A Complete Bibliography of Publications in the Journal of Functional Programming

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

27 May 2024 Version 1.09

Title word cross-reference

$37.99 [Rag13]. $60.00 [Mil13]. $70 [Gam09]. $AU [Tho04a]. λ [AMR09, BFG97, BS96, Cro99, CH94b, CD96, DM05, GK96, Gol00, HD97, Hue94, KR97, Kam05, ØP97]. λν [BBLRD96]. λs_{e} [Gui00]. ω [HL96]. p [JJ23]. Π [KN96, DG02, GK96, Jef01, RH03].

-calculi [Kam05]. -calculus [AMR09, Cro99, CD96, DG02, DM05, GK96, Gol00, Gui00, HD97, Hue94, KR97, ØP97, RH03, Jef01]. -complexity [JJ23]. -conversion [KN96]. -cube [BFG97]. -RED [GK96].

//www.pragprog.com [Hag11].

0 [Bha09, Boi03, Boi12, Bud08, Ceb11, Ceb12, Der04, Hut94, Jon92b, Kar03, Kel03, Klu01, McK99, Mit95, San06, Sim07, Tho02, Tho04b, Vil07, dP01, vS00]. 0-13-184854-2 [Fou96]. 0-13-288788-6 [Fou96]. 0-13-722075-8 [Fou96].

0-19-511306-3 [Tho02]. 0-201-39820-6 [Klu01]. 0-201-59604-0 [Rus03]. 0-262-06217-8 [Tah03a]. 0-262-19455-4 [Kar03]. 0-262-22069-5 [Gam09]. 0-262-51087-1 [Woo01]. 0-333-99285-7 [Han04]. 0-387-95349-3 [Der04]. 0-387-95391-4 [Tho04b]. 0-471-93844-0 [Mit95]. 0-471-99886-9 [Kel03]. 0-521-39022-2 [Fou96]. 0-521-41497-0 [Jon92b]. 0-521-44189-7 [Hut94]. 0-521-45520-0 [Klo01]. 0-521-48089-2 [O’D01]. 0-521-54310-X [Bud08]. 0-521-57681-4 [Har99b]. 0-521-58775-1 [McK99, dP01]. 0-521-61410-4 [Ceb11]. 0-521-63168-8 [Kah98]. 0-521-64408-9 [vS00]. 0-521-69269-5 [Bha09]. 0-521-74758-9 [Ceb12]. 0-521-78098-5 [Boi03]. 0-521-89556-1 [Boi12]. 0-9543006-3-7 [Sim07]. 0-9543006-9-6 [San06]. 0-9576711-0-5 [Rag13].

1 [Boi12, De 11, Hud15, Jad14, Jon07, MF08, McK99, Mic07, Pat14, San09, Slo09, dP01]. 1-107-02957-0 [Wat14]. 1-4039-0772-2 [Han04]. 1-4302-6250-8 [Gre14]. 1-4398-9666-6 [Pat14]. 1-59059-757-5 [De 11]. 1-59327-283-9 [Mic13]. 1-59327-491-2 [Jad14]. 1-74009-404-2 [Tho04a]. 1-934356-00-X [San09]. 150 [Tho04a]. 1994 [Gor96]. 1995 [Bai99, Bur99, Pag97, Pag98, Ric98]. 1996 [App97a, Rea97, Ten97, Tho99]. 1997 [Jef98, Oka99a]. 1998 [Han01, Ten00]. 1999 [Jef01, Lin00]. 1ML [Ros18].

2 [Jad14]. 2.0 [BR04]. 2000 [Rea04]. 2001 [Hai03, Tah03a]. 2002 [Ang03, Sha06]. 2003 [Kah05, Lou04, Sha04a]. 2004 [Sha04b]. 2008 [Chi09, TN10]. 2009 [TL11]. 2014 [CS16, DS16, Gru15]. 2015 [Tan20]. 2018 [Rag20]. 2019 [Hac20]. 2020 [Gru20]. 20652 [Cas13]. 28-30081991 [Joh92]. 2nd [Rea97, Tah03a, Woo01].

3 [Aba94, BAD^{+}01, Der04, Kin07, Mic07, Tho02, Vil07]. 3-4 [Ano09d, Ano09c, Ano10f, Ano10e]. 3-540-00296-0 [Vil07]. 3-540-20234-X [Kin07]. 3-540-40174-1 [Mic07]. 3-540-43092-X [Wal04]. 359 [Wal04]. 3rd [Tho02].

4 [Car11, Ceb11, Cla15, Kar03, Spe14, Tho04b]. 4th [KR02].

5 [Bha09, Boi03, De 11]. 5-6 [Ano10h, Ano10g]. 50-year [HB17].

6 [Hag11, Klu01, Pat14, San06].

7 [Hut94, Sim07, Tho16].

8 [Spe20].

9 [Ceb12, Kel03, vS00]. ’91 [Joh92]. 97 [Cas13]. 97-81-4493-20652 [Cas13]. 978 [Car11, Cla15, Hag11, Hud15, Slo09, Spe14, Spe20, Tho16].

978-0-262-06279-4 [Chi09]. 978-0-262-10114-1 [Slo09]. 978-0-262-52664-7 [Tho16]. 978-0-521-51338-8 [Mil13]. 978-0-521-89957-4 [Car11]. 978-1-107-03660-4 [Spe14]. 978-1-4493-1103-2 [Bro12]. 978-1-4665-0455-4 [Cla15]. 978-1-59327-283-8 [Mic13]. 978-1-59327-435-1 [Hud15]. 978-1-617-29224-8 [Spe20]. 978-1-84882-617-5 [Kol12]. 978-1-934356-45-6 [Hag11].

= [MNVS99, THLJ98].

A-translation [CH94a]. Abelson [Woo01]. Abrial [Boi12]. Abstract [Bou92, GAM18, Har93a, Bur91b, BC93, DRR01, DP12, GPN19, JSE^{+}14, Ses97, SR98, VPD21, VM12, WW08, Mit95]. Abstraction [Bro12, Sch03a, VM12]. Abstractions [Slo09, Jac06]. Abstracts [Hut14, Hut15a, Hut15b, Hut16a, Hut16b, Hut17a, Hut17b, Hut18, Hut19a, Hut19b, Hut20a, Hut20b, Hut21a, Hut21b, Hut22a, Hut22b, Hut22c, Hut23a, Hut23b, Hut15c]. access [BN05, Swi20]. accesses [BKP01]. accidents [DD03b]. account [Fil21, SSTH18]. accumulating [IB99, KN08, VK04]. accumulative [MKHT12]. accumulators [Dan23b]. acids [TLM95]. across [DM14]. act [Ada93, Hin03]. action [BP96]. activations [SHMJ16]. actor [AMST97]. actors [CGJF20]. Acute [SLW^{+}07]. ad [GZND13]. Ada [TO98]. Addis [Kol12]. Addison [Hag12, Klu01, Ric98, Rus03]. addition [OR04]. adequacy [Bre18]. adequate [Gol00, RH03]. adjusting [ABD13, CDHA14]. administrative [DD03a]. Advanced [Bur99, Sha06, Tho99]. advice [BJI17]. again [STW21]. Agda [AAS17, MKJ09, Swi11, VMA21]. Agility [Bro12]. Alejandro [Gre14]. Algebra [MKJ09, Feg17, HMGW96, Spe24, PT99]. Algebra-Driven [Spe24]. Algebraic [KN08, SW17, BFG97, KP17, LP20, Mor21, STI24, Sel02]. algebraic- [BFG97]. Algebras [Spi09]. Algol [O’H96]. Algorithm [Bir10, Gru20, Mil13, THLJ98, AFB92, ABR97, AD15, Bir13, Bir20, BM21]. Algorithmic [NF94, Oka01, SDF^{+}09, SHMB05]. Algorithms [App97a, BCMT01, Edm08, Sch04, Bir92a, Bir92b, Bir01a, Erw01, Joh98, LP99, Oka97, RL99, Rag20, Ceb11, Rus03]. aliasing [BHY07]. Alice [GHM24]. allocation [GAM18]. almost [BR08]. alone [KMS21]. alternative [Ach21]. always [SH12]. Amadio [dP01]. amortized [HS15]. Analogical [Sch03a]. analyses [BJ93, GAM18, PS05]. analysis [AA02, BFGJ04, Ben01, BHY07, BL96, CGMS06, CKP05, DG02, DD03b, DHW05, EM95, FS03, Han98, HS15, HY05, Jac06, JSE^{+}14, MS08, Pal93, QGG21, RH03, SSSS08, SVJB17, Wan93, Slo09]. analytic [AS96]. analytics [Feg17]. Andrew [Kah98, McK99]. ANF [KRAB22]. Animated [MJ98a]. animation [Ary94, Tho00]. Annie [Spe14]. antiquote [Hin11]. Antoni [Gor96]. anyone [Oka98]. Apostolos [Ceb12]. Appel [McK99]. application [Bru92, FF99, Hin23, Hue05, Jon92a, Swi22]. Applications [Cas13, HP93, SB99, TW00, Tah03b, HP95a, Kes12, Lis94, Pag97, Pag98].

Applicative [MW18, MP08]. applied [GK96]. apply [MP06]. Applying [BSZL^{+}21]. Approach [Har99b, CCP01, CM98, GAM18, Men14, MS01, RL99, SJ92, VYH^{+}10, Xi07, Rus03]. Apress [De 11, Gre14]. April [Mic13]. arbitrary [PVWS07]. arbitrary-rank [PVWS07]. Architecture [Joh92]. Armstrong [Dup95, San09]. array [Gre05, Sch03b, WC01]. Arrays [Pey03h, HS15, Joh98, MBG^{+}20, OB97]. arrow [AVDP13, LWY10]. arrow-based [AVDP13]. arrows [JHH09, LCH11]. Art [HP93, HP95a, Lew13, Pat14]. Arvind [Hai03]. Aspects [Wal04, AHK08, KR02, Oka01, Ram03]. Asperti [Mic00]. assembly [MCGW02, MCGW03]. assertions [BCG05]. Assignment [Sch03b, Di 93]. assistant [Chl08, Dan22, Dan23a]. associated [Gol00]. associative [KM01]. associative-commutative [KM01]. assumptions [VPSS11]. Asymptotic [HLL24]. asynchronous [GV10]. augment [GJ07]. Author [Ano91a, Ano92a, Ano93a, Ano94a, Ano95a, Ano96a]. authoring [GS05]. AutoBayes [FS03]. Automata [Kri06]. Automated [Ben01, Car11, GKM^{+}08, Har09b]. Automatic [App97a, DG00, IB99, Lap98]. automation [GZND13]. avoidability [BJI17]. avoidance [Cra20]. aware [SZJ14]. axiomatic [AGNS21]. axioms [CD18].

B [Boi12, Lou04, Abr10]. Baader [Klo01]. Baby [Aba94]. Back [Ano91d, Ano91f, Ano91h, Ano91j, Ano92b, Ano92d, Ano92f, Ano92h, Ano93b, Ano93d, Ano93f, Ano93h, Ano94b, Ano94d, Ano94f, Ano94h, Ano95b, Ano95d, Ano95f, Ano95h, Ano96b, Ano96d, Ano96f, Ano96h, Ano96j, Ano96l, Ano09a, Ano09c, Ano09e, Ano09g, Ano09i, Ano10a, Ano10c, Ano10e, Ano10g, Ano11a, Ano11c, Ano11e, Ano11g, Ano11i, Ano12a, Ano12c, Ano12e, Ano12g, Ano12i, Ano13a, Ano13c, Ano13e, Ano13g, Ano13i, Ano13k, Ano14a, Ano14c, Ano14e, Ano14g, Ano14i, HM96, PP22]. backend [TMK^{+}19]. Backhouse [Sha04b]. Backtracking [PS17]. backward [BB21]. balanced [FL11, HY11, MC21]. Balancing [HY11, Ada93]. bananas [WW08]. Baral [Sha04a]. Barendregt [KN96]. bargain [Nee05]. Barron [Kel03]. Barski [Jad14]. base [LP09]. Based [AMS05, AVDP13, BN05, Bla18, BL96, CG96, Cla96, HMM18, Han98, HS15, KSK08, LG05, Mac94, Men14, Mor18, MCGW02, MCGW03, SHMB05, SFL17, SDK18, SGM^{+}03, TN15, ZPV06]. Basic [McK99, Pey03y, App97b, IRK01]. Basics [Tho16, HM96]. Basis [Jon06, CL91, Cla96]. Batcher [HM18a]. be [Gre96]. bee [CLMM20, Ser00]. Beginner [Mic13, Hud15, Lip12, TUmI15]. Beginning [Men14, Whi13, Rag13, Gre14]. behaviour [ANN97]. behind [CA20]. being [DP22, Nil01]. Ben [Gor96]. benchmark [HSB97, HFA^{+}96]. Benchmarking [HFA^{+}96, MR01]. benefit [AH08]. benefits [CK04]. Berger [Dyc01]. Better [Bro12]. between [LTJ94]. Bibliography [Pey03-28]. Bice [Jad14]. bidirectional [MW18, ZTX^{+}23]. bidirectionalization [MW18, VHMW13]. bidirectionalizer [VHMW13]. Big [AC09, CG18, NH09, Feg17]. Big-step [AC09, NH09]. binary [Gol00, GL15, GS06, Hin02, KV12, Pad17, Swi20]. binder [Kam05]. binders [LP13, PP12]. binding [AAC^{+}21, Pal93, SW16, Wan93]. binding-safe [SW16]. binding-time [Pal93, Wan93]. Bindings [Pey03v, FCDF12]. binomial [Hin99b]. Biological [Wak06]. Bird [Gru15, Gru20, Hin09a, Mil13, PT99]. bisimulation [FHJ98]. bit [Hin00b, KV12]. bit-reversal [Hin00b]. black [GM14, Kah01, Oka99b]. Blame [STW21, SCYZ24]. Bob [GHM24]. Böhm [Tro95]. Böhm-out [Tro95]. Bologna [AGN96]. Book [Ang03, App97a, Bai99, Bha09, Boi03, Boi12, Bro12, Bud08, Bur99, Car11, Cas13, Ceb11, Ceb12, Chi09, De 11, Der04, Dup95, Dyc01, Fou96, Gam09, Gam11a, Gor96, Gre14, Gru15, Hag11, Hag12, Hai03, Han04, Han01, Har99b, Hin99a, Hut94, Jad14, Jef98, Jef01, Jon92b, Jon06, Jon07, Kah98, Kah05, Kar03, Kel03, Kin07, Klo01, Klu01, Kol12, Lin00, Lou04, McK99, Mic00, Mic07, Mic13, Mil13, Mit95, O’D01, Oka99a, Orc14, Pag97, Pag98, Pat14, PT99, Rag13, Rea97, Rea04, Ric98, Rus03, San09, San06, Sha04a, Sha04b, Sha06, Sim07, Slo09, Spe14, Tah03a, Ten97, Ten00, Tho02, Tho04a, Tho04b, Tho99, Vil07, Wal04, Wat14, Woo01, dP01, vS00]. Bookreview [Hag12]. Books [Ang03]. Bookshelf [Hag11, San09]. Boolean [JJ23]. Bosworth [Rea97]. boundaries [BAD^{+}01]. bounded [SC09]. bounds [AGLK20]. Boxes [WW08]. bracketed [SDB21]. Braun [Oka97]. breadth [Spi00]. breadth-first [Spi00]. Bristol [Ang03]. Brother [Hin09b]. Brouwer [Ste21]. Bruce [Tho02]. Bruijn [BP99]. Build [MMJ20]. Building [Bir11, FS08, Kes12, Kre04, Bir97a, Cas13]. Bulterman [Kin07]. bunched [O’H03]. Burrows [BM04]. business [Mol93]. busy [Nil01, SJCA17].

C [Boi03, Der04, Fou96, Kin07, Pat14, Rea97, Sha04a, Sha04b, Ten97, Ten00, Tho04a, Wal04, BJZS10, DP09, HMG04, RA19, Sch03b, Waz10]. C-Rules [Waz10]. CA [Jad14]. caching [BKP01]. CakeML [TMK^{+}19]. Calculating [BH15, BH20, CM15, MG06, Mee04, Sha04b]. calculation [ZTX^{+}23]. calculi [CC19, HM94, IN97, Kam05, SC21, Vas05, dP01]. Calculus [WC03, AZ02, AMR09, AF97, BB96, Bar91b, Bar92b, BG00, Bar91c, BBLRD96, Ben95, Cou07, Cro99, CD96, DG02, DM05, DA18, FD04, GPN19, GK96, GN91, Gol00, GJ91, GL15, Gui00, HMP98, HD97, Hin10a, Hof99, Hue94, KN95, KR97, LWY10, Mai04, MOW98, Mog92, Mor21, ØP97, PI11, RH03, Sel02, Stu17, Sur07, Swi23, WDMT02, Jef01, Gam11a]. Call [BDT01, JFDA17, MK97, Mel97, AF97, MOW98, MKHT12, NH09, SF06]. Call-by-name [JFDA17]. call-by-need [AF97, MOW98, NH09]. call-time [MKHT12]. CAM [Vas05]. Cambridge [Bha09, Boi03, Boi12, Car11, Ceb11, Ceb12, Fou96, Gru15, Gru20, Har99b, Hin99a, Hut94, Jef01, Joh92, Jon92b, Jon06, Jon07, Kah98, Kah05, Klo01, McK99, Mic00, Mil13, O’D01, Rea97, Sha04a, Spe14, Ten00, Wat14, vS00]. Caml [MW08, Lou04]. can [Gre96]. Canonical [KN96]. capabilities [SDB21, VPD21]. Capability [BSO20, SFL17, Pot13]. Capability-based [SFL17]. Capability-passing [BSO20]. cardinality [MS08, SVJB17]. Carl [Tan20]. Carlo [HLB95]. carry [OR04]. carte [MMJ20, Swi08]. case [BM05, BRS94]. cast [SC21, Wei04]. catch [Cro99]. catch/throw [Cro99].

Categorical [JHH09, Asp92]. categorically [HD00]. Causal [LCH11]. celebrities [BC06b]. centric [IEM^{+}22]. Certified [Hin23, Chl08]. certifying [OCR^{+}21]. Chailloux [Lou04]. Chakravarty [Tho04a]. changers [HM96]. channel [PI09]. channels [PI09]. Chapman [Pat14]. Character [Pey03k]. characterization [Par16]. characterizations [IRK01]. chart [Lju04]. checking [Bar99, Bet00, DO02, JM00]. checkpointing [ZJ10]. chessboard [Bir04]. choice [BB96]. choices [AH04]. choreographies [GHM24]. Chris [Fou96, Hin99a]. Christina [Orc14]. Christopher [Tah03a]. Christos [Ceb12]. Church [Hin05]. CIC [ZS17]. circuit [OR04]. City [Gre14]. Clack [Fou96]. Clarity [Spe14, Liu13]. class [HDM93, Hin03, JK09]. Classes [NP95, Pey03x, BJZS10, DO02, Jon92a, Jon95, Läu96]. Classical [DA23, Kar03, SR98, DA18, SWM01]. classification [CCP01]. clean [DRvE95, AP95]. client [CC19]. client-server [CC19]. Closed [CMS03, SFG14, SB99]. closures [AS96]. CLP [SS08]. CMC [LTJ94]. CML [FHJ98]. co [CA20, DA23, Gib21]. co-programs [Gib21]. COCHIS [SOWM19]. code [AD15, BSZL^{+}21, CX05, DG00, HLB95, RA19, SGM^{+}03, VPD21]. coercion [STW21]. coercions [BEJW16]. Cogent [OCR^{+}21]. cognitively [TI15]. Coherence [Sch09, BKP01, Hed98]. Coherent [Rag13, SOWM19]. coin [Bir19]. Colin [Fou96, Pag97, Pag98, Rea04]. Collard [Tho04b]. Collection [App97a, EH21, JSE^{+}14]. collections [PTW22]. College [San06, Sim07]. Column [Fel03, PW03b]. combinator [Str11]. combinatorial [Spi09]. Combinators [Gam11a, Hil96, Spi00, Thi99, AS21, Bou99, BD97, CH94a, Ken04, NFFM17, Oka03, PW96, Rhi09]. combinatory [BD97, Swi23]. Commercial [FSNB09, SM13, HNSSH97]. Communicating [Jef01]. Communication [VH95, CE18, Hay00, SSTH18]. communications [BCG05]. commutative [KM01, LCH11]. comparative [GJL^{+}07]. Comparing [HLB95, OG10]. comparison [BJZS10]. competitive [ABR97]. Compilation [FM00, GHL99, OCR^{+}21, SM21, VPD21, Wak98]. Compile [FCDF12, Hin11]. Compile-time [FCDF12, Hin11]. compiled [AHN12]. Compiler [McK99, Pey03c, App97b, BL96, Cle16, KM01, RA19, SWD05, SHMB05, TMK^{+}19, PM02]. compilers [BH15, BH20, BP96]. Compiling [EFD03, HB17, SGM^{+}03, Wak99, GK96, FBH17]. complete [BM06, BD97]. completeness [Mai04]. Complex [Pey03e]. Complexity [MK97, Oka99a, Ben01, EHBR14, HM94, JJ23]. components [RT91]. Composable [Pet21, SHMJ16]. Composing [Jon04, FS02]. Composition [VK04, DHLB16]. compositional [BR08]. comprehensible [ZS17]. comprehensions [Rea93]. compression [MC13]. Computability [Oka99a]. Computation [BRL17, ABD13, AMST97, Atk09, CDHA14, HMM18, MW24, MS03, NP05, Nie00, RJ17, SLW^{+}07, SMT16, VH95]. Computational [BBD98, MK97, BJI17, CGR96, DA14, DA18, Dyc01, Orc14]. computations [MKHT12]. Computer [Bud08, Bur99, Gam09, Hic04, Joh92, Jon92b, Tho99, Woo01, ASS96, ASS00, FFFK04, Har93b, HR04, VH04]. computers [FM00].

Computing [Jon92a, Rea04, Tho04a, CK02, Cla15]. Concepts [Mit03, VH04, BJZS10, Gam09, Boi03]. Concrete [Hin10a]. Concurrency [Bro12, Cla99, Jef98]. Concurrent [AH04, KH18, Rep99, San09, ANN97, Arm07, BCG05, JKJ^{+}18, Nie00, SM21, ZJ10, Mar02, Dup95, O’D01]. conference [Joh92]. confinement [PNCB06, ZPV06]. confluence [AMR09, JFDA17, Nie00]. confluent [Cro99, CD96, KR97]. conformations [TLM95]. Connections [DV19]. conquer [HL99]. Conrad [Jad14]. CONS [Jon01]. consensus [Bir11]. conservatively [O’H96]. Consistency [BHM^{+}06]. consistent [ABD13]. constant [HW07, Stä98]. constrained [CDP15, HG96]. Constraint [NT01, MLC91, SSW09, SDPJS07]. Constraints [Han01, Nar07, Sur07, WC01]. Constructed [BFGJ04]. constructing [Hin18]. Construction [Sha04b, Swi23]. Constructions [WC03, BB96, GN91]. Constructive [DV19, DD95]. constructively [UV17]. constructor [Jon95]. constructors [AMR09]. Container [HD00]. containers [AGH^{+}15]. context [BC15, GM04, Lju04]. context-free [BC15, Lju04]. context-sensitive [GM04]. contexts [BHM^{+}06, FCDF12]. Contextual [LP13]. continuation [SR98]. continuations [Dan23b, DPS07, FS02, HDM93, HLA20]. Contract [ZTX^{+}23, NTHV17, SSTH18]. Contracts [GPW12, BM06, MF23]. Contributions [BJI17]. Control [AH08, Thi09, BN05, BD06, DA14, DNB12, GAM18, HL96, KKS11, SDB21, SMMR17]. control-flow [GAM18]. controlling [ACR16]. conversion [KN96, SH12, Swi23]. converting [ZMCN18]. copatterns [AP16]. coq [BSZL^{+}21, AMNS22, CLB23, Dan22, Dan23a, ZDK^{+}15]. Core [FHJ98, Ros18]. coroutine [PS20]. correct [BH15, BH20, EAKC06, Gam11b, Joh03, Swi23]. correct-by-construction [Swi23]. Correctness [FBH17, Pag07, Pal93, AF04, BB21, BL96, CGMS06, CK95, FH03, Wan93]. Correspondence [BCG05]. corresponding [GPN19]. Corrigendum [Har09a]. cost [AS96, BEJW16, DL22, QGG21]. countdown [Hut02a, BM05]. counterexample [CH94b]. Counting [GL13, GL15, CGR96]. counts [KV12]. Course [Rea97]. Cousineau [Har99b]. Cover [Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano92c, Ano92b, Ano92e, Ano92d, Ano92g, Ano92f, Ano92i, Ano92h, Ano93c, Ano93b, Ano93e, Ano93d, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95e, Ano95d, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96i, Ano96h, Ano96k, Ano96j, Ano96m, Ano96l, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano09j, Ano09i, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano11j, Ano11i, Ano12b, Ano12a, Ano12d, Ano12c]. Cover [Ano12f, Ano12e, Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i]. covering [Bir11]. CPS [DD03a, DD03b, DMN07, FS02, PW03a]. CPU [Pey03s]. Craft [Ric98]. CRAY [BH92]. CRC [Cla15, Pat14]. Create [Hag11, Par10]. critique [App93]. CS1 [Mor18]. cube [BFG97, Jon13, KN96]. Cubical [VMA21]. cubism [Jac96]. CUFP [MME12]. CUFP’13 [ESM15]. Cuneiform [BRL17]. Curien [dP01]. curriculum [FFFK04]. curry [MP06]. curse [GM14]. Custom [SFG14]. cut [BG00, GJ07, Joh03, PS17].

D [Fou96, Kah98, Oka99a]. Daniel [Chi09, Slo09, Tah03a, Tan20]. d’applications [Lou04]. Data [CG18, Hin99a, Mit95, Swi08, All05, AJ15, Bur91b, BC93, CDP15, Col03, FS03, GS06, Har93a, HP06, KV12, MR01, NO99, NS93, Pet21, SW17, SNS16, Tro95, Feg17, Tho04b]. data-parallel [NO99]. dataspace [CGJF20]. datatype [BP99, VYH^{+}10]. datatype-generic [VYH^{+}10]. datatypes [Hin01, Mat09]. dated [PT99]. Dates [Pey03q]. David [Gor96, Jad14, Kel03, Pag97, Pag98, Rag20]. Dean [Bro12]. Debasish [Spe20]. Debugger [TA95, Nil01]. debugging [Har99a, Har09a, NF94, Yi06]. December [Wat14]. Deciding [CCG05, Di 93]. decision [TN15]. Declarations [Pey03v]. Declarative [Sha04a, Wal04, KR02]. Dedicated [AP13, SF14, TN10, TL11, ACW12, DS16]. Deduction [LS96, BG00]. defined [MBG^{+}20, NO99]. Defining [Tro95]. definition [FCDF12]. definitional [HP99]. definitions [LP13]. Deletion [GM14]. Delimited [DA14, BD06, DPS07, KKS11]. DeMaio [Jad14]. Demand [GBF05, HK05, LB96]. Denotational [DL22]. densest [CM15]. densest-segment [CM15]. dependencies [SDPJS07]. Dependent [AF04, BDT01, BDT04, CD18, MKJ09, SD17, Xi07, BJP12, Bet00, CDP16, CA20, DTT18, FI00, HJP23, KRAB22, McB02, Stu17, SCF^{+}13]. dependently [Bra13, VMA21]. deques [Oka95]. Derivation [OR04, DP09, Gib99, MKJ09]. derivatives [ORT09, TN15]. Derived [Pey03b]. Deriving [Gib96a, HM96, Ses97]. description [HW97, PI11]. Design [Bir10, Bra13, Bru94, Gru20, Spe14, Tho02, Ber93, Bru92, Gib21, GMO06, Liu13, LW99, Mac99, SLW^{+}07, She11, Str11, SNS16, Spe24, Mil13]. design-space [SNS16]. destruct [SSH^{+}08]. destruct-time [SSH^{+}08]. destructive [WC01]. determinacy [Bur91b]. determinate [Bur91b]. determinism [PS05]. deterministic [KM01]. developed [AGLK20]. Developers [Bro12, Wam11]. Developing [GBF05, Mar02]. Development [De 20, Chl08, Hue94, Kre04, Ser00]. Développement [Lou04]. devices [BR04, Kin07]. dialects [BLH16]. Dick [Kin07]. dictionary [SW23]. dictionary-passing [SW23]. Diets [Erw98, FL11]. digital [Bru92]. digits [BM21]. Diller [Gor96]. dimensions [RR96]. Direct [Dan23b]. directed [Fil21, Har99a, Har09a, KH95, SW23, Yi06]. Directions [Lin00]. Directory [Pey03o]. disciplinary [LG05]. discrimination [Hen12]. Dissertations [Jon92b]. distinct [AH04]. Distinguished [Jon92b]. Distributed [BP92, GBF05, Hay00, TLP02, ABR97, AS21, DHW05, Feg17, HMM18, Mor18, SFL17, SLW^{+}07, SCF^{+}13]. distribution [CDP15, DLP10]. distributive [PS17]. divide [HL99]. divide-and-conquer [HL99]. Do [FI00]. documents [Thi02]. does [Gui00]. Doets [San06]. Domain [BS00, Hag11, Hag12, MLC91, Spe20, IEM^{+}22, Par10, SJ92, WE09]. Domain-free [BS00]. Domain-Specific [Hag11, Hag12, IEM^{+}22, Par10, WE09]. domain-theoretic [SJ92]. domains [CG97, dP01]. Doo [CLMM20]. Double [Hut02b]. down [Hen12]. Drawing [AA10, Ken96, Kol12]. drawings [Gib96a]. Driven [De 20, Spe24]. dropping [BM21]. DrScheme [FCF^{+}02]. duality [LP09]. duplicates [Bir91a]. Dutch [Swi11]. Dynamic [ACPR95, App97a, Lap98, SJW18, SSH^{+}08, Bou92, EK21, FBH17, MG06, She11, SMMR17, SV13, Wak98]. Dynamics [LM93].

eager [BJD97]. Eastlund [Tan20]. easy [Hut23c]. Ed [Joh92, Woo01, Jef98, Tah03a, Tho02]. Eden [DHLB16, LOMPM05]. edited [Bur99, Han04, Kah98, Pag98, Tho99]. Edition [Gam11a, Bud08, Chi09, Gor96, Rea97]. editor [AVDP13, Ang03]. editor-based [AVDP13]. EditorArrow [AVDP13]. Editorial [Ano91b, AP13, BHH^{+}93, Bir99, CG18, DS16, Fel12, Fel14, Gib10, GK22, Hin04, Hin10b, HM04, KH18, Law08, LC01, Ler09, LF09, SF14, SA14, SD17, TN10, TL11, PW98, Wad97, ACW12]. editors [Ang03, Bai99, Der04, Dyc01, Jon06, Lin00, Pag97]. Edmonds [Ceb11]. eds [Hut94, Rea04, Wal04]. education [Kri21, SWD05, TW93, Bai99]. Educational [Fel03, Kri06, PW03b, SWD05, Wak06, Yi06]. Effect [HLA20, BGT16, BSO20, HLL24, KPS^{+}20, LP20, TJ92]. effect-safe [BSO20]. Effective [SNO^{+}10, TI15]. effectiveness [HSB97]. effects [AJ15, DA14, DNB12, KP17, KMS21, MT91, MP08, MS01, PM00, PRS20, STI24, SSV08, Thi09]. Effekt [BSO20]. Efficiency [SF16, Liu13, Spe14]. Efficient [Ada93, BC15, BJ93, GKM^{+}08, GS06, Joh98, Mog92, Fil21, Gol00, Hoo92, Oka95, Sch03b]. Eijck [Orc14, San06]. Elaborating [CA20, Dun14]. Elastic [MBG^{+}20]. Elements [Ull94, Fou96]. Eliminating [CDP16]. elimination [BG00, DG00, Mai92]. eliminations [Stu17]. Ellen [Fou96]. Embedded [Ben05, SFG14, EFD03]. Embedding [Ram11, RH03]. Emden [Jon06]. Empirical [AS96]. Encapsulating [Bur91b]. encapsulation [MS01, SDB21]. Encoding [WW08, Hin23]. encodings [SF16]. Enforcing [SDB21]. engine [SFG14]. Engineering [Pag07, Abr10, Boi12]. engines [SDK18]. English [Wat14]. Enhancing [VHMW13]. Enno [Sha06]. enough [Spi12]. enriched [LP09]. ensure [AF04]. enter [MP06]. Enumerating [GLB06, McI04]. enumeration [NFFM17]. Enumerators [Bar92a]. environment [Asp92, FCF^{+}02, GM93, GKM^{+}08, HNSSH97, Hut94, RTF93, Ser00]. equality [ACD^{+}20, Ada06, BBJR21, EHBR14, KRAB22, VW10]. Equals [KRRS97]. equational [BC06a]. equations [KMS21]. Equivalence [MT91, Bie00, GHL99, LP13, LTJ94]. erasure [CWM02]. Eratosthenes [Mee04, Nyk11, O’N09]. Eric [Jad14]. Erik [Bur99, Tho99]. Erlang [Hud15, San09, Arm07, Kes12, Cas13, Dup95]. Errataum [Ano91c]. Erratum [Dan23a, Jon93]. errors [CE18, PW96, SJW18]. Escape [Erw04, Han98]. Essays [Rea04]. essence [Dag17, GdSO09]. Essentials [FWH01, FW08, Chi09, Tah03a]. eta [JG95]. eta-expansion [JG95]. eval [MP06]. eval/apply [MP06]. evaluate [GTN^{+}19]. evaluated [CKS09, NR12]. Evaluation [Sch15, Tho02, AHN12, BJD97, BJ93, BP96, Bre18, Bur91a, CK95, DRR01, HJP23, Lau91, Mac99, PM00, Jon92b]. evaluator [GJ91]. Even [Oka98, HM18a]. Event [Boi12, Abr10]. Event-B [Boi12, Abr10]. events [DF08]. ever [Nil01, Oka98, Pie13]. Every [KV12]. everything [Pie13]. examined [ORT09]. example [TI15]. Exceptional [BK01]. exceptions [DM05, SMMR17]. excluded [BB96]. excluded-middle [BB96]. execution [ABR97, AS21, GKM^{+}08, KH95, NTHV17, Thi03]. exercise [Erw04, Lüt03, MC21]. existential [Läu96]. exotic [FSNB09]. expanding [Bir13]. Expansion [Pol98, FCDF12, JG95]. Experience [Waz10, HMG04]. experiences [MW08]. experimental [WE09]. Expert [HNSSH97]. Explainable [EK21]. Explaining [Hin99b]. Explicit [ACCL91, KPS^{+}20, BBLRD96, FP15, KR97]. Exploiting [MS08]. exploration [SNS16]. exploratory [RTF93]. explore [BAD^{+}01]. Exploring [TLM95]. expressed [CG96]. Expression [vS00, Fil21, Hos06, HP03, Hud00, ORT09, Jon07]. Expressions [Pey03u, AJ97, Chi94, Hil96, PM00, TN15]. expressive [Jon01]. expressiveness [Hut99]. Extended [Han93, HW16, Mor21, GJL^{+}07, Hin10a, MS01, PG20, SSTH18]. Extending [KR97, O’H96]. extensible [BSO20, KMS21]. extension [HL96]. Extensional [BBJR21, KRAB22]. extensionality [JFDA17]. extensions [Tho16]. extracted [Ben01]. Extracting [AMNS22]. extraction [DL22].

F [HL96, RRD14]. F# [Pic07, De 11]. F-ing [RRD14]. Factorisations [Jon92b, Lau91]. Factorising [HJG10]. factorizable [BS23]. Fair [NFFM17]. fake [Bir19]. Faking [McB02]. family [SIV08]. fast [KRRS97, MP06]. Faster [PS20, HJG10]. fat [Erw98]. Fault [PG20, SMT16]. Fault-tolerant [PG20]. FC [MS04]. feasible [IRK01]. Featherweight [PNCB06, SW23]. features [HSB97]. Felleisen [Jad14]. Feng [Jad14]. Feng-Yun [Jad14]. Fernández [Sim07]. Fethi [Rus03]. fifth [Joh92]. filters [Hos06]. Finally [CKS09, DP12]. find [Bir19]. Finding [BC06b, CG97]. Finger [HP06]. Finite [Sch03a, CG97]. Finiteness [UV17]. First [JK09, Mcb03, RW97, BAD^{+}01, CK04, CIZ22, HDM93, JVV93, LLR93, STW21, Spi00, Yi06]. First-class [JK09, HDM93]. First-order [Mcb03, RW97, BAD^{+}01, Yi06]. first-year [JVV93]. fix [SH14]. Fixed [HH91, CG97, VH95]. Flag [Swi11]. Flagship [Kea94]. Flattening [Oka03]. Flemming [Jef98]. Flexible [SMMR17, SW16]. float [HFA^{+}96]. float-intensive [HFA^{+}96]. FLOPS [CS16]. Florence [Jad14]. Flow [Tho04b, BN05, Col03, CKP05, DD03a, GAM18, HY05, JSE^{+}14, MS08, PP01, PW03a, SDB21, SMMR17, WDMT02]. focused [Cra20]. Fold [Dan22, Dan23a, Hut99, SH14]. Folding [Dan19, Dan23b, Sch03a, Zhu94]. folding/unfolding [Zhu94]. folds [HJG10]. folklore [BD06]. forest [FP03]. Formal [Cla96, Gor96, Sch04, VDH10, Bru92, DP09, Hue94, LP03]. Formalisation [AD15]. format [TI15]. Formatting [Hin03]. Forms [BAD^{+}01]. Forms/3 [BAD^{+}01]. formula [HW97]. Forrest [Jad14]. Fortifying [Cul12]. foundation [AMST97, DL22, HMM18, JKJ^{+}18]. Foundations [DTT18, Pic07, Rea04, Wat14, Har13, PT94, De 11, Ten97]. founded [AP16]. four [PW96]. Fowler [Hag12]. FPCA [Joh92]. Fractal [MC13]. fractals [Jon04]. fragment [HPR98]. Framework [Ran04, Di 93, DPS07, HL07, HP95b, LG05, Luo03, SWD05]. Frameworks [DH00, Hue03]. Fran [Tho00]. Francisco [Jad14]. Francois [Tho04b]. Franz [Klo01]. Fred [Hud15]. free [BS00, BJP12, BC15, EH21, Lju04, Luo03]. Fregel [IEM^{+}22]. fresh [Hin02]. Friedman [Chi09, Tah03a, Tan20]. Front [Ano91e, Ano91g, Ano91i, Ano91k, Ano92c, Ano92e, Ano92g, Ano92i, Ano93c, Ano93e, Ano93g, Ano93i, Ano94c, Ano94e, Ano94g, Ano94i, Ano95c, Ano95e, Ano95g, Ano95i, Ano96c, Ano96e, Ano96g, Ano96i, Ano96k, Ano96m, Ano09b, Ano09d, Ano09f, Ano09h, Ano09j, Ano10b, Ano10d, Ano10f, Ano10h, Ano11b, Ano11d, Ano11f, Ano11h, Ano11j, Ano12b, Ano12d, Ano12f, Ano12h, Ano12j, Ano13b, Ano13d, Ano13f, Ano13h, Ano13j, Ano13l, Ano14b, Ano14d, Ano14f, Ano14h, Ano14j]. frontiers [HH91]. FRP [Bah22]. FT [HSB97]. full [GPN19]. full-reducing [GPN19]. fully [AGLK20, RH03, VPD21]. Fun [Ach11, Han04, GdM03]. function [Oka98, Rit91, SF06, Swi22]. Functional [Ada93, Ang03, AH04, ARS94, BLS04, Bah22, Bir91b, Bir91a, Bir92a, Bir92b, Bir93, Bir97a, Bir97b, Bir01a, BM04, Bir04, BC06b, Bir06, Bir10, Bur99, Ceb12, Cla04, Cur06, Dan98, EK06, Gib96a, Gib96b, GLB06, GBF05, HMP98, Har09a, HP93, Har99b, HMG04, Hin99a, Hin09b, Hoo92, HM96, Jon04, KH18, Ken96, Ken04, Kol12, LS96, Lin00, Lju04, LW99, MK97, Mai04, Mar04, MS04, Mee04, Mel97, Mic00, Mic07, Mil13, Mol93, OR04, Orc14, Pag97, Pag98, Rag20, Rea97, Ric98, Run97, Rus03, Sch03a, She11, SA14, Spe24, TW93, Tho99, Tri05, Wam11, Waz10, Wei04, vS00, AA10, AFB92, ABR97, All05, AMNS22, AD15, Ary94, BKP01, Bie00, BDH96, Bir20, BH92, BRL17, BO96, BRS94, CK04, CDHA14, Chi94, Cla15, CM98, DG00, DL22, DRvE95]. functional [DLP10, Erw01, Fil21, Fil03, FF06, FKS11, FM00, FSNB09, GS05, GG99, GSWZ96, Gre05, GMO06, HSB97, Han99, Han98, Har93b, HP95a, HFA^{+}96, HG96, HW97, Hin04, Hud00, Hue05, HNSSH97, IEM^{+}22, JVV93, LP13, LOMPM05, LS10, Lüt03, Mac94, MT91, MW08, Mor18, MJ98b, MR01, NF94, Nil01, NO99, O’H96, OB97, Oka95, Oka99b, Oka01, PS05, PG20, PN20, Pey92, QGG21, RL99, RW97, RW99, RW93b, RTF93, Sab98, Sch03b, SMZ21, SJ92, SM13, SBHG10, Stä98, SMT16, SDPJS07, SC09, Tho00, TL13, WS92, Wak98, Wak99, Wak07, Won00, ZSW95, Joh92, Spe20, Bai99, Bro12]. functional-imperative [QGG21]. Functionally [Gru15, FP03, HM96]. functionals [IRK01, JO15]. Functions [Pey03o, Pey03f, Pey03p, BHY07, BR08, BD92, CG97, DM14, Hut92, HJG10, IB99, JJ23, KN08, MLC91, MBG^{+}20, MKHT12, NN91, Oka98, Ram11, Ste21, VK04]. Functorial [JBM98]. Funser [ZSW95]. Further [Chi94]. fusion [Chi94, Gam11b, GJ07, Joh03, KN08, SDK18]. future [FF99]. futures [PP22]. fuzzy [MJ98a].

G [Pey92]. G-machine [Pey92]. Gabriel [LP09]. Gabriel-Ulmer [LP09]. Gabrielle [Tho04a]. GADTs [ZMCN18]. Galois [DV19]. game [F^{+}13, Lüt03, WE09]. games [Mor18]. Gansner [Jon06]. garbage [EH21, JSE^{+}14, App97a]. Garsia [Bir20]. Gauss [AD15]. Gencel [EAKC06]. General [Hag11, Bra13, HP06, LP09, Par10]. general-purpose [Bra13, HP06]. generalised [GJ07, HLA20]. Generalising [MBG^{+}20]. generalization [GG99]. Generalized [PRS20, Bar91a, Hin00a]. Generalizing [Hin00a]. Generating [BP96, CDP15, ARS94, FS03, GL13, GL15, MLC91]. Generation [TW00, Tah03b, AD15, BP92, Glü94, SGM^{+}03, Thi99]. generational [EH21]. generativity [Dre07, Ler96]. generator [EAKC06]. generators [Ram03, Sch15]. Generic [BJZS10, BDH96, Hen12, Hin10b, SW23, BBJR21, GJL^{+}07, OG10, PNCB06, VYH^{+}10]. Generics [Hin06]. Genuine [O’N09, Nyk11]. geometry [DD95]. Gerald [Kar03]. Ghosh [Spe20]. Ghostbuster [ZMCN18]. Gibbons [Gru20, Han04]. Gilmore [Ang03]. Girard [Has00]. Glasgow [PM02]. Global [Hug04, BL96]. Go [SJW18, WW08, SW23]. Going [FSNB09]. Good [Mic13, Hud15, Lip12, Spi12]. Gordon [Hut94, Kah98, Rea04]. GPUs [SNS16]. Gradual [BGT16, NLA21, GTN^{+}19, ICEW22, SCYZ24]. gradually [SC21]. Graham [Bha09]. grammar [DP12]. grammars [Ben17, Lju04]. Grammatical [Ran04]. granularity [ACR16, AFB92]. graph [Bir13, Erw01, GK96, IEM^{+}22, Joh98]. graphical [AAS17, GS03, LW99]. graphs [Erw01]. Great [Mic13, Hud15, Lip12]. greedy [Bir92a, Bir92b, BM21]. Greg [Ang03, Lin00]. ground [JKJ^{+}18]. groups [DG02]. Guerrini [Mic00]. guide [Lip12, ZS17, Mic13, Hud15]. guided [ACR16]. Guy [Har99b, Rus03].

H [Bai99, Jon06, Klu01, O’D01]. Hall [Dup95, Fou96, Gor96, Pat14, PT99]. Hall/CRC [Pat14]. Hammond [Lin00]. Handbook [Har09b, Car11]. handlers [BSO20, HLA20, HLL24, KP17, PRS20]. handling [SDPJS07]. Hans [Ang03]. Hans-Wolfgang [Ang03]. Hansen [Klu01]. hardback [Kah98, Klo01, Mil13]. Hardcover [Wat14]. hardware [Bru92, GMO06, PI11, Pey92]. Haridi [Gam09]. Harper [Wat14, DP22]. Harrison [Car11, Mit95]. Hartel [Bai99]. Haskell [Gru20, Hud00, Jon07, Mic13, PM02, San06, vS00, Bah22, BFGJ04, BJZS10, BLH16, BEJW16, BSZL^{+}21, CG96, CK02, DvJE04, EK06, Fax02, HLB95, HK05, Hug04, HM98, Hut00, Hut02b, Hut07, Jon13, Lip12, Lüt03, Mar02, McB02, Mei00, Men14, OG10, SFG14, SHMJ16, Thi02, TUmI15, Bha09, Gre14, Gru15, Tho04a, Ric98]. Haskell-Embedded [SFG14]. Haskells [TLP02]. Haskore [HMGW96]. haste [BJD97]. Haynes [Tah03a]. HB [Han04]. Heap [RW93b, AS96, RR96]. heaps [Hin99b]. Hébert [Hud15]. height [Bir97a]. Helmut [Dyc01]. Herman [Cla15]. Heterogeneous [Swi20, Abe09, SFL17]. hidden [Pot13]. hierarchical [SNS16]. High [SLW^{+}07, Tri05, Kre04, Mar02, Sch03b]. High-level [SLW^{+}07, Sch03b]. high-performance [Kre04, Mar02]. Higher [HP99, Hut92, NTHV17, RW99, Ste21, ANN97, AD15, AGN96, BHY07, Bla18, CG97, DNB12, GM93, HM94, Hut94, Jon95, Jon01, JKJ^{+}18, LP13, MP06, MS08, MO14, NN91, Oka98, PI09, Ram11, SVJB17, SSV08, Tof94, VMA21, VW10, WW08, GP98, Kah98]. Higher-order [HP99, Hut92, RW99, ANN97, AD15, AGN96, BHY07, Bla18, CG97, DNB12, HM94, Jon95, Jon01, JKJ^{+}18, MP06, MS08, MO14, NN91, Oka98, PI09, Ram11, Tof94, VW10, WW08]. Hill [Rea97]. Hindley [Gam11a]. histories [BGP11]. HM [SS08]. Hoare [BGP11, NMB08]. hoc [GZND13]. HOL [Hut94, GM93, MNVS99, RH03]. HOLCF [MNVS99]. Homomorphism [Gib96b]. Homotopical [AMLH16]. Honour [Rea04]. Horn [Jad14]. HOT [DG00, Wad97]. Hough [Fok11]. hs [BSZL^{+}21]. hs-to-coq [BSZL^{+}21]. HtDP [Mor18]. HtDP-based [Mor18]. HTML [Thi02]. http [Hag11]. Hudak [Jon07, vS00]. Huffman [Hin23]. Hughes [Joh92]. Huth [Bud08]. Hutton [Bha09].

I/O [AP95]. ICFP [ACW12, AP13, DS16, SF14, TN10, TL11]. Id [HLB95]. ideals [FP03]. identification [CE18]. identity [ACD^{+}20]. Idris [Bra13, De 20]. II [BH20, Chi94, DD03a]. Ill [SJW18]. Ill-Typed [SJW18]. illustrated [Moy24]. image [MC13]. impact [DD03b, DNB12]. Imperative [Tho04b, BHY07, CMS03, Col03, GHL99, QGG21]. Implementation [Hag11, McK99, Mic00, Nil01, TW00, Tah03b, AHN12, App97b, Bir20, Bra13, KRRS97, LW99, Mac99, NN91, Pad17, Par10, Tho16, Tho02]. Implementations [Sha04b, HFA^{+}96]. implemented [HG96]. Implementing [Abe09, Bur91a, DRvE95, Han99, Pey92]. Implicit [CDHA14, Jon95, Hai03]. Implicitly [FRRS10, ACR16]. implicits [SOWM19]. improvement [KN08]. improvements [Chi94]. Improving [JB93]. in-situ [Bir13]. including [ZS17]. incremental [BC15, OSC12, RTF93]. indentation [Lap98]. Index [Ano91a, Ano92a, Ano93a, Ano94a, Ano95a, Ano96a]. Indexed [AGH^{+}15]. Indexing [Pey03g]. induction [BB21, Mat09]. Inductive [Erw01, MR01, LP13, VMA21, Sch03a, Mic07]. inference [HM94, Mit91, PVWS07, SS08, TJ92, VPSS11]. information [BN05, CKP05, DD03a, PP01, PW03a, SMMR17, ZSW95]. Infusing [Mor18]. ing [RRD14]. inhabitation [BS96]. injective [Bar99]. inliner [PM02]. Inlining [MS03]. Input [Pey03y, Pey03n, MBG^{+}20]. Input/Output [Pey03y, Pey03n]. ins [AP95, VHMW13]. insider [Pie13]. inspection [BDJ05]. Instances [Pey03b]. integers [HS15]. integrated [Ser00].

Integrating [EH21]. integration [SW17]. Intellect [Ang03]. Intensional [CWM02, Mat09, RW97, RW99]. intensive [HFA^{+}96]. Interaction [Rea04]. Interactive [AAS17, Kin07, Aug93, BR04, GK96, Lüt03]. Interfaces [BDJ05, AAS17, GS03, LW99]. interference [OSC12]. Interleaving [AJ15]. International [Fou96, KR02]. interoperability [DTT18, TO98]. Interpretation [Kar03, Woo01, ASS96, ASS00, Bar91b, BCMT01, Bou92, CGR96, FFFK04, Mog92, SWM01]. interpretations [Fil03, HL96]. interpreted [Ram11]. interpreters [Ben05, CKS09]. interruptions [HW07]. intersection [Dun14, FD04, Nee05, PP01]. intervals [JB93]. Introduction [Bar91a, BDT04, Fis06, GM93, Gor96, HR99, Lew13, Pey03a, Tho04a, TW93, Wan94, Ceb12, CK02, Fer04, Fer06, LS10, Gam11a, Han01, Hut94, Klu01, Pat14]. intuitionistic [DM05]. invariants [Zei16]. inversion [MC21]. invertible [MW24]. Inverting [BM04]. Iris [JKJ^{+}18]. irrelevance [BB96]. Isabelle [RH03]. Isabelle/HOL [RH03]. ISBN [Bha09, Boi03, Boi12, Bro12, Bud08, Car11, Cas13, Ceb11, Ceb12, Chi09, Cla15, De 11, Der04, Fou96, Gam09, Gre14, Hag11, Han04, Har99b, Hud15, Hut94, Jad14, Jon92b, Kah98, Kar03, Kel03, Kin07, Klo01, Klu01, Kol12, McK99, Mic07, Mic13, Mil13, Mit95, O’D01, Pat14, Rag13, Rus03, San09, San06, Sim07, Slo09, Spe14, Spe20, Tah03a, Tho02, Tho04a, Tho04b, Tho16, Vil07, Wal04, Wat14, Woo01, dP01, vS00]. ISBN-10 [Boi12, Gre14, Jad14, Mic13, Pat14, Rag13, San09, Wat14]. ISBN-13 [Bro12, Mic13]. isomorphic [ACD^{+}20]. isomorphisms [Di 93]. Issue [AH92, AH93, AP13, BDT04, CG18, HP93, HP95a, Hin10b, Hut02b, KH18, MK97, Mel97, SF14, SA14, TN10, TL11, Tri05, ACW12, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano92c, Ano92b, Ano92e, Ano92d, Ano92g, Ano92f, Ano92i, Ano92h, Ano93c, Ano93b, Ano93e, Ano93d, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95e, Ano95d, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96i, Ano96h, Ano96k, Ano96j, Ano96m, Ano96l, Ano04, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano09j, Ano09i, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e]. issue [Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano11j, Ano11i, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i, BDT01, DH00, DS16, Hin04, Hue03, Hut00, Oka01, SD17, TW00]. Iterating [PTW22]. Iterator [GdSO09].

J [Gam11a, Han01, Tho02]. Jack [Kar03]. Jackson [Slo09]. Jan [Kol12, Orc14, San06]. Jane [Gor96]. January [KR02]. Java [Bro12, Cle16, Wak99, Wam11]. Jay [Kar03]. Jean [Boi12, Tho04b]. Jean-Francois [Tho04b]. Jean-Raymond [Boi12]. Jeff [Ceb11]. Jeffrey [Fou96]. Jeremy [Gru20, Han04]. Jersey [Fou96]. Jeuring [Bur99]. JFP [Hag12, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano92c, Ano92b, Ano92e, Ano92d, Ano92g, Ano92f, Ano92i, Ano92h, Ano93c, Ano93b, Ano93e, Ano93d, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95e, Ano95d, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96i, Ano96h, Ano96k, Ano96j, Ano96m, Ano96l, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano09j, Ano09i, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano11j, Ano11i, Ano12b, Ano12a, Ano12d]. JFP [Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i]. jigsaw [Nak13]. Joe [Dup95, San09]. Johan [Bur99]. John [App97a, Boi03, Car11, Gor96, Joh92, Jon92b, Jon06, Kel03, Mit95, Rag13, Rag20, Sha04b, Ten97, Ten00, Tho99]. Jones [App97a, Oka99a]. Jordan [AD15]. Journal [MK97, Mel97, Spe24]. journey [HB17]. judgemental [Ada06]. July [Bro12].

Kansas [SFG14]. Kaufmann [Hai03]. Kees [San06]. Keller [Tho04a]. kernels [GY13]. Kessin [Cas13]. Kevin [Lin00]. keys [Rit91]. Kim [Han01]. King [San06, Sim07]. kit [Ary94]. Klaus [Vil07]. Kleisli [Won00]. KN [GPN19]. know [Pie13]. Knowledge [Sha04a]. known [Swi22]. Knuth [Moy24]. Krishnamurthi [Wal04]. Krivine [GPN19].

L [Fou96, Rea97, dP01]. lambda [AJ97, AF97, Bar91b, Bar92a, Bar92b, Bar91c, FH03, FD04, GJ91, GL13, GL15, HMP98, HM94, Hof99, KN95, Luo03, Mai04, MOW98, Mog92, Par16, Sel02, SC21, SF16, Stu17, Swi23, Zei16, BG00, Mor21, Vas05, dP01, Gam11a]. Lambda-Calculi [dP01]. lambda-calculus [FD04, GJ91, Gam11a]. lambda-encodings [SF16]. lambda-free [Luo03]. lambda-sigma [HMP98]. lambda-terms [Par16]. LAML [Nør05]. Language [Ano04, Par10, Rea04, SFG14, Slo09, SDF^{+}09, Bie00, Bra13, BRL17, Bru94, BAD^{+}01, CF10, FS08, FM00, GS05, GJL^{+}07, GSWZ96, GMO06, HSB97, HG96, HPR98, Hut23c, IEM^{+}22, Jac06, KRRS97, KM01, LLR93, LP03, Mac94, MW24, MCGW02, MCGW03, PS05, Ram11, Sab98, Ser00, SLW^{+}07, SNS16, TO98, VMA21, WS92, WE09, AMS05, Hag11]. Language-Based [AMS05]. Languages [Bai99, CG18, Chi09, Joh92, Sim07, SA14, Tah03a, Tho02, ACPR95, ACR16, AS96, AD15, Bar00, BC15, CKS09, DL22, EFD03, Fer04, Fer06, FWH01, FW08, Han98, Har13, Har93b, HP93, HP95a, HFA^{+}96, KR02, Mac99, Mai92, MP06, MT91, McI04, Mit03, NF94, Oka01, Par10, Pey92, RW97, RW99, Stä98, TL13, Boi03, Hag11, Hag12, Kel03, Mic00, Ten97, Ten00, Wal04, Wat14].

Lapalme [Rus03]. large [IEM^{+}22, MJ98b]. large-scale [IEM^{+}22, MJ98b]. last [Bir93]. lattice [CG96]. lattice-based [CG96]. lattices [Jon92a]. Lauchbury [Tho99]. Launchbury [Jon92b, Bre18]. law [PS17]. Lawvere [LP09]. layout [HW97, Lap98]. layouts [Str11]. Lazy [Aug93, BFR^{+}12, Run97, BJD97, Bre18, FKS11, HG96, Joh98, KRRS97, MG06, MJ98b, NF94, Nil01, NT01, Pey92, RW93b, RTF93, Ses97, Wak98, Wak99]. LCF [MNVS99]. leaks [Bah22, Jon92c]. LeanPub [Hac20, Spe24]. Learn [Lip12, F^{+}13, Hud15, Mic13]. Learning [vS00, All05, Hud00]. least [CG97]. Lecture [Bur99, Joh92, Tho99]. left [CA20, Dan19, Dan23b, Gol00, MM04, PS17]. left-associated [Gol00]. left-zero [PS17]. Leftmost [IN97]. Leibniz [ACD^{+}20]. lemmas [Dan22, Dan23a]. lenses [MW18, ZTX^{+}23]. less [BJD97, GZND13]. lesson [BC06b]. Lessons [Ber93]. Level [JJ23, Sch03b, SLW^{+}07, SP04]. Level- [JJ23]. Lewis [Pat14]. Lexical [CCP95, Pey03l]. LF [Pie13]. libraries [Rit91]. library [Ber93, DD95, MS04, Pad17, Str11, SV13, Jon06]. life [Jon01]. lift [Tho00]. lifting [FH03, VH95]. Lightweight [SIV08, ZJ10, VYH^{+}10]. Lilac [Mac94]. Limited [Pag98]. Lin [Jad14]. Lindeman [Jad14]. line [BJ93, CK95]. lineage [HMM18]. lineage-based [HMM18]. Linear [GV10, Mai04, VPD21, SC09, Zei16, AJ97, Ben95, Bie00, CGR96, CM15, Hen12, Hof99, KSK08, Mac94, SDB21, WS92]. linear-time [CM15]. linear/modal [Hof99]. Linearization [FD04]. lines [RA19]. Lins [App97a]. Linux [DLP10]. Lipovaca [Mic13]. List [Pey03i, Hoo92]. lists [Swi20]. literacy [Hic04]. Little [Tan20, FS02, CF10]. Liu [Spe14]. Lloyd [Kin07]. LNCS [Wal04]. Local [LP20, DNB12, VPSS11]. Locales [Pey03r]. locality [PI09]. localization [SFL17]. locally [Pit11]. Loci [LG05]. lockstep [GPN19]. Löf [ACD^{+}20, Hed98]. logarithmic [OR04]. Logic [Car11, HR04, San06, Slo09, AD15, CCG05, CGR96, DvJE04, DA18, Erw04, GM93, HK05, Har09b, Jac06, JKJ^{+}18, Mac94, MJ98a, MO14, RW97, RW99, SJ92, SR98, VPD21, Swi23, Dyc01, Hut94, Bud08]. Logical [DH00, Hue03, BBD98, BHY07, HL07, Has00, Luo03]. Logics [Stä98]. Loidl [Ang03]. Longest [MC21]. look [Hin02, Nil01, Pie13]. lookahead [OR04]. loop [SDK18]. looping [CH94a]. loops [HL99]. Loverdos [Ceb12].

M [Kah98, Klu01, Tho04a, dP01]. MA [Joh92]. Machine [Wak99, All05, AGN96, GPN19, Ses97, Cle16, Pey92]. machines [Asp92, BH20, Bur91a, DP09, SR98, VM12, Vas05]. MacLennan [Tho02]. Macmillan [Han04]. Macros [FCDF12, Cul12, Kri06]. made [GPW12]. Mads [Rea04]. Maguire [Hac20, Spe24]. make [Gam11b, GZND13, LP99]. Making [MP06]. man [Cla99]. management [EH21, App97a]. managing [AFB92, ABR97]. manifest [GPW12]. Manipulating [MKHT12]. manipulation [GS06]. Manning [Spe20]. Manoury [Lou04]. Manticore [FRRS10]. Manuel [Tho04a]. Manufacturing [Hin01]. Many [Kis20]. maps [Zei16]. Marble [Cur06]. Maribel [Sim07]. Mark [Bud08, Pat14]. marking [Bir01b]. Marriott [Han01]. marshalling [SSH^{+}08]. Martin [Hag12, ACD^{+}20, Hed98]. Martin-Löf [ACD^{+}20, Hed98]. masses [Hin06]. massively [FM00]. matching [BC93, CDP16, CD18, CA20, Fil21, GS06, HP03, Mar07]. Maths [San06, DvJE04]. matrix [GSWZ96, LP99, WS92]. matter [Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano92c, Ano92b, Ano92e, Ano92d, Ano92g, Ano92f, Ano92i, Ano92h, Ano93c, Ano93b, Ano93e, Ano93d, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95e, Ano95d, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96i, Ano96h, Ano96k, Ano96j, Ano96m, Ano96l, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano09j, Ano09i, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano11j, Ano11i, Ano12b, Ano12a, Ano12d, Ano12c]. matter [Ano12f, Ano12e, Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i]. matters [Dan23b]. Matthias [Jad14]. Mauny [Har99b]. Maximum [Bir01b]. May [Boi12]. Maybe [Pey03j, Spi12]. Mayer [Kar03]. McGraw [Rea97]. McIver [Der04]. MD [Jad14]. me [CIZ22]. meaning [HW07]. Mechanics [DA23, SWM01, Kar03]. mechanism [Cro99]. mechanized [RH03]. Mechanizing [HL07]. Media [Bro12, Cas13]. Meertens [Bir98]. Meets [BDT01, BDT04]. Meijer [Bur99, Tho99]. Meinhard [Kar03]. Melham [Hut94]. memoization [JJ23]. Memory [App97a, EH21, Gre05]. Mena [Gre14]. Merge [HZO21]. merging [Bir97b, HM18a]. message [CIZ22]. message-passing [CIZ22]. Meta [CX05, SC21]. Meta-programming [CX05]. meta-theory [SC21]. metalanguage [Hue03]. Metalanguages [DH00]. MetaML [CMS03]. Metamorph [Bru92]. Metamorphism [Nak13]. metasystem [Tur93]. metatheory [HL07]. method [CG97, OB97, SW23]. methodology [MM03, Der04]. Methods [Sch04, Bru92, Gor96]. Michael [Bud08]. Michaelson [Ang03, Lin00]. Michel [Har99b]. middle [BB96]. Migrating [ICEW22]. Mike [Dup95, Hut94]. Milner [Jef01, Rea04]. Mimi [Jad14]. mingle [Bir15]. mingling [Cur06]. minimum [Bir97a]. mining [All05]. Minout [Bir91b]. Miran [Mic13]. Miranda [LLR93, Ric98]. mistakes [TUmI15]. Mitchell [Boi03, Chi09, Tah03a, Ten97]. mixed [SM21]. mixed-sensitivity [SM21]. Mixing [MW18]. ML [Fou96, Jon06, McK99, AH92, App93, AH93, App97b, Aug93, Ber93, Fou96, HDM93, Har93a, Hay00, JBM98, LM93, MS95, MCP93, MO14, Pau91, Rea97, Rep99, SMZ21, SZJ14, TA95, TO98, Ull94, Xi07, ZJ10, Fou96, Mit95, O’D01, Jef98, Rea97]. MLton [SMZ21]. Mobile [Jef01, Kin07, BR04]. Modal [Bah22, Hof99, HJP23]. modality [Thi09]. model [AVDP13, Bur91a, CC19, DP12, HMM18]. Modeling [Abr10, AS21, Spe20, Boi12]. Modelling [BD92, Bud08, HR04]. Models [All05, Gam09, BM06, FS03, VH04]. Modern [App97b, McK99]. Modula [Aba94]. Modula-3 [Aba94]. Modular [Chl08, GN91, GY13, NT01, OSC12, SVJB17, VPSS11, JKJ^{+}18, Ler00]. Modularity [BFG97]. module [AZ02, Cou07, Ler00]. Modules [Pey03w, CF10, RRD14, Ros18, SP04, Tof94]. Monad [Pey03m, AGNS21, Cla99, ZDK^{+}15]. Monadic [FM06, GJ07, HM98, MS01, SSW09, BB21, CKP05, DPS07, Sch09]. monitoring [KH95, SSTH18, Thi03]. monoidal [PRS20]. monoids [PS17, RJ17, Swi22]. monolithic [Joh98]. Monte [HLB95]. Moor [Han04, PT99]. Morgan [Der04, Hai03]. morphisms [SB99]. morphological [Hue05]. Morris [Moy24]. MP [BH92]. MRI [OSC12]. MSO [TN15]. Mtac [ZDK^{+}15]. Multi [SF06, DO02, IT16, LG05]. multi-disciplinary [LG05]. multi-parameter [DO02]. Multi-return [SF06]. multi-stage [IT16]. multicore [SZJ14]. multicore-aware [SZJ14]. Multimedia [Kin07, BR04, Hud00, vS00]. MultiMLton [SZJ14]. multiple [PTW22]. multiprocessor [Gre05]. music [HMGW96]. mutable [CKP05]. Myers [Fou96].

n [Pie13]. name [JFDA17]. names [ARS94, NP05, Pit11]. Nanopass [SWD05]. narrowing [HP99, IN97]. NAS [HSB97]. National [Swi11]. Natural [HNSSH97, BG00, Bre18, DP09, Han93]. naturals [HR22]. necessity [NP05]. need [AF97, FI00, MOW98, NH09, PW96]. needed [KP17]. Neil [Oka99a]. nested [BP99, HL99, Mat09]. network [HM18a, HM18b]. networks [Kre04, She11]. Nicole [Jad14]. Nielson [Jef98]. Nikhil [Hai03]. Nipkow [Klo01]. NixOS [DLP10]. No [CA20, Hud15, Jad14, KP17, Mic13]. Nöcker [SSSS08]. Non [NDR11, PS05, BC06a, Bur91b, Fil03, KM01, NS93, Stä98]. non-determinacy [Bur91b]. Non-determinism [PS05]. non-deterministic [KM01]. non-functional [Fil03]. non-normalizing [BC06a]. Non-parametric [NDR11]. non-recursive [NS93]. non-strict [Stä98]. nondeterminism [AGNS21]. nondeterministic [FKS11]. Noninterference [HY05, HLSZ^{+}16]. Normal [Ben17, GPN19]. Normal-order [Ben17, GPN19]. normalisation [AHN12, AC09, BBLRD96, Ben95, Gui00, KR97]. normalising [Pol98]. Normalization [DRR01, HJP23, BFG97, GN91, Nee05]. normalizer [Abe09]. normalizing [BC06a]. notation [BP99, HMGW96]. Note [O’H96, Gam11b, Nyk11]. Notes [Bur99, Joh92, Tho99]. Notions [RJ17, Atk09]. nucleic [TLM95]. number [Bir98, BP92]. Numbers [Pey03e, Pey03d, Pey03t, Dan19]. numeral [Bar91c, Gol00]. numerals [Hin05, Par16]. Numeric [Pey03f]. Nuprl [Ben01, Kre04]. Nussbaum [Jad14]. NY [Gre14].

O [AP95]. Object [Ceb12, Bru94, CD96, LS10, PT94, Sur07, Wad94]. Object-Functional [Ceb12, LS10]. object-oriented [Bru94, PT94, Sur07, Wad94]. Objective [Lou04]. Objects [AAS17, Aba94, Bou04, GHL99, HP95b]. observable [DP12]. OCaml [Cle16, Whi13, Rag13]. OCaml-Java [Cle16]. odd [HM18a]. Oege [Han04, PT99]. off [BJ93, CK95]. off-line [BJ93, CK95]. Ohlebusch [Sha06]. Okasaki [Hin99a]. On-line [CK95]. one [DMN07, F^{+}13, Kam05]. one-pass [DMN07]. online [RW93a]. only [CD18]. open [KR97]. operating [BD92]. Operational [HL96, HPR98, Kah98, Sim07, Fer04, Fer06, GP98, MF08]. Operations [Pey03g, Hoo92, Sch03b]. Operator [HZO21]. operators [HL96]. opportunity [Wad97]. Optimal [BO96, Mic00, AGN96, Bir20]. optimisation [Bou99]. optimisations [BL96]. optimization [WC01]. Oracle [ACR16]. Oracle-guided [ACR16]. Order [Kah98, ANN97, AD15, AGN96, Ben17, BHY07, Bla18, BAD^{+}01, CG97, DNB12, GPN19, GM93, GP98, HP99, HM94, Hut92, Hut94, JO15, Jon95, Jon01, JKJ^{+}18, LP13, MP06, Mcb03, MS08, MO14, NTHV17, NN91, Oka98, PM00, PI09, Ram11, RW97, RW99, SVJB17, SSV08, Ste21, Tof94, VW10, WW08, Yi06]. ordered [KSK08]. O’Reilley [Lou04]. O’Reilly [Bro12, Cas13]. oriented [Bru94, PT94, Sur07, Wad94]. origami [BM05]. ornaments [DM14, Dag17, KG17]. Ott [SNO^{+}10]. Output [Pey03y, Pey03n]. outs [AP95]. outside-in [IN97]. OutsideIn [VPSS11]. overloading [Jon95, SW23, ZS17]. overview [Kea94]. own [Hag11, Par10]. Oxford [Tho02].

P [Bai99, BS96, Chi09, Lou04, Tah03a, Tan20, dP01]. P. [Gam11a]. P.-L [dP01]. PADL [KR02]. PAEAN [BLH16]. Pagano [Lou04]. pairing [CD96]. pairwise [HM18b]. PAL [Luo03]. Palgrave [Han04]. paper [Kri21]. paperback [Hud15, Pat14, Rag13]. Papers [MK97, Mel97, BDT01]. paradigm [BAD^{+}01]. paradigmatic [Bru94]. Parallel [Cla04, KH18, Lin00, LOMPM05, NO99, TLP02, Tri05, ACR16, AFB92, BC15, BLH16, Bur91a, BRS94, FM00, Hai03, KRRS97, LG05, MS95, PS05, SHMB05, She11, SBHG10, SNS16]. parallel-functional [PS05]. parallelisation [Mor21]. parallelism [FRRS10, THLJ98, VH95]. Parallelization [HL99]. Parallelizing [GG99]. parameter [DO02, KN08]. Parameterised [Atk09]. Parameterized [SC21, SP04]. parameters [IB99, VK04]. parametric [Mai92, NDR11, WW08]. parametrically [ACD^{+}20]. Parametricity [BJP12, VW10, NDR11]. Parberry [HM18b]. parentheses [MC21, Oka03]. Parr [Hag11]. parser [PW96]. Parsing [BLS04, Cla04, BC15, Hil96, Hin11, Hut92, HM98, Lju04, Oka98]. Part [DD03a, IRK01]. Partial [Jon92b, AF04, BJ93, BP96, CK95, FCDF12, FS02, GJ91, Lau91, LB96, Stä98]. Partially [NS93, CKS09, MW24]. partitioning [Bou92, Hen12]. partitions [SB99]. pass [DMN07]. passing [BSO20, CIZ22, SW23]. patch [AMLH16]. path [CGMS06, HG96]. Pattern [BC93, BR08, CDP16, CD18, CA20, GdSO09, GS06, HP03, Mar07, Rhi09]. Patterns [Hag11, JK09, Par10, SSTH18]. Paul [Jon07, vS00]. Paulson [Fou96, Rea97]. pbk [Har99b]. Peano [Dan19]. Pearl [AH04, ARS94, BLS04, BD06, Bir97a, Bir97b, Bir01a, BM04, Bir04, Bir06, Cla04, CH94b, Cur06, GLB06, Har09a, Hin05, HM96, Jon04, Ken04, Kri06, LS96, Lju04, Mai04, Mar04, Mee04, OR04, Run97, SWD05, Wak06, Wei04, Yi06]. Pearls [Ada93, Bar91b, Bar92a, Bar92b, Bar91c, Bir91b, Bir91a, Bir92a, Bir92b, Bir93, BC06b, Bir10, EK06, Fel03, Gib96a, Gib96b, Hoo92, Ken96, Nee05, Oka03, PW03b, Hin04, Mil13]. Pearson [Tho04a]. Perfect [Hin00b]. Performance [Tri05, GTN^{+}19, Kre04, Mar02]. permutation [BLS04]. permutations [Hin00b]. Persistent [BBC^{+}01]. perspective [BBD98, HH91, Oka99a]. Peter [Han01, Tho16]. Peterson [Jad14]. pH [Hai03]. Phantom [FP06]. PhD [DP22, Hut14, Hut15a, Hut15b, Hut15c, Hut16a, Hut16b, Hut17a, Hut17b, Hut18, Hut19a, Hut19b, Hut20a, Hut20b, Hut21a, Hut21b, Hut22a, Hut22b, Hut22c, Hut23a, Hut23b]. Phil [Ang03]. phonological [Hue05]. photon [HLB95]. phrases [BLS04]. pi [Vas05]. Pickering [De 11]. Pickler [Ken04]. pieces [FS02]. Pipelined [BKP01]. pipelines [PS20]. Pitts [Kah98]. Planning [Sch03a]. Plasmeijer [Bai99]. Platforms [SA14]. playing [LP03]. Plessner [Jad14]. Plotkin [Rea04, SB99]. Plotkin-terms [SB99]. plug [VHMW13]. plug-ins [VHMW13]. point [VH95]. pointer [Bir01a]. pointless [Gib99]. points [CG97, HH91]. policy [BJI17]. Polymorphic [TJ92, ACPR95, DD95, JM00, RT91, STI24, SV13, WDMT02]. polymorphism [CWM02, GHM24, Gre96, FBH17, Jon95, Mai92, NMB08, SIV08, VW10, WW08, ZS17]. polynomial [BCMT01]. polynomials [JJ23]. Polytypic [JJ98, VDH10, Wei06]. polyvariance [GAM18]. Polyvariant [PP01, WDMT02]. Poon [Fou96]. poor [Cla99]. Portable [BLH16]. Portland [KR02]. positive [SGJ96]. postponement [Pol98]. Potter [Gor96]. Power [Mci99, Jon01]. powerful [Zhu94]. Pp [Tho04a, Bha09, Ceb12, Fou96, Kin07, Mic07, Mit95, San09, Slo09, Vil07, Wal04, vS00]. pp. [Bro12, Dup95, Dyc01, Gam09, Gre14, Har99b, Hin99a, Joh92, Jon06, Klo01, Kol12, Mic00, Mic13, Mil13, O’D01, Rag13, Rus03, Wat14, Woo01]. Practical [BDT04, Car11, Har13, PVWS07, Rea97, BC15, BH92, CLB23, Har09b, LP99, Xi07, KR02, Wal04, Wat14]. Practice [BDT01, Kol12, AA10, CCP95, MMJ20, SVJB17]. Pragmas [Pey03c]. Pragmatic [Hag11, Ram03, San09]. Pratt [Moy24]. predecessors [Kis20]. Predefined [Pey03x]. predicates [Has00]. predicative [AA02, CCP01]. Predictive [PW96]. preemptive [GY13]. Preface [CS16, FP07, MU09]. prefix [She11]. Prelude [Pey03z]. Prentice [Dup95, Fou96, Gor96, PT99]. Prentice-Hall [Gor96]. presence [SMMR17]. presentations [FF06]. preservation [BBJR21]. preserve [Gui00]. preserves [BBLRD96, KR97, KRAB22]. Press [Bha09, Boi03, Boi12, Car11, Ceb11, Ceb12, Chi09, Cla15, Fou96, Gam09, Gru15, Gru20, Han01, Har99b, Hin99a, Hud15, Hut94, Jad14, Jef01, Jon92b, Jon06, Jon07, Kah98, Kah05, Kar03, Klo01, McK99, Mic00, Mic13, Mil13, O’D01, Oka99a, Pag97, Pag98, Pat14, Rag13, Rea97, Rea04, Sha04a, Slo09, Spe14, Tah03a, Tan20, Ten97, Ten00, Tho02, Tho16, Wat14, Woo01, vS00]. pretty [SC09]. pretty-printing [SC09]. Price [Gam09, Tho04a, Wal04]. primes [Run97]. Principal [Tof94]. principle [Mat09]. Principles [Mac99, Tho02]. printing [SC09]. Priority [CIZ22, BO96]. Probabilistic [EK06]. probabilities [MG06]. probability [AGNS21]. Problem [Sha04a, Swi11, Ach21, Bir91b, Bir11, CCP01, Cra20, CM15, HG96, Hut02a, MC21]. Problems [NT01, Bir01b, CG96, MLC91]. procedures [TN15]. proceedings [Joh92, KR02]. process [BCG05, DHLB16, GHM24]. Processes [Cla04]. processing [BD92, Gre05, Hue05, IEM^{+}22, KSK08]. Producing [FP03, MO14]. product [BFGJ04]. Profiling [MJ98b, CCP95, RW93b, RR96, SBHG10]. Program [Bie00, Spe14, TW00, Tah03b, Thi03, Tho04b, Tur93, Bir06, BH92, Chl08, Col03, DD03b, EAKC06, F^{+}13, Fil21, KH95, Liu13, MKJ09, MC21, Ram03, RW93a, Thi99, Tof94, Sha04b]. programmer [Pau91, SJCA17, Fou96, Rea97]. programmers [LS96, OG10]. Programming [Arm07, Bai99, BDT01, BDT04, Boi03, Bro12, CG18, Chi09, Dup95, Fer04, Fer06, GS03, GBF05, Hag11, HP93, Hin10b, Hut07, Hut23c, Joh92, KH18, Klu01, KG17, MK97, MM03, Mel97, Mic00, MCP93, O’D01, Oka99a, Rea97, Rus03, SFG14, SD17, Tah03a, Ten97, Ten00, Tho02, Tho04b, Tri05, Wat14, vS00, AAS17, AVDP13, AA10, All05, Bah22, BJZS10, BDH96, BM05, BC06b, BBJR21, Bra13, Bru94, CK04, CX05, Col03, CM98, DvJE04, Erw04, EK06, EK21, FCF^{+}02, FKS11, FWH01, FW08, GJL^{+}07, GdM03, Hai03, HMM18, HR99, Har13, HP95a, Hic04, Hud00, HNSSH97, JVV93, KM01, LP13, LLR93, Lew13, LOMPM05, LS10, Lüt03, Mac99, Mac94, Mai92, MG06, MW18, MP08, MS04, MW08, Mit03, Mol93, Mor18, MKJ09, NO99, Nør05, O’H96, Oka01]. programming [OG10, OSC12, Par10, Pat14, PG20, PT94, RL99, Rep99, RTF93, SSW09, Ser00, SLW^{+}07, She11, SJ92, SM13, SW16, Stä98, SCF^{+}13, TW93, Ull94, VH04, VMA21, Wad94, Wak07, Wam11, Waz10, Wei06, Xi07, ZDK^{+}15, Ano04, Fou96, Gam09, Rag20, Ric98, Ang03, Bur99, Ceb12, Han04, Har99b, Kol12, Lin00, Orc14, Pag97, Pag98, PT99, Ric98, San06, Tho99, Bha09, Der04, Fou96, Han01, MK97, Mel97, San09, Sim07, Spe24]. Programs [Kol12, Sch03a, SJW18, ASS96, ASS00, AA10, AFB92, ABR97, ANN97, AMNS22, Ben01, BS23, BRS94, CDHA14, DG00, Dan22, Dan23a, Fil03, FS03, GG99, Gib21, IT16, KV12, KM01, KSK08, MJ98b, RW93b, SMZ21, SM21, SSV08, SBHG10, VDH10, Wak98, Wak99, ZTX^{+}23, Mic07, Woo01]. project [Ach11, CF10, Men14]. project-based [Men14]. Projection [Lau91, Jon92b]. projections [LB96]. Promoting [KM01]. Proof [BB96, CD18, Fil21, Har99a, Har09a, MO14, Yi06, BS96, BD06, BCMT01, Chl08, Dan22, Dan23a, GPN19, GN91, GZND13, Kre04, Nee05, Pot13, Rea04]. Proof-directed [Fil21, Har99a, Har09a, Yi06]. Proof-irrelevance [BB96]. Proof-producing [MO14]. Proof-relevant [CD18]. Proofs [BJP12, AAC^{+}21, CK95, VDH10]. proper [BD97]. Propositions [Wad14]. Protocol [AS21]. Prototyping [MS95]. Prover [Tan20]. Provers [Mel97, AF04, Han99, LW99]. Proving [BL96, GM93, GMO06, Hut94]. pseudo [Sch15]. pseudo-random [Sch15]. Pseudoknot [HFA^{+}96]. PSSM [MG06]. PTIME [Mai04]. PTIME-completeness [Mai04]. Publications [San06, Sim07]. pull [SDK18]. pull-based [SDK18]. Pure [Ada06, Cou07, FP15, HP93, SH12, Bar99, BS00, BC06a, CH94a, GPN19, HP95a, MO14, Pol98]. Purely [FKS11, Hin09b, Bir20, BO96, CK04, CDHA14, DLP10, Han99, MR01, Oka95, Sab98, Hin99a]. purpose [Bra13, HP06]. Push [SDK18, MP06]. push/enter [MP06]. Pushdown [JSE^{+}14]. puzzle [Jon13]. Python [Cla15].

Quantifier [Mai92]. Quantitative [DHW05]. queries [CGMS06, SFL17]. query [SDK18, Won00]. Quest [CL91]. queues [BO96, Oka95]. Quick [SJCA17]. quickly [HLSZ^{+}16]. quote [Hin11]. quote/antiquote [Hin11].

R [Bai99, Gam11a, Hai03, Jon06, Klu01, Rea97, Sha04b, Wal04, dP01]. Rabhi [Rus03]. Rachel [Mit95]. Racket [F^{+}13, Jad14]. radix [Gib99]. Rafael [App97a]. railway [Str11]. rainfall [Ach21]. Ramakrishnan [Wal04]. Random [Pey03t, BP92, CDP15, Sch15, Swi20]. random-access [Swi20]. randomization [LP99]. rank [PVWS07]. Rational [Pey03d, UV17]. rationals [GLB06]. Raymond [Boi12]. re [NR12, ORT09]. re-evaluated [NR12]. re-examined [ORT09]. reachability [MS08]. Reactive [Sch04, Spe20, Vil07, Bah22, PG20, PN20, Tho00]. Read [BS23]. Read/write [BS23]. Ready [BSZL^{+}21]. Real [SMZ21, BSZL^{+}21]. Real-time [SMZ21]. real-world [BSZL^{+}21]. realistic [BJ93]. Realm [F^{+}13, Jad14]. Reasoning [Bud08, Col03, IT16, KMS21, Sha04a, ZTX^{+}23, AGNS21, Dan22, Dan23a, DNB12, GKM^{+}08, Har09b, HR04, OSC12, Sch03a, Car11, Tho04b]. rebinding [SSH^{+}08]. reconfigured [NR12]. Reconstruction [NP95, ANN97, Pie13, PS20]. record [Bet00, Bou04, KD14]. rectangle [Bir11]. recurrence [DL22]. recursion [AA02, AP16, DP12, DA23, HW16, Jon92c, Mcb03, NO99, Pit11]. Recursive [Dre07, GLP02, Bou04, Dan22, Dan23a, LP99, NS93]. recursively [NO99]. Red [Kah01, Oka99b, GM14, GK96]. Red-black [Kah01, Oka99b, GM14]. redex [SSH^{+}08]. redex-time [SSH^{+}08]. Reduceron [NR12]. reducing [Bar92a, GPN19]. reduction [AH08, Ben17, Bur91a, CD96, GPN19, GK96, KN95, Mor21, SSH^{+}08]. reductions [DD03a]. Refactoring [TL13]. Reference [CGR96, Pey03-27]. refinement [Cla96]. Refining [KN95]. reflective [GMO06]. refutation [NTHV17]. Region [DG02, EH21, TJ92]. regions [FM06]. register [BH20]. Regular [Hos06, HP03, ORT09, Fil21, McI04, TN15]. Regular-expression [ORT09]. relation [FD04]. Relational [QGG21, DNB12, MKJ09]. relations [BDH96]. relevant [CD18]. reliable [Kre04]. Remarks [BC06a]. removing [Bir91a]. Report [MME12, SDF^{+}09, ESM15, PW96, SM13, Waz10]. Reppy [Jon06, O’D01]. Representation [Sha04a, CX05, HM96, JO15, KD14, KV12, Kis20, Swi22, Thi02]. representations [GSWZ96, WS92]. Representing [Bar92b, LB96]. Research [Lin00]. Residual [Hue94]. resource [HG96, HS15]. restriction [KP17, STI24]. result [BFGJ04]. retrieval [ZSW95]. Retrieving [RT91]. Return [BH20, MKHT12, SF06]. return-time [MKHT12]. reusable [Ram03, RT91, SC21]. revealed [GLP02, HM18a, HM18b]. reversal [Hin00b]. reverse [Swi22]. Review [Ang03, App97a, Bai99, Bha09, Boi03, Boi12, Bro12, Bud08, Bur99, Car11, Cas13, Ceb11, Ceb12, Chi09, Cla15, De 11, Der04, Dup95, Dyc01, Gam09, Gam11a, Gre14, Gru15, Gru20, Hac20, Hag11, Hag12, Hai03, Han04, Han01, Har99b, Hin99a, Hud15, Hut94, Jad14, Jef98, Jef01, Jon92b, Jon06, Jon07, Kah98, Kah05, Kar03, Kel03, Kin07, Klo01, Klu01, Kol12, Lin00, Lou04, McK99, Mic00, Mic07, Mic13, Mil13, Mit95, O’D01, Oka99a, Orc14, Pag97, Pag98, Pat14, PT99, Rag13, Rag20, Rea04, Rus03, San09, San06, Sha04a, Sha04b, Sha06, Sim07, Slo09, Spe14, Spe20, Tah03a, Tan20, Ten97, Ten00, Tho02, Tho04a, Tho04b, Tho99, Tho16, Vil07, Wal04, Wat14, Woo01, dP01, vS00, Joh92, Spe24]. Reviews [Fou96, Gor96, Rea97, Ric98]. Revised [SDF^{+}09, MS01]. revisited [Bou04, Yi06]. rewrite [GM04, KM01, SFG14]. Rewriting [BN98, Kah05, Klo01, Bla18, Bou99, KM01, VYH^{+}10, WC03, Sha06]. Reynolds [Ten00]. Richard [App97a, Gru15, Gru20, Mil13, PT99]. right [Dan19, Dan23b]. Rischel [Klu01]. risks [CK04]. Road [San06, DvJE04]. Robert [De 11, Dup95, Wat14, DP22]. Robin [Jef01, Rea04]. Roles [BGP11]. Romeo [SW16]. rooted [Zei16]. Rose [Jad14]. Roy [Gam09]. RPC [CC19]. rule [Cla96, Lap98, LG05]. rule-based [LG05]. rules [SDPJS07, Waz10]. Run [SA14, AFB92, KD14, SGM^{+}03, Wei06]. Run-Time [SA14, AFB92, KD14, SGM^{+}03, Wei06]. Runciman [Pag97, Pag98]. Runtime [PN20, BLH16, HMP98, SMZ21, SZJ14]. Rutledge [Kin07]. Ryan [Bud08, Jad14].

S [Hai03, Mic00]. S1 [Ano09j, Ano09i]. SAC [Gre05]. Safe [BEJW16, Chi94, AAC^{+}21, BSO20, CMS03, PI09, Rhi09, SW16, Wei04, Wei06]. Safety [SSSS08]. San [Jad14]. Sandy [Hac20, Spe24]. Sanskrit [Hue05]. Satisfaction [NT01, MLC91, Mar04]. SB [Han04]. Scala [Ceb12, BSO20, Lew13, LS10, OG10, Pat14]. scalable [BLH16, SMT16]. scale [IEM^{+}22, MJ98b]. scheduler [SHMJ16]. scheduling [ACR16, BKP01, BRS94]. Schema [Sch03a]. Schematic [Kol12, AA10]. Scheme [FCF^{+}02, GS03, Hic04, HPR98, KD14, MF08, Nør05, Ser00, SDF^{+}09]. schemes [HW16]. Schmid [Mic07]. Schneider [Vil07]. School [Jon07, vS00, Hud00]. Schwichtenberg [Dyc01]. Science [Bud08, Bur99, Joh92, Jon92b, Tho99, FFFK04, Har93b, HR04]. scientific [BRL17, CG96]. scope [AAC^{+}21]. scope-safe [AAC^{+}21]. scoped [Pit11]. Scott [Jad14]. scribe [ESM15]. Scripting [Kel03, Bar00, Mei00]. search [Hin02, NT01, Rit91, Spi00, Spi09]. SECD [Vas05]. Second [Gam11a, Bud08, Gor96, JO15]. second-order [JO15]. Secrets [PM02]. section [CS16]. Secure [SCF^{+}13, BN05, SM21]. Security [AMS05, CKP05]. segment [CM15, MC21]. Segments [Ach21]. Seif [Gam09]. Seldin [Gam11a]. selection [Bir97b]. Self [Bar91b, ABD13, CDHA14, Mog92]. self-adjusting [ABD13, CDHA14]. Self-interpretation [Bar91b, Mog92]. Semantic [BGHL12, CL91, CK95, MW18, VHMW13]. semanticist [SNO^{+}10]. Semantics [Hof99, KH95, Orc14, TW00, Tah03b, ABD13, AAC^{+}21, Bou04, BRL17, Bre18, Bru94, Bur91b, CWM02, DL22, DP09, Fax02, Fer04, Fer06, FF99, Han93, Hut23c, JHH09, MF08, NH09, PM00, SR98, HPR98, GP98, Kah98, Sim07]. Send [CIZ22]. sensitive [GM04]. sensitivity [SM21]. separation [AMR09, JKJ^{+}18, NMB08, VPD21]. separation-logic-verified [VPD21]. September [Mil13]. sequence [Wak06]. sequences [UV17]. sequent [BG00, DA18]. serialisers [Hin23]. Series [Rea04, Mci99, Hag12]. serious [Mci99]. Serrano [Gre14]. Server [GBF05, Mei00, CC19, Mar02, ZSW95]. session [GV10]. sessions [Pad17, Wad14]. Sestoft [Tho16]. Set [BSZL^{+}21, WC01, Hoo92]. Setoids [BCP03]. sets [Ada93, BD97, Erw98]. setting [Oka99b, QGG21, Sch03b]. shallow [RH03]. shape [VHMW13]. Shared [Gre05]. sharing [ABR97, Ler96, Nar07]. Sheard [Tho99]. sheet [MBG^{+}20]. sheet-defined [MBG^{+}20]. Shifting [KKS11]. Short [Gam11b, Joh03, GJ07]. shortest [HG96]. Shrinking [AJ97]. Shriram [Wal04]. side [Mei00, PM00]. Sieve [Mee04, Nyk11, O’N09]. sieves [Run97]. sigma [HMP98]. Signature [Hag12, STI24]. signatures [Tof94]. similarity [Wak06]. Simon [Ric98]. Simple [Oka95, PT94, BS96, BD06, HP06, Mit91, Nee05, Pad17]. simpler [CKS09]. simplification [Mor21]. simplifying [ZMCN18]. simulates [GPN19]. Simulating [McB02]. simulation [LG05]. Sinclair [Gor96]. Single [Sch03b]. situ [Bir13]. sixth [Oka98]. sixth-order [Oka98]. Size [Bla18, MBG^{+}20]. Size-based [Bla18]. sized [Abe09, AP16, CLB23]. Skeleton [DHLB16]. skeletons [SHMB05]. Skribe [GS05]. Slideshow [FF06]. Small [NH09]. Small-step [NH09]. SMIL [BR04, Kin07]. SML [HR99, SHMB05, Klu01]. SMT [BGHL12]. snake [Jon13]. Soccer [Ach11]. Soccer-Fun [Ach11]. Software [Boi12, Jac06, Pag07, San09, Abr10, Arm07, RT91, Slo09]. solid [DD95]. solution [Cra20, CM15]. solve [Bir06]. solver [BGHL12]. Solving [Jon13, MLC91, SS08, WS92, Sha04a]. Some [CG96, Hud15]. Sons [App97a, Gor96, Kel03, Mit95, Sha04b]. sort [Gib99]. Sorted [Swi11]. sorting [Hen12, HM18b]. Sound [BM06, Gre96]. soundness [Pot13]. source [TO98]. SP [CH94b]. Space [Lüt03, SBHG10, Bah22, Jon92c, SNS16]. Sparcl [MW24]. Sparse [GSWZ96, WS92]. spatial [CCG05]. Special [ACW12, Ano04, AH92, AH93, AP13, BDT01, BDT04, CG18, DH00, DS16, HP93, HP95a, Hin04, Hin10b, Hue03, Hut00, Hut02b, KH18, MK97, Mel97, Oka01, SF14, SA14, SD17, TW00, TN10, TL11, Tri05, CS16]. specialization [RW93a, Thi03]. specialized [FM00]. specializers [Glü94, RW93a]. Specific [Hag11, Hag12, IEM^{+}22, Par10, WE09]. Specification [FH03, Gor96, Pey03b]. specifications [CK95, SJCA17, Sha04b]. Specifying [Wan93]. speed [BJD97]. speedup [HLL24]. Spencer [Jad14]. Spineless [Pey92]. spirals [Run97]. split [AGLK20]. splittable [Sch15]. splitting [BFR^{+}12]. Spreadsheet [Wak07, BAD^{+}01, DRvE95, MBG^{+}20, Tho16]. spreadsheets [EAKC06]. Springer [Bai99, Bur99, Der04, Dyc01, Jef98, Joh92, Kin07, Kol12, Lin00, Mic07, Rag20, Sha06, Tho04b, Tho99, Vil07, Wal04, dP01]. Springer-Verlag [Bai99, Bur99, Der04, Dyc01, Jef98, Joh92, Lin00, Mic07, Rag20, Tho04b, Tho99, Vil07, Wal04, dP01]. SprintPrint [Tho04a]. SQL [RA19]. Stable [SOWM19, DHLB16]. Stack [BN05, MCGW02, MCGW03, AS96, BDJ05, SDB21]. Stack-based [BN05, MCGW02, MCGW03]. stacks [BGP11]. stage [IT16, KKS11]. Staged [NP05, CKS09, MS03]. Staging [KKS11]. Standard [Fou96, Jon06, Mit95, Pey03z, Rea97, SMZ21, Har93a, SZJ14, App93, Ber93, MS95, MCP93, TA95]. Starch [Hud15, Jad14, Mic13]. start [Cla15]. starter [Ary94]. starter-kit [Ary94]. State [HP93, HP95a, CKP05, DNB12, Pot13]. State-of-The-Art [HP93, HP95a]. stateful [MO14]. Static [CGMS06, Bru94, DHW05, Fax02, SJW18, SCYZ24]. statistical [FS03]. step [AC09, NH09]. Stephen [Ang03]. Steps [LS10, Ceb12]. Sterling [Rea04]. STG [DP09]. StkTokens [SDB21]. stock [Pey92]. Stone [Rag20]. Strategic [SFG14]. strategies [SSTH18]. strategy [THLJ98]. stream [BD92, EHBR14, Hin10a, KSK08]. stream-processing [KSK08]. streams [Bir15]. Street [MW08]. Strict [Gam11b, NS93, Stä98]. strictness [BL96, EM95, SSSS08]. strings [McI04]. Strong [Ben95, BFG97, BBLRD96, GN91, Gui00, KR97, Nee05]. Strongly [TO98]. Strongly-typed [TO98]. Structural [Pit11, AA02, AH08, Mcb03, SW23]. Structure [ASS96, ASS00, Pey03l, SWM01, FFFK04, HP06, Kar03, Woo01]. structured [HW16, HPR98]. structures [BD92, MR01, Tro95, Hin99a]. Stuckey [Han01]. student [DP22]. students [JVV93, Mol93]. Study [HW16, AS96, BM05, BL96, GJL^{+}07, Hin10a, Mor21, PM00, WS92]. style [BSO20, Dan23b, Han99]. substitution [AH08, KR97]. substitutions [ACCL91, BBLRD96, FP15]. subsumption [Sch09]. subtypes [Mit91]. subtyping [Bet00, BGHL12, FP06, GLP02, KPS^{+}20, SW23]. Sudoku [Bir06]. Super [HR22]. Super-naturals [HR22]. supercompiler [SGJ96]. support [BLH16, GJL^{+}07, Gre05, OG10, Sch03b, SNO^{+}10]. surjective [CD96]. surviving [Oka03]. Sussman [Kar03, Woo01]. swapping [MKHT12]. symbolic [NTHV17]. symmetric [Hoo92]. Symposium [KR02]. synchronization [BKP01, BCG05]. synchronous [CIZ22]. synchrony [PTW22]. Syntactic [DD03b, Pot13, CG97, Ler96]. Syntax [AMR09, Pey03-27, BK01, DRR01, PP12, RH03, WW08]. syntaxes [AAC^{+}21]. Synthesis [Mic07, Sch03a, LG05]. Syropoulos [Ceb12]. System [Boi12, Pey03p, SH12, Abr10, AHK08, Aug93, Bar91c, BD92, FS03, GK96, Gol00, Jon95, Kea94, Kre04, Ler00, MS95, Pot13, SW16, Won00]. Systematic [CE18, Liu13, VM12, Spe14]. Systems [Cou07, Jef01, Sch04, SA14, Ada06, AZ02, AS21, BGT16, Bar91a, Bar99, BS00, BC06a, Cla96, CH94a, DHW05, DHLB16, FD04, FP15, GM04, GTN^{+}19, HMP98, HR04, LP03, MMJ20, PN20, Pol98, Wad94, WS92, Bud08, Kah05, Vil07].

T [Tah03a, Tho04a]. tabbing [Lap98]. tactic [ZDK^{+}15]. tag [EH21]. tag-free [EH21]. tagger [Hue05]. Tagless [Pey92, CKS09, DP12]. Tail [Jon92c, Bir93]. Taming [HZO21]. Target [SA14, Cle16]. Teaching [JVV93, TI15, CK04, Har93b]. Techniques [Gam09, Kah98, App97b, GM04, GP98, VH04, McK99]. Technologies [Ano04]. technology [Tho16]. Ted [Cla15]. template [SGM^{+}03]. template-based [SGM^{+}03]. Terence [Hag11]. Terese [Kah05]. Term [BN98, Sha06, Ben95, Sur07, Kah05, Klo01]. terminal [CD96]. Terminating [Rea93]. Termination [WC03, Bla18, BCMT01]. terms [Bar92a, BG00, GL13, GL15, KR97, Par16, Stä98, SB99, Zei16]. Testing [HLSZ^{+}16, AS21]. textual [ZSW95]. their [AAC^{+}21]. Theorem [Gib96b, Mel97, AF04, BD06, GM93, GMO06, Han99, Hed98, Hut94, JO15, LW99]. theoretic [HP95b, PT94, SJ92]. Theoretical [Bar91b, Bar92a, Bar92b, Bar91c, BD06, CH94b, Hin05, Nee05, Oka03]. theories [AH08, KM01, LP09, LP20, Ten00]. Theory [BDT01, BDT04, Kol12, MMJ20, Aba94, AA10, AMLH16, BCP03, BC06a, BJI17, BHM^{+}06, CC19, CCP95, CD18, FHJ98, Fil03, GV10, Hed98, HJP23, Hue94, Ler96, Lis94, Mat09, NMB08, NLA21, SVJB17, SC21, SF16, WE09]. these [HW07]. thesis [Ste21]. Think [Edm08, Ceb11]. Thinking [Hac20, Gru15]. thinning [JJ23]. Third [Gib96b, Chi09]. Thompson [Ric98]. threaded [FRRS10]. Three [Oka97]. throw [Cro99]. Thunks [HD97]. tidy [Gib96a]. Tight [AGLK20]. tiling [Bir04]. Till [Gor96]. Tim [Tho99, LTJ94]. Time [Pey03s, SA14, AFB92, AJ97, CM15, F^{+}13, FCDF12, Hen12, Hin11, KD14, MKHT12, OR04, Pal93, SSH^{+}08, SMZ21, STW21, SGM^{+}03, Wan93, Wei06]. timed [BD92]. Times [Pey03q]. TinkerType [LP03]. Tobias [Klo01]. Tofte [Rea04]. Together [FCDF12, STW21]. tolerance [SMT16]. tolerant [PG20]. Tom [Hut94, Kol12]. tool [SNO^{+}10, ZMCN18]. toolkit [Bru92, Hue05]. Tools [Bro12, TL13]. top [Hen12]. top-down [Hen12]. Topics [Sha06]. Total [Lis94, Hin23, SF16]. totally [Gam11b]. towers [FS08]. Trace [MF23, SSV08]. track [Str11]. trades [FSNB09]. trading [MW08]. Transactional [DF08]. transform [BM04, Fok11]. Transformation [GM04, DD03a, DD03b, Fil21, GH09, PW03a, Tur93]. transformational [RTF93]. Transformations [Tho04b, Col03, DMN07, NN91, Zhu94]. transformer [Bur91a]. transforming [Par16]. transition [Cla96]. transitions [Tur93]. Translation [KSK08, CH94a, Has00, HL99, MO14, SW23, TO98]. Transparent [Bou99, SMT16]. transport [HLB95]. Transporting [DM14]. treatment [PP12]. Tree [Hin09a, BKP01, BFR^{+}12, GM14, KSK08]. tree-processing [KSK08]. Trees [Hin09b, Bir97a, CCG05, Gib96a, HP99, Hin00b, Hin02, HP06, Hin18, HY11, Kah01, Ken96, Oka97, Oka99b]. Trends [Ang03]. triangulations [BBC^{+}01]. tries [Hin00a]. Trinder [Ang03]. triple [BGP11]. trivalent [Zei16]. Trust [ØP97]. trustful [AGNS21]. tutorial [DA18, Hut99]. twice [Hin05]. Two [Bir92a, SP04, DP09]. Two-level [SP04]. typable [SH12]. Type [ANN97, Bar99, BDT01, BDT04, Bet00, Cou07, DO02, HS15, Läu96, Mit91, De 20, Nar07, NP95, PI09, Rhi09, SH12, Wad94, Wei04, Wei06, ZPV06, Ada06, AAC^{+}21, AHK08, BGT16, Bar91a, BS00, BCP03, BC06a, BJZS10, BSO20, Bur91b, CE18, CD18, CH94a, CWM02, Di 93, Dre07, Fil03, FD04, FP15, GV10, GTN^{+}19, Han98, Hed98, HM94, HP95b, HJP23, Jon92a, JM00, KSK08, Ler96, Mat09, NMB08, NLA21, PVWS07, Pie13, PT94, Pol98, Pot13, RT91, SJW18, SF16, SS08, SW23, TJ92, VW10, VPSS11]. type- [AAC^{+}21, BSO20]. type-and-capability [Pot13]. type-and-effect [BGT16]. type-assignment [Di 93]. Type-based [HS15, ZPV06, Han98]. Type-checking [Bar99, DO02]. type-directed [SW23]. Type-Driven [De 20]. type-erasure [CWM02]. Type-safe [PI09, Rhi09, Wei04, Wei06]. type-theoretic [HP95b, PT94]. Typed [CGJF20, Hin11, Kam05, SJW18, Bra13, CKS09, DRR01, HM94, KV12, LP13, MCGW02, MCGW03, SC21, Thi02, TO98, VMA21, ZDK^{+}15]. typeful [CX05]. Types [Mit95, Pey03x, SSV08, SD17, Abe09, AP16, AF04, BBD98, BJP12, Bet00, BDH96, BC93, CMS03, Dun14, FP06, FI00, GV10, Har93a, HD00, ICEW22, Jon01, Kah01, KD14, KRAB22, Läu96, McB02, MKJ09, Nee05, NS93, OCR^{+}21, PP01, PVWS07, Ram11, Rit91, Sch09, SP04, SCF^{+}13, Swi08, TI15, VMA21, WDMT02, Xi07, Hac20]. Typing [HDM93, ACPR95, Bru94, CLB23, FBH17, KN96, O’H03, SCYZ24, SV13]. typings [AGLK20].

U [Mic07]. UCL [Pag97, Pag98]. Ullman [Fou96]. Ulmer [LP09]. Ulrich [Dyc01]. undecidability [BS96, Nar07]. undefined [Bar92b, Stä98]. Understanding [SDPJS07, TUmI15, Asp92]. unfold [Dan22, Dan23a]. Unfolding [Bir01a, Lis94, Zhu94]. Unger [Orc14]. unification [CD18, JJ98, Mcb03]. unified [GAM18, PP12]. unifier [ZS17]. Uniform [Nie00, CDP15]. Unifying [HW16, HP95b]. union [Dun14, PP01]. unique [ARS94]. uniqueness [OCR^{+}21]. united [Ros18]. univalence [VMA21]. Univerity [Jon07]. Universal [Sch03a]. universality [Hut99]. universe [AAC^{+}21, ZS17]. Universes [SFG14]. University [Bha09, Boi03, Boi12, Car11, Ceb11, Ceb12, Fou96, Gru15, Gru20, Har99b, Hin99a, Hut94, Jef01, Jon92b, Jon06, Kah98, Kah05, Klo01, McK99, Mic00, Mil13, O’D01, Rea97, Sha04a, Spe14, Ten00, Tho02, Wat14, vS00, SFG14]. unparsing [Dan98]. Unravelling [Bir92b]. unsolvable [Bar91c]. Untyped [EM95, GJ91]. unwraps [Gam11b]. update [SSH^{+}08, WC01]. updates [BKP01]. USA [Fou96, KR02]. usage [AHK08]. use [Har93b, Oka98]. useless [DG00]. useless-code [DG00]. user [AAS17, GS03, LW99]. users [SM13]. uses [FSNB09]. Using [Klu01, LLR93, LP99, NN91, Rea97, Rit91, Abe09, Dan22, Dan23a, Fil03, GBF05, GS06, HR99, JJ23, Joh98, Lew13, Mor18, Pat14, Ram11, SDB21, Tho00]. Usually [SJW18]. Utilities [Pey03k, Pey03i, Pey03j, Pey03m].

validation [PN20]. validity [CCG05]. value [KP17, SCF^{+}13]. value-dependent [SCF^{+}13]. values [PW96]. variable [MBG^{+}20]. variable-size [MBG^{+}20]. variables [Hug04]. vector [Swi22]. Verification [Fil03, Sch04, GY13, IB99, NTHV17, PN20, Vil07]. Verified [SM21, TN15, BBJR21, VPD21, TMK^{+}19]. verifiers [Chl08]. Verify [BSZL^{+}21]. Verifying [Swi11]. Verlag [Bai99, Bur99, Der04, Dyc01, Jef98, Joh92, Lin00, Mic07, Rag20, Sha06, Tho04b, Tho99, Vil07, Wal04, dP01]. version [MS01, PG20, Yi06]. versus [AS96, BJD97, DHLB16, SDK18]. vertex [IEM^{+}22]. vertex-centric [IEM^{+}22]. Very [Rag13, Whi13]. via [GPN19, Hic04, HLA20, HLL24, Kri06, NO99, PS17, SSH^{+}08, SDPJS07, TO98, Tro95, VHMW13, ZTX^{+}23]. Video [GBF05, Mor18]. Video-on-Demand [GBF05]. view [MM04]. Virding [Dup95]. Virtual [Cle16, Wak99]. virtues [JG95]. vision [MS95]. visual [BAD^{+}01]. visualisation [JM00]. visualizations [Pet21]. VoDKA [GBF05]. Vol [Jon92b]. Volume [Ano91a, Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, Joh92, Ang03, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano92c, Ano92b, Ano92e, Ano92d, Ano92g, Ano92f, Ano92i, Ano92h, Ano93c, Ano93b, Ano93e, Ano93d, Ano93g, Ano93f, Ano93i, Ano93h, Ano94c, Ano94b, Ano94e, Ano94d, Ano94g, Ano94f, Ano94i, Ano94h, Ano95c, Ano95b, Ano95e, Ano95d, Ano95g, Ano95f, Ano95i, Ano95h, Ano96c, Ano96b, Ano96e, Ano96d, Ano96g, Ano96f, Ano96i, Ano96h, Ano96k, Ano96j, Ano96m, Ano96l, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano09j, Ano09i, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h]. volume [Ano11g, Ano11j, Ano11i, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano12j, Ano12i, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano13j, Ano13i, Ano13l, Ano13k, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano14j, Ano14i]. volumes [Ang03]. vs [MP06].

W [Kel03, McK99]. Wachs [Bir20]. Wakeling [Pag97, Pag98]. Wall [MW08]. Wampler [Bro12]. Wand [Chi09, Tah03a]. want [Oka98]. wanted [Pie13]. Warnings [Mar07]. Weak [Gre96, FHJ98]. Weaving [HJ01]. Web [Cas13, Kin07, BR04, Hic04, HJ01, Kes12, Mar02, Mei00, Nør05]. weight [HY11]. weight-balanced [HY11]. Welcome [Fel03]. Well [AP16, SDB21, Swi22]. well-bracketed [SDB21]. Well-founded [AP16]. well-known [Swi22]. Wesley [Hag12, Klu01, Ric98, Rus03]. wheel [Run97]. Wheeler [BM04]. which [BBLRD96, KR97]. while [Nil01]. Whitington [Rag13]. Wiley [App97a, Gor96, Kel03, Mit95, Sha04b]. Williams [Dup95]. Wisdom [Kar03]. within [HMP98]. without [Bah22, CDP16, Jon92c, Jon01, Oka03]. witnesses [SJW18]. Wolfgang [Ang03]. words [TN15]. Work [FCDF12, SH14]. worked [TI15]. worker [GH09]. worker/wrapper [GH09]. workerg [Gam11b]. workerg/wrapper [Gam11b]. workflow [BRL17]. Working [Fou96, Rea97, Pau91, SNO^{+}10]. workout [Kis20]. Workshop [MME12, SM13]. World [Kel03, Arm07, Bar00, BSZL^{+}21, San09]. Worst [BRS94]. would [Oka98]. wrap [SH14]. wrapper [Gam11b, GH09]. write [BS23]. Wrong [SJW18]. www.pragprog.com [Hag11].

X [SS08, VPSS11, Wal04, BH92, Bud08, Kin07, San09]. X-MP [BH92]. X.media.publishing [Kin07]. xi [Rus03]. xiv [Har99b]. XML [Ano04, CGMS06, Hos06, HP03, Thi02]. xv [O’D01].

Yanhong [Spe14]. year [CK04, HB17, JVV93]. York [Gre14, Wat14]. Yun [Jad14].

Z [Gor96, Gor96]. Zachary [Cas13]. zero [BEJW16, PS17]. zero-cost [BEJW16]. Ziggurat [FS08]. Zipper [Hue97]. Zurg [Erw04].

References

Abel:2002:PAS

[AA02] Andreas Abel and Thorsten Altenkirch. A predicative analysis of structural recursion. Journal of Functional Programming, 12(1): 1–41, January 2002. CODEN JFPRES. ISSN 0956-7968 (print),

          1469-7653 (electronic). URL https://www.cambridge.org/core/
          product/D587A7DC2E95DFA4C8E1F977C73B92FE.

Addis:2010:DPT

[AA10] Tom R. Addis and Jan Addis. Drawing programs: the theory and practice of schematic functional programming. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84882-617-6, 1-84882-618-4 (e). xvi + 395 pp. LCCN QA76.62 .A375 2010. URL http:

          //bvbr.bib-bvb.de:8991/F?func=service\%26doc\%5Flibrary=
          BVB01\%26doc\%5Fnumber=017646367\%26line\%5Fnumber=0001\
          %26func\%5Fcode=DB\%5FRECORDS\%26service\%5Ftype=MEDIA;
           http://www.loc.gov/catdir/enhancements/fy1208/2009927007-
          d.html.

Allais:2021:TSS

[AAC^{+}21] Guillaume Allais, Robert Atkey, James Chapman, Conor Mcbride, and James Mckinna. A type- and scope-safe universe of syntaxes with binding: their semantics and proofs. Journal of Functional Programming, 31(??):e22, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/type-and-scopesafe-universe-of-syntaxes-
with-binding-their-semantics-and-proofs/8A0865F34313BA65F4FE46D4522B4568.

Abel:2017:IPA

[AAS17] Andreas Abel, Stephan Adelsberger, and Anton Setzer. Interactive programming in Agda — objects and graphical user interfaces. Journal of Functional Programming, 27:e8, ???? 2017. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/interactive-programming-
in-agda-objects-and-graphical-user-interfaces/56ECE95C3A0C208D5ABCD072643BC8FB.

Abadi:1994:BMT

[Aba94] Martin Abadi. Baby Modula-3 and a theory of objects. Journal of Functional Programming, 4(2):249–283, April 1994. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/7091DC3F4FC43DC753EE1D5372D8C38C.

Acar:2013:CSS

[ABD13] Umut A. Acar, Matthias Blume, and Jacob Donham. A consistent semantics of self-adjusting computation. Journal of Functional Programming, 23(3):249–292, May 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/441A28C813BDA23B57F1ED2BB1A7E36E.

Abel:2009:INU

[Abe09] Andreas Abel. Implementing a normalizer using sized heterogeneous types. Journal of Functional Programming, 19(3–4):287–310, July 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
150DEA4898B591CAA5FA031289C09571.

Aharoni:1997:CAM

[ABR97] Gad Aharoni, Amnon Barak, and Amir Ronen. A competitive algorithm for managing sharing in the distributed execution of functional programs. Journal of Functional Programming, 7(4):421–440, July 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4474555DB08311874146D1752D2DEA5E.

Abrial:2010:MEB

[Abr10] Jean-Raymond Abrial. Modeling in Event-B: system and software engineering. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-89556-1 (hardcover). xxvi + 586 pp. LCCN QA76.9.F67

A27 2010. URL http://assets.cambridge.org/97805218/95569/
cover/9780521895569.jpg.

Altenkirch:2009:BSN

[AC09] Thorsten Altenkirch and James Chapman. Big-step normalisation. Journal of Functional Programming, 19(3–4):311–333, July 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
749A35BCD063A03AE16897256E08DD09.

Abadi:1991:ES

[ACCL91] M. Abadi, L. Cardelli, P.-L. Curien, and J.-J. Lévy. Explicit substitutions. Journal of Functional Programming, 1(4):375–416, October 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C1B1AFAE8F34C953C1B2DF3C2D4C2125.

Abel:2020:LEI

[ACD^{+}20] Andreas Abel, Jesper Cockx, Dominique Devriese, Amin Timany, and Philip Wadler. Leibniz equality is isomorphic to Martin-Löf identity, parametrically. Journal of Functional Programming, 30(??):e17, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
leibniz-equality-is-isomorphic-to-martinlof-identity-parametrically/
50D76A9C314AB24E7CD46DA4A0A766EB.

Achten:2011:SFP

[Ach11] Peter Achten. The soccer-fun project. Journal of Functional Programming, 21(1):1–19, January 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/CA981494BF6064CEA47E819C69603626.

Achten:2021:SAR

[Ach21] Peter Achten. Segments: an alternative rainfall problem. Journal of Functional Programming, 31(??):e23, ???? 2021. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/segments-an-alternative-
rainfall-problem/ABEA634EB9763953CBCC8D2AC58FE710.

Abadi:1995:DTP

[ACPR95] M. Abadi, L. Cardelli, B. Pierce, and D. Rémy. Dynamic typing in polymorphic languages. Journal of Functional Programming, 5(1): 111–130, January 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/E9028529B70C236AAF53C6136A28ACA0.

Acar:2016:OGS

[ACR16] Umut A. Acar, Arthur Charguéraud, and Mike Rainey. Oracleguided scheduling for controlling granularity in implicitly parallel languages. Journal of Functional Programming, 26:e23, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
oracleguided-scheduling-for-controlling-granularity-in-
implicitly-parallel-languages/0316B0F89CD4C2006713CDA608146300.

Acar:2012:ESI

[ACW12] Umut A. Acar, James Cheney, and Stephanie Weirich. Editorial: Special issue dedicated to ICFP 2010. Journal of Functional Programming, 22(4–5):379–381, September 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/08DC440E00041469B2A815F5E330CBC5.

Aransay:2015:FHO

[AD15] Jesús Aransay and Jose Divasón. Formalisation in higher-order logic and code generation to functional languages of the Gauss– Jordan algorithm. Journal of Functional Programming, 25:e9, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/formalisation-
in-higherorder-logic-and-code-generation-to-functional-
languages-of-the-gaussjordan-algorithm/2B1780CA4588D531821AFD950C782848.

Adams:1993:FPE

[Ada93] Stephen Adams. Functional pearls: Efficient sets — a balancing act. Journal of Functional Programming, 3(4):553–561, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0CAA1C189B4F7C15CE9B8C02D0D4B54E.

Adams:2006:PTS

[Ada06] Robin Adams. Pure type systems with judgemental equality. Journal of Functional Programming, 16(2):219–246, March 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
43C76AF58785F4F6EC41F4D94FFE74C6.

Ariola:1997:CNL

[AF97] Zena M. Ariola and Matthias Felleisen. The call-by-need lambda calculus. Journal of Functional Programming, 7(3):265–301, May 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F4FC3C34E9CAE3F4326503E254FCF6F2.

Appel:2004:DTE

[AF04] Andrew W. Appel and Amy P. Felty. Dependent types ensure partial correctness of theorem provers. Journal of Functional Programming, 14(1):3–19, January 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C968F39F6E1649C722A4572FAB862F0C.

Aharoni:1992:RTA

[AFB92] Gad Aharoni, Dror G. Feitelson, and Amnon Barak. A run-time algorithm for managing the granularity of parallel functional programs. Journal of Functional Programming, 2(4):387–405, October 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4BC37BC8A62FAAEAE672A699179AA086.

Altenkirch:2015:IC

[AGH^{+}15] Thorsten Altenkirch, Neil Ghani, Peter Hancock, Conor Mcbride, and Peter Morris. Indexed containers. Journal of Functional Programming, 25:e5, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
indexed-containers/FB9C7DC88A65E7529D39554379D9765F.

Accattoli:2020:TTS

[AGLK20] Beniamino Accattoli, Stéphane Graham-Lengrand, and Delia Kesner. Tight typings and split bounds, fully developed. Journal of Functional Programming, 30(??):e14, ???? 2020. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/tight-typings-and-split-
bounds-fully-developed/95BD54E1AF44234025F3462836B7D83D.

Asperti:1996:BOH

[AGN96] Andrea Asperti, Cecilia Giovannetti, and Andrea Naletto. The Bologna optimal higher-order machine. Journal of Functional Programming, 6(6):763–810, November 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/1F2763B0F931680F9B15BDC750BEB343.

Affeldt:2021:TMA

[AGNS21] Reynald Affeldt, Jacques Garrigue, David Nowak, and Takafumi Saikawa. A trustful monad for axiomatic reasoning with probability and nondeterminism. Journal of Functional Programming, 31(??):e17, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
trustful-monad-for-axiomatic-reasoning-with-probability-
and-nondeterminism/315018412FBB338E9B47BB55B2E55107.

Appel:1992:SIM

[AH92] Andrew W. Appel and Robert Harper. Special issue on ML. Journal of Functional Programming, 2(1):i, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
899F7BB7F3038E11CA6BF76E01A856D6.

Appel:1993:SIM

[AH93] Andrew W. Appel and Robert Harper. Special issue on ML. Journal of Functional Programming, 3(4):389, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B3F49FA29A27A6A856F4766418ACE209.

Antoy:2004:FPC

[AH04] Sergio Antoy and Michael Hanus. Functional pearl: Concurrent distinct choices. Journal of Functional Programming, 14(6):657– 668, November 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/679F409229981E36DF95372B66BD6EC5.

Ariola:2008:CRT

[AH08] Zena M. Ariola and Hugo Herbelin. Control reduction theories: the benefit of structural substitution. Journal of Functional Programming, 18(3):373–419, May 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/A983A64C551FEB771405F9FEBAD09F65.

Aspinall:2008:TSU

[AHK08] David Aspinall, Martin Hofmann, and Michal Konecný. A type system with usage aspects. Journal of Functional Programming, 18(2):141–178, March 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/D0D72E86D1C1CEBB7475F8EB1C25E2D4.

Aehlig:2012:CIN

[AHN12] Klaus Aehlig, Florian Haftmann, and Tobias Nipkow. A compiled implementation of normalisation by evaluation. Journal of Functional Programming, 22(1):9–30, January 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/599BCD246F47455F1A804C41F6A83C21.

Appel:1997:SLE

[AJ97] Andrew W. Appel and Trevor Jim. Shrinking lambda expressions in linear time. Journal of Functional Programming, 7(5):515– 540, September 1997. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/3D535161F9664C7CBA184D5EE901D14C.

Atkey:2015:IDE

[AJ15] Robert Atkey and Patricia Johann. Interleaving data and effects. Journal of Functional Programming, 25:e20, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic).  URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/interleaving-
data-and-effects/FCDD7EDCF7E7A9484C59A13B53F75742.

Allison:2005:MML

[All05] Lloyd Allison. Models for machine learning and data mining in functional programming. Journal of Functional Programming, 15(1): 15–32, January 2005. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0E075DBAAD53F0868560C3DD93698440.

Angiuli:2016:HPT

[AMLH16] Carlo Angiuli, Edward Morehouse, Daniel R. Licata, and Robert Harper. Homotopical patch theory. Journal of Functional Programming, 26:e18, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
homotopical-patch-theory/42AD8BB8A91688BCAC16FD4D6A2C3FE7.

Annenkov:2022:EFP

[AMNS22] Danil Annenkov, Mikkel Milo, Jakob Botsch Nielsen, and Bas Spitters. Extracting functional programs from Coq, in Coq. Journal of Functional Programming, 32:e11, ???? 2022. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/extracting-functional-programs-from-
coq-in-coq/ACA1A3F43DD4EA96646F705BC2F8E370.

Arbiser:2009:CCS

[AMR09] Ariel Arbiser, Alexandre Miquel, and Alejandro Ríos. The λcalculus with constructors: Syntax, confluence and separation. Journal of Functional Programming, 19(5):581–631, September 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BEDEC96FA93DA749AB8C595A023F008B.

Abadi:2005:LBS

[AMS05] Martín Abadi, Greg Morrisett, and Andrei Sabelfeld. “Languagebased security”. Journal of Functional Programming, 15(2):129, March 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
28F1B5E40F21B5B3C1AF525FCDB623C5.

Agha:1997:FAC

[AMST97] Gul A. Agha, Ian A. Mason, Scott F. Smith, and Carolyn L. Talcott. A foundation for actor computation. Journal of Functional Programming, 7(1):1–72, January 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/E9A5266BA5D37A1856D50C939679F31C.

Angus:2003:BRT

[Ang03] Chris Angus. Book review: Trends in Functional Programming (volumes 1 & 2) by Greg Michaelson, Phil Trinder and Hans-Wolfgang Loidl (editors volume 1), and Stephen Gilmore (editor volume 2). Intellect Books, Bristol, 2001, 2002. Journal of Functional Programming, 13(4):823–824, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/FF4F5DD1D38A1EE30F5BA641BC21F0E8.

Amtoft:1997:TBR

[ANN97] Torben Amtoft, Flemming Nielson, and Hanne Riis Nielson. Type and behaviour reconstruction for higher-order concurrent programs. Journal of Functional Programming, 7(3):321–347, May 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
8F4E3EC214A87119B3C94E00FCE636FE.

Anonymous:1991:AIV

[Ano91a] Anonymous. Author index to volume 1. Journal of Functional Programming, 1(4):495, October 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/1E3F38DB456FF1B271C27919F3E0B468.

Anonymous:1991:Ea

[Ano91b] Anonymous. Editorial. Journal of Functional Programming, 1(1): 1–2, January 1991. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/80CF4ECDEFB3B3454E0AD35200B5DA85.

Anonymous:1991:Eb

[Ano91c] Anonymous. Errataum. Journal of Functional Programming, 1 (3):373, July 1991. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/220A820B0E326ABCEDDFBD8B0BAAEF96.

Anonymous:1991:JVIb

[Ano91d] Anonymous. JFP volume 1 issue 1 cover and back matter. Journal of Functional Programming, 1(1):b1–b6, January 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
306D549E961E36117E7D9E9957047BB1.

Anonymous:1991:JVIa

[Ano91e] Anonymous. JFP volume 1 issue 1 cover and front matter. Journal of Functional Programming, 1(1):f1–f2, January 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2A531AA8AB7885633115BC00A83F4AFE.

Anonymous:1991:JVId

[Ano91f] Anonymous. JFP volume 1 issue 2 cover and back matter. Journal of Functional Programming, 1(2):b1–b2, April 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BA9A1ED222FACCD3B24554C0BC316264.

Anonymous:1991:JVIc

[Ano91g] Anonymous. JFP volume 1 issue 2 cover and front matter. Journal of Functional Programming, 1(2):f1–f2, April 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
726AFBB7468E620EB9D8689A825CEF50.

Anonymous:1991:JVIf

[Ano91h] Anonymous. JFP volume 1 issue 3 cover and back matter. Journal of Functional Programming, 1(3):b1–b4, July 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
120031C6DFBBBCDA229925B1D6B3EE6A.

Anonymous:1991:JVIe

[Ano91i] Anonymous. JFP volume 1 issue 3 cover and front matter. Journal of Functional Programming, 1(3):f1–f2, July 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D4EA01EECB66B9830F8D034F30A5A8B5.

Anonymous:1991:JVIh

[Ano91j] Anonymous. JFP volume 1 issue 4 cover and back matter. Journal of Functional Programming, 1(4):b1–b7, October 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2548710DB08DE323F3215EC6ABE8450D.

Anonymous:1991:JVIg

[Ano91k] Anonymous. JFP volume 1 issue 4 cover and front matter. Journal of Functional Programming, 1(4):f1–f2, October 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
25C4D8EDA27889757D6C2485E347A238.

Anonymous:1992:AIV

[Ano92a] Anonymous. Author index to volume 2. Journal of Functional Programming, 2(4):515, October 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/FFCC72D11DECDEC984C88AA62DC10BA5.

Anonymous:1992:JVIb

[Ano92b] Anonymous. JFP volume 2 issue 1 cover and back matter. Journal of Functional Programming, 2(1):b1–b3, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
105EF62E27A41601CFDF851224DF4409.

Anonymous:1992:JVIa

[Ano92c] Anonymous. JFP volume 2 issue 1 cover and front matter. Journal of Functional Programming, 2(1):f1–f2, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
658393D8742F5B78BD2BFAC2DAB0AA6E.

Anonymous:1992:JVId

[Ano92d] Anonymous. JFP volume 2 issue 2 cover and back matter. Journal of Functional Programming, 2(2):b1–b4, April 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C91356084B14E5B9B78B5A7FF2C13C05.

Anonymous:1992:JVIc

[Ano92e] Anonymous. JFP volume 2 issue 2 cover and front matter. Journal of Functional Programming, 2(2):f1–f2, April 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9DDAAFCADFAD5BF0F6BEB51C6DE6F54A.

Anonymous:1992:JVIf

[Ano92f] Anonymous. JFP volume 2 issue 3 cover and back matter. Journal of Functional Programming, 2(3):b1–b4, July 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E62DE76A992240B79655C3B91D4825D4.

Anonymous:1992:JVIe

[Ano92g] Anonymous. JFP volume 2 issue 3 cover and front matter. Journal of Functional Programming, 2(3):f1–f2, July 1992. CODEN JFPRES.

ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/98E9DCE8E7E917147C2E78B9D107345F.

Anonymous:1992:JVIh

[Ano92h] Anonymous. JFP volume 2 issue 4 cover and back matter. Journal of Functional Programming, 2(4):b1–b3, October 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
40928BB55459F38F4097AEA764AE2E5C.

Anonymous:1992:JVIg

[Ano92i] Anonymous. JFP volume 2 issue 4 cover and front matter. Journal of Functional Programming, 2(4):f1–f2, October 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
04688C64C5B84F8800AB6EE3E0740779.

Anonymous:1993:AIV

[Ano93a] Anonymous. Author index to volume 3. Journal of Functional Programming, 3(4):563, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/EDE3B2B689AECE1FB130156AC36340A2.

Anonymous:1993:JVIb

[Ano93b] Anonymous. JFP volume 3 issue 1 cover and back matter. Journal of Functional Programming, 3(1):b1–b4, January 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E3EF5CEDBB269233D1B5819615E0E0D0.

Anonymous:1993:JVIa

[Ano93c] Anonymous. JFP volume 3 issue 1 cover and front matter. Journal of Functional Programming, 3(1):f1–f2, January 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
ED5D8FDB9A869C0284B734782815AB9E.

Anonymous:1993:JVId

[Ano93d] Anonymous. JFP volume 3 issue 2 cover and back matter. Journal of Functional Programming, 3(2):b1–b2, April 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
34D18B462A0EC0FF4B21FD1AC083C2BE.

Anonymous:1993:JVIc

[Ano93e] Anonymous. JFP volume 3 issue 2 cover and front matter. Journal of Functional Programming, 3(2):f1–f2, April 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
76B6D0CE428C671C6BDB060B653C1BED.

Anonymous:1993:JVIf

[Ano93f] Anonymous. JFP volume 3 issue 3 cover and back matter. Journal of Functional Programming, 3(3):b1–b2, July 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
AB15602677A6658F8C5CC8324E5658C4.

Anonymous:1993:JVIe

[Ano93g] Anonymous. JFP volume 3 issue 3 cover and front matter. Journal of Functional Programming, 3(3):f1–f2, July 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7EEFC0B6AAFAD945F1F481939C7F56CA.

Anonymous:1993:JVIh

[Ano93h] Anonymous. JFP volume 3 issue 4 cover and back matter. Journal of Functional Programming, 3(4):b1–b3, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
800F985FBB186ECEA4EA0F06AAAD1A30.

Anonymous:1993:JVIg

[Ano93i] Anonymous. JFP volume 3 issue 4 cover and front matter. Journal of Functional Programming, 3(4):f1–f2, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
46159BFF9821BBBE1CE88448760D83D5.

Anonymous:1994:AIV

[Ano94a] Anonymous. Author index to volume 4. Journal of Functional Programming, 4(4):561, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4AB8656771DAE4908679CD034E4D103F.

Anonymous:1994:JVIb

[Ano94b] Anonymous. JFP volume 4 issue 1 cover and back matter. Journal of Functional Programming, 4(1):b1–b4, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D3DB7278C1E6BDD760BA31EC36527795.

Anonymous:1994:JVIa

[Ano94c] Anonymous. JFP volume 4 issue 1 cover and front matter. Journal of Functional Programming, 4(1):f1–f2, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BF28AC16B29EB3BB959485F16770DF36.

Anonymous:1994:JVId

[Ano94d] Anonymous. JFP volume 4 issue 2 cover and back matter. Journal of Functional Programming, 4(2):b1–b2, April 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6CF24CBF46D704648FC37757FC80F541.

Anonymous:1994:JVIc

[Ano94e] Anonymous. JFP volume 4 issue 2 cover and front matter. Journal of Functional Programming, 4(2):f1–f2, April 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
22F19A32E680929D1C24D52A9C085BC6.

Anonymous:1994:JVIf

[Ano94f] Anonymous. JFP volume 4 issue 3 cover and back matter. Journal of Functional Programming, 4(3):b1–b3, July 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
43C9A8E5BBF7F5A066B3BCB8897B14C2.

Anonymous:1994:JVIe

[Ano94g] Anonymous. JFP volume 4 issue 3 cover and front matter. Journal of Functional Programming, 4(3):f1–f2, July 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/21C92EC22FBE340B6190C7A98F1F4595.

Anonymous:1994:JVIh

[Ano94h] Anonymous. JFP volume 4 issue 4 cover and back matter. Journal of Functional Programming, 4(4):b1–b2, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B067EBD582C763E93278FECE8F49DB48.

Anonymous:1994:JVIg

[Ano94i] Anonymous. JFP volume 4 issue 4 cover and front matter. Journal of Functional Programming, 4(4):f1–f2, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E1F8A54036C3B73797AB9C374D8215B5.

Anonymous:1995:AIV

[Ano95a] Anonymous. Author index to volume 5. Journal of Functional Programming, 5(4):661–662, October 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/B90DF8FDC92301088D5026D239149807.

Anonymous:1995:JVIb

[Ano95b] Anonymous. JFP volume 5 issue 1 cover and back matter. Journal of Functional Programming, 5(1):b1–b2, January 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EE67D0B74F35BADDA49CA8EC6270C365.

Anonymous:1995:JVIa

[Ano95c] Anonymous. JFP volume 5 issue 1 cover and front matter. Journal of Functional Programming, 5(1):f1–f2, January 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
995D288CCD012B12EBA2FB9446A675C3.

Anonymous:1995:JVId

[Ano95d] Anonymous. JFP volume 5 issue 2 cover and back matter. Journal of Functional Programming, 5(2):b1–b3, April 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
020A7A56B52855E9F008576F660EA1D4.

Anonymous:1995:JVIc

[Ano95e] Anonymous. JFP volume 5 issue 2 cover and front matter. Journal of Functional Programming, 5(2):f1–f2, April 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2137D237C90F844E9623B148D5FE5CD6.

Anonymous:1995:JVIf

[Ano95f] Anonymous. JFP volume 5 issue 3 cover and back matter. Journal of Functional Programming, 5(3):b1–b4, July 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
5FEF97875B112C66E80B38A1321792CF.

Anonymous:1995:JVIe

[Ano95g] Anonymous. JFP volume 5 issue 3 cover and front matter. Journal of Functional Programming, 5(3):f1–f2, July 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/1AB3AB176193F00CB79CE8A8395C7A4A.

Anonymous:1995:JVIh

[Ano95h] Anonymous. JFP volume 5 issue 4 cover and back matter. Journal of Functional Programming, 5(4):b1–b5, October 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
937005EE721C3260BA97ABBDD31E5D3A.

Anonymous:1995:JVIg

[Ano95i] Anonymous. JFP volume 5 issue 4 cover and front matter. Journal of Functional Programming, 5(4):f1–f2, October 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
5B4C685765870A98B8135A36E7F82870.

Anonymous:1996:AIV

[Ano96a] Anonymous. Author index to volume 6. Journal of Functional Programming, 6(6):859–861, November 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/16F8BEEE294BB88AE215DB5C30E3577E.

Anonymous:1996:JVIb

[Ano96b] Anonymous. JFP volume 6 issue 1 cover and back matter. Journal of Functional Programming, 6(1):b1–b5, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E13A8A5806B51DDCD980A60F6A40FE03.

Anonymous:1996:JVIa

[Ano96c] Anonymous. JFP volume 6 issue 1 cover and front matter. Journal of Functional Programming, 6(1):f1–f2, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C6CCF392B68049AF093AC45134F8E475.

Anonymous:1996:JVId

[Ano96d] Anonymous. JFP volume 6 issue 2 cover and back matter. Journal of Functional Programming, 6(2):b1–b3, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6DC3FB49813FA8C539A730C01E9E1B8E.

Anonymous:1996:JVIc

[Ano96e] Anonymous. JFP volume 6 issue 2 cover and front matter. Journal of Functional Programming, 6(2):f1–f2, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E48B15E6E336D35502F74FADCA9A4EB7.

Anonymous:1996:JVIf

[Ano96f] Anonymous. JFP volume 6 issue 3 cover and back matter. Journal of Functional Programming, 6(3):b1–b2, May 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2F8A3AAF3014695167743E3D4C1169CE.

Anonymous:1996:JVIe

[Ano96g] Anonymous. JFP volume 6 issue 3 cover and front matter. Journal of Functional Programming, 6(3):f1–f2, May 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/40261842A8C32416DFEEEAF4C80EAD59.

Anonymous:1996:JVIh

[Ano96h] Anonymous. JFP volume 6 issue 4 cover and back matter. Journal of Functional Programming, 6(4):b1–b3, July 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3EB0CC6131DE72748E223E9475C4D7D7.

Anonymous:1996:JVIg

[Ano96i] Anonymous. JFP volume 6 issue 4 cover and front matter. Journal of Functional Programming, 6(4):f1–f2, July 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D6E2495CBEECBA3F6F3032FD514196BE.

Anonymous:1996:JVIj

[Ano96j] Anonymous. JFP volume 6 issue 5 cover and back matter. Journal of Functional Programming, 6(5):b1–b3, September 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
5AF48DD323991C011878EB53AC9257D5.

Anonymous:1996:JVIi

[Ano96k] Anonymous. JFP volume 6 issue 5 cover and front matter. Journal of Functional Programming, 6(5):f1–f2, September 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
03D6C5AA87690B4138937A6B71B71A2D.

Anonymous:1996:JVIl

[Ano96l] Anonymous. JFP volume 6 issue 6 cover and back matter. Journal of Functional Programming, 6(6):b1–b3, November 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9EDC77B4150B421FC314FA2DDFE32820.

Anonymous:1996:JVIk

[Ano96m] Anonymous. JFP volume 6 issue 6 cover and front matter. Journal of Functional Programming, 6(6):f1–f2, November 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2FF9953B349723B1276EDBEC719BCD5D.

Anonymous:2004:SIP

[Ano04] Anonymous. Special issue on “Programming Language Technologies for XML”. Journal of Functional Programming, 14(4):473, July 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1FB130234AAACC67FE431FE92BFBED88.

Anonymous:2009:JVIb

[Ano09a] Anonymous. JFP volume 19 issue 2 cover and back matter. Journal of Functional Programming, 19(2):b1–b2, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2214A553E81ED56DCD67591F95C2FEBC.

Anonymous:2009:JVIa

[Ano09b] Anonymous. JFP volume 19 issue 2 cover and front matter. Journal of Functional Programming, 19(2):f1–f2, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
CCF2A1E28D143A2C9567BC1AA4D0B6EA.

Anonymous:2009:JVId

[Ano09c] Anonymous. JFP volume 19 issue 3-4 cover and back matter. Journal of Functional Programming, 19(3–4):b1–b8, July 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
949BB6996BD42079FFC60726501DC8F9.

Anonymous:2009:JVIc

[Ano09d] Anonymous. JFP volume 19 issue 3-4 cover and front matter. Journal of Functional Programming, 19(3–4):f1–f2, July 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
73BB6230DA2BA46A19FCC202558B8026.

Anonymous:2009:JVIf

[Ano09e] Anonymous. JFP volume 19 issue 5 cover and back matter. Journal of Functional Programming, 19(5):b1–b3, September 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
A093F301A0D9164C38D1D33BFB672B66.

Anonymous:2009:JVIe

[Ano09f] Anonymous. JFP volume 19 issue 5 cover and front matter. Journal of Functional Programming, 19(5):f1–f2, September 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BFC508D20D9F9714A48AA17E385125E9.

Anonymous:2009:JVIh

[Ano09g] Anonymous. JFP volume 19 issue 6 cover and back matter. Journal of Functional Programming, 19(6):b1–b5, November 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
81EBE94A4A8E1DF19D935823AE5535A8.

Anonymous:2009:JVIg

[Ano09h] Anonymous. JFP volume 19 issue 6 cover and front matter. Journal of Functional Programming, 19(6):f1–f2, November 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
47A4D1A3CFDBED52AE0ADF1C4D0E7C7C.

Anonymous:2009:JVIj

[Ano09i] Anonymous. JFP volume 19 issue S1 cover and back matter. Journal of Functional Programming, 19(S1):b1–b2, August 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
CB73556BFF33F882F59518144E6576C5.

Anonymous:2009:JVIi

[Ano09j] Anonymous. JFP volume 19 issue S1 cover and front matter. Journal of Functional Programming, 19(S1):f1–f2, August 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
934A7AD08A1352E730CE000FC39424BB.

Anonymous:2010:JVIb

[Ano10a] Anonymous. JFP volume 20 issue 1 cover and back matter. Journal of Functional Programming, 20(1):b1–b8, January 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0736351948D60CF636E802F9F62D3017.

Anonymous:2010:JVIa

[Ano10b] Anonymous. JFP volume 20 issue 1 cover and front matter. Journal of Functional Programming, 20(1):f1–f2, January 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E7BF720012E19F74576EC17C3ABD7D7E.

Anonymous:2010:JVId

[Ano10c] Anonymous. JFP volume 20 issue 2 cover and back matter. Journal of Functional Programming, 20(2):b1–b10, March 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1F9E0C6A7BFCF29738C4A1423A7687F7.

Anonymous:2010:JVIc

[Ano10d] Anonymous. JFP volume 20 issue 2 cover and front matter. Journal of Functional Programming, 20(2):f1–f2, March 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
02F8CD0A456EC0E07423E94CD660F40F.

Anonymous:2010:JVIf

[Ano10e] Anonymous. JFP volume 20 issue 3-4 cover and back matter. Journal of Functional Programming, 20(3–4):b1–b7, July 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2F217FC44B081E84B646FB72B5121C6D.

Anonymous:2010:JVIe

[Ano10f] Anonymous. JFP volume 20 issue 3-4 cover and front matter. Journal of Functional Programming, 20(3–4):f1–f2, July 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
FCF76C4033D2A9A8A9388000F2D524B0.

Anonymous:2010:JVIh

[Ano10g] Anonymous. JFP volume 20 issue 5-6 cover and back matter. Journal of Functional Programming, 20(5–6):b1–b8, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
49F6932EB019F410190A0F2F09D9E6C2.

Anonymous:2010:JVIg

[Ano10h] Anonymous. JFP volume 20 issue 5-6 cover and front matter. Journal of Functional Programming, 20(5–6):f1–f2, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D77EF8288C179B0F3C0D634487C6FD87.

Anonymous:2011:JVIb

[Ano11a] Anonymous. JFP volume 21 issue 1 cover and back matter. Journal of Functional Programming, 21(1):b1–b5, January 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
41187A86C66A7F0793064A28618B169F.

Anonymous:2011:JVIa

[Ano11b] Anonymous. JFP volume 21 issue 1 cover and front matter. Journal of Functional Programming, 21(1):f1–f2, January 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EF679DA8B56C0FE5D51652AB9E49A2DC.

Anonymous:2011:JVId

[Ano11c] Anonymous. JFP volume 21 issue 2 cover and back matter. Journal of Functional Programming, 21(2):b1–b6, March 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7314D0958EE1F943A0335ECA6DFB38F4.

Anonymous:2011:JVIc

[Ano11d] Anonymous. JFP volume 21 issue 2 cover and front matter. Journal of Functional Programming, 21(2):f1–f2, March 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4C4F55251D6966E4D3C09462935150DF.

Anonymous:2011:JVIf

[Ano11e] Anonymous. JFP volume 21 issue 3 cover and back matter. Journal of Functional Programming, 21(3):b1–b3, May 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E5BDBED4A7C54E07A8D1478255C51640.

Anonymous:2011:JVIe

[Ano11f] Anonymous. JFP volume 21 issue 3 cover and front matter. Journal of Functional Programming, 21(3):f1–f2, May 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C40D63F55ADD45C5E4FB322953BA3BF6.

Anonymous:2011:JVIh

[Ano11g] Anonymous. JFP volume 21 issue 4–5 cover and back matter. Journal of Functional Programming, 21(4–5):b1–b2, September 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B804891859A0C279736C260EC3A14445.

Anonymous:2011:JVIg

[Ano11h] Anonymous. JFP volume 21 issue 4–5 cover and front matter. Journal of Functional Programming, 21(4–5):f1–f2, September 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
63A9A2265078435A03E9E1F5A95715E1.

Anonymous:2011:JVIj

[Ano11i] Anonymous. JFP volume 21 issue 6 cover and back matter. Journal of Functional Programming, 21(6):b1–b2, November 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2927AA6F33E2B0D228E9B22A4D1B00C1.

Anonymous:2011:JVIi

[Ano11j] Anonymous. JFP volume 21 issue 6 cover and front matter. Journal of Functional Programming, 21(6):f1–f2, November 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
10EF8CC032B42257FCF979C009C3099D.

Anonymous:2012:JVIb

[Ano12a] Anonymous. JFP volume 22 issue 1 cover and back matter. Journal of Functional Programming, 22(1):b1–b9, January 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4DD247946E2BE5509F438FDA5365288A.

Anonymous:2012:JVIa

[Ano12b] Anonymous. JFP volume 22 issue 1 cover and front matter. Journal of Functional Programming, 22(1):f1–f2, January 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1C31EF1F2A5E52B7B3888D7973A17410.

Anonymous:2012:JVId

[Ano12c] Anonymous. JFP volume 22 issue 2 cover and back matter. Journal of Functional Programming, 22(2):b1–b7, March 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D18E9BFDA6EBA04B86D03569855EDA14.

Anonymous:2012:JVIc

[Ano12d] Anonymous. JFP volume 22 issue 2 cover and front matter. Journal of Functional Programming, 22(2):f1–f2, March 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
81D98E25828637CED536B86C804B893D.

Anonymous:2012:JVIf

[Ano12e] Anonymous. JFP volume 22 issue 3 cover and back matter. Journal of Functional Programming, 22(3):b1–b7, May 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9F53A644E37DD9AD471942FB9A1AA27D.

Anonymous:2012:JVIe

[Ano12f] Anonymous. JFP volume 22 issue 3 cover and front matter. Journal of Functional Programming, 22(3):f1–f2, May 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
00D13D739FBC8D70A111BBF8F4CEF6AA.

Anonymous:2012:JVIh

[Ano12g] Anonymous. JFP volume 22 issue 4–5 cover and back matter. Journal of Functional Programming, 22(4–5):b1–b2, September 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2378D43ECE1E51508DEE8FE68293A70D.

Anonymous:2012:JVIg

[Ano12h] Anonymous. JFP volume 22 issue 4–5 cover and front matter. Journal of Functional Programming, 22(4–5):f1–f2, September 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EA51E3057B86331E4B702230F4208230.

Anonymous:2012:JVIj

[Ano12i] Anonymous. JFP volume 22 issue 6 cover and back matter. Journal of Functional Programming, 22(6):b1–b5, November 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E5C5212218B6BB69F8CC891F03A256C2.

Anonymous:2012:JVIi

[Ano12j] Anonymous. JFP volume 22 issue 6 cover and front matter. Journal of Functional Programming, 22(6):f1–f2, November 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3FC1AEFFCC0474290A1A2E145676A131.

Anonymous:2013:JVIb

[Ano13a] Anonymous. JFP volume 23 issue 1 cover and back matter. Journal of Functional Programming, 23(1):b1–b2, January 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
8253F6C2B88683656E5BF8E5A504E410.

Anonymous:2013:JVIa

[Ano13b] Anonymous. JFP volume 23 issue 1 cover and front matter. Journal of Functional Programming, 23(1):f1–f2, January 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B9C5D1FB0EA29C98E1F36D2CCA396A1E.

Anonymous:2013:JVId

[Ano13c] Anonymous. JFP volume 23 issue 2 cover and back matter. Journal of Functional Programming, 23(2):b1–b3, March 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F9357E3251014CB78E8D97B989C2C121.

Anonymous:2013:JVIc

[Ano13d] Anonymous. JFP volume 23 issue 2 cover and front matter. Journal of Functional Programming, 23(2):f1–f2, March 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4A0D83937782C78AC3C1CCAC7CA9E351.

Anonymous:2013:JVIf

[Ano13e] Anonymous. JFP volume 23 issue 3 cover and back matter. Journal of Functional Programming, 23(3):b1–b4, May 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
46A266FDDD77018FFB2E0C61B2E894DB.

Anonymous:2013:JVIe

[Ano13f] Anonymous. JFP volume 23 issue 3 cover and front matter. Journal of Functional Programming, 23(3):f1–f2, May 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
79AA2F988CE6B847BCD92026024672A1.

Anonymous:2013:JVIh

[Ano13g] Anonymous. JFP volume 23 issue 4 cover and back matter. Journal of Functional Programming, 23(4):b1–b2, July 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
135D0080323EA46FA3FC553E3D194825.

Anonymous:2013:JVIg

[Ano13h] Anonymous. JFP volume 23 issue 4 cover and front matter. Journal of Functional Programming, 23(4):f1–f2, July 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7839ACCC5F30D14C5D68EC94CBAF3060.

Anonymous:2013:JVIj

[Ano13i] Anonymous. JFP volume 23 issue 5 cover and back matter. Journal of Functional Programming, 23(5):b1–b8, September 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
8B3F8576CA644C2C9C9783FFF01636AE.

Anonymous:2013:JVIi

[Ano13j] Anonymous. JFP volume 23 issue 5 cover and front matter. Journal of Functional Programming, 23(5):f1–f2, September 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
011E556AFDF70089AA059707760ED391.

Anonymous:2013:JVIl

[Ano13k] Anonymous. JFP volume 23 issue 6 cover and back matter. Journal of Functional Programming, 23(6):b1–b6, November 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C7852E481C37F98BE115E83AD49D67A4.

Anonymous:2013:JVIk

[Ano13l] Anonymous. JFP volume 23 issue 6 cover and front matter. Journal of Functional Programming, 23(6):f1–f2, November 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
716F863590A784AEEC633045924CE84F.

Anonymous:2014:JVIb

[Ano14a] Anonymous. JFP volume 24 issue 1 cover and back matter. Journal of Functional Programming, 24(1):b1–b8, January 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
01549E117978A3871D9B337561477933.

Anonymous:2014:JVIa

[Ano14b] Anonymous. JFP volume 24 issue 1 cover and front matter. Journal of Functional Programming, 24(1):f1–f2, January 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
95855F19AB96866ADC116A3A27269408.

Anonymous:2014:JVId

[Ano14c] Anonymous. JFP volume 24 issue 2–3 cover and back matter. Journal of Functional Programming, 24(2–3):b1–b6, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
380A795B3A0ED6A35E860658B67C90BD.

Anonymous:2014:JVIc

[Ano14d] Anonymous. JFP volume 24 issue 2–3 cover and front matter. Journal of Functional Programming, 24(2–3):f1–f2, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
331E03C4690950AC35AE4685408BFA1C.

Anonymous:2014:JVIf

[Ano14e] Anonymous. JFP volume 24 issue 4 cover and back matter. Journal of Functional Programming, 24(4):b1–b9, July 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2FF1CDA9A3AEFC2D2D021B7CA67B8056.

Anonymous:2014:JVIe

[Ano14f] Anonymous. JFP volume 24 issue 4 cover and front matter. Journal of Functional Programming, 24(4):f1–f2, July 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
71F2DB8DCE7753CCFC90A91303598D9F.

Anonymous:2014:JVIh

[Ano14g] Anonymous. JFP volume 24 issue 5 cover and back matter. Journal of Functional Programming, 24(5):b1–b9, September 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2EA3F245ED7227A5767A432B4BFA6295.

Anonymous:2014:JVIg

[Ano14h] Anonymous. JFP volume 24 issue 5 cover and front matter. Journal of Functional Programming, 24(5):f1–f2, September 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B1DB25FCC662CE53D07A7AE070813215.

Anonymous:2014:JVIj

[Ano14i] Anonymous. JFP volume 24 issue 6 cover and back matter. Journal of Functional Programming, 24(6):b1–b8, November 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
31C05FF2178FDA4F93605D5BA8B468DB.

Anonymous:2014:JVIi

[Ano14j] Anonymous. JFP volume 24 issue 6 cover and front matter. Journal of Functional Programming, 24(6):f1–f2, November 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1B173E75FFF61F2A4196865A73CC7332.

Achten:1995:IOC

[AP95] Peter Achten and Rinus Plasmeijer. The ins and outs of Clean I/O. Journal of Functional Programming, 5(1):81–110, January 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2EFAEBBE3A19EA03A8D6D75A5348E194.

Asai:2013:SID

[AP13] Kenichi Asai and Benjamin C. Pierce. Special issue dedicated to ICFP 2011 editorial. Journal of Functional Programming, 23(4): 355–356, July 2013. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/17B5E9856C72935F39A121CB00747E1D.

Abel:2016:WFR

[AP16] Andreas Abel and Brigitte Pientka. Well-founded recursion with copatterns and sized types. Journal of Functional Programming, 26:e2, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
wellfounded-recursion-with-copatterns-and-sized-types/39794AEA4D0F5003C8E9F88E564DA8DD.

Appel:1993:CSM

[App93] Andrew W. Appel. A critique of Standard ML. Journal of Functional Programming, 3(4):391–429, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/8BFFACE1EFA4E2AEDB93B7EEA5F08D54.

Appel:1997:BRG

[App97a] Andrew W. Appel. Book review: Garbage Collection: Algorithms for Automatic Dynamic Memory Management, by Richard Jones and Rafael Lins, John Wiley & Sons, 1996. Journal of Functional Programming, 7(2):227–229, March 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/DE15A080BD3879A62CD078A7B5AA466C.

Appel:1997:MCI

[App97b] Andrew W. Appel. Modern compiler implementation in ML: basic techniques. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-58275-X (hardcover), 0-521-58775-1 (paperback). x + 390 pp. LCCN QA76.76.C65 A6753 1997. URL http://www.

loc.gov/catdir/description/cam028/96050393.html;   http:
//www.loc.gov/catdir/toc/cam022/96050393.html.

Armstrong:2007:PES

[Arm07] Joe Armstrong. Programming ERLANG: software for a concurrent world. Pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2007. ISBN 1-934356-00-X (paperback). xiii + 515 pp. LCCN QA76.73.E75 A765 2007. URL http://www.oreilly.com/

catalog/.

Augustsson:1994:FPG

[ARS94] Lennart Augustsson, Mikael Rittri, and Dan Synek. Functional pearl: On generating unique names. Journal of Functional Programming, 4(1):117–123, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/763DE73EB4761FDF681A613BE0E98443.

Arya:1994:FAS

[Ary94] Kavi Arya. A functional animation starter-kit. Journal of Functional Programming, 4(1):1–18, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/36B0FB7D4DB4B90BDE721E0A36BE5226.

Appel:1996:EAS

[AS96] Andrew W. Appel and Zhong Shao. Empirical and analytic study of stack versus heap cost for languages with closures. Journal of Functional Programming, 6(1):47–74, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
30303C7D7A9ACCC12AAA130855B7E6CF.

Andersen:2021:PCM

[AS21]  Kristoffer Just Arndal Andersen and Ilya Sergey.  Proto-

col combinators for modeling, testing, and execution of distributed systems. Journal of Functional Programming, 31(??): e3, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
protocol-combinators-for-modeling-testing-and-execution-
of-distributed-systems/BACA56805598C43F1B6AD36E40EF4B82.

Asperti:1992:CUE

[Asp92] Andrea Asperti. A categorical understanding of environment machines. Journal of Functional Programming, 2(1):23–59, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B9C0B6276C7F560A638902270459CA58.

Abelson:1996:SIC

[ASS96] Harold Abelson, Gerald Jay Sussman, and Julie Sussman. Structure and interpretation of computer programs. MIT electrical engineering and computer science. MIT Press, Cambridge, MA, USA, second edition, 1996. ISBN 0-262-01153-0 (MIT Press hardcover), 0-07-000484-6 (McGraw-Hill hardcover), 0-262-51087-1 (paperback). xxiii + 657 pp. LCCN QA76.6 .A255 1996. URL http://catdir.

loc.gov/catdir/description/mh022/96017756.html;   http:/
/catdir.loc.gov/catdir/toc/mh022/96017756.html. Foreword

by Alan J. Perlis.

Abelson:2000:SIC

[ASS00] Harold Abelson, Gerald Jay Sussman, and Julie Sussman. Structure and interpretation of computer programs. The MIT electrical engineering and computer science series. MIT Press, Cambridge, MA, USA, second edition, 2000. ISBN 0-07-000484-6, 0-262-01153-0, 0-262-51087-1. xxiii + 657 pp. LCCN ????

Atkey:2009:PNC

[Atk09] Robert Atkey. Parameterised notions of computation. Journal of Functional Programming, 19(3–4):335–376, July 2009. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/82CE5F0583C3390BBBD305830255FAA0.

Augustsson:1993:ILM

[Aug93] Lennart Augustsson. The interactive lazy ML system. Journal of Functional Programming, 3(1):77–92, January 1993. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/D76BABE75A9AD9902ECED03A5FFC0318.

Achten:2013:EAB

[AVDP13] Peter Achten, Marko Van Eekelen, Maarten De Mol, and Rinus Plasmeijer. EditorArrow: An arrow-based model for editor-based programming. Journal of Functional Programming, 23(2):185–224, March 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0415C28274BEF2D6C9124C1EEECC690D.

Ancona:2002:CMS

[AZ02] Davide Ancona and Elena Zucca. A calculus of module systems. Journal of Functional Programming, 12(2):91–132, March 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9FDD4C5B6EE93F4AC673D96BDE2DEBEA.

Burnett:2001:FFO

[BAD^{+}01] Margaret Burnett, John Atwood, Rebecca Walpole Djang, James Reichwein, Herkimer Gottfried, and Sherry Yang. Forms/3: A first-order visual language to explore the boundaries of the spreadsheet paradigm. Journal of Functional Programming, 11(2):155–206, March 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
66F1ACF22742EA28DD52608A5122E70F.

Bahr:2022:MFA

[Bah22] Patrick Bahr. Modal FRP for all: Functional reactive programming without space leaks in Haskell. Journal of Functional Programming, 32:e15, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
modal-frp-for-all-functional-reactive-programming-without-
space-leaks-in-haskell/9BE20E8D61E9B74811CF3CF97B5D10C7.

Bailes:1999:BRF

[Bai99] Paul A. Bailes. Book review: Functional Programming Languages in Education, by P. H. Hartel and R. Plasmeijer, editors, Springer-Verlag, 1995. Journal of Functional Programming, 9(3):347–354, May 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
677F2FF7A45978FCEE6512FC5116B2BE.

Barendregt:1991:IGT

[Bar91a] Henk Barendregt. Introduction to generalized type systems. Journal of Functional Programming, 1(2):125–154, April 1991. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/869991BA6A99180BF96A616894C6D710.

Barendregt:1991:TPS

[Bar91b] Henk Barendregt. Theoretical pearls: Self-interpretation in lambda calculus. Journal of Functional Programming, 1(2):229–233, April 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D16FE1FAB38CD06AD831C845156E10D5.

Barendsen:1991:TPU

[Bar91c] Erik Barendsen. Theoretical pearls: An unsolvable numeral system in lambda calculus. Journal of Functional Programming, 1(3): 367–372, July 1991. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/00483DAE6FF74553E2C896CDD001ABD3.

Barendregt:1992:TPE

[Bar92a] Henk Barendregt. Theoretical pearls: Enumerators of lambda terms are reducing. Journal of Functional Programming, 2(2):233–236, April 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F3145D4DE8AD9734951632E985EAF908.

Barendregt:1992:TPR

[Bar92b] Henk Barendregt. Theoretical pearls: Representing ‘undefined’ in lambda calculus. Journal of Functional Programming, 2(3):367–374, July 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
593885802DF6EFD4D7EB8F2E3D8FBB18.

Barthe:1999:TCI

[Bar99] Gilles Barthe. Type-checking injective pure type systems. Journal of Functional Programming, 9(6):675–698, November 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7F96FF1C8E04E3C4932AA574FFBD219A.

Barron:2000:WSL

[Bar00] D. W. (David William) Barron. The world of scripting languages. Worldwide series in computer science. John Wiley, New York, NY, USA, 2000. ISBN 0-471-99886-9. xii + 492 pp. LCCN QA76.6

.B3715 2000. URL http://www.loc.gov/catdir/description/
wiley034/99089450.html;   http://www.loc.gov/catdir/toc/
onix06/99089450.html.

Barbanera:1996:PIE

[BB96] Franco Barbanera and Stefano Berardi. Proof-irrelevance out of excluded-middle and choice in the calculus of constructions. Journal of Functional Programming, 6(3):519–526, May 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
698316A70788A331B3E10B2D3235255C.

Brede:2021:CMB

[BB21] Nuria Brede and Nicola Botta. On the correctness of monadic backward induction. Journal of Functional Programming, 31 (??):e26, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
on-the-correctness-of-monadic-backward-induction/0F71FFE6AE1222E77F52341C695E366A.

Blelloch:2001:PT

[BBC^{+}01] Guy Blelloch, Hal Burch, Karl Crary, Robert Harper, Gary Miller, and Noel Walkington. Persistent triangulations. Journal of Functional Programming, 11(5):441–466, September 2001. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/C7F680713ADE671FB51850F1850B9B2D.

Benton:1998:CTL

[BBD98] P. N. Benton, G. M. Bierman, and V. C. V. De Paiva. Computational types from a logical perspective. Journal of Functional Programming, 8(2):177–193, March 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/37B1EAE149C3EE88BE5A90EF9B56FD4F.

Botta:2021:EEP

[BBJR21] Nicola Botta, Nuria Brede, Patrik Jansson, and Tim Richter. Extensional equality preservation and verified generic programming. Journal of Functional Programming, 31(??):e24, ???? 2021. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/extensional-equality-preservation-and-
verified-generic-programming/BF60B28E6CAC815B83D3D94D8103133F.

Benaissa:1996:CES

[BBLRD96] Zine-El-Abidine Benaissa, Daniel Briaud, Pierre Lescanne, and Jocelyne Rouyer-Degli. λν, a calculus of explicit substitutions which preserves strong normalisation. Journal of Functional Programming, 6(5):699–722, September 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/3231CD0A276A27ED82C778CB262B75F7.

Burton:1993:PMA

[BC93] F. Warren Burton and Robert D. Cameron. Pattern matching with abstract data types. Journal of Functional Programming, 3(2): 171–190, April 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/04DD26A0E6CA3A1E87E0E6AE8BC02EED.

Barthe:2006:RET

[BC06a] Gilles Barthe and Thierry Coquand. Remarks on the equational theory of non-normalizing pure type systems. Journal of Functional Programming, 16(2):137–155, March 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7802BAE4C650D853790158D63E0CC4FE.

Bird:2006:FPF

[BC06b] Richard Bird and Sharon Curtis. Functional pearls: Finding celebrities: A lesson in functional programming. Journal of Functional Programming, 16(1):13–20, January 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3E79673B3E7EA223532194DFBFF69569.

Bernardy:2015:EPI

[BC15] Jean-Philippe Bernardy and Koen Claessen. Efficient parallel and incremental parsing of practical context-free languages. Journal of Functional Programming, 25:e10, ???? 2015. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/efficient-parallel-and-incremental-
parsing-of-practical-contextfree-languages/4D620F0BFADE2B588F854AAAEA252F5C.

Bonelli:2005:CAP

[BCG05] Eduardo Bonelli, Adriana Compagnoni, and Elsa Gunter. Correspondence assertions for process synchronization in concurrent communications. Journal of Functional Programming, 15(2):219–247, March 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BCA1F400E23BC6E584EC6844182B639F.

Bonfante:2001:API

[BCMT01] G. Bonfante, A. Cichon, J.-Y. Marion, and H. Touzet. Algorithms with polynomial interpretation termination proof. Journal of Functional Programming, 11(1):33–53, January 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/B4F33C45A883BF3C875699EAC091F86E.

Barthe:2003:STT

[BCP03] Gilles Barthe, Venanzio Capretta, and Olivier Pons. Setoids in type theory. Journal of Functional Programming, 13(2):261–293, March 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6A223F72737E421BD9D642C14EB5600B.

Broy:1992:MOS

[BD92] Manfred Broy and Claus Dendorfer. Modelling operating system structures by timed stream processing functions. Journal of Functional Programming, 2(1):1–21, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3A13783378428EA08B1AE8A4BCC8021E.

Broda:1997:CCS

[BD97] Sabine Broda and Luís Damas. On combinatory complete sets of proper combinators. Journal of Functional Programming, 7(6):593– 612, November 1997. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0303D13A5B4E8160B9FEC85D514B18F9.

Biernacki:2006:TPS

[BD06] Dariusz Biernacki and Olivier Danvy. Theoretical pearl: A simple proof of a folklore theorem about delimited control. Journal of Functional Programming, 16(3):269–280, May 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7F5EF8CF2330282D403FC22CA47CCF17.

Bird:1996:GFP

[BDH96] Richard Bird, Oege De Moor, and Paul Hoogendijk. Generic functional programming with types and relations. Journal of Functional Programming, 6(1):1–28, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/67C5B097533659132953E973698B81A6.

Besson:2005:ISI

[BDJ05] Frédéric Besson, Thomas De Grenier De Latour, and Thomas Jensen. Interfaces for stack inspection. Journal of Functional Programming, 15(2):179–217, March 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/E528AA7AE8E1A030A95886E5C7D835AE.

Barthe:2001:SID

[BDT01] Gilles Barthe, Peter Dybjer, and Peter Thiemann. Special issue on dependent type theory meets programming practice: Call for papers. Journal of Functional Programming, 11(4):437, July 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
67A3CF74B38FD242DD8ADBCF8EDD1901.

Barthe:2004:ISI

[BDT04] Gilles Barthe, Peter Dybjen, and Peter Thiemann. Introduction to the special issue on dependent type theory meets practical programming. Journal of Functional Programming, 14(1):1–2, January 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B4C8A7E3FD276535C317B175EA3C9E8B.

Breitner:2016:SZC

[BEJW16] Joachim Breitner, Richard A. Eisenberg, Simon Peyton Jones, and Stephanie Weirich. Safe zero-cost coercions for Haskell. Journal of Functional Programming, 26:e15, ???? 2016. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/safe-zerocost-coercions-for-haskell/
FFB002A34C6B40EF150B5F357AA865DC.

Benton:1995:SNL

[Ben95] P. N. Benton. Strong normalisation for the linear term calculus. Journal of Functional Programming, 5(1):65–80, January 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
DDBBD2A74BC4F0D5F9DDB1326DCD70E9.

Benzinger:2001:ACA

[Ben01] Ralph Benzinger. Automated complexity analysis of Nuprl extracted programs. Journal of Functional Programming, 11(1):3–31, January 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
14847C50A5F64488246A9D19B62BFB50.

Benton:2005:EI

[Ben05] Nick Benton. Embedded interpreters. Journal of Functional Programming, 15(4):503–542, July 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/34E3F8995878604F6E1CC9B332F65424.

Bendkowski:2017:NOR

[Ben17] Maciej Bendkowski. Normal-order reduction grammars. Journal of Functional Programming, 27:e6, ???? 2017. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/normalorder-reduction-grammars/B043EABD00457D588EA10B449CDF91AB.

Berry:1993:LDS

[Ber93] Dave Berry. Lessons from the design of a Standard ML library. Journal of Functional Programming, 3(4):527–552, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0CECFD1D61A6547F4AE5716BCCAAEDE5.

Betarte:2000:TCD

[Bet00] Gustavo Betarte. Type checking dependent (record) types and subtyping. Journal of Functional Programming, 10(2):137–166, March 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1793E1F504A8B156B7A3EF9F17A42549.

Barbanera:1997:MSN

[BFG97] Franco Barbanera, Maribel Fernández, and Herman Geuvers. Modularity of strong normalization in the algebraic-λ-cube. Journal of Functional Programming, 7(6):613–660, November 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
A4D2DA3CE83133C249BFCF3C6C444404.

Baker-Finch:2004:CPR

[BFGJ04] Clem Baker-Finch, Kevin Glynn, and Simon Peyton Jones. Constructed product result analysis for Haskell. Journal of Functional Programming, 14(2):211–245, March 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/53F8937E9686B7CB261AE727124FE1D0.

Bergstrom:2012:LTS

[BFR^{+}12] Lars Bergstrom, Matthew Fluet, Mike Rainey, John Reppy, and Adam Shaw. Lazy tree splitting. Journal of Functional Programming, 22(4–5):382–438, September 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/EF3F7E296F505383FA01B64E7E299F72.

Barendregt:2000:LTN

[BG00] Henk Barendregt and Silvia Ghilezan. Lambda terms for natural deduction, sequent calculus and cut elimination. Journal of Functional Programming, 10(1):121–134, January 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/C94BB5ECFC2E2B3F3EF4F8920A2840BC.

Bierman:2012:SSS

[BGHL12] Gavin M. Bierman, Andrew D. Gordon, Catalin Hritcu, and David Langworthy. Semantic subtyping with an SMT solver. Journal of Functional Programming, 22(1):31–105, January 2012. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/5093B3D8253A1A47E50357AFC7A321EF.

Borgstrom:2011:RSH

[BGP11] Johannes Borgström, Andrew D. Gordon, and Riccardo Pucella. Roles, stacks, histories: A triple for Hoare. Journal of Functional Programming, 21(2):159–207, March 2011. CODEN JFPRES.

ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/23F4063DD70B051277741864576BF603.

BanadosSchwerter:2016:GTE

[BGT16] Felipe Bañados Schwerter, Ronald Garcia, and Éric Tanter. Gradual type-and-effect systems. Journal of Functional Programming, 26:e19, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
gradual-typeandeffect-systems/5FB15D90345693369960654137B3C8D0.

Boyle:1992:PFP

[BH92] James M. Boyle and Terence J. Harmer. A practical functional program for the CRAY X-MP. Journal of Functional Programming, 2(1):81–126, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/64C00F14FB1E6B1A549A8F499BA0C576.

Bahr:2015:CCC

[BH15] Patrick Bahr and Graham Hutton. Calculating correct compilers. Journal of Functional Programming, 25:e14, ???? 2015. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/calculating-correct-compilers/70AA17724EBCA4182B1B2B522362A9AF.

See also Part II [BH20].

Bahr:2020:CCC

[BH20] Patrick Bahr and Graham Hutton. Calculating correct compilers II: Return of the register machines. Journal of Functional Programming, 30(??):e25, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
calculating-correct-compilers-ii-return-of-the-register-
machines/194A3C46A13A0BFDDE85976E96232AEC. See also Part I

[BH15].

Bhamidipati:2009:BRP

[Bha09] Saketh Bhamidipati. Book review: Programming in Haskell, by Hutton Graham, Cambridge University Press, 2007, 184 pp., ISBN 0-521-69269-5. Journal of Functional Programming, 19(2):256–259, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
A2E94CADF97E06E6EE1591311D4EFF6D.

Barendregt:1993:E

[BHH^{+}93] Henk Barendregt, Paul Hudak, John Hughes, Simon Peyton Jones, and Philip Wadler. Editorial. Journal of Functional Programming, 3(1):1–2, January 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/EBA8D8A91A51C0433550B1B29240C1E5.

Bucalo:2006:CTC

[BHM^{+}06] Anna Bucalo, Furio Honsell, Marino Miculan, Ivan Scagnetto, and Martin Hoffman. Consistency of the theory of contexts. Journal of Functional Programming, 16(3):327–372, May 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
CB41667D0A5A007107777F6FD1E9ADCA.

Berger:2007:LAA

[BHY07] Martin Berger, Kohei Honda, and Nobuko Yoshida. A logical analysis of aliasing in imperative higher-order functions. Journal of Functional Programming, 17(4–5):473–546, July 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/6A98863EFA4BE6D0AA14242CEC1715D5.

Bierman:2000:PEL

[Bie00] G. M. Bierman. Program equivalence in a linear functional language. Journal of Functional Programming, 10(2):167–190, March 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0A99518FCBBB490A21FF29018A32C97E.

Bird:1991:FPR

[Bir91a] Richard S. Bird. Functional pearls: On removing duplicates. Journal of Functional Programming, 1(2):235–243, April 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2CC831E0F77C77E86DD1E9AF165DECEE.

Bird:1991:FPM

[Bir91b] Richard S. Bird. Functional pearls: The minout problem. Journal of Functional Programming, 1(1):121–124, January 1991. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/64738655BA2DD3BB86B8416397D40837.

Bird:1992:FPT

[Bir92a] R. S. Bird. Functional pearls: Two greedy algorithms. Journal of Functional Programming, 2(2):237–244, April 1992. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/E55EFDCCB071BBE9FFE0F6543C1BCBA1.

Bird:1992:FPU

[Bir92b] Richard S. Bird. Functional pearls: Unravelling greedy algorithms. Journal of Functional Programming, 2(3):375–385, July 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F477E874FBCB0C2F828A7F2C152DBB5A.

Bird:1993:FPL

[Bir93] R. S. Bird. Functional pearls: The last tail. Journal of Functional Programming, 3(1):117–122, January 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/CC6079CCEC88883F6C10DFB12908795F.

Bird:1997:FPB

[Bir97a] Richard S. Bird. Functional pearl: On building trees with minimum height. Journal of Functional Programming, 7(4):441–445, July 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
23D5C1B90A6B00A5D87239054FEDC8CF.

Bird:1997:FPM

[Bir97b] Richard S. Bird. Functional pearl: On merging and selection. Journal of Functional Programming, 7(3):349–354, May 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
DB0CF71E6B3D975D1F171C02336F3FA3.

Bird:1998:MN

[Bir98] Richard S. Bird. Meertens number. Journal of Functional Programming, 8(1):83–88, January 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/991084CDC1A348354224F893E2BD1D0C.

Bird:1999:E

[Bir99] Richard Bird. Editorial. Journal of Functional Programming, 9 (3):311, May 1999. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/1080708F356AC4CFDAA0F382103D82D5.

Bird:2001:FPU

[Bir01a] Richard S. Bird. Functional pearl: Unfolding pointer algorithms. Journal of Functional Programming, 11(3):347–358, May 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6C58105CDEFA4B5B46D59F275575996B.

Bird:2001:MMP

[Bir01b] Richard S. Bird. Maximum marking problems. Journal of Functional Programming, 11(4):411–424, July 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/63E57511542AE88B1A6EE66C959F3AAE.

Bird:2004:FPT

[Bir04] Richard S. Bird. Functional pearl: On tiling a chessboard. Journal of Functional Programming, 14(6):613–622, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
44A9D4F4D9AF0D8380D5E573E7E7F885.

Bird:2006:FPP

[Bir06] Richard Bird. Functional pearl: A program to solve Sudoku. Journal of Functional Programming, 16(6):671–679, November 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2A2C12F0CAA9A48C1D6697CBFC6A559C.

Bird:2010:PFA

[Bir10] Richard Bird. Pearls of Functional Algorithm Design. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-51338-3 (hardcover). xi + 277 pp. LCCN QA76.62 .B57 2010.

Bird:2011:BCR

[Bir11] Richard S. Bird. Building a consensus: A rectangle covering problem. Journal of Functional Programming, 21(2):119–128, March 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
561C117A97FE0388AF96155469348B6B.

Bird:2013:SAE

[Bir13] Richard S. Bird. An in-situ algorithm for expanding a graph. Journal of Functional Programming, 23(2):174–184, March 2013. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/EF06351157803712CCD8F24F75511FC7.

Bird:2015:HMS

[Bir15] Richard S. Bird. How to mingle streams. Journal of Functional Programming, 25:e4, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
how-to-mingle-streams/F4F6E6D0EC3671D5F8D6CAAF9408E9E2.

Bird:2019:HFF

[Bir19] Richard S. Bird. How to find a fake coin. Journal of Functional Programming, 29:e8, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
how-to-find-a-fake-coin/BD43BF3F22072B224171E9A0B961DF22.

Bird:2020:OPF

[Bir20] Richard S. Bird. An optimal, purely functional implementation of the Garsia–Wachs algorithm. Journal of Functional Programming, 30(??):e3, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
an-optimal-purely-functional-implementation-of-the-garsiawachs-
algorithm/6CC5120546FAE30DF5F7F16A9784E136.

Bondorf:1993:EAR

[BJ93] Anders Bondorf and Jesper Jørgensen. Efficient analyses for realistic off-line partial evaluation. Journal of Functional Programming, 3 (3):315–346, July 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/85CDDF0A659BC825CFF35D29AD4B99C3.

Bird:1997:MHL

[BJD97] Richard Bird, Geraint Jones, and Oege De Moor. More haste, less speed: lazy versus eager evaluation. Journal of Functional Programming, 7(5):541–547, September 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/162B391CBCD864794C766CA2A2EC7CBE.

Botta:2017:CCT

[BJI17] Nicola Botta, Patrik Jansson, and Cezar Ionescu. Contributions to a computational theory of policy advice and avoidability. Journal of Functional Programming, 27:e23, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic).  URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/contributions-
to-a-computational-theory-of-policy-advice-and-avoidability/
CDB4C9601702AAB336A2FB2C34B8F49B.

Bernardy:2012:PFP

[BJP12] Jean-Philippe Bernardy, Patrik Jansson, and Ross Paterson. Proofs for free: Parametricity for dependent types. Journal of Functional Programming, 22(2):107–152, March 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/FB3D0FC91F42AF027A2EB09969F0E759.

Bernardy:2010:GPC

[BJZS10] Jean-Philippe Bernardy, Patrik Jansson, Marcin Zalewski, and Sibylle Schupp. Generic programming with C++ concepts and Haskell type classes — a comparison. Journal of Functional Programming, 20(3–4):271–302, July 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/2BF57A31E8173929FCBEF6E3E9AB76BE.

Benton:2001:ES

[BK01] Nick Benton and Andrew Kennedy. Exceptional syntax. Journal of Functional Programming, 11(4):395–410, July 2001. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/58206FB399EDC9F197A0D53BC46E4667.

Bennett:2001:PFT

[BKP01] Andrew J. Bennett, Paul H. J. Kelly, and Ross A. Paterson. Pipelined functional tree accesses and updates: scheduling, synchronization, caching and coherence. Journal of Functional Programming, 11(4):359–393, July 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/21248D1EB7DD91D3D0810D2BB3690A3A.

Burn:1996:PCC

[BL96] Geoffrey Burn and Daniel Le Métayer. Proving the correctness of compiler optimisations based on a global analysis: a study of strictness analysis. Journal of Functional Programming, 6(1):75– 109, January 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/9F5B8B249126E3088B0E0998AE92AEB9.

Blanqui:2018:SBT

[Bla18] Frédéric Blanqui. Size-based termination of higher-order rewriting. Journal of Functional Programming, 28:e11, ???? 2018. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/sizebased-termination-of-
higherorder-rewriting/2134D9160988448FA62DD693D337892D.

Berthold:2016:PPS

[BLH16] Jost Berthold, Hans-Wolfgang Loidl, and Kevin Hammond. PAEAN: Portable and scalable runtime support for parallel Haskell dialects. Journal of Functional Programming, 26:e10, ???? 2016. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/paean-portable-and-scalable-runtime-
support-for-parallel-haskell-dialects/FFAF6CAE496B45F215755B6CB25584A1.

Baars:2004:FPP

[BLS04] Arthur I. Baars, Andres Löh, and S. Doaitse Swierstra. Functional pearl: Parsing permutation phrases. Journal of Functional Programming, 14(6):635–646, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/DB7B6AFE506CF84BBDBBF54306F28D38.

Bird:2004:FPI

[BM04] Richard S. Bird and Shin-Cheng Mu. Functional pearl: Inverting the Burrows–Wheeler transform. Journal of Functional Programming, 14(6):603–612, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/EB01B27756846A88CD5C2787138B143D.

Bird:2005:CCS

[BM05] Richard Bird and Shin-Cheng Mu. Countdown: A case study in origami programming. Journal of Functional Programming, 15(5): 679–702, September 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/F7469590A11FDC55ED1C5401B0BD642D.

Blume:2006:SCM

[BM06] Matthias Blume and David McAllester. Sound and complete models of contracts. Journal of Functional Programming, 16(4–5):375–414, July 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B41BD2BF7E986641E5FACDE9C6F45340.

Bird:2021:GAD

[BM21] Richard Bird and Shin-Cheng Mu. A greedy algorithm for dropping digits. Journal of Functional Programming, 31(??): e29, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
greedy-algorithm-for-dropping-digits/A11B56DDB98C8B789AD0F80E96CCB3AD.

Baader:1998:TRA

[BN98] Franz Baader and Tobias Nipkow. Term Rewriting and All That. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-45520-0 (hardcover). xii + 301 pp. LCCN QA267 .B314

1998. URL http://www.loc.gov/catdir/description/cam028/
97028286.html;   http://www.loc.gov/catdir/toc/cam027/97028286.
html.

Banerjee:2005:SBA

[BN05] Anindya Banerjee and David A. Naumann. Stack-based access control and secure information flow. Journal of Functional Programming, 15(2):131–177, March 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/27C5A3345D3CE01E733207F1BF9BF363.

Brodal:1996:OPF

[BO96] Gerth Stølting Brodal and Chris Okasaki. Optimal purely functional priority queues. Journal of Functional Programming, 6(6):839– 857, November 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/1A799209383FD6535F90715192863C13.

Boiten:2003:BRC

[Boi03] Eerke Boiten. Book review: Concepts in Programming Languages, by John C. Mitchell, Cambridge University Press, 2002, ISBN 0-521-78098-5. Journal of Functional Programming, 13(6):1087– 1088, November 2003. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/1865415A15871C20697B2B6A6554C335.

Boiten:2012:BRM

[Boi12] Eerke Boiten. Book review: Modeling in Event-B — System and Software Engineering, Abrial Jean-Raymond, Cambridge University Press, May 2010 ISBN-10: 0-521-89556-1. Journal of Functional Programming, 22(2):217–219, March 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/35DE12E3B2E217067A8A3C316666B102.

Bourdoncle:1992:AID

[Bou92] François Bourdoncle. Abstract interpretation by dynamic partitioning. Journal of Functional Programming, 2(4):407–435, October 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4A8CB4DB6C477E8844D89B28CF0CE5C1.

Boulton:1999:TOR

[Bou99] Richard J. Boulton. Transparent optimisation of rewriting combinators. Journal of Functional Programming, 9(2):113–146, March 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B1CBB51C9905A56327BB68336A17303C.

Boudol:2004:RRS

[Bou04] Gérard Boudol. The recursive record semantics of objects revisited. Journal of Functional Programming, 14(3):263–315, May 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D198B38143E6DB7AA13954BF390FF553.

Burton:1992:DRN

[BP92] F. Warren Burton and Rex L. Page. Distributed random number generation. Journal of Functional Programming, 2(2):203–212, April 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6D10F1D0A2FB7E66D5F746F6D0822D78.

Bondorf:1996:GAC

[BP96] Anders Bondorf and Jens Palsberg. Generating action compilers by partial evaluation. Journal of Functional Programming, 6(2): 269–298, March 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/DEB7DA3B2CBD20990EF1BF18241CBCA6.

Bird:1999:BNN

[BP99] Richard S. Bird and Ross Paterson. de Bruijn notation as a nested datatype. Journal of Functional Programming, 9(1):77–91, January 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D8BFA383FDA7EA3DC443B4C42A168F30.

Bulterman:2004:SIM

[BR04] Dick C. A. Bulterman and Lloyd Rutledge. SMIL 2.0: interactive multimedia for web and mobile devices. X. media.publishing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20234-X (hardcover). ISSN 1612-1449. xxx + 439 pp. LCCN QA76.76.H94 B858 2004. URL http:/

/aleph.unisg.ch/hsgscan/hm00130755.pdf;   http://catdir.
loc.gov/catdir/enhancements/fy0826/2004041320-b.html;
http://catdir.loc.gov/catdir/enhancements/fy0826/2004041320-
d.html;   http://catdir.loc.gov/catdir/enhancements/fy0826/
2004041320-t.html;   http://opac.nebis.ch/cgi-bin/showAbstract.
pl?u20=354020234X;   http://swbplus.bsz-bw.de/bsz107852276cov.
htm;   http://swbplus.bsz-bw.de/bsz107852276inh.htm;   http:
//swbplus.bsz-bw.de/bsz107852276kap.htm;   http://swbplus.
bsz-bw.de/bsz107852276vlg.htm;   http://www.gbv.de/du/services/
toc/bs/37256514x;   http://www.XmediaSMIL.net.

Bringert:2008:PAC

[BR08] Björn Bringert and Aarne Ranta. A pattern for almost compositional functions. Journal of Functional Programming, 18(5):567– 598, September 2008. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/682674057A68EFA9D46A8201B18C80E6.

Brady:2013:IGP

[Bra13] Edwin Brady. Idris, a general-purpose dependently typed programming language: Design and implementation. Journal of Functional Programming, 23(5):552–593, September 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/418409138B4452969AC0736DB0A2C238.

Breitner:2018:ALN

[Bre18] Joachim Breitner. The adequacy of Launchbury’s natural semantics for lazy evaluation. Journal of Functional Programming, 28:e1, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
adequacy-of-launchburys-natural-semantics-for-lazy-evaluation/
74FA9327F749042DC9C824261547F94B.

Brandt:2017:CSF

[BRL17] Jörgen Brandt, Wolfgang Reisig, and Ulf Leser. Computation semantics of the functional scientific workflow language Cuneiform. Journal of Functional Programming, 27:e22, ???? 2017. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/computation-semantics-of-
the-functional-scientific-workflow-language-cuneiform/1A3B8AB825939117C5BD9F850F63ADCC.

Brown:2012:BRF

[Bro12] Neil Brown. Book review: Functional Programming for Java Developers — Tools for Better Concurrency, Abstraction, and Agility, By Dean Wampler, O’Reilly Media, July 2011, ISBN-13: 978-1-4493-1103-2, 90 pp. Journal of Functional Programming, 22(6):853– 854, November 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/02755329E5B068C2D403EE6722FCBD06.

Burton:1994:WCS

[BRS94] F. Warren Burton and V. J. Rayward-Smith. Worst case scheduling for parallel functional programs. Journal of Functional Programming, 4(1):65–75, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/9A0ED3A53A5EBEED1D4900F718E892C4.

Brumfitt:1992:MFM

[Bru92] P. J. Brumfitt. Metamorph — a formal methods toolkit with application to the design of digital hardware. Journal of Functional Programming, 2(4):437–473, October 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/144E154639185A74FC4F98EBCE67E970.

Bruce:1994:POO

[Bru94] Kim B. Bruce. A paradigmatic object-oriented programming language: Design, static typing and semantics. Journal of Functional Programming, 4(2):127–206, April 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/DBA72A6F6E3E0235CA2321C822BC000E.

Bezem:1996:SPU

[BS96] Marc Bezem and Jan Springintveld. A simple proof of the undecidability of inhabitation in λp. Journal of Functional Programming, 6(5):757–761, September 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C5867A407284956845F36FC53BE7078F.

Barthe:2000:DFP

[BS00] Gilles Barthe and Morten Heine Sørensen. Domain-free pure type systems. Journal of Functional Programming, 10(5):417–452, September 2000. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/5B8D3E23AD63A508EDF6AF706123E56E.

Bhaskar:2023:RWF

[BS23] Siddharth Bhaskar and Jakob Grue Simonsen. Read/write factorizable programs. Journal of Functional Programming, 33: ??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
readwrite-factorizable-programs/E00B881B91D79D7806634C2B306DEF06.

Brachthauser:2020:ECP

[BSO20] Jonathan Immanuel Brachthäuser, Philipp Schuster, and Klaus Ostermann. Effekt: Capability-passing style for type- and effectsafe, extensible effect handlers in Scala. Journal of Functional Programming, 30(??):e8, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/effekt-capabilitypassing-style-for-
type-and-effectsafe-extensible-effect-handlers-in-scala/
A19680B18FB74AD95F8D83BC4B097D4F.

Breitner:2021:RSV

[BSZL^{+}21] Joachim Breitner, Antal Spector-Zabusky, Yao Li, Christine Rizkallah, John Wiegley, Joshua Cohen, and Stephanie Weirich. Ready, set , verify! Applying hs-to-coq to real-world Haskell code. Journal of Functional Programming, 31(??):e5, ???? 2021. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/ready-set-verify-applying-
hstocoq-to-realworld-haskell-code/88AC7FD74F828740A1CD961559147F70.

Buday:2008:BRL

[Bud08] Gergely Buday. Book review: Logic in Computer Science: Modelling and Reasoning About Systems, by Huth Michael and Ryan Mark, second edition. ISBN 0-521-54310-X. Journal of Functional Programming, 18(3):421–422, May 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/B71B4E6ABE3ACA010DDB546DBA3F808E.

Burn:1991:IET

[Bur91a] G. L. Burn. Implementing the evaluation transformer model of reduction on parallel machines. Journal of Functional Programming, 1 (3):329–366, July 1991. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/56557FB8F959D8464B5327D8043D7512.

Burton:1991:END

[Bur91b] F. Warren Burton. Encapsulating non-determinacy in an abstract data type with determinate semantics. Journal of Functional Programming, 1(1):3–20, January 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/16828BC4A60520B4ED8AA8924AE40540.

Burton:1999:BRA

[Bur99] F. Warren Burton. Book review: Advanced Functional Programming: Lecture Notes in Computer Science 925, edited by Johan Jeuring and Erik Meijer, Springer-Verlag, 1995. Journal of Functional Programming, 9(1):105–111, January 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/F1EEA48458E15F0D5929E41543BB2848.

Cockx:2020:EDC

[CA20] Jesper Cockx and Andreas Abel. Elaborating dependent (co)pattern matching: No pattern left behind. Journal of Functional Programming, 30(??):e2, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
elaborating-dependent-copattern-matching-no-pattern-left-
behind/F13CECDAB2B6200135D45452CA44A8B3.

Carette:2011:BRH

[Car11] Jacques Carette. Book review: Handbook of Practical Logic and Automated Reasoning, by Harrison John, Cambridge University Press, 2009 ISBN 978-0-521-89957-4. Journal of Functional Programming, 21(6):663–664, November 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/EB9225C2C5AE94808B5DCF3AAC0CD200.

Castro:2013:BRB

[Cas13] Laura Castro. Book review: Building Web Applications with Erlang, by Zachary Kessin, O’Reilly Media, Inc., 2012, ISBN 97-81-4493-20652. Journal of Functional Programming, 23(2):225–226, March 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
CF6F62636DA83333AB52CF127D4FA576.

Choi:2019:TRC

[CC19] Kwanghoon Choi and Byeong-Mo Chang. A theory of RPC calculi for client-server model. Journal of Functional Programming, 29:e5, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/theory-of-rpc-
calculi-for-clientserver-model/15DC9096F78E604ABD5F34A96F277EFE.

Calcagno:2005:DVS

[CCG05] Cristiano Calcagno, Luca Cardelli, and Andrew D. Gordon. Deciding validity in a spatial logic for trees. Journal of Functional Programming, 15(4):543–572, July 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/A651D556570DF9C23054A6CB74C69CD2.

Clack:1995:LPT

[CCP95] Chris Clack, Stuart Clayman, and David Parrott. Lexical profiling: theory and practice. Journal of Functional Programming, 5(2): 225–277, April 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0E578D77A794F3E48A8D848AF7E3F60F.

Caporaso:2001:PAC

[CCP01] Salvatore Caporaso, Emanuele Covino, and Giovanni Pani. A predicative approach to the classification problem. Journal of Functional Programming, 11(1):95–116, January 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/BBD626FEFD6A1DDC3A607FFF08676C95.

Curien:1996:CRC

[CD96] Pierre-Louis Curien and Roberto Di Cosmo. A confluent reduction for the λ-calculus with surjective pairing and terminal object. Journal of Functional Programming, 6(2):299–327, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
325956525D5F9286604E8BA3C197AFA0.

Cockx:2018:PRU

[CD18] Jesper Cockx and Dominique Devriese. Proof-relevant unification: Dependent pattern matching with only the axioms of your type theory. Journal of Functional Programming, 28:e12, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/proofrelevant-
unification-dependent-pattern-matching-with-only-the-axioms-
of-your-type-theory/E54D56DC3F5D5361CCDECA824030C38E.

Chen:2014:ISA

[CDHA14] Yan Chen, Joshua Dunfield, Matthew A. Hammer, and Umut A. Acar. Implicit self-adjusting computation for purely functional programs. Journal of Functional Programming, 24(1):56–112, January 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
13F0A8D40493757FFB93C9AB676659EB.

Claessen:2015:GCR

[CDP15] Koen Claessen, Jonas Duregård, and Michal H. Palka. Generating constrained random data with uniform distribution. Journal of Functional Programming, 25:e8, ???? 2015. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/generating-constrained-random-data-
with-uniform-distribution/567438B9A7FABDD0F191FF65DAEA7005.

Cockx:2016:EDP

[CDP16] Jesper Cockx, Dominique Devriese, and Frank Piessens. Eliminating dependent pattern matching without K. Journal of Functional Programming, 26:e16, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
eliminating-dependent-pattern-matching-without-k/4BC4EA2D02D801E5ABED264FE5FB177A.

Chen:2018:SIC

[CE18] Sheng Chen and Martin Erwig. Systematic identification and communication of type errors. Journal of Functional Programming, 28:e2, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
systematic-identification-and-communication-of-type-errors/
40EF6949DF63E97ABA353999C72AB6C5.

Cebrian:2011:BRH

[Ceb11] Toni Cebrián. Book review: How to think about algorithms, by Edmonds Jeff, Cambridge University Press, ISBN 0-521-61410-4. Journal of Functional Programming, 21(6):664–666, November 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6FEDE8C30866A960255C49CAF21647A3.

Cebrian:2012:BRS

[Ceb12] Toni Cebrián. Book review: Steps in Scala: An introduction to Object-Functional Programming, by Christos K. K. Loverdos, Apostolos Syropoulos, Cambridge University Press, 2010, 504 pp., ISBN 0-521-74758-9. Journal of Functional Programming, 22(6):854– 855, November 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/92ACCF4945864075EA64FA6265B1F747.

Clements:2010:LLP

[CF10] John Clements and Kathi Fisler. “Little language” project modules. Journal of Functional Programming, 20(1):3–18, January 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0C29454A298651F53AE9378EB33573A7.

Carpenter:1996:SLB

[CG96] D. B. Carpenter and H. Glaser. Some lattice-based scientific problems, expressed in Haskell. Journal of Functional Programming, 6 (3):419–444, May 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/043C15CC02663588B429E0999050E61C.

Chuang:1997:SMF

[CG97] Tyng-Ruey Chuang and Benjamin Goldberg. A syntactic method for finding least fixed points of higher-order functions over finite domains. Journal of Functional Programming, 7(4):357–394, July 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4E5CCB9053B3D45425F348BBF17F4142.

Cheney:2018:SIP

[CG18] James Cheney and Torsten Grust. Special issue on programming languages for big data editorial. Journal of Functional Programming, 28:e8, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
special-issue-on-programming-languages-for-big-data-editorial/
4525FDEF1AE0B07F0D54B904696ADC62.

Caldwell:2020:TDA

[CGJF20] Sam Caldwell, Tony Garnock-Jones, and Matthias Felleisen. Typed dataspace actors. Journal of Functional Programming, 30(??):e27, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
typed-dataspace-actors/95C1CBE3CF43FDDEA3724EEDE67A53EF.

Colazzo:2006:SAP

[CGMS06] Dario Colazzo, Giorgio Ghelli, Paolo Manghi, and Carlo Sartiani. Static analysis for path correctness of XML queries. Journal of Functional Programming, 16(4–5):621–661, July 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/E18D9D538130677C3CECCE63D2AC24B2.

Chirimar:1996:RCC

[CGR96] Jawahar Chirimar, Carl A. Gunter, and Jon G. Riecke. Reference counting as a computational interpretation of linear logic. Journal of Functional Programming, 6(2):195–244, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
57AE85B618932EE0D716556774E73904.

Coquand:1994:TLC

[CH94a] Thierry Coquand and Hugo Herbelin. A-translation and looping combinators in pure type systems. Journal of Functional Programming, 4(1):77–88, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/A041A93B5965D84CBBE3D2CBD477AA5B.

Curien:1994:TPY

[CH94b] Pierre-Louis Curien and Thérèse Hardin. Theoretical pearl: Yet yet a counterexample for λ +SP. Journal of Functional Programming, 4(1):113–115, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/9C225C5A15E450FA268A2BA476622598.

Chin:1994:SFF

[Chi94] Wei-Ngan Chin. Safe fusion of functional expressions II: Further improvements. Journal of Functional Programming, 4(4):515–555, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3388FFCCA872572087B8D1678D752F44.

Chitil:2009:BRE

[Chi09] Olaf Chitil. Book review: Essentials of Programming Languages (third edition), by Friedman Daniel P. and Wand Mitchell, MIT Press, ISBN 978-0-262-06279-4, 2008. Journal of Functional Programming, 19(6):723–725, November 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/0BFF41EE0E99729831ADA16E47881520.

Chlipala:2008:MDC

[Chl08] Adam Chlipala. Modular development of certified program verifiers with a proof assistant 1, 2. Journal of Functional Programming, 18 (5):599–647, September 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/90ECD60D060DCCEBC81521A36FC565A0.

Chuang:2022:SMF

[CIZ22] Cheng-En Chuang, Grant Iraci, and Lukasz Ziarek. Send to me first: Priority in synchronous message-passing. Journal of Functional Programming, 32:e16, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
send-to-me-first-priority-in-synchronous-messagepassing/
F59BD7EF0FD493515F9AC3890E40A40B.

Consel:1995:LLP

[CK95] Charles Consel and Siau Cheng Khoo. On-line and off-line partial evaluation: semantic specifications and correctness proofs. Journal of Functional Programming, 5(4):461–500, October 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/297804C171F96758B178B748358B9672.

Chakravarty:2002:ICH

[CK02] Manuel M. T. Chakravarty and Gabriele C. Keller. An introduction to computing: with Haskell. French Forest, ????, 2002. ISBN 1-74009-404-2 (paperback). viii + 145 pp. LCCN BLI YK.2004.b.1796.

Chakravarty:2004:RBT

[CK04] Manuel M. T. Chakravarty and Gabriele Keller. The risks and benefits of teaching purely functional programming in first year. Journal of Functional Programming, 14(1):113–123, January 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
39F929A1793B67BCEF316DBDB717F273.

Crary:2005:MAI

[CKP05] Karl Crary, Aleksey Kliger, and Frank Pfenning. A monadic analysis of information flow security with mutable state. Journal of Functional Programming, 15(2):249–291, March 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/01A8F47B757460768168CAF18989F83A.

Carette:2009:FTP

[CKS09] Jacques Carette, Oleg Kiselyov, and Chung-Chieh Shan. Finally tagless, partially evaluated: Tagless staged interpreters for simpler typed languages. Journal of Functional Programming, 19(5):509– 543, September 2009. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/7B2DC44A2127EBBA71ADE63809D9425F.

Cardelli:1991:SBQ

[CL91] Luca Cardelli and Giuseppe Longo. A semantic basis for quest. Journal of Functional Programming, 1(4):417–458, October 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E42A0E3A3A5D903F017FA06AC5190E5C.

Clark:1996:FBR

[Cla96] A. N. Clark. Formal basis for the refinement of rule based transition systems. Journal of Functional Programming, 6(2):329–354, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6F70510B7B47DB7B3A54232BDC655EA5.

Claessen:1999:PMC

[Cla99] Koen Claessen. A poor man’s concurrency monad. Journal of Functional Programming, 9(3):313–323, May 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A369E310ADAE4455020C918FC1D47958.

Claessen:2004:FPP

[Cla04] Koen Claessen. Functional pearl: Parallel parsing processes. Journal of Functional Programming, 14(6):741–757, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0AF17481A41F2007752F530F07698139.

Clare:2015:RFS

[Cla15] Amanda Clare. Review of “A functional start to computing with Python”, Ted Herman, CRC Press, 2014, ISBN 978-1-4665-0455-4. Journal of Functional Programming, 25:e15, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
review-of-a-functional-start-to-computing-with-python-ted-
herman-crc-press-2014-isbn-9781466504554/CDD2895726478A185EDA78AE98C30BB3.

Chan:2023:STC

[CLB23] Jonathan Chan, Yufeng Li, and William J. Bowman. Is sized typing for Coq practical? Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
is-sized-typing-for-coq-practical/9B55C4ABE4CB011CBD6405CA291FD920.

Clerc:2016:OJJ

[Cle16] Xavier Clerc. OCaml-Java: The Java Virtual Machine as the target of an OCaml compiler. Journal of Functional Programming, 26:e7, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/ocamljava-the-
java-virtual-machine-as-the-target-of-an-ocaml-compiler/
EDA83983550B6025B3705E2F8D97EB81.

Convent:2020:DBD

[CLMM20] Lukas Convent, Sam Lindley, Conor Mcbride, and Craig Mclaughlin. Doo bee doo bee doo. Journal of Functional Programming, 30(??):e9, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
doo-bee-doo-bee-doo/DEC5F8FDABF7DE3088270E07392320DD.

Cousineau:1998:FAP

[CM98] Guy Cousineau and Michel Mauny. The functional approach to programming. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-57183-9 (hardcover), 0-521-57681-4 (paperback). xiv + 445 pp. LCCN QA76.62 .C68 1998. URL http://www.

loc.gov/catdir/description/cam028/97042892.html;   http:
//www.loc.gov/catdir/toc/cam026/97042892.html.

Curtis:2015:CLT

[CM15] Sharon Curtis and Shin-Cheng Mu. Calculating a linear-time solution to the densest-segment problem. Journal of Functional Programming, 25:e22, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
calculating-a-lineartime-solution-to-the-densestsegment-
problem/B7A7F3DB6A6220FB95F819F91BA94DCE.

Calcagno:2003:CTS

[CMS03] C. Calcagno, E. Moggi, and T. Sheard. Closed types for a safe imperative MetaML. Journal of Functional Programming, 13(3): 545–571, May 2003. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/155E678C81DAE8E2E945180BE177D414.

Collard:2003:RAP

[Col03] Jean-François Collard. Reasoning about program transformations: imperative programming and flow of data. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95391-4. xviii + 237 pp. LCCN QA76.7.C645 2003.

Courant:2007:MCP

[Cou07] Judicaël Courant. A module calculus for pure type systems. Journal of Functional Programming, 17(3):287–352, May 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
252EC49AA63D8439BAF5A74CDE42A1CE.

Crary:2020:FSA

[Cra20] Karl Crary. A focused solution to the avoidance problem. Journal of Functional Programming, 30(??):e24, ???? 2020. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/focused-solution-to-the-
avoidance-problem/F6F764A18625588897968919C2B2F4CC.

Crolard:1999:CCC

[Cro99] Tristan Crolard. A confluent λ-calculus with a catch/throw mechanism. Journal of Functional Programming, 9(6):625–647, November 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
69855F4EC7E940B0FB7B038832033C4F.

Codish:2016:PSS

[CS16] Michael Codish and Eijiro Sumii. Preface for special section from FLOPS 2014. Journal of Functional Programming, 26:e14, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
preface-for-special-section-from-flops-2014/3956976D0A10E8715C843581BC69F1E7.

Culpepper:2012:FM

[Cul12] Ryan Culpepper. Fortifying macros. Journal of Functional Programming, 22(4–5):439–476, September 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/2AC116FC3DD38075BDF61C464ED35507.

Curtis:2006:FPM

[Cur06] S. A. Curtis. Functional pearl: Marble mingling. Journal of Functional Programming, 16(2):129–136, March 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D5BFB0A762ABF9973398081F373515EF.

Crary:2002:IPT

[CWM02] Karl Crary, Stephanie Weirich, and Greg Morrisett. Intensional polymorphism in type-erasure semantics. Journal of Functional Programming, 12(6):567–600, November 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/B46353534BE6D97919C442C4BD326845.

Chen:2005:MPT

[CX05] Chiyan Chen and Hongwei Xi. Meta-programming through typeful code representation. Journal of Functional Programming, 15(6):797– 835, November 2005. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0CBD5BD0E394EA5981ECB8CF0F004129.

Downen:2014:DCC

[DA14] Paul Downen and Zena M. Ariola. Delimited control and computational effects. Journal of Functional Programming, 24(1):1–55, January 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
84C685AB5E4BD71EACD2E42A6594DF83.

Downen:2018:TCC

[DA18] Paul Downen and Zena M. Ariola. A tutorial on computational classical logic and the sequent calculus. Journal of Functional Programming, 28:e3, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/tutorial-on-computational-classical-
logic-and-the-sequent-calculus/4C5D37F74D95ED7CCFC0BB3E37F342A5.

Downen:2023:CCR

[DA23] Paul Downen and Zena M. Ariola. Classical (co)recursion: Mechanics. Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/classical-corecursion-mechanics/
9C61F6814C3CA0A70D4922B0DE40849E.

Dagand:2017:EO

[Dag17] Pierre-Evariste Dagand. The essence of ornaments. Journal of Functional Programming, 27:e9, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/essence-of-ornaments/4D2DF6F4FE23599C8C1FEA6C921A3748.

Danvy:1998:FU

[Dan98] Olivier Danvy. Functional unparsing. Journal of Functional Programming, 8(6):621–625, November 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/789945109AD2AB168B504472E6B786A0.

Danvy:2019:FLR

[Dan19] Olivier Danvy. Folding left and right over Peano numbers. Journal of Functional Programming, 29:e6, ???? 2019. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/folding-left-and-right-over-peano-numbers/
BF89C1D1283424DBF295ECFAB2ABE66C.

Danvy:2022:FUL

[Dan22] Olivier Danvy. Fold–unfold lemmas for reasoning about recursive programs using the Coq proof assistant. Journal of Functional Programming, 32:e13, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
foldunfold-lemmas-for-reasoning-about-recursive-programs-
using-the-coq-proof-assistant/CE55ACA989A03E9073540DCCEBA4E89C.

See erratum [Dan23a].

Danvy:2023:FUL

[Dan23a] Olivier Danvy. Fold–unfold lemmas for reasoning about recursive programs using the Coq proof assistant — erratum. Journal of Functional Programming, 33:??, ???? 2023. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/foldunfold-lemmas-for-reasoning-about-
recursive-programs-using-the-coq-proof-assistant-erratum/
240CE5C1CADFED0F92E1B5FD8E23D5DD. See [Dan22].

Danvy:2023:FLR

[Dan23b] Olivier Danvy. Folding left and right matters: Direct style, accumulators, and continuations. Journal of Functional Programming, 33: ??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/folding-left-
and-right-matters-direct-style-accumulators-and-continuations/
55377D4E137AFCF13A7922F3EFBCDCBF.

Davy:1995:PLC

[DD95] J. R. Davy and P. M. Dew. A polymorphic library for constructive solid geometry. Journal of Functional Programming, 5(3):415–442, July 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
77A754FA5CBA6583C70F9CB45572DEB5.

Damian:2003:CTF

[DD03a] Daniel Damian and Olivier Danvy. CPS transformation of flow information, part II: administrative reductions. Journal of Functional Programming, 13(5):925–933, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3DA1271C07DBA366BE4D997C96138FC2.

Damian:2003:SAP

[DD03b] Daniel Damian and Olivier Danvy. Syntactic accidents in program analysis: on the impact of the CPS transformation. Journal of Functional Programming, 13(5):867–904, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9D1BE850131ED8EC1E93FF086E84C86A.

DeJong:2011:BRF

[De 11] Willem De Jong. Book review: Foundations of F#, Pickering Robert, Apress, 2007 ISBN 10: 1-59059-757-5. Journal of Functional Programming, 21(2):217–218, March 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/AA9107CA673B6072D9CB7AB7320BAEE2.

Muijnck-Hughes:2020:TDD

[De 20] Jan De Muijnck-Hughes. Type-driven development with Idris. Journal of Functional Programming, 30(??):e16, ???? 2020. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/typedriven-development-with-idris/90AF5662A02C45CCDAFF8980F19A41D5.

Derrick:2004:BRP

[Der04] John Derrick. Book review: Programming Methodology. A. McIver and C. Morgan, editors, Springer-Verlag, 2002, ISBN 0-387-95349-

3. Journal of Functional Programming, 14(5):597–598, September 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B3360A8BF8ED76EC4E3932751B3EE8B7.

Donnelly:2008:TE

[DF08] Kevin Donnelly and Matthew Fluet. Transactional events. Journal of Functional Programming, 18(5):649–706, September 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
84D5EB6061CC0B2A2C73AC148A26EB6C.

Damiani:2000:AUC

[DG00] Ferruccio Damiani and Paola Giannini. Automatic useless-code elimination for HOT functional programs. Journal of Functional Programming, 10(6):509–559, November 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/C78E57AB79A5EF9C3B0C7AE2DD9D8CE9.

DalZilio:2002:RAC

[DG02] Silvano Dal Zilio and Andrew D. Gordon. Region analysis and a π-calculus with groups. Journal of Functional Programming, 12(3): 229–292, May 2002. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/47E03F3644192C26732CDBD1AA210952.

Despeyroux:2000:SIL

[DH00] Joëlle Despeyroux and Robert Harper. Special issue on logical frameworks and metalanguages. Journal of Functional Programming, 10(1):135–136, January 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL http//www-sop.inria.

fr/certilab/LFM00/cfp-jfp.html;   https://www.cambridge.
org/core/product/1D1E430960CA01946A231DD0C8D44BC1.

Dieterle:2016:SCV

[DHLB16] M. Dieterle, T. Horstmeyer, R. Loogen, and J. Berthold. Skeleton composition versus stable process systems in Eden. Journal of Functional Programming, 26:e11, ???? 2016. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/skeleton-composition-versus-stable-
process-systems-in-eden/0F108B29BA5AC5CA666DB8D2639A3017.

DiPierro:2005:QSA

[DHW05] Alessandra Di Pierro, Chris Hankin, and Herbert Wiklicky. Quantitative static analysis of distributed systems. Journal of Functional Programming, 15(5):703–749, September 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D50D655CD96D4DB9400563B9BE7F1D82.

DiCosmo:1993:DTI

[Di 93] Roberto Di Cosmo. Deciding type isomorphisms in a typeassignment framework. Journal of Functional Programming, 3(4): 485–525, October 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/5FEA30D55D009A2AA04036A30D5382F8.

Danner:2022:DSF

[DL22] Norman Danner and Daniel R. Licata. Denotational semantics as a foundation for cost recurrence extraction for functional languages. Journal of Functional Programming, 32:e8, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/denotational-
semantics-as-a-foundation-for-cost-recurrence-extraction-
for-functional-languages/BEC8FAB23FECD752D2968FD000E7C3C8.

Dolstra:2010:NPF

[DLP10] Eelco Dolstra, Andres Löh, and Nicolas Pierron. NixOS: A purely functional Linux distribution. Journal of Functional Programming, 20(5–6):577–615, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/C1ACBA2A51D2E5466820F5B5086EA2CE.

David:2005:ICE

[DM05] R. David and G. Mounier. An intuitionistic λ-calculus with exceptions. Journal of Functional Programming, 15(1):33–52, January 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C1E24D3CCF38A980E6E4166D468BE920.

Dagand:2014:TFA

[DM14] Pierre-Évariste Dagand and Conor McBride. Transporting functions across ornaments. Journal of Functional Programming, 24(2– 3):316–383, May 2014. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/86A539E953412C85452ADA9E39E67EB7.

Danvy:2007:OPC

[DMN07] Olivier Danvy, Kevin Millikin, and Lasse R. Nielsen. On onepass CPS transformations. Journal of Functional Programming, 17 (6):793–812, November 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/FA023612B1FAB59194D6AD8884135FDC.

Dreyer:2012:IHO

[DNB12] Derek Dreyer, Georg Neis, and Lars Birkedal. The impact of higher-order state and control effects on local relational reasoning. Journal of Functional Programming, 22(4–5):477–528, September 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
DA921680527426F49ED79644C8C6A565.

Duggan:2002:TCM

[DO02] Dominic Duggan and John Ophel. Type-checking multi-parameter type classes. Journal of Functional Programming, 12(2):133–158, March 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EABE479DF71BA3B31843E1518E5FA265.

dePaiva:2001:BRD

[dP01] Valeria de Paiva. Book review: Domains and Lambda-Calculi, by R. M. Amadio and P.-L. Curien, Springer-Verlag, 1999, ISBN 0-521-58775-1. Journal of Functional Programming, 11(5):589, September 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
FF3CC1C45998ABE60941BAA6D9B7E97B.

DeLaEncina:2009:NSC

[DP09] Alberto De La Encina and Ricardo Peña. From natural semantics to C: A formal derivation of two STG machines. Journal of Functional Programming, 19(1):47–94, January 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/0713809339C4CCFFAFB8E793F9F350D3.

Devriese:2012:FTO

[DP12] Dominique Devriese and Frank Piessens. Finally tagless observable recursion for an abstract grammar model. Journal of Functional Programming, 22(6):757–796, November 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7FED2F29DBD2740E00C796241BFBF94E.

Dreyer:2022:BPS

[DP22] Derek Dreyer and Benjamin C. Pierce. On being a PhD student of Robert Harper. Journal of Functional Programming, 32(??):e3, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
on-being-a-phd-student-of-robert-harper/B68BCF31384084D9FDCB17A7DF1DE7A0.

Dyvbig:2007:MFD

[DPS07] R. Kent Dyvbig, Simon Peyton Jones, and Amr Sabry. A monadic framework for delimited continuations. Journal of Functional Programming, 17(6):687–730, November 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D99D1394370DFA8EA8428D552B5D8E7E.

Dreyer:2007:RTG

[Dre07] Derek Dreyer. Recursive type generativity. Journal of Functional Programming, 17(4–5):433–471, July 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/897191BEE5243EEE18FE5738728350D3.

Danvy:2001:NET

[DRR01] Olivier Danvy, Morten Rhiger, and Kristoffer H. Rose. Normalization by evaluation with typed abstract syntax. Journal of Functional Programming, 11(6):673–680, November 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/EB9C84EDC74834E940111BC94D2E0481.

DeHoon:1995:IFS

[DRvE95] Walter A. C. A. J. De Hoon, Luc M. W. J. Rutten, and Marko C. J. D. van Eekelen. Implementing a functional spreadsheet in clean. Journal of Functional Programming, 5(3):383–414, July 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9B7355B606D0F77AD03C308C70B4FF9D.

Dreyer:2016:SID

[DS16] Derek Dreyer and Mary Sheeran. Special issue dedicated to ICFP 2014: Editorial. Journal of Functional Programming, 26:e20, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/special-issue-
dedicated-to-icfp-2014-editorial/014390426F6D74E82AC2D68EFDEF4DD3.

Dagand:2018:FDI

[DTT18] Pierre-Évariste Dagand, Nicolas Tabareau, and Éric Tanter. Foundations of dependent interoperability. Journal of Functional Programming, 28:e9, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
foundations-of-dependent-interoperability/801E5D733DE9BC30B1A34A5FDC621242.

Dunfield:2014:EIU

[Dun14] Joshua Dunfield. Elaborating intersection and union types. Journal of Functional Programming, 24(2–3):133–165, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/F5922703222B5F9C04C46BAE7AB8F7EB.

Duponcheel:1995:BRC

[Dup95] Luc Duponcheel. Book review: Concurrent Programming in Erlang, by Armstrong Joe, Virding Robert and Williams Mike, Prentice Hall, 1993, 281 pp. Journal of Functional Programming, 5(4): 653–660, October 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/3E436B38A9A927DC045CD0B1AB80DA46.

Darais:2019:CGC

[DV19] David Darais and David Van Horn. Constructive Galois connections. Journal of Functional Programming, 29(??):e11, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/constructive-
galois-connections/216F5350F896AB024486361287C08489.

Doets:2004:HRL

[DvJE04] Kees Doets and J. van (Jan) Eijck. The Haskell road to logic, maths and programming, volume 4 of Texts in computing. King’s College Publications, London, UK, 2004. ISBN 0-9543006-9-6. x + 426 pp. LCCN QA9 DOE.

Dyckhoff:2001:BRC

[Dyc01] Roy Dyckhoff. Book review: Computational Logic, by Ulrich Berger and Helmut Schwichtenberg, editors, Springer-Verlag, 1999, 444 pp. Journal of Functional Programming, 11(4):433–436, July 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2C9DC025F44D11832888C7B21E8F446E.

Erwig:2006:GPG

[EAKC06] Martin Erwig, Robin Abraham, Steve Kollmansberger, and Irene Cooperstein. Gencel: a program generator for correct spreadsheets. Journal of Functional Programming, 16(3):293–325, May 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D61E245196C3F56071F89B07F904E0EF.

Edmonds:2008:HTA

[Edm08] Jeff Edmonds. How to Think about Algorithms. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-84931-4 (hardcover), 0-521-61410-4 (paperback), 0-511-41278-9 (e-book), 0-511-41370-X (e-book). xiii + 448 pp. LCCN QA9.58 .E36

2008. URL http://www.loc.gov/catdir/enhancements/fy0808/
2008001238-b.html;   http://www.loc.gov/catdir/enhancements/
fy0808/2008001238-d.html;   http://www.loc.gov/catdir/enhancements/
fy0808/2008001238-t.html.

Elliott:2003:CEL

[EFD03] Conal Elliott, Sigbjørn Finne, and Oege De Moor. Compiling embedded languages. Journal of Functional Programming, 13(3):455–481, May 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4B0A7526CC16907F445CCF27277E9B9B.

Elsman:2021:IRM

[EH21] Martin Elsman and Niels Hallenberg. Integrating region memory management and tag-free generational garbage collection. Journal of Functional Programming, 31(??):e4, ???? 2021. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/integrating-region-memory-management-
and-tagfree-generational-garbage-collection/782D317A9B811CD99FA0E924A35B6A58.

Endrullis:2014:CSE

[EHBR14] Jörg Endrullis, Dimitri Hendriks, Rena Bakhshi, and Grigore Rosu. On the complexity of stream equality. Journal of Functional Programming, 24(2–3):166–217, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/34CF8DEA07BCC5FEF11008CD79E07626.

Erwig:2006:FPP

[EK06] Martin Erwig and Steve Kollmansberger. Functional pearls: Probabilistic functional programming in Haskell. Journal of Functional Programming, 16(1):21–34, January 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/8DF2F2B191C68260A961BA3A9F50B332.

Erwig:2021:EDP

[EK21] Martin Erwig and Prashant Kumar. Explainable dynamic programming. Journal of Functional Programming, 31(??):e10, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
explainable-dynamic-programming/E8DC46B3363847A2BA18B345C8591DAA.

Ernoult:1995:USA

[EM95] Christine Ernoult and Alan Mycroft. Untyped strictness analysis. Journal of Functional Programming, 5(1):37–49, January 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6F10D7549B2CBB39789970907D954380.

Erwig:1998:DFS

[Erw98] Martin Erwig. Diets for fat sets. Journal of Functional Programming, 8(6):627–632, November 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/A4574A4130665EE6CF4FAE35FD302469.

Erwig:2001:IGF

[Erw01] Martin Erwig. Inductive graphs and functional graph algorithms. Journal of Functional Programming, 11(5):467–492, September 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2210F7C31A34EA4CF5008ED9E7B4EF62.

Erwig:2004:EZE

[Erw04] Martin Erwig. Escape from Zurg: an exercise in logic programming. Journal of Functional Programming, 14(3):253–261, May 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1444B442C9BEB3E705691E711D50ADAE.

Eriksen:2015:CSR

[ESM15] Marius Eriksen, Michael Sperber, and Anil Madhavapeddy. CUFP’13 scribe’s report. Journal of Functional Programming, 25:e3, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
cufp13-scribes-report/F38AAE60DA9AD95E1737E3F863075C13.

Felleisen:2013:RRL

[F^{+}13] Matthias Felleisen et al. Realm of Racket: learn to program, one game at a time! No Starch Press, San Francisco, CA, USA, 2013. ISBN 1-59327-491-2. LCCN QA76.73.R33 R43 2013eb.

Faxen:2002:SSH

[Fax02] Karl-Filip Faxén. A static semantics for Haskell. Journal of Functional Programming, 12(4–5):295–357, July 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/9D90E0C7DE8DA7D6BAEAC5143E658E1D.

Hou:2017:CCP

[FBH17] Kuen-Bang Hou (Favonia), Nick Benton, and Robert Harper. Correctness of compiling polymorphism to dynamic typing. Journal of Functional Programming, 27:e1, ???? 2017. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/correctness-of-compiling-
polymorphism-to-dynamic-typing/AE529D08033550814B59D305CCE5A6A2.

Flatt:2012:MWT

[FCDF12] Matthew Flatt, Ryan Culpepper, David Darais, and Robert Bruce Findler. Macros that work together: Compile-time bindings, partial expansion, and definition contexts. Journal of Functional Programming, 22(2):181–216, March 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/375043C6746405B22014D235FA4C90C3.

Findler:2002:DPE

[FCF^{+}02] Robert Bruce Findler, John Clements, Cormac Flanagan, Matthew Flatt, Shriram Krishnamurthi, Paul Steckler, and Matthias Felleisen. DrScheme: a programming environment for Scheme. Journal of Functional Programming, 12(2):159–182, March 2002. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/A3F8F65C27FF86D9CEB63B61DBB3EEC6.

Florido:2004:LLC

[FD04] Mário Florido and Luís Damas. Linearization of the lambda-calculus and its relation with intersection type systems. Journal of Functional Programming, 14(5):519–546, September 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/6D82E41D080B2E762AD82FC810315F62.

Fegaras:2017:ADB

[Feg17] Leonidas Fegaras. An algebra for distributed Big Data analytics. Journal of Functional Programming, 27:e27, ???? 2017. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/an-algebra-for-distributed-big-data-
analytics/580E6A327FD3AC8E4F9B0EA0408B434B.

Felleisen:2003:WEP

[Fel03] Matthias Felleisen. Welcome to the educational pearls column. Journal of Functional Programming, 13(5):835–838, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
DC05D831A73F2E216E8927837CFBB3F1.

Felleisen:2012:E

[Fel12] Matthias Felleisen. Editorial. Journal of Functional Programming, 22(3):223–224, May 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/F590A9F6C4FA5235E6B1798802FEBE49.

Felleisen:2014:E

[Fel14] Matthias Felleisen. Editorial. Journal of Functional Programming, 24(6):611, November 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/02392626D7555E55B6A29BCFA3474E1C.

Fernandez:2004:PLO

[Fer04] Maribel Fernández. Programming languages and operational semantics: an introduction, volume 1 of Texts in computing. King’s College Publications, London, UK, 2004. ISBN 0-9543006-3-7 (paperback). vi + 142 pp. LCCN QA76.7.

Fernandez:2006:PLO

[Fer06] Maribel Fernández. Programming languages and operational semantics: an introduction. King’s College Publications, London, UK, 2006. ISBN 0-9543006-3-7. vi + 142 pp. LCCN ????

Flanagan:1999:SFA

[FF99] C. Flanagan and M. Felleisen. The semantics of future and an application. Journal of Functional Programming, 9(1):1–31, January 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F41EA83BB29440A7B3350DA0A94A9F41.

Findler:2006:SFP

[FF06] Robert Bruce Findler and Matthew Flatt. Slideshow: functional presentations. Journal of Functional Programming, 16(4–5):583–619, July 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
8B86402C62CCC27E4A8ED347E2C5155B.

Felleisen:2004:SIC

[FFFK04] Matthias Felleisen, Robert Bruce Findler, Matthew Flatt, and Shriram Krishnamurthi. The structure and interpretation of the computer science curriculum. Journal of Functional Programming, 14 (4):365–378, July 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/49C3FE2372A32C9B27BD60767F8CB05C.

Fischbach:2003:SCL

[FH03] Adam Fischbach and John Hannan. Specification and correctness of lambda lifting. Journal of Functional Programming, 13(3):509–543, May 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1788816553C0B29C2C9BB0E075D2F99D.

Ferreira:1998:TWB

[FHJ98] William Ferreira, Matthew Hennessy, and Alan Jeffrey. A theory of weak bisimulation for core CML. Journal of Functional Programming, 8(5):447–491, September 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7DE2F97B777344441D6DEDBD525780FB.

Fridlender:2000:DWN

[FI00] Daniel Fridlender and Mia Indrika. Do we need dependent types? Journal of Functional Programming, 10(4):409–415, July 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
374DA8854D388E986FE35859390C1C4D.

Filliatre:2003:VNF

[Fil03] Jean-Christophe Filliâtre. Verification of non-functional programs using interpretations in type theory. Journal of Functional Programming, 13(4):709–745, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/DC4EC7D2F45D032135ADB8432933D0FD.

Filinski:2021:PDP

[Fil21] Andrzej Filinski. Proof-directed program transformation: a functional account of efficient regular expression matching. Journal of Functional Programming, 31(??):e12, ???? 2021. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/proofdirected-program-transformation-
a-functional-account-of-efficient-regular-expression-matching/
454BB5CD9B0B056FA91957F2F9CC3EC5.

Fisher:2006:I

[Fis06] Kathleen Fisher. Introduction. Journal of Functional Programming, 16(4–5):373–374, July 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C53F79B43E3509D738F1BE290BC7C1F1.

Fischer:2011:PFL

[FKS11] Sebastian Fischer, Oleg Kiselyov, and Chung-Chieh Shan. Purely functional lazy nondeterministic programming. Journal of Functional Programming, 21(4–5):413–465, September 2011. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/1E8BA117E549A9612BC4AF9804E5507A.

Friedmann:2011:MBD

[FL11] Oliver Friedmann and Martin Lange. More on balanced diets. Journal of Functional Programming, 21(2):135–157, March 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C1B1BDEE6823CCD3190B1C7A4855127A.

Fradet:2000:CSF

[FM00] Pascal Fradet and Julien Mallet. Compilation of a specialized functional language for massively parallel computers. Journal of Functional Programming, 10(6):561–605, November 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BF4B9B8C427AEB52A57377DECD78BEC6.

Fluet:2006:MR

[FM06] Matthew Fluet and Greg Morrisett. Monadic regions. Journal of Functional Programming, 16(4–5):485–545, July 2006. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/DD56B542BFFEB2F3F8206EBB95FE777B.

Fokkinga:2011:HT

[Fok11] Maarten Fokkinga. The Hough transform. Journal of Functional Programming, 21(2):129–133, March 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/DEF0A0FBC42F6782F92120CADA72437A.

Fourman:1996:BRP

[Fou96] Michael Fourman. Book reviews: Programming with Standard ML, by Myers Colin, Clack Chris and Poon Ellen, Prentice Hall International, Inc., New Jersey, 301 pp., 1993, ISBN 0-13-722075-8. ML for the Working Programmer, by Paulson, L. C., Cambridge University Press, 1991, 429 pp., ISBN 0-521-39022-2. Elements of ML Programming, by Ullman, Jeffrey D., Prentice Hall International, Inc., New Jersey, 1994, 320 pp., ISBN 0-13-288788-6, 0-13-184854-2 (USA). Journal of Functional Programming, 6(1):189–193, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D30599E898CC0B7A66C70330A6ABD6E4.

Filliatre:2003:PAI

[FP03] Jean-Christophe Filliâtre and François Pottier. Producing all ideals of a forest, functionally. Journal of Functional Programming, 13 (5):945–956, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/153FB8EA081D57838EE48E7183F47F9A.

Fluet:2006:PTS

[FP06] Matthew Fluet and Riccardo Pucella. Phantom types and subtyping. Journal of Functional Programming, 16(6):751–791, November 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
08E1C18BA8C61F0EDF70EFD4051604E5.

Flatt:2007:P

[FP07] Matthew Flatt and Benjamin C. Pierce. Preface. Journal of Functional Programming, 17(4–5):431, July 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/E7DD9AD5D7E14E9190C284F746D96E6B.

Fridlender:2015:PTS

[FP15] Daniel Fridlender and Miguel Pagano. Pure type systems with explicit substitutions. Journal of Functional Programming, 25: e19, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
journals/journal-of-functional-programming/article/pure-
type-systems-with-explicit-substitutions/DB9536503A77C1B2D8357588D8DF31DE.

Fluet:2010:ITP

[FRRS10] Matthew Fluet, Mike Rainey, John Reppy, and Adam Shaw. Implicitly threaded parallelism in Manticore. Journal of Functional Programming, 20(5–6):537–576, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/9F8A993EE51C782F80AB993F05DCA215.

Friedman:2002:CLP

[FS02] Daniel P. Friedman and Amr Sabry. CPS in little pieces: composing partial continuations. Journal of Functional Programming, 12(6):617–622, November 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/A704CEDFCFBC2DF259D269B584FC038F.

Fischer:2003:ASG

[FS03] Bernd Fischer and Johann Schumann. AutoBayes: a system for generating data analysis programs from statistical models. Journal of Functional Programming, 13(3):483–508, May 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
00956DF4D7E5C4362C86FDF8B28F3D7B.

Fisher:2008:BLT

[FS08] David Fisher and Olin Shivers. Building language towers with Ziggurat. Journal of Functional Programming, 18(5):707–780, September 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0930B7E8DC1B3CF55E5701814E1A590E.

Frankau:2009:CUG

[FSNB09] Simon Frankau, Diomidis Spinellis, Nick Nassuphis, and Christoph Burgard. Commercial uses: Going functional on exotic trades. Journal of Functional Programming, 19(1):27–45, January 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/C151F30DBC7391CE0E4D7091DD9784E9.

Friedman:2008:EPL

[FW08] Daniel P. Friedman and Mitchell Wand. Essentials of programming languages. MIT Press, Cambridge, MA, USA, third edition, 2008. ISBN 0-262-06279-8 (hardcover). xxii + 410 pp.

LCCN QA76.7 .F73 2008. URL http://www.loc.gov/catdir/toc/
ecip081/2007039723.html.

Friedman:2001:EPL

[FWH01] Daniel P. Friedman, Mitchell Wand, and Christopher Thomas Haynes. Essentials of programming languages. MIT Press, Cambridge, MA, USA, second edition, 2001. ISBN 0-262-06217-8 (hardcover). xviii + 389 pp. LCCN QA76.7 .F73 2001.

Gammie:2009:BRR

[Gam09] Peter Gammie. Book review: Roy Peter Van and Haridi Seif. Concepts, Techniques, and Models of Computer Programming. The MIT Press, 2004. ISBN: 0-262-22069-5 Price $70. 930 pp. Journal of Functional Programming, 19(2):254–256, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7612DCD7AF920631E2A4A56740C6244B.

Gammie:2011:BRL

[Gam11a] Peter Gammie. Book review: Lambda-Calculus and Combinators: An Introduction, Second Edition, by Hindley, J. R. and Seldin, J. P. Journal of Functional Programming, 21(1):115–117, January 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
09F380E186CC6DE4A1CE8F6A4F3EF1AD.

Gammie:2011:SNS

[Gam11b] Peter Gammie. Short note: Strict unwraps make workerg/wrapper fusion totally correct. Journal of Functional Programming, 21(2): 209–213, March 2011. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/63A83354E6AD110597493C50692DCD81.

Gilray:2018:AAU

[GAM18] Thomas Gilray, Michael D. Adams, and Matthew Might. Abstract allocation as a unified approach to polyvariance in control-flow analyses. Journal of Functional Programming, 28:e18, ???? 2018. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/abstract-allocation-as-a-
unified-approach-to-polyvariance-in-controlflow-analyses/
1555B0130B5B00DFFCEDA327DF0EB232.

Gulias:2005:VDV

[GBF05] Victor M. Gulias, Miguel Barreiro, and Jose L. Freire. VoDKA: Developing a video-on-demand server using distributed functional programming. Journal of Functional Programming, 15(3):403–430, May 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F2728700B4F86CBF2110DF6DE8C2F265.

Gibbons:2003:FP

[GdM03] Jeremy Gibbons and Oege de Moor, editors. The fun of programming. Cornerstones of computing. Palgrave Macmillan, Basingstoke, 2003. ISBN 1-4039-0772-2, 0-333-99285-7 (PBK.). viii + 277 pp. LCCN QA76.62.F858 2003. £24.99.

Gibbons:2009:EIP

[GdSO09] Jeremy Gibbons and Bruno C. d. S. Oliveira. The essence of the iterator pattern. Journal of Functional Programming, 19(3–4):377–402, July 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3FC26EB2A63E6A2B29E07B9F0D5C5BCD.

Geser:1999:PFP

[GG99] Alfons Geser and Sergei Gorlatch. Parallelizing functional programs by generalization. Journal of Functional Programming, 9(6):649– 673, November 1999. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/5459217AA3E3850B8BAB23F884C5B1FB.

Gill:2009:WWT

[GH09] Andy Gill and Graham Hutton. The worker/wrapper transformation. Journal of Functional Programming, 19(2):227–251, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
75629BBEDB11D8463553A09BF5DEA235.

Gordon:1999:CEI

[GHL99] Andrew D. Gordon, Paul D. Hankin, and Søren B. Lassen. Compilation and equivalence of imperative objects. Journal of Functional Programming, 9(4):373–426, July 1999. CODEN JFPRES.

ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/1AA6AF3D5AB7A4270503884EFF380211.

Graversen:2024:ABP

[GHM24] Eva Graversen, Andrew K. Hirsch, and Fabrizio Montesi. Alice or Bob?: Process polymorphism in choreographies. Journal of Functional Programming, 34:??, ???? 2024. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/alice-or-bob-process-polymorphism-in-
choreographies/382AD3B58F86FF95AB59DDF0EDE96F65.

Gibbons:1996:FPD

[Gib96a] Jeremy Gibbons. Functional pearls: Deriving tidy drawings of trees. Journal of Functional Programming, 6(3):535–562, May 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F3F2023B7C2416D2FA6CB04485B18B4E.

Gibbons:1996:FPT

[Gib96b] Jeremy Gibbons. Functional pearls: The third homomorphism theorem. Journal of Functional Programming, 6(4):657–665, July 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
8DB1295DC8EB3C4942FB635C439B016E.

Gibbons:1999:PDR

[Gib99] Jeremy Gibbons. A pointless derivation of radix sort. Journal of Functional Programming, 9(3):339–346, May 1999. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/398024F932C04F26CD69ACF9E1E6E734.

Gibbons:2010:E

[Gib10] Jeremy Gibbons. Editorial. Journal of Functional Programming, 20 (1):1, January 2010. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/D55C4C2B9EEF2B3328255F72FACFDA50.

Gibbons:2021:HDC

[Gib21] Jeremy Gibbons. How to design co-programs. Journal of Functional Programming, 31(??):e15, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/how-to-design-coprograms/02B37487DB71F1CFD8B46BBE7823E440.

Gomard:1991:PEU

[GJ91] Carsten K. Gomard and Neil D. Jones. A partial evaluator for the untyped lambda-calculus. Journal of Functional Programming, 1(1): 21–69, January 1991. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/EE324F936F0A009B6766B13FF6755DFC.

Ghani:2007:MAG

[GJ07] Neil Ghani and Patricia Johann. Monadic augment and generalised short cut fusion. Journal of Functional Programming, 17(6):731– 776, November 2007. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/B9F2ABDD010E6521AC500CE66BA14FA9.

Garcia:2007:ECS

[GJL^{+}07] Ronald Garcia, Jaakko Jarvi, Andrew Lumsdaine, Jeremy Siek, and Jeremiah Willcock. An extended comparative study of language support for generic programming. Journal of Functional Programming, 17(2):145–205, March 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C97D5964ECC2E651EEF9A70BC50600A6.

Gartner:1996:RIC

[GK96] Dietmar Gärtner and Werner E. Kluge. π-RED + an interactive compiling graph reduction system for an applied λcalculus. Journal of Functional Programming, 6(5):723–756, September 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
62EC6AD594AB3E2892594304C5F9B032.

Gibbons:2022:E

[GK22] Jeremy Gibbons and Shriram Krishnamurthi. Editorial. Journal of Functional Programming, 32(??):e1, ???? 2022. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/editorial/6AB193BF2D5CF1B72328A2AF0D92C5C0.

Greve:2008:EEA

[GKM^{+}08] David A. Greve, Matt Kaufmann, Panagiotis Manolios, J. Strother Moore, Sandip Ray, José Luis Ruiz-Reina, Rob Sumners, Daron Vroon, and Matthew Wilding. Efficient execution in an automated reasoning environment. Journal of Functional Programming, 18(1):

15–46, January 2008. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/CA24027BF9359A7B0DF2A06CCED49DEA.

Grygiel:2013:CGL

[GL13] Katarzyna Grygiel and Pierre Lescanne. Counting and generating lambda terms. Journal of Functional Programming, 23(5):594– 628, September 2013. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/9A0F5D0A36BBDC81914117DC30AD5A53.

Grygiel:2015:CGT

[GL15] Katarzyna Grygiel and Pierre Lescanne. Counting and generating terms in the binary lambda calculus. Journal of Functional Programming, 25:e24, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
counting-and-generating-terms-in-the-binary-lambda-calculus/
47DE83E8BD697326F0FFD43351E083E3.

Gibbons:2006:FPE

[GLB06] Jeremy Gibbons, David Lester, and Richard Bird. Functional pearl: Enumerating the rationals. Journal of Functional Programming, 16 (3):281–291, May 2006. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/23B95751CFFFDA20CD810F561E23CB08.

Gapeyev:2002:RSR

[GLP02] Vladimir Gapeyev, Michael Y. Levin, and Benjamin C. Pierce. Recursive subtyping revealed. Journal of Functional Programming, 12 (6):511–548, November 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/E4A5C810FDEB5B865381D8F890A1623F.

Gluck:1994:GS

[Glü94] Robert Glück. On the generation of specializers. Journal of Functional Programming, 4(4):499–514, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/0A9174548FCE5383C67CCAC1D542D284.

Gordon:1993:IHT

[GM93] Michael J. C. Gordon and T. F. (Tom F.) Melham, editors. Introduction to HOL: a theorem proving environment for higher order logic. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-44189-7 (spiral). xx + 472 pp.

LCCN QA76.9.A96 I6 1993. URL http://www.loc.gov/catdir/
description/cam026/94140201.html;   http://www.loc.gov/
catdir/toc/cam029/94140201.html.

Giesl:2004:TTC

[GM04] Jürgen Giesl and Aart Middeldorp. Transformation techniques for context-sensitive rewrite systems. Journal of Functional Programming, 14(4):379–427, July 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C2B83A3F89C5E6ED2E3C064438336048.

Germane:2014:DCR

[GM14] Kimball Germane and Matthew Might. Deletion: The curse of the red-black tree. Journal of Functional Programming, 24(4):423–433, July 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
471C92AF3D431403FEE6C66FE070C492.

Grundy:2006:RFL

[GMO06] Jim Grundy, Tom Melham, and John O’Leary. A reflective functional language for hardware design and theorem proving. Journal of Functional Programming, 16(2):157–196, March 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
FB955BDD939520270BA0B07796CD08FC.

Geuvers:1991:MPS

[GN91] Herman Geuvers and Mark-Jan Nederhof. Modular proof of strong normalization for the calculus of constructions. Journal of Functional Programming, 1(2):155–189, April 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/348B6914C707F5282ED91E08AE47BDB8.

Goldberg:2000:AEL

[Gol00] Mayer Goldberg. An adequate and efficient left-associated binary numeral system in the λ-calculus. Journal of Functional Programming, 10(6):607–623, November 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/FD3368D7BE39F3E82A171766AB7189D1.

Gordon:1996:BRI

[Gor96] Andrew D. Gordon. Book reviews: An Introduction to Formal Specification and Z, by Potter Ben, Sinclair Jane and Till David, Prentice-Hall, 1991. Z: An Introduction to Formal Methods (second edition) by Diller Antoni, John Wiley & Sons, 1994. Journal of Functional Programming, 6(2):375–377, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F30892982A4DC7797F3E5396D58D5869.

Gordon:1998:HOO

[GP98] Andrew D. Gordon and A. M. (Andrew M.) Pitts, editors. Higher order operational techniques in semantics, volume 12 of Publications of the Newton Institute. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63168-8 (hardcover).

LCCN QA76 .H479 1998. URL http://www.loc.gov/catdir/
description/cam028/97032339.html;   http://www.loc.gov/
catdir/toc/cam027/97032339.html.

Garcia-Perez:2019:FRK

[GPN19] Álvaro García-Pérez and Pablo Nogueira. The full-reducing Krivine abstract machine KN simulates pure normal-order reduction in lockstep: A proof via corresponding calculus. Journal of Functional Programming, 29:e7, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
fullreducing-krivine-abstract-machine-kn-simulates-pure-
normalorder-reduction-in-lockstep-a-proof-via-corresponding-
calculus/F9D6AC47F4C2FC0903CD28AB451B37EC.

Greenberg:2012:CMM

[GPW12] Michael Greenberg, Benjamin C. Pierce, and Stephanie Weirich. Contracts made manifest. Journal of Functional Programming, 22 (3):225–274, May 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/223F83D633FC9075D12274FC086F2EFD.

Greiner:1996:WPC

[Gre96] John Greiner. Weak polymorphism can be sound. Journal of Functional Programming, 6(1):111–141, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4FA7CFD20FDD996954155E84846AC1A5.

Grelck:2005:SMM

[Gre05] Clemens Grelck. Shared memory multiprocessor support for functional array processing in SAC. Journal of Functional Programming, 15(3):353–401, May 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/1B90594F5E1399E6A3756134F3BDB63D.

Green:2014:BRB

[Gre14] Robin Green. Book review: Beginning Haskell, by Alejandro Serrano Mena, Apress, New York City, NY, 2014, ISBN-10: 1-4302-6250-8, 428 pp. Journal of Functional Programming, 24(2–3):421–422, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EE1E701FB3EE176FDC0C4B37DCFA2558.

Grust:2015:TFH

[Gru15] Torsten Grust. Book review: Thinking Functionally with Haskell by Richard Bird, Cambridge University Press, 2014. Journal of Functional Programming, 25:e2, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/thinking-functionally-with-haskell-
by-richard-bird-cambridge-university-press-2014/9E3EB0C8A858BE50E1E67CBA75C50BEC.

Grust:2020:RAD

[Gru20] Torsten Grust. Review of “Algorithm Design with Haskell” by Richard Bird and Jeremy Gibbons, Cambridge University Press, 2020. Journal of Functional Programming, 30(??):e29, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
review-of-algorithm-design-with-haskell-by-richard-bird-
and-jeremy-gibbons-cambridge-university-press-2020/60310B41A83CFB91E7C6A3E4312B98B0.

Gallesio:2003:PGU

[GS03] Erick Gallesio and Manuel Serrano. Programming graphical user interfaces with Scheme. Journal of Functional Programming, 13 (5):839–866, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/73D29423791429E4ED5B161A425489D8.

Gallesio:2005:SFA

[GS05] Erick Gallesio and Manuel Serrano. Skribe: a functional authoring language. Journal of Functional Programming, 15(5):751–770, September 2005. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/689EF8960521587B23203A56C46D2592.

Gustafsson:2006:EMB

[GS06] Per Gustafsson and Konstantinos Sagonas. Efficient manipulation of binary data using pattern matching. Journal of Functional Programming, 16(1):35–74, January 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/5FC43829285903225B3BD0A52B1AD975.

Grant:1996:SMR

[GSWZ96] P. W. Grant, J. A. Sharp, M. F. Webster, and X. Zhang. Sparse matrix representations in a functional language. Journal of Functional Programming, 6(1):143–170, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/669431E9C12EDC16F02603D833FAC31B.

Greenman:2019:HEP

[GTN^{+}19] Ben Greenman, Asumu Takikawa, Max S. New, Daniel Feltey, Robert Bruce Findler, Jan Vitek, and Matthias Felleisen. How to evaluate the performance of gradual type systems. Journal of Functional Programming, 29:e4, ???? 2019. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/how-to-evaluate-the-performance-of-
gradual-type-systems/DC765724C52A3A462F16C7FB3AD18697.

Guillaume:2000:CDP

[Gui00] Bruno Guillaume. The λs_{e}-calculus does not preserve strong normalisation. Journal of Functional Programming, 10(4):321–325, July 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E331421A97E0CAF4BA06790C2D1FE488.

Gay:2010:LTT

[GV10] Simon J. Gay and Vasco T. Vasconcelos. Linear type theory for asynchronous session types. Journal of Functional Programming, 20 (1):19–50, January 2010. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/605DA26DDDE9B1CCD35D33D0D62DF20C.

Gotsman:2013:MVP

[GY13] Alexey Gotsman and Hongseok Yang. Modular verification of preemptive OS kernels. Journal of Functional Programming, 23(4): 452–514, July 2013. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/B7B82B068880AD05DD5067F0A2994558.

Gonthier:2013:HMA

[GZND13] Georges Gonthier, Beta Ziliani, Aleksandar Nanevski, and Derek Dreyer. How to make ad hoc proof automation less ad hoc. Journal of Functional Programming, 23(4):357–401, July 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
30171361444D577470B3F1E306501E2A.

Hackett:2020:RTT

[Hac20] J. Hackett. Review of “Thinking with Types” by Sandy Maguire, LeanPub, 2019. Journal of Functional Programming, 30(??): e4, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
review-of-thinking-with-types-by-sandy-maguire-leanpub-
2019/151CBE6431DDB2C828F77CECDDAA828B.

Hage:2011:BRL

[Hag11] Jurriaan Hage. Book review: Language Implementation Patterns: Create your own Domain-Specific and General Programming Languages, by Parr Terence, Pragmatic Bookshelf, http://www.pragprog.com, ISBN 978-1-934356-45-6. Journal of Functional Programming, 21(2):215–217, March 2011. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/D18C6F8334E9E97AC8DF593BF45A90EB.

Hage:2012:BRB

[Hag12] Jurriaan Hage. Book review: Bookreview JFP: Domain-Specific Languages, by Martin Fowler The Addison Wesley Signature Series. Journal of Functional Programming, 22(3):375–377, May 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4B239180E572D00396517AA0075C3039.

Hains:2003:BRI

[Hai03] Gaétan Hains. Book review: Implicit parallel programming in pH, by R. S. Nikhil and Arvind, Morgan Kaufmann, 2001. Journal of Functional Programming, 13(4):824–826, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/2915113E74DEF3C74DF1CFD62FD6A0C6.

Hannan:1993:ENS

[Han93] John Hannan. Extended natural semantics. Journal of Functional Programming, 3(2):123–152, April 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/92F625091815AFC17724C201944C9158.

Hannan:1998:TBE

[Han98] John Hannan. A type-based escape analysis for functional languages. Journal of Functional Programming, 8(3):239–273, May 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
04423A17BA9A127ADE9BDD149ABF79B2.

Hanna:1999:ITP

[Han99] Keith Hanna. Implementing theorem provers in a purely functional style. Journal of Functional Programming, 9(2):147–166, March 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
5F6C9B0FB71F479BF44B513495DB820D.

Hanus:2001:BRP

[Han01] Michael Hanus. Book review: Programming with Constraints: An Introduction, by Kim Marriott and Peter J. Stuckey, MIT Press, 1998. Journal of Functional Programming, 11(2):253–262, March 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
DEEA0CCD23684416A44744A8908C95C3.

Hancock:2004:BRF

[Han04] Peter Hancock. Book review: The Fun of Programming, edited by Jeremy Gibbons and Oege de Moor, Palgrave Macmillan, 2003, ISBN 1-4039-0772-2 (HB), 0-333-99285-7 (SB). Journal of Functional Programming, 14(5):594–597, September 2004. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/25C76BB328E8180BA2087EC3BA0F529F.

Harrison:1993:ADT

[Har93a] Rachel Harrison. Abstract data types in standard ML. Wiley professional computing. John Wiley, New York, NY, USA, 1993. ISBN 0-471-93844-0 (paperback). x + 212 pp. LCCN QA76.9.A23 H37 1993. US$44.75 (U.S.). URL http://www.loc.

gov/catdir/description/wiley031/92039975.html;   http://
www.loc.gov/catdir/toc/onix01/92039975.html.

Harrison:1993:UFL

[Har93b] Rachel Harrison. The use of functional languages in teaching computer science. Journal of Functional Programming, 3(1):67–75, January 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
80589BD5EAF49F31018129364BD5C054.

Harper:1999:PDD

[Har99a] Robert Harper. Proof-directed debugging. Journal of Functional Programming, 9(4):463–469, July 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/C33E27D7A70190D0A265935CAF9938A8.

See corrigendum [Har09a].

Hartel:1999:BRF

[Har99b] Pieter Hartel. Book review: The Functional Approach to Programming, by Guy Cousineau and Michel Mauny, Cambridge University Press, 1998, ISBN 0-521-57681-4 pbk, xiv+445 pp. Journal of Functional Programming, 9(4):479–482, July 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1F0BE3028E4FAF42EF9BEF79A5755869.

Harper:2009:FPP

[Har09a] Robert Harper. Functional pearl: Proof-directed debugging — corrigendum. Journal of Functional Programming, 19(2):262, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3962675105136FE490AF6481583CCE3B. See [Har99a].

Harrison:2009:HPL

[Har09b] J. (John) Harrison. Handbook of practical logic and automated reasoning. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-89957-5 (hardcover). xix + 681 pp. LCCN QA76.9.L63 H375

Harper:2013:PFP

[Har13] Robert Harper. Practical foundations for programming languages. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-02957-0 (hardcover). xviii + 471 pp. LCCN QA76.7 .H377 2013.

Hasegawa:2000:GTL

[Has00] Masahito Hasegawa. Girard translation and logical predicates. Journal of Functional Programming, 10(1):77–89, January 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
33BFABF75573B9D81D9BBE51E0D6C524.

Hayden:2000:DCM

[Hay00] Mark Hayden. Distributed communication in ML. Journal of Functional Programming, 10(1):91–120, January 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A6222A3C76C0B6C086AB7404A1929E1D.

Hutton:2017:CYJ

[HB17] Graham Hutton and Patrick Bahr. Compiling a 50-year journey. Journal of Functional Programming, 27:e20, ???? 2017. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/compiling-a-50year-journey/31A66593E940CEF13253C34714F78041.

Hatcliff:1997:TC

[HD97] John Hatcliff and Olivier Danvy. Thunks and the λ-calculus. Journal of Functional Programming, 7(3):303–319, May 1997. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/1EEAC7B043CC6D6EF06AE23A8022D7CF.

Hoogendijk:2000:CTC

[HD00] Paul Hoogendijk and Oege De Moor. Container types categorically. Journal of Functional Programming, 10(2):191–225, March 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F65973701316677A02E2F03981043AB8.

Harper:1993:TFC

[HDM93] Robert Harper, Bruce F. Duba, and David Macqueen. Typing firstclass continuations in ML. Journal of Functional Programming, 3(4): 465–484, October 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/8E466C17B3C1EB69FF426EDE1BE750FC.

Hedberg:1998:CTM

[Hed98] Michael Hedberg. A coherence theorem for Martin-Löf’s type theory. Journal of Functional Programming, 8(4):413–436, July 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
37DC581B6D915FE83B74AC5984899319.

Henglein:2012:GTD

[Hen12] Fritz Henglein. Generic top-down discrimination for sorting and partitioning in linear time. Journal of Functional Programming, 22 (3):300–374, May 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/B85E48EFC0B4D2BDDDE9A3885094FDD7.

Hartel:1996:BIF

[HFA^{+}96] Pieter H. Hartel, Marc Feeley, Martin Alt, Lennart Augustsson, Peter Baumann, Marcel Beemster, Emmanuel Chailloux, Christine H. Flood, Wolfgang Grieskamp, John H. G. Van Groningen, Kevin Hammond, Bogumil Hausman, Melody Y. Ivory, Richard E. Jones, Jasper Kamperman, Peter Lee, Xavier Leroy, Rafael D. Lins, Sandra Loosemore, Niklas Röjemo, Manuel Serrano, Jean-Pierre Talpin, Jon Thackray, Stephen Thomas, Pum Walters, Pierre Weis, and Peter Wentworth. Benchmarking implementations of functional languages with ‘pseudoknot’, a float-intensive benchmark. Journal of Functional Programming, 6(4):621–655, July 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
39682AA1BC05B5F1C511BB939839F158.

Hartel:1996:RCS

[HG96] Pieter H. Hartel and Hugh Glaser. The resource constrained shortest path problem implemented in a lazy functional language. Journal of Functional Programming, 6(1):29–46, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
A087385C294D6E2F79D82FA1FF49B64F.

Hunt:1991:FPF

[HH91] Sebastian Hunt and Chris Hankin. Fixed points and frontiers: a new perspective. Journal of Functional Programming, 1(1):91–120, January 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
CCB4CE0A8FFA15B5BA4A0EFACB08CE54.

Hickey:2004:CLS

[Hic04] Timothy J. Hickey. Computer literacy via Scheme and web programming. Journal of Functional Programming, 14(5):475–488, September 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
30482CE8C54B73BCF98DA96CC9AF006C.

Hill:1996:CPE

[Hil96] Steve Hill. Combinators for parsing expressions. Journal of Functional Programming, 6(3):445–464, May 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/59AA145C7980FA88F4FED5D0B474860A.

Hinze:1999:BRP

[Hin99a] Ralf Hinze. Book review: Purely Functional Data Structures, by Chris Okasaki, Cambridge University Press, 1998, 220 pp. Journal of Functional Programming, 9(5):577–578, September 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F95E7BBCF0DFE43815A52B26E6D07914.

Hinze:1999:EBH

[Hin99b] Ralf Hinze. Explaining binomial heaps. Journal of Functional Programming, 9(1):93–104, January 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D53DC543F1EF73B158D1006AD25842E4.

Hinze:2000:GGT

[Hin00a] Ralf Hinze. Generalizing generalized tries. Journal of Functional Programming, 10(4):327–351, July 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/03C839ABDC2CE3326B73CDDD35DD568E.

Hinze:2000:PTB

[Hin00b] Ralf Hinze. Perfect trees and bit-reversal permutations. Journal of Functional Programming, 10(3):305–317, May 2000. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/AD88A8EA0B024153C58AC1C30C6E278E.

Hinze:2001:MD

[Hin01] Ralf Hinze. Manufacturing datatypes. Journal of Functional Programming, 11(5):493–524, September 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/5F23278226F73D01110E6E5A130D36A3.

Hinze:2002:FLB

[Hin02] Ralf Hinze. A fresh look at binary search trees. Journal of Functional Programming, 12(6):601–607, November 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/95DB17CADC4CBA335334C2122D564299.

Hinze:2003:FCA

[Hin03] Ralf Hinze. Formatting: a class act. Journal of Functional Programming, 13(5):935–944, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/74ABD300574892B9763EEB7735D03357.

Hinze:2004:SIF

[Hin04] Ralf Hinze. Special issue on functional pearls: Editorial. Journal of Functional Programming, 14(6):601, November 2004. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/5C49635D13554D11604B780CEE59F38A.

Hinze:2005:TPC

[Hin05] Ralf Hinze. Theoretical pearl: Church numerals, twice! Journal of Functional Programming, 15(1):1–13, January 2005. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/9C64E44B1410C56A4361F7E7F33E732B.

Hinze:2006:GM

[Hin06] Ralf Hinze. Generics for the masses. Journal of Functional Programming, 16(4–5):451–483, July 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A0CECF4B43C6B6EC8EF431D61230F0BB.

Hinze:2009:BT

[Hin09a] Ralf Hinze. The bird tree. Journal of Functional Programming, 19 (5):491–508, September 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/E32A68A3F9D1129BAC6E8A835BAF9C46.

Hinze:2009:PFB

[Hin09b] Ralf Hinze. Purely functional 1–2 brother trees. Journal of Functional Programming, 19(6):633–644, November 2009. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/874BB3A6CB664718275D40FA9ECBDFB4.

Hinze:2010:CSC

[Hin10a] Ralf Hinze. Concrete stream calculus: An extended study. Journal of Functional Programming, 20(5–6):463–535, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F5DADE7E7819BB96D9091F3404AFB323.

Hinze:2010:SIG

[Hin10b] Ralf Hinze. Special issue on generic programming editorial. Journal of Functional Programming, 20(3–4):211–212, July 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7AFE2E70F95A1714D2E3811253691420.

Hinze:2011:TQA

[Hin11] Ralf Hinze. Typed quote/antiquote or: Compile-time parsing. Journal of Functional Programming, 21(3):219–234, May 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EAF6284F3BAC130CA4915E0D91FA100C.

Hinze:2018:CT

[Hin18] Ralf Hinze. On constructing 2–3 trees. Journal of Functional Programming, 28:e19, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
on-constructing-23-trees/129759AE03543B461536F78B386D93B5.

Hinze:2023:CTS

[Hin23] Ralf Hinze. Certified, total serialisers with an application to Huffman encoding. Journal of Functional Programming, 33: ??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
certified-total-serialisers-with-an-application-to-huffman-
encoding/A8A4B67E800CEECFC434C8F69B96F35F.

Hinze:2001:WW

[HJ01] Ralf Hinze and Johan Jeuring. Weaving a web. Journal of Functional Programming, 11(6):681–689, November 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/CD429AE301A49C755941BD867E133766.

Hutton:2010:FFF

[HJG10] Graham Hutton, Mauro Jaskelioff, and Andy Gill. Factorising folds for faster functions. Journal of Functional Programming, 20(3–4): 353–373, July 2010. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/1AC94646C45744F2474475CE5B3C7F90.

Hu:2023:NEM

[HJP23] Jason Z. S. Hu, Junyoung Jang, and Brigitte Pientka. Normalization by evaluation for modal dependent type theory. Journal of Functional Programming, 33:??, ???? 2023. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/normalization-by-evaluation-for-modal-
dependent-type-theory/F2803F6894AE7E39D51B2CFC23CA6874.

Harrison:2005:LDH

[HK05] William L. Harrison and Richard B. Kieburtz. The logic of demand in Haskell. Journal of Functional Programming, 15(6):837– 891, November 2005. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/F4BF326F9A085D338B0E4F70F008B742.

Harper:1996:OIE

[HL96] Robert Harper and Mark Lillibridge. Operational interpretations of an extension of f ω with control operators. Journal of Functional Programming, 6(3):393–418, May 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/B6AD5301F8BAD6D4DB93D8CAB9036801.

Herrmann:1999:PDC

[HL99] Christoph A. Herrmann and Christian Lengauer. Parallelization of divide-and-conquer by translation to nested loops. Journal of Functional Programming, 9(3):279–310, May 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/1FCD1927A7690BC9E370C847BF3E804C.

Harper:2007:MML

[HL07] Robert Harper and Daniel R. Licata. Mechanizing metatheory in a logical framework. Journal of Functional Programming, 17(4–5): 613–673, July 2007. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/E4EC2F77CEF3DC56EABCF6F6658B179E.

Hillerstrom:2020:EHG

[HLA20] Daniel Hillerström, Sam Lindley, and Robert Atkey. Effect handlers via generalised continuations. Journal of Functional Programming, 30(??):e5, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
effect-handlers-via-generalised-continuations/DF590482FEE2F6888CD68B4B446E31D5.

Hammes:1995:CIH

[HLB95] Jeffrey Hammes, Olaf Lubeck, and Wim Böhm. Comparing Id and Haskell in a Monte Carlo photon transport code. Journal of Functional Programming, 5(3):283–316, July 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/8429771C4EAC0792C18B8BE4A5F9723E.

Hillerstrom:2024:ASE

[HLL24] Daniel Hillerström, Sam Lindley, and John Longley. Asymptotic speedup via effect handlers. Journal of Functional Programming, 34:??, ???? 2024. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
asymptotic-speedup-via-effect-handlers/296879DE2FD96FB6CF388F27978C76E4.

Hritcu:2016:TNQ

[HLSZ^{+}16] Catalin Hritcu, Leonidas Lampropoulos, Antal Spector-Zabusky, Arthur Azevedo De Amorim, Maxime Dénès, John Hughes, Benjamin C. Pierce, and Dimitrios Vytiniotis. Testing noninterference, quickly. Journal of Functional Programming, 26:e4, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/testing-noninterference-
quickly/00D792464D8E0CAD7F6B30417F121729.

Henglein:1994:CTI

[HM94] Fritz Henglein and Harry G. Mairson. The complexity of type inference for higher-order typed lambda calculi. Journal of Functional Programming, 4(4):435–477, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/9CFBB776DD580AFE6BB33F0B1BBA5AA4.

Hutton:1996:FPB

[HM96] Graham Hutton and Erik Meijer. Functional pearl: Back to basics: Deriving representation changers functionally. Journal of Functional Programming, 6(1):181–188, January 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/59D4690A72161FB983EEBBC8722A22B4.

Hutton:1998:MPH

[HM98] Graham Hutton and Erik Meijer. Monadic parsing in Haskell. Journal of Functional Programming, 8(4):437–444, July 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/E557DFCCE00E0D4B6ED02F3FB0466093.

Hudak:2004:E

[HM04] Paul Hudak and Greg Morrisett. Editorial. Journal of Functional Programming, 14(4):i–ii, July 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/1D716A1AD146A9FCC1F8C92CE66460C4.

Hinze:2018:BOE

[HM18a] Ralf Hinze and Clare Martin. Batcher’s odd–even merging network revealed. Journal of Functional Programming, 28:e14, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/batchers-oddeven-merging-
network-revealed/14DB825CE81B605C919746923FFF8BA2.

Hinze:2018:PPS

[HM18b] Ralf Hinze and Clare Martin. Parberry’s pairwise sorting network revealed. Journal of Functional Programming, 28:e21, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
parberrys-pairwise-sorting-network-revealed/C9761D61FD85D2E1A053BF48A6CB1399.

Hartel:2004:FCE

[HMG04] Pieter Hartel, Henk Muller, and Hugh Glaser. The functional “C” experience. Journal of Functional Programming, 14(2):129–135, March 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B6FECA5C02015052FF9ADF07022E20F7.

Hudak:1996:HMN

[HMGW96] Paul Hudak, Tom Makucevich, Syam Gadde, and Bo Whong. Haskore music notation — an algebra of music. Journal of Functional Programming, 6(3):465–484, May 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/89A16D59034EF5C649F7846635989E2B.

Haller:2018:PMF

[HMM18] Philipp Haller, Heather Miller, and Normen Müller. A programming model and foundation for lineage-based distributed computation. Journal of Functional Programming, 28:e7, ???? 2018. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/programming-model-and-foundation-for-
lineagebased-distributed-computation/B410CE79B21E33462843B408B716E1E5.

Hardin:1998:FRS

[HMP98] Thérèse Hardin, Luc Maranget, and Bruno Pagano. Functional runtime systems within the lambda-sigma calculus. Journal of Functional Programming, 8(2):131–176, March 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D52C011DF690ACA5C30A3E32FA524AD5.

Hutchison:1997:NEC

[HNSSH97] Nigel W. O. Hutchison, Ute Neuhaus, Manfred Schmidt-Schauss, and Cordy V. Hall. Natural Expert: a commercial functional programming environment. Journal of Functional Programming, 7(2): 163–182, March 1997. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/AAD5A16CB8E3C148BE5D7CB5F36C1D93.

Hofmann:1999:SLM

[Hof99] Martin Hofmann. Semantics of linear/modal lambda calculus. Journal of Functional Programming, 9(3):247–277, May 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C093CB42F73825CD9E1F2472DEE272AE.

Hoogerwoord:1992:FPS

[Hoo92] Rob R. Hoogerwoord. Functional pearls: A symmetric set of efficient list operations. Journal of Functional Programming, 2(4):

505–513, October 1992. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/A77930CC2C68694EBB93964AF16D06D0.

Hosaya:2006:REF

[Hos06] Haruo Hosaya. Regular expression filters for XML. Journal of Functional Programming, 16(6):711–750, November 2006. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/F27B4BABC2160A54AF14D9B76D6CC999.

Hartel:1993:SIS

[HP93] Pieter Hartel and Rinus Plasmeijer. Special issue on stateof-the-art applications of pure functional programming languages. Journal of Functional Programming, 3(4):565–566, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3EA8E94B90C550A603FA3C3DC5907D7D.

Hartel:1995:SIS

[HP95a] Pieter Hartel and Rinus Plasmeijer. Special issue on stateof-the-art applications of pure functional programming languages. Journal of Functional Programming, 5(3):279–281, July 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1AB4C77DC65438D7D3EC97EDA1885A8D.

Hofmann:1995:UTT

[HP95b] Martin Hofmann and Benjamin Pierce. A unifying type-theoretic framework for objects. Journal of Functional Programming, 5(4): 593–635, October 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/803C5BEB16FCF29A7420F660405AB653.

Hanus:1999:HON

[HP99] Michael Hanus and Christian Prehofer. Higher-order narrowing with definitional trees. Journal of Functional Programming, 9(1): 33–75, January 1999. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/964ED12C8B25E04C1AAAA9E86C1CF532.

Hosoya:2003:REP

[HP03] Haruo Hosoya and Benjamin C. Pierce. Regular expression pattern matching for XML. Journal of Functional Programming, 13(6):961– 1004, November 2003. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/C845B41E6B150FBD7731EE396FBCB911.

Hinze:2006:FTS

[HP06] Ralf Hinze and Ross Paterson. Finger trees: a simple generalpurpose data structure. Journal of Functional Programming, 16(2): 197–217, March 2006. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/BF419BCA07292DCAAF2A946E6BDF573B.

Honsell:1998:SOS

[HPR98] Furio Honsell, Alberto Pravato, and Simona Ronchi Della Rocca. Structured Operational Semantics of a fragment of the language Scheme. Journal of Functional Programming, 8(4):335–365, July 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
90C43B581A42F4F94C45B7E7A09768FC.

Hansen:1999:IPU

[HR99] Michael R. Hansen and Hans Rischel. Introduction to programming using SML. International computer science series. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-39820-6. xvi + 355 pp. LCCN QA76.73.M6 H36 1999.

Huth:2004:LCS

[HR04] Michael Huth and Mark Ryan. Logic in computer science: modelling and reasoning about systems. Cambridge University Press, Cambridge, UK, second edition, 2004. ISBN 0-521-54310-X (paperback). xiv + 427 pp. LCCN QA76.9.L63 H88 2004. URL http://www.loc.

gov/catdir/description/cam041/2004045921.html;   http://
www.loc.gov/catdir/samples/cam041/2004045921.html;   http:
//www.loc.gov/catdir/toc/cam041/2004045921.html.

Hinze:2022:SN

[HR22] Ralf Hinze and Colin Runciman. Super-naturals. Journal of Functional Programming, 32(??):e5, ???? 2022. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/supernaturals/9D3E021A2B1717B186ABD500CD34AD75.

Hoffmann:2015:TBA

[HS15] Jan Hoffmann and Zhong Shao. Type-based amortized resource analysis with integers and arrays. Journal of Functional Programming, 25:e17, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
typebased-amortized-resource-analysis-with-integers-and-
arrays/B226DB9379DAF00F7BD7992283CEA906.

Hammes:1997:EFL

[HSB97] J. Hammes, S. Sur, and W. Böhm. On the effectiveness of functional language features: NAS benchmark FT. Journal of Functional Programming, 7(1):103–123, January 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/72D1F22F1E0E488EA6214D12A0104FDE.

Hudak:2000:HSE

[Hud00] Paul Hudak, editor. The Haskell school of expression: learning functional programming through multimedia. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-64338-4 (hardcover), 0-521-64408-9 (paperback). LCCN QA76.62 .H83 2000. URL http:/

/www.loc.gov/catdir/description/cam0210/99045529.html;
http://www.loc.gov/catdir/toc/cam023/99045529.html.

Hudson:2015:RLY

[Hud15] Shane Hudson. Review of “Learn you some Erlang for great good! A beginner’s guide”, by Fred Hébert, No Starch Press, 2013, £26.80 (paperback), ISBN: 978-1-59327-435-1. Journal of Functional Programming, 25:e16, ???? 2015. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/review-of-learn-you-some-
erlang-for-great-good-a-beginners-guide-by-fred-hebert-
no-starch-press-2013-2680-paperback-isbn-9781593274351/
63A0A56C14B429CA2F748D163E7FF07D.

Huet:1994:RTC

[Hue94] Gérard Huet. Residual theory in λ-calculus: a formal development. Journal of Functional Programming, 4(3):371–394, July 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
10C9E95ABFCEEFD4F1CBAF2C800647AA.

Huet:1997:Z

[Hue97] Gérard Huet. The zipper. Journal of Functional Programming, 7 (5):549–554, September 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/0C058890B8A9B588F26E6D68CF0CE204.

Huet:2003:SIL

[Hue03] Gérard Huet. Special issue on ‘logical frameworks and metalanguages’. Journal of Functional Programming, 13(2):257–260, March 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
62D0801AECDE360EAC6DF16078B69B4E.

Huet:2005:FTM

[Hue05] Gérard Huet. A functional toolkit for morphological and phonological processing, application to a Sanskrit tagger. Journal of Functional Programming, 15(4):573–614, July 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/52F91F2AB03C5A222F1E85AFD0899629.

Hughes:2004:GVH

[Hug04] John Hughes. Global variables in Haskell. Journal of Functional Programming, 14(5):489–502, September 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/37A6F7551C3A84120D658CE2D2C55E6E.

Hutton:1992:HOF

[Hut92] Graham Hutton. Higher-order functions for parsing. Journal of Functional Programming, 2(3):323–343, July 1992. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/0490F2C8511F7625F9FC15BFFEDBB0AA.

Hutton:1994:BRI

[Hut94] Graham Hutton. Book review: Introduction to HOL: a theorem proving environment for higher order logic, by Gordon Mike and Melham Tom (eds.), Cambridge University Press, 1993, ISBN 0-521-44189-7. Journal of Functional Programming, 4(4):557–559, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
682CAD7058D7014549AE3F9580D0220B.

Hutton:1999:TUE

[Hut99] Graham Hutton. A tutorial on the universality and expressiveness of fold. Journal of Functional Programming, 9(4):355–372, July 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
CDBAA53C7120E23CBBBE206FD47FDBAA.

Hutton:2000:SIH

[Hut00] Graham Hutton. Special issue on Haskell. Journal of Functional Programming, 10(3):319, May 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/450EE1A5F92FD090872D016121C4889B.

Hutton:2002:CP

[Hut02a] Graham Hutton. The countdown problem. Journal of Functional Programming, 12(6):609–616, November 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/FCF81E290B194D1ACBF699842D4D4CA5.

Hutton:2002:SDI

[Hut02b] Graham Hutton. Special double issue on Haskell. Journal of Functional Programming, 12(4–5):293–294, July 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A139AA548F13B28FCFA9D8D38CBDEC66.

Hutton:2007:PH

[Hut07] Graham Hutton. Programming in Haskell. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87172-7 (hardcover), 0-521-69269-5 (paperback). xii + 171 pp. LCCN QA76.62 .H88

2007. URL http://www.loc.gov/catdir/enhancements/fy0729/
2007274987-b.html;   http://www.loc.gov/catdir/enhancements/
fy0729/2007274987-d.html;   http://www.loc.gov/catdir/enhancements/
fy0729/2007274987-t.html.

Hutton:2014:PA

[Hut14] Graham Hutton. PhD abstracts. Journal of Functional Programming, 24(4):474–522, July 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/746B45DF269A7746FF42CCC1DE5A2748.

Hutton:2015:PAa

[Hut15a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 25:e1, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/A571A5D4D001F8FE31C4672AC184671F.

Hutton:2015:PAb

[Hut15b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 25:e7, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/E150B65762A51870356263EA0F052152.

Hutton:2015:PAc

[Hut15c] Graham Hutton. PhD abstracts. Journal of Functional Programming, 25:e21, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/26EDA64896CD456D3BA13F03298E248A.

Hutton:2016:PAa

[Hut16a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 26:e8, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/A39BE6465AD4FD4744B8689F01BEF1C9.

Hutton:2016:PAb

[Hut16b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 26:e24, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/BAF394984CB4971A0EA778D176382CC6.

Hutton:2017:PAa

[Hut17a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 27:e16, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/880BC4CBC71DA9AFF193B616E1C637D8.

Hutton:2017:PAb

[Hut17b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 27:e25, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/BA901528CBC710EB1FAC68EAFDA70836.

Hutton:2018:PA

[Hut18] Graham Hutton. PhD abstracts. Journal of Functional Programming, 28:e17, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/755F5652224B435080756DD21623839C.

Hutton:2019:PAa

[Hut19a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 29:e1, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/06E7785894BD9EAC9903940267B9A4DF.

Hutton:2019:PAb

[Hut19b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 29:e10, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/A0B4D8FFC69D131EA44B6B1A71A66A1B.

Hutton:2020:PAa

[Hut20a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 30(??):e1, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/F66CECFC31BDB7D2790F39DE25426EEA.

Hutton:2020:PAb

[Hut20b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 30(??):e18, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/3A33B8690F953DDD3AFD44F25B69EF00.

Hutton:2021:PAa

[Hut21a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 31(??):e1, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/86D39E70D5D89E4C3E6D38434D10ADAC.

Hutton:2021:PAb

[Hut21b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 31(??):e14, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/7BA112E2114AF9F97A9C1B8D74A5EE7A.

Hutton:2022:PAa

[Hut22a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 32(??):e2, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/5540C7C053B5E5FD73713D82B7A60915.

Hutton:2022:PAb

[Hut22b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 32:e7, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/FA3E743E00B155C2A5AB4946EF4C6A4F.

Hutton:2022:PAc

[Hut22c] Graham Hutton. PhD abstracts. Journal of Functional Programming, 32:e17, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/D0E41859562B6733AED9979BB67017C1.

Hutton:2023:PA

[Hut23a] Graham Hutton. PhD abstracts. Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/A32EE036508B4F3B309230D342DA7C2C.

Hutton:2023:PAb

[Hut23b] Graham Hutton. PhD abstracts. Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
phd-abstracts/BDC0F1227031BFB67C816ED3FCC0CD96.

Hutton:2023:PLS

[Hut23c] Graham Hutton. Programming language semantics: It’s easy as 1, 2, 3. Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/programming-language-semantics-
its-easy-as-123/EC2C046CF94382B3B408036B84475DC7.

Heckmann:1997:FDT

[HW97] Reinhold Heckmann and Reinhard Wilhelm. A functional description of TEX’s formula layout. Journal of Functional Programming, 7(5):451–485, September 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/18A13702FF264BBA3894B55E01832976.

Hutton:2007:WMC

[HW07] Graham Hutton and Joel Wright. What is the meaning of these constant interruptions? Journal of Functional Programming, 17 (6):777–792, November 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/F8D5045B9F5DAAD8460ED1F8A57E7E74.

Hinze:2016:USR

[HW16] Ralf Hinze and Nicolas Wu. Unifying structured recursion schemes: An extended study. Journal of Functional Programming, 26:e1, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
unifying-structured-recursion-schemes/2CF2B29B85890E893F9761EA1C3B709E.

Honda:2005:NTF

[HY05] Kohei Honda and Nobuko Yoshida. Noninterference through flow analysis. Journal of Functional Programming, 15(2):293–349, March 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
316515BBDF23C6F0948CB529D7654E68.

Hirai:2011:BWB

[HY11] Yoichi Hirai and Kazuhiko Yamamoto. Balancing weight-balanced trees. Journal of Functional Programming, 21(3):287–307, May 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7281C4DE7E56B74F2D13F06E31DCBC5B.

Huang:2021:TMO

[HZO21] Xuejing Huang, Jinxu Zhao, and Bruno C. D. S. Oliveira. Taming the Merge Operator. Journal of Functional Programming, 31(??):e28, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
taming-the-merge-operator/B8A3A5D16B7AB175D44DB6F980450863.

Ireland:1999:AVF

[IB99] Andrew Ireland and Alan Bundy. Automatic verification of functions with accumulating parameters. Journal of Functional Programming, 9(2):225–245, March 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/38920FB596F373F0EAF3D19243FA77F2.

Iii:2022:MGT

[ICEW22] John Peter Campora Iii, Sheng Chen, Martin Erwig, and Eric Walkingshaw. Migrating gradual types. Journal of Functional Programming, 32:e14, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
migrating-gradual-types/BD262271D2A9694DCF5212735E50CC33.

Iwasaki:2022:FFD

[IEM^{+}22] Hideya Iwasaki, Kento Emoto, Akimasa Morihata, Kiminori Matsuzaki, and Zhenjiang Hu. Fregel: a functional domain-specific language for vertex-centric large-scale graph processing. Journal of Functional Programming, 32(??):e4, ???? 2022. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/fregel-a-functional-domainspecific-
language-for-vertexcentric-largescale-graph-processing/
A6D4FC0B42C87B752E9AF88F681C5F13.

Ida:1997:LON

[IN97] Tetsuo Ida and Koichi Nakahara. Leftmost outside-in narrowing calculi. Journal of Functional Programming, 7(2):129–161, March 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
87DB5DF908052E51DE36792332422603.

Irwin:2001:CBF

[IRK01] Robert J. Irwin, James S. Royer, and Bruce M. Kapron. On characterizations of the basic feasible functionals, Part I. Journal of Functional Programming, 11(1):117–153, January 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9FECA793B66F66EC01C17ED56823CC38.

Inoue:2016:RAM

[IT16] Jun Inoue and Walid Taha. Reasoning about multi-stage programs. Journal of Functional Programming, 26:e22, ???? 2016. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/reasoning-about-multistage-programs/
60E3E040633DD97C4B61766123F1D639.

Jacobs:1996:C

[Jac96] Bart Jacobs. On cubism. Journal of Functional Programming, 6 (3):379–392, May 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/9AB78ED619AB80A60D82CED62DE430D3.

Jackson:2006:SAL

[Jac06] Daniel Jackson. Software abstractions: logic, language and analysis. MIT Press, Cambridge, MA, USA, 2006. ISBN 0-262-10114-9, 0-262-10115-9. xvi + 350 pp. LCCN QA76.76.D47 J29 2006.

Jadud:2014:BRR

[Jad14] Matt Jadud. Book review: Realm of Racket, by Forrest Bice, Rose DeMaio, Spencer Florence, Feng-Yun Mimi Lin, Scott Lindeman, Nicole Nussbaum, Eric Peterson, Ryan Plessner, David Van Horn, Matthias Felleisen and Conrad Barski, MD, No Starch Press, San Francisco, CA, 2013, £27.49. ISBN-10: 1-59327-491-2. Journal of Functional Programming, 24(2–3):419–421, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
AFBE1E4602FCE8B93CEE8DB789163CDC.

Jackson:1993:II

[JB93] W. Ken Jackson and F. Warren Burton. Improving intervals. Journal of Functional Programming, 3(2):153–169, April 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6945A96DCF281D77C6E8E9690F551DC8.

Jay:1998:FM

[JBM98] C. B. Jay, G. Bellè, and E. Moggi. Functorial ML. Journal of Functional Programming, 8(6):573–619, November 1998. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/DE496E31AB5D5BC0D62DCA207E98E996.

Jeffrey:1998:BRM

[Jef98] Alan Jeffrey. Book review: ML with Concurrency, by Flemming Nielson (ed.), Springer-Verlag, 1997. Journal of Functional Programming, 8(5):537–542, September 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/B7C9632B5AC533BBC31669A4D31F52AA.

Jeffrey:2001:BRC

[Jef01] Alan Jeffrey. Book review: Communicating and Mobile Systems: the π-calculus, by Robin Milner, Cambridge University Press, 1999. Journal of Functional Programming, 11(4):433–436, July 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BD49BCAF28ADE3CA15C9083B19B5AD9B.

Johnson-Freyd:2017:CNE

[JFDA17] Philip Johnson-Freyd, Paul Downen, and Zena M. Ariola. Callby-name extensionality and confluence. Journal of Functional Programming, 27:e12, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
callbyname-extensionality-and-confluence/65772ED558B918E54B32A70C012AE78D.

Jay:1995:VEE

[JG95] C. Barry Jay and Neil Ghani. The virtues of eta-expansion. Journal of Functional Programming, 5(2):135–154, April 1995. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/FAAA27786A519A85C8DFB650AC8864AD.

Jacobs:2009:CSA

[JHH09] Bart Jacobs, Chris Heunen, and Ichiro Hasuo. Categorical semantics for arrows. Journal of Functional Programming, 19(3–4): 403–438, July 2009. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/35C603B99F43CBD8109C97FA51BDAB23.

Jansson:1998:PU

[JJ98] Patrik Jansson and Johan Jeuring. Polytypic unification. Journal of Functional Programming, 8(5):527–536, September 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/FD51DC3B623D490DE78A479648E05D25.

Jansson:2023:LCB

[JJ23] Julia Jansson and Patrik Jansson. Level-p-complexity of Boolean functions using thinning, memoization, and polynomials. Journal of Functional Programming, 33:??, ???? 2023. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/levelpcomplexity-of-boolean-functions-
using-thinning-memoization-and-polynomials/58122B71C40F99E0D19ACD0FAFF867A9.

Jay:2009:FCP

[JK09] Barry Jay and Delia Kesner. First-class patterns. Journal of Functional Programming, 19(2):191–225, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/968C982CA9B727A2C04D216EEF4E6CFC.

Jung:2018:IGM

[JKJ^{+}18] Ralf Jung, Robbert Krebbers, Jacques-Henri Jourdan, Ales Bizjak, Lars Birkedal, and Derek Dreyer. Iris from the ground up: A modular foundation for higher-order concurrent separation logic. Journal of Functional Programming, 28:e20, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-

7653 (electronic).  URL https://www.cambridge.org/core/
journals/journal-of-functional-programming/article/iris-
from-the-ground-up-a-modular-foundation-for-higherorder-
concurrent-separation-logic/26301B518CE2C52796BFA12B8BAB5B5F.

Jung:2000:VPT

[JM00] Yang Jung and Greg Michaelson. A visualisation of polymorphic type checking. Journal of Functional Programming, 10(1):57–75, January 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3234142DE64E7B8508F8C803012A4FFA.

Jaskelioff:2015:RTS

[JO15] Mauro Jaskelioff and Russell O’Connor. A representation theorem for second-order functionals. Journal of Functional Programming, 25:e13, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
representation-theorem-for-secondorder-functionals/4B782B0BB6EB53E53935D7A0F8432E8E.

Johnson:1992:RFP

[Joh92] Thomas Johnson. A review of the FPCA ’91 proceedings — Hughes John (Ed); Functional Programming Languages and Computer Architecture. Proceedings of the fifth conference (Cambridge, MA, 28-30081991), Volume 523 of Lecture Notes in Computer Science, Springer-Verlag. 666 pp. Journal of Functional Programming, 2(2): 227–231, April 1992. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/154D49CD0319B167E6D13C19D7A7C0F6.

Johnsson:1998:EGA

[Joh98] Thomas Johnsson. Efficient graph algorithms using lazy monolithic arrays. Journal of Functional Programming, 8(4):323–333, July 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2853872041BCA4DD4C58A672369FBAFE.

Johann:2003:SCF

[Joh03] Patricia Johann. Short cut fusion is correct. Journal of Functional Programming, 13(4):797–814, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/35A6587372FCE50ED52D279054BE5C3E.

Jones:1992:CLA

[Jon92a] Mark P. Jones. Computing with lattices: An application of type classes. Journal of Functional Programming, 2(4):475–503, October 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EA22C34575C9C9B0CB1264B3D9680950.

Jones:1992:BRP

[Jon92b] Neil D. Jones. Book review: Projection Factorisations in Partial Evaluation. By John Launchbury. Distinguished Dissertations in Computer Science, Vol. 1. Cambridge University Press, 1991. ISBN 0-521-41497-0. Journal of Functional Programming, 2(3):365, July 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2615182311586CDC7E215624CDCCD50F.

Jones:1992:TRS

[Jon92c] Richard Jones. Tail recursion without space leaks. Journal of Functional Programming, 2(1):73–79, January 1992. CODEN JFPRES.

ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3F1FE7625B1633B8D8BEC0955F0F4906.

Jones:1993:E

[Jon93] Neil Jones. Erratum. Journal of Functional Programming, 3(4): 389, October 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/582907D1CF55A2005DDA13FA85E9505B.

Jones:1995:SCC

[Jon95] Mark P. Jones. A system of constructor classes: overloading and implicit higher-order polymorphism. Journal of Functional Programming, 5(1):1–35, January 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/D2D05B479E365B913F7F0CD3543CB7EE.

Jones:2001:EPH

[Jon01] Neil D. Jones. The expressive power of higher-order types or, life without CONS. Journal of Functional Programming, 11(1):55–94, January 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
21B766EA77D6646049BCF312279B1B17.

Jones:2004:FPC

[Jon04] Mark P. Jones. Functional pearl: Composing fractals. Journal of Functional Programming, 14(6):715–725, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/EBF5B1EBB717AFEC0CA84FE8A07C5C24.

Jones:2006:BRS

[Jon06] Isaac Jones. Book review: The Standard ML Basis Library, by Emden R. Gansner and John H. Reppy, editors, Cambridge University Press, 2004, 406 pp. Journal of Functional Programming, 16(2): 250–251, March 2006. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/AF4F4E64538F286D9D24DC3D84D622D5.

Jones:2007:BRH

[Jon07] Isaac Jones. Book review: The Haskell School of Expression, by Paul Hudak, Cambridge Univerity Press, 2000 1. Journal of Functional Programming, 17(3):426–428, May 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/AD9FD992E8EF22E5171BE9DF3E2368DC.

Jones:2013:SSC

[Jon13] Mark P. Jones. Solving the snake cube puzzle in Haskell. Journal of Functional Programming, 23(2):145–160, March 2013. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/D0C6BD42DB4BAD20F0866F162D723137.

Johnson:2014:PFA

[JSE^{+}14] J. Ian Johnson, Ilya Sergey, Christopher Earl, Matthew Might, and David Van Horn. Pushdown flow analysis with abstract garbage collection. Journal of Functional Programming, 24(2–3):218–283, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
92567231A16859F51056AB6F6781ACAE.

Joosten:1993:TFP

[JVV93] Stef Joosten, Klaas Van Den Berg, and Gerrit Van Der Hoeven. Teaching functional programming to first-year students. Journal of Functional Programming, 3(1):49–65, January 1993. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/041D6A27806B921685CD562695270216.

Kahrs:1998:BRH

[Kah98] Stefan Kahrs. Book review: Higher Order Operational Techniques in Semantics edited by Andrew D. Gordon and Andrew M. Pitts, Cambridge University Press 1998, ISBN 0-521-63168-8 (hardback). Journal of Functional Programming, 8(6):633–637, November 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
AF45458996477BB29BBD1E8CFD3B3E15.

Kahrs:2001:RBT

[Kah01] Stefan Kahrs. Red-black trees with types. Journal of Functional Programming, 11(4):425–432, July 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/632BDF7BE8CD1F6EAEDEA37E6046E4A1.

Kahrs:2005:BRT

[Kah05] Stefan Kahrs. Book review: Term Rewriting Systems, by “Terese”, Cambridge University Press, 2003. Journal of Functional Programming, 15(4):651, July 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/DA6187CE28D2702FAAF6B520EC65EAA3.

Kamareddine:2005:TCO

[Kam05] Fairouz Kamareddine. Typed λ-calculi with one binder. Journal of Functional Programming, 15(5):771–796, September 2005. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/4B1A0F23966F5E95A5D6658C176AFF99.

Karczmarczuk:2003:BRS

[Kar03] Jerzy Karczmarczuk. Book review: Structure and Interpretation of Classical Mechanics, by Gerald Jay Sussman and Jack Wisdom with Meinhard E. Mayer, The MIT Press; 2001, ISBN 0-262-19455-4. Journal of Functional Programming, 13(4):826–828, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
59F058C2CAAB96A215C472858F4A3E7C.

Keep:2014:RTR

[KD14] Andrew W. Keep and R. Kent Dybvig. A run-time representation of Scheme record types. Journal of Functional Programming, 24 (6):675–716, November 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/B8FE043B9B0D8865EBB8ABBF58865E5C.

Keane:1994:OFS

[Kea94] John A. Keane. An overview of the Flagship system. Journal of Functional Programming, 4(1):19–45, January 1994. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/A1FAEBEDCE3D10D07F328481BDFC97A9.

Keller:2003:BRW

[Kel03] Bryn Keller. Book review: The World of Scripting Languages, by David W. Barron, John Wiley & Sons, 2000, ISBN 0-471-99886-9. Journal of Functional Programming, 13(4):831–832, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
85C1B5AC12B6D233A7F99B34D8D60D80.

Kennedy:1996:FPD

[Ken96] Andrew J. Kennedy. Functional pearls: Drawing trees. Journal of Functional Programming, 6(3):527–534, May 1996. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/535113105064F7354260FF55841D529E.

Kennedy:2004:FPP

[Ken04] Andrew J. Kennedy. Functional pearl: Pickler combinators. Journal of Functional Programming, 14(6):727–739, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
FFE741C63CC891BE2CAF9CD82AAB5B4D.

Kessin:2012:BWA

[Kes12] Zachary Kessin. Building web applications with Erlang. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2012. ISBN 1-4493-2066-X (e-book), 1-4493-2065-1 (ebook), 1-4493-2062-7, 1-4493-0996-8. xvi + 133 pp. LCCN

QA76.73.E75. URL http://cdn.totalboox.com/static/covers/
ORL/798004101cf496f1-b.jpg;   http://excerpts.cdn.overdrive.
com/FormatType-410/2858-1/433/F8B/F8/BuildingWebApplicationswithErlangWork9781449320652.
epub;   http://images.contentreserve.com/ImageType-100/
2858-1/{433F8BF8-1081-406E-AC43-B29E5DCB23B9}Img100.jpg;
 http://proquest.safaribooksonline.com/9781449320621;
http://public.eblib.com/choice/publicfullrecord.aspx?p=
939411;   http://site.ebrary.com/id/10758595;   http://
www.totalboox.com/book/id-8755002142860023537.

Ko:2017:PO

[KG17] Hsiang-Shang Ko and Jeremy Gibbons. Programming with ornaments. Journal of Functional Programming, 27:e2, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/programming-with-ornaments/
97C63D7C90556ACD2EC1482E63026A74.

Kishon:1995:SDP

[KH95] Amir Kishon and Paul Hudak. Semantics directed program execution monitoring. Journal of Functional Programming, 5(4):501– 547, October 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/5B283376131AEA9A1DF6722F140AC3AB.

Keller:2018:ESI

[KH18] Gabriele Keller and Fritz Henglein. Editorial for the special issue on parallel and concurrent functional programming. Journal of Functional Programming, 28:e6, ???? 2018. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-

programming/article/editorial-for-the-special-issue-on-
parallel-and-concurrent-functional-programming/8C1315D41C9264978A08BD6EC1DE91FB.

King:2007:BRS

[Kin07] Peter King. Book review: SMIL 2.0 Interactive Multimedia for Web and Mobile Devices, by Dick C. A. Bulterman and Lloyd Rutledge, Springer X.media.publishing, 2004, 440 pp., ISBN 3-540-20234-X. Journal of Functional Programming, 17(3):423–424, May 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1EF269E7304B3A7A920DE7307E562106.

Kiselyov:2020:MMP

[Kis20] Oleg Kiselyov. Many more predecessors: A representation workout. Journal of Functional Programming, 30(??):e7, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/many-more-predecessors-a-
representation-workout/72C397C5CEC70ACB379658A431447500.

Kameyama:2011:SSS

[KKS11] Yukiyoshi Kameyama, Oleg Kiselyov, and Chung-Chieh Shan. Shifting the stage: Staging with delimited control. Journal of Functional Programming, 21(6):617–662, November 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/2D16264FDEB84B745599C095174675B9.

Klop:2001:BRT

[Klo01] Jan Willem Klop. Book review: Term Rewriting and All That, by Franz Baader and Tobias Nipkow, Cambridge University Press, 1998, ISBN 0-521-45520-0 (hardback), 301 pp. Journal of Functional Programming, 11(2):253–262, March 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3053BC8F59F25E200EEA9D58CAE2F557.

Kluge:2001:BRI

[Klu01] Werner Kluge. Book review: Introduction to Programming Using SML, by M. R. Hansen and H. Rischel, Addison Wesley, 1999, ISBN 0-201-39820-6. Journal of Functional Programming, 11(2): 253–262, March 2001. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0722E35459563C17C96059B8E28E0A69.

Kirchner:2001:PRP

[KM01] Hélène Kirchner and Pierre-Etienne Moreau. Promoting rewriting to a programming language: a compiler for non-deterministic rewrite programs in associative-commutative theories. Journal of Functional Programming, 11(2):207–251, March 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/668DC798B3CB263206CE9FB50CAAAF02.

Kiselyov:2021:EAR

[KMS21] Oleg Kiselyov, Shin-Cheng Mu, and Amr Sabry. Not by equations alone: Reasoning with extensible effects. Journal of Functional Programming, 31(??):e2, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
not-by-equations-alone-reasoning-with-extensible-effects/
8E5B14BA0105D056F0ED48564E79BEBC.

Kamareddine:1995:RRL

[KN95] Fairouz Kamareddine and Rob Nederpelt. Refining reduction in the lambda calculus. Journal of Functional Programming, 5(4):637– 651, October 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/E35C06DB3BCAD433E4BADE2743678C63.

Kamareddine:1996:CTC

[KN96] Fairouz Kamareddine and Rob Nederpelt. Canonical typing and Πconversion in the Barendregt Cube. Journal of Functional Programming, 6(2):245–267, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/534C8EEBBB83AA415CBCCE5C46CFB195.

Katsumata:2008:AFF

[KN08] Shin-Ya Katsumata and Susumu Nishimura. Algebraic fusion of functions with an accumulating parameter and its improvement. Journal of Functional Programming, 18(5):781–819, September 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2942891EAC1C2029C767CAC2E96E3EB2.

Kolar:2012:BRD

[Kol12] Dusan Kolàr. Book review: Drawing Programs: The Theory and Practice of Schematic Functional Programming, by Addis, Tom and Addis, Jan, Springer, 2010, ISBN 978-1-84882-617-5, 379 pp. Journal of Functional Programming, 22(2):219–221, March 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9DF7D4C29D4390EE93C7E136C8CD7F0F.

Kammar:2017:NVR

[KP17] Ohad Kammar and Matija Pretnar. No value restriction is needed for algebraic effects and handlers. Journal of Functional Programming, 27:e7, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/no-value-restriction-is-needed-for-
algebraic-effects-and-handlers/19DD87AEDAEABBB45A06D5EA21F03428.

Karachalias:2020:EES

[KPS^{+}20] Georgios Karachalias, Matija Pretnar, Amr Hany Saleh, Stien Vanderhallen, and Tom Schrijvers. Explicit effect subtyping. Journal of Functional Programming, 30(??):e15, ???? 2020. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/explicit-effect-subtyping/4851B1C994BEAAB7F04A58B60F11D0AF.

Kamareddine:1997:ECE

[KR97] Fairouz Kamareddine and Alejandro Ríos. Extending a λ-calculus with explicit substitution which preserves strong normalisation into a confluent calculus on open terms. Journal of Functional Programming, 7(4):395–420, July 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/E3DEB2EE26074D37AD38F522FDD7652E.

Krishnamurthi:2002:PAD

[KR02] Shriram Krishnamurthi and C. R. Ramakrishnan, editors. Practical aspects of declarative languages: 4th International Symposium, PADL 2002, Portland, OR, USA, January 19–20, 2002: proceedings, volume 2257 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-43092-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7

.P32 2002. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t2257.htm;   http://www.loc.gov/catdir/
enhancements/fy0817/2002283456-d.html;   http://www.springerlink.
com/content/978-3-540-43092-6;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=2257.

Koronkevich:2022:APD

[KRAB22] Paulette Koronkevich, Ramon Rakow, Amal Ahmed, and William J. Bowman. ANF preserves dependent types up to extensional equality. Journal of Functional Programming, 32:e12, ???? 2022. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/anf-preserves-dependent-types-up-to-
extensional-equality/73FC888A23E5E87BAE16B158ABE349C8.

Kreitz:2004:BRH

[Kre04] Christoph Kreitz. Building reliable, high-performance networks with the Nuprl proof development system. Journal of Functional Programming, 14(1):21–68, January 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/C4B3E502A3028EF2157204C59AC0DEAD.

Krishnamurthi:2006:EPA

[Kri06] Shriram Krishnamurthi. Educational pearl: Automata via macros. Journal of Functional Programming, 16(3):253–267, May 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
CE394A89B92EF0F1193426C2D1731AB1.

Krishnamurthi:2021:WEP

[Kri21] Shriram Krishnamurthi. What is an education paper? Journal of Functional Programming, 31(??):e16, ???? 2021. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/what-is-an-education-paper/22F85A4207E40020C96FCFB6E989B089.

Kaser:1997:EFP

[KRRS97] Owen Kaser, C. R. Ramakrishnan, I. V. Ramakrishnan, and R. C. Sekar. Equals — a fast parallel implementation of a lazy language. Journal of Functional Programming, 7(2):183–217, March 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
80E1A2C835A01963E05B125B41A7AC9E.

Kodama:2008:TTP

[KSK08] Koichi Kodama, Kohei Suenaga, and Naoki Kobayashi. Translation of tree-processing programs into stream-processing programs based on ordered linear type. Journal of Functional Programming, 18(3): 333–371, May 2008. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/8BB6F3EC40DF255B72255B7B2D947D34.

Kennedy:2012:EBC

[KV12] Andrew J. Kennedy and Dimitrios Vytiniotis. Every bit counts: The binary representation of typed data and programs. Journal of Functional Programming, 22(4–5):529–573, September 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
A9A5BA8A2385066935ADA3989D26DE5A.

Lapalme:1998:DTA

[Lap98] Guy Lapalme. Dynamic tabbing for automatic indentation with the layout rule. Journal of Functional Programming, 8(5):493– 502, September 1998. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0A13628FF5D9B759B12BFB488D91C128.

Launchbury:1991:PFP

[Lau91] John Launchbury. Projection factorisations in partial evaluation. Distinguished dissertations in computer science. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-41497-0. xii + 163 pp. LCCN QA76.6 .L3486 1991. URL http://www.

loc.gov/catdir/description/cam025/92165640.html;   http:
//www.loc.gov/catdir/toc/cam029/92165640.html.

Laufer:1996:TCE

[Läu96] Konstantin Läufer. Type classes with existential types. Journal of Functional Programming, 6(3):485–518, May 1996. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/E5FA27DB3DFD8CF92F018BC6884B3478.

Lawall:2008:E

[Law08] Julia Lawall. Editorial. Journal of Functional Programming, 18 (5):565–566, September 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/16CC1ECDC150C37F57C35A39EE4A4A22.

Launchbury:1996:RDP

[LB96] John Launchbury and Gebreselassie Baraki. Representing demand by partial projections. Journal of Functional Programming, 6(4): 563–585, July 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/DE6A25A1F8220AB06C0C6A7A8BD3F39B.

Leivant:2001:E

[LC01] Daniel Leivant and Bob Constable. Editorial. Journal of Functional Programming, 11(1):1, January 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/E97F5976941A1F90B1BDBDF15D1931C8.

Liu:2011:CCA

[LCH11] Hai Liu, Eric Cheng, and Paul Hudak. Causal commutative arrows. Journal of Functional Programming, 21(4–5):467–496, September 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
68A0B61A360B4A2BFA38512DD661667C.

Leroy:1996:STT

[Ler96] Xavier Leroy. A syntactic theory of type generativity and sharing. Journal of Functional Programming, 6(5):667–698, September 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
2270FCE3983541C446AE69D231B35C8D.

Leroy:2000:MMS

[Ler00] Xavier Leroy. A modular module system. Journal of Functional Programming, 10(3):269–303, May 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A8D022C76CBFB0DD9EEA05458D5C662D.

Leroy:2009:Ea

[Ler09] Xavier Leroy. Editorial. Journal of Functional Programming, 19 (2):143, March 2009. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/BB5634045A4DB569217FAB218D34D166.

Lewis:2013:IAP

[Lew13] Mark C. Lewis. Introduction to the art of programming using Scala. Chapman and Hall/CRC textbooks in computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4398-9666-6 (paperback). ???? pp. LCCN QA76.73.S28 L49 2013.

Leroy:2009:Eb

[LF09] Xavier Leroy and Matthias Felleisen. Editorial. Journal of Functional Programming, 19(5):489–490, September 2009. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/834A00EA963E854901165AD6B4DED87B.

Luke:2005:LRB

[LG05] Edward A. Luke and Thomas George. Loci: a rule-based framework for parallel multi-disciplinary simulation synthesis. Journal of Functional Programming, 15(3):477–502, May 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/DA6C7EA3C26CA48B5E8248B1318F9B16.

Lins:2000:BRR

[Lin00] Rafael D. Lins. Book review: Research Directions in Parallel Functional Programming, by Kevin Hammond and Greg Michaelson, editors, Springer-Verlag, 1999. Journal of Functional Programming, 10 (5):501–508, September 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/3BC7EEFC025142E799170BA58B53CCA0.

Lipovaca:2012:LYH

[Lip12] Miran Lipovača. Learn you a Haskell for great good!: a beginner’s guide. No Starch Press, San Francisco, CA, USA, 2012. ISBN 1-59327-283-9. xviii + 375 pp. LCCN QA76.73.H37 L69

2012. URL http://www.loc.gov/catdir/enhancements/fy1107/
2011000790-b.html;   http://www.loc.gov/catdir/enhancements/
fy1107/2011000790-d.html;   http://www.loc.gov/catdir/enhancements/
fy1108/2011000790-t.html.

Lisper:1994:TUT

[Lis94] Björn Lisper. Total unfolding: theory and applications. Journal of Functional Programming, 4(4):479–498, October 1994. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/672EFB8243199A8431472F3F5AD3EE64.

Liu:2013:SPD

[Liu13] Yanhong Annie Liu. Systematic program design: from clarity to efficiency. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-33691-0 (e-book), 1-139-56787-X (e-book), 1-299-83844-8 (e-book), 1-107-33359-8 (e-book), 1-107-03660-7 (hardcover), 1-107-61079-6 (paperback), 1-107-33525-6. xv + 240 pp. LCCN

QA76.6 .L578 2013. URL http://proquest.safaribooksonline.
com/9781107327153.

Ljunglof:2004:FPF

[Lju04] Peter Ljunglöf. Functional pearl: Functional chart parsing of context-free grammars. Journal of Functional Programming, 14 (6):669–680, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/FDEC977AD0E2A5067F2A865F2BE72740.

Lambert:1993:UMF

[LLR93] Tim Lambert, Peter Lindsay, and Ken Robinson. Using Miranda as a first programming language. Journal of Functional Programming, 3 (1):5–34, January 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/2FB7DC08AA6A4FE4A65E40DEBA82EF5C.

Leroy:1993:DM

[LM93] Xavier Leroy and Michel Mauny. Dynamics in ML. Journal of Functional Programming, 3(4):431–463, October 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4732BA16D60F8E525E56938D4404047B.

Loogen:2005:PFP

[LOMPM05] Rita Loogen, Yolanda Ortega-Mallén, and Ricardo Peña-Marí. Parallel functional programming in Eden. Journal of Functional Programming, 15(3):431–475, May 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/F60BD0DF77CC25B7ABFC37A98EB1C607.

Loulergue:2004:BRD

[Lou04] Frèdèric Loulergue. Book review: Développement d’applications avec Objective CAML, by E. Chailloux, P. Manoury and B. Pagano, O’Reilley, 2003. Journal of Functional Programming, 14(5):592– 594, September 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/975F7772B4710DBA30EBC8804E382658.

Le:1999:URM

[LP99] Dinh Lê and D. Stott Parker. Using randomization to make recursive matrix algorithms practical. Journal of Functional Programming, 9 (6):605–624, November 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/5F45C741CF61E5345A6C8174DA9E120D.

Levin:2003:TLP

[LP03] Michael Y. Levin and Benjamin C. Pierce. TinkerType: a language for playing with formal systems. Journal of Functional Programming, 13(2):295–316, March 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/48B7A773D0B8FDB7D1A3E644F8BEA9CA.

Lack:2009:GUD

[LP09] Stephen Lack and John Power. Gabriel-ulmer duality and Lawvere theories enriched over a general base. Journal of Functional Programming, 19(3–4):265–286, July 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/15991DB3A44153A628968282D8287C20.

Lakin:2013:CEI

[LP13] Matthew R. Lakin and Andrew M. Pitts. Contextual equivalence for inductive definitions with binders in higher order typed functional programming. Journal of Functional Programming, 23(6): 658–700, November 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/D6A89CB54DE7C5CBF1314DDD2B206C12.

Luksic:2020:LAE

[LP20] Ziga Luksic and Matija Pretnar. Local algebraic effect theories. Journal of Functional Programming, 30(??):e13, ???? 2020. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/local-algebraic-effect-theories/
670D4567BB7D87DA956E1EAD7F1DFD6F.

Leifer:1996:FPD

[LS96] J. J. Leifer and B. A. Sufrin. Functional pearl deduction for functional programmers. Journal of Functional Programming, 6(2):365– 373, March 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/BDAF47F832165FC6855534DB0025E085.

Loverdos:2010:SSI

[LS10] Christos K. K. Loverdos and Apostolos Syropoulos. Steps in Scala: an introduction to object-functional programming. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76217-0, 0-521-74758-9 (paperback). xvii + 485 pp. LCCN QA76.73.S28

L68 2010. URL http://assets.cambridge.org/97805217/62175/
cover/9780521762175.jpg;   http://catdir.loc.gov/catdir/
enhancements/fy1012/2010033849-b.html;   http://catdir.
loc.gov/catdir/enhancements/fy1012/2010033849-d.html;
http://catdir.loc.gov/catdir/enhancements/fy1012/2010033849-
t.html.

Lins:1994:EBC

[LTJ94] Rafael D. Lins, Simon J. Thompson, and Simon Peyton Jones. On the equivalence between CMC and TIM. Journal of Functional Programming, 4(1):47–63, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4EEF3A8AC974156CDF2A588CF83537F3.

Luo:2003:PLF

[Luo03] Zhaohui Luo. PAL+: a lambda-free logical framework. Journal of Functional Programming, 13(2):317–338, March 2003. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/C2F3220804821E7AA0D57FF1B48F8AE4.

Luth:2003:HSI

[Lüt03] Christoph Lüth. Haskell in space: An interactive game as a functional programming exercise. Journal of Functional Programming, 13(6):1077–1085, November 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/5FC8F1953AFD28FA7C69A8AC1EB5F9A7.

Luth:1999:FDI

[LW99] C. Lüth and B. Wolff. Functional design and implementation of graphical user interfaces for theorem provers. Journal of Functional Programming, 9(2):167–189, March 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/CCF0F17A6E0B3A38BF7FADA195D2CD17.

Lindley:2010:AC

[LWY10] Sam Lindley, Philip Wadler, and Jeremy Yallop. The arrow calculus. Journal of Functional Programming, 20(1):51–69, January 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
06F0C85EA019CB355B6E89932BBF28F7.

Mackie:1994:LFP

[Mac94] Ian Mackie. Lilac: a functional programming language based on linear logic. Journal of Functional Programming, 4(4):395–433, October 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
67CAB4EBFCC3B9861BEFA108793C699A.

MacLennan:1999:PPL

[Mac99] Bruce J. MacLennan. Principles of programming languages: design, evaluation, and implementation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, third edition, 1999. ISBN 0-19-511306-3 (hardcover). xvii + 509 pp. LCCN QA76.7 .M33 1999. URL

http://www.loc.gov/catdir/enhancements/fy0602/98027755-
d.html;   http://www.loc.gov/catdir/enhancements/fy0602/
98027755-t.html.

Mairson:1992:QEP

[Mai92] Harry G. Mairson. Quantifier elimination and parametric polymorphism in programming languages. Journal of Functional Programming, 2(2):213–226, April 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/64FDA6F4CEB7A0EDCB39F98407309623.

Mairson:2004:FPL

[Mai04] Harry G. Mairson. Functional pearl: Linear lambda calculus and PTIME-completeness. Journal of Functional Programming, 14(6): 623–633, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/E8F3718983561B265BB6BAECC1974DDB.

Marlow:2002:DHP

[Mar02] Simon Marlow. Developing a high-performance web server in Concurrent Haskell. Journal of Functional Programming, 12(4–5): 359–374, July 2002. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/69170C1B55537BEE303E846C87A33CFD.

Maranget:2004:FPF

[Mar04] Luc Maranget. Functional pearl: Functional satisfaction. Journal of Functional Programming, 14(6):647–656, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/DC5E2CA1D666AB43CA19E0CF72A02948.

Maranget:2007:WPM

[Mar07] Luc Maranget. Warnings for pattern matching. Journal of Functional Programming, 17(3):387–421, May 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3165B75113781E2431E3856972940347.

Matthes:2009:IPN

[Mat09] Ralph Matthes. An induction principle for nested datatypes in intensional type theory. Journal of Functional Programming, 19(3–4): 439–468, July 2009. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/7B1D02B3287CD508F83545E1B5157FDF.

Mccutchen:2020:ESD

[MBG^{+}20] Matt Mccutchen, Judith Borghouts, Andrew D. Gordon, Simon Peyton Jones, and Advait Sarkar. Elastic sheet-defined functions: Generalising spreadsheet functions to variable-size input arrays. Journal of Functional Programming, 30(??):e26, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
elastic-sheetdefined-functions-generalising-spreadsheet-
functions-to-variablesize-input-arrays/A85E83284ADAFF27B8607BBF20175D7B.

Martin:2013:FIC

[MC13] C. E. Martin and S. A. Curtis. Fractal image compression. Journal of Functional Programming, 23(6):629–657, November 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EFC967F9A094A9D43075789B2D12A8D1.

Mu:2021:LSB

[MC21] Shin-Cheng Mu and Tsung-Ju Chiang. Longest segment of balanced parentheses: an exercise in program inversion in a segment problem. Journal of Functional Programming, 31 (??):e31, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
longest-segment-of-balanced-parentheses-an-exercise-in-
program-inversion-in-a-segment-problem/219E51626DE1E48DF2FBEA0070483A99.

McBride:2002:FIS

[McB02] Conor McBride. Faking it simulating dependent types in Haskell. Journal of Functional Programming, 12(4–5):375–392, July 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
A904B84CA962F2D75578445B703F199A.

Mcbride:2003:FOU

[Mcb03] Conor Mcbride. First-order unification by structural recursion. Journal of Functional Programming, 13(6):1061–1075, November 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
91476025EE07F5E0F81F8F61A0C06C3B.

Morrisett:2002:SBT

[MCGW02] Greg Morrisett, Karl Crary, Neal Glew, and David Walker. Stackbased typed assembly language. Journal of Functional Programming, 12(1):43–88, January 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/FAA86C307845C6E28B88F57EE64C6F3B.

Morrisett:2003:SBT

[MCGW03] Greg Morrisett, Karl Crary, Neal Glew, and David Walker. Stackbased typed assembly language. Journal of Functional Programming, 13(5):957–959, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C2436AF12AEDEADC0CD2AB40A1EA185F.

Mcilroy:1999:PSP

[Mci99] M. Douglas Mcilroy. Power series, power serious. Journal of Functional Programming, 9(3):325–337, May 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/19863F4EAACC33E1E01DE2A2114EC7DF.

McIlroy:2004:ESR

[McI04] M. Douglas McIlroy. Enumerating the strings of regular languages. Journal of Functional Programming, 14(5):503–518, September 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL http://www.cs.dartmouth.edu/~doug/nfa.ps.gz;
 https://www.cambridge.org/core/product/1D46239B6CC6299AA385B3094EBC80E1.

McKenzie:1999:BRM

[McK99] Bruce McKenzie. Book review: Modern Compiler Implementation in ML: Basic Techniques, by Andrew W. Appel, Cambridge University Press, 1997, ISBN 0-521-58775-1. Journal of Functional Programming, 9(1):105–111, January 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A0EAB2711910493EF8D2F068798B5E45.

Myers:1993:PSM

[MCP93] Colin Myers, Chris Clack, and Ellen Poon. Programming with Standard ML. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-722075-8 (paperback). x + 301 pp. LCCN QA76.73.M6 M94 1993. US$40.00.

Meertens:2004:FPC

[Mee04] Lambert Meertens. Functional pearl: Calculating the Sieve of Eratosthenes. Journal of Functional Programming, 14(6):759–763, November 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/703B6CED6FA81741408572C1399376B0.

Meijer:2000:SSW

[Mei00] Erik Meijer. Server side web scripting in Haskell. Journal of Functional Programming, 10(1):1–18, January 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/F34DEFF8D123F59DB19AAF76DA5241F0.

Melham:1997:CPJ

[Mel97] Tom Melham. Call for papers: Journal of Functional Programming special issue on theorem provers and functional programming. Journal of Functional Programming, 7(1):125–126, January 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E07EBE06FB1FA1296F0EEF5A8A6DC9F4.

Mena:2014:BHP

[Men14] Alejandro Serrano Mena. Beginning Haskell: a project-based approach. The expert’s voice in programming. Apress, Berkeley, CA, USA, 2014. ISBN 1-4302-6250-8. xxiii + 402 pp. LCCN

QA76.73.H37 M46 2014. URL http://www.loc.gov/catdir/
enhancements/fy1513/2014466576-b.html;   http://www.loc.
gov/catdir/enhancements/fy1513/2014466576-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1513/2014466576-t.html.

Matthews:2008:OSS

[MF08] Jacob Matthews and Robert Bruce Findler. An operational semantics for Scheme 1. Journal of Functional Programming, 18(1): 47–86, January 2008. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/8C698C51D281C4E22DA1917DCCFE5EDA.

Moy:2023:TC

[MF23] Cameron Moy and Matthias Felleisen. Trace contracts. Journal of Functional Programming, 33:??, ???? 2023. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/trace-contracts/4AF1C7361751839FF7E2DEBC65A050EE.

Malde:2006:CPP

[MG06] Ketil Malde and Robert Giegerich. Calculating PSSM probabilities with lazy dynamic programming. Journal of Functional Programming, 16(1):75–81, January 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/90FB7BFBED26FB865F87D45BDEB947E8.

Michaelson:2000:BRO

[Mic00] Greg Michaelson. Book review: The Optimal Implementation of Functional Programming Languages, by A. Asperti and S. Guerrini, Cambridge University Press, 1998, 392 pp. Journal of Functional Programming, 10(5):501–508, September 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7822F5E4AF914276C9AC0CF5062D54AF.

Michaelson:2007:BRI

[Mic07] Greg Michaelson. Book review: Inductive Synthesis of Functional Programs, by Schmid U., Springer-Verlag, 2003, 420 pp., ISBN 3-540-40174-1. Journal of Functional Programming, 17(3):428–429, May 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
57162A67231632A36C491075606DD299.

Michaelson:2013:BRL

[Mic13] Greg Michaelson. Book review: Learn You a Haskell for Great Good! A Beginner’s Guide, by Miran Lipovaca, No Starch Press, April 2011, ISBN-10: 1-59327-283-9; ISBN-13: 978-1-59327-283-8, 376 pp. Journal of Functional Programming, 23(3):351–352, May 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D535063849F031C4E4371D6EB4A45967.

Miljenovic:2013:BRP

[Mil13] Ivan Lazar Miljenovic. Book review: Pearls of Functional Algorithm Design, by Richard Bird, Cambridge University Press, September 2010, £35.00, US $60.00. ISBN: 978-0-521-51338-8 (hardback), 286 pp. Journal of Functional Programming, 23(2):226–227, March 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D2D1D7A35DAFB5FC4343475901F61CF6.

Mitchell:1991:TIS

[Mit91] John C. Mitchell. Type inference with simple subtypes. Journal of Functional Programming, 1(3):245–285, July 1991. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/4A68C2794A117954C960CFA777815573.

Mitchell:1995:BRA

[Mit95] K. Mitchell. Book review: Abstract Data Types in Standard ML, by Rachel Harrison, John Wiley & Sons, 1993, 212 pp., ISBN 0-471-93844-0. Journal of Functional Programming, 5(1):131–134, January 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E4DD1AF5CFE6441B2D1E85B7323E812E.

Mitchell:2003:CPL

[Mit03] John C. Mitchell. Concepts in programming languages. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-78098-5. x + 529 pp. LCCN QA76.6 .M5293 2003. URL http://www.loc.

gov/catdir/description/cam0210/2002031217.html;   http:/
/www.loc.gov/catdir/samples/cam033/2002031217.html;   http:
//www.loc.gov/catdir/toc/cam031/2002031217.html.

Meehan:1998:AFL

[MJ98a] Gary Meehan and Mike Joy. Animated fuzzy logic. Journal of Functional Programming, 8(5):503–525, September 1998. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/D6B160DF985798F1B6D2531CF310EC29.

Morgan:1998:PLS

[MJ98b] R. G. Morgan and S. A. Jarvis. Profiling large-scale lazy functional programs. Journal of Functional Programming, 8(3):201–237, May 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E92505FBBC05DC6A8F31FB66BB6609D9.

Mairson:1997:CPJ

[MK97] Harry Mairson and Bruce Kapron. Call for papers: Journal of Functional Programming special issue on functional programming and computational complexity. Journal of Functional Programming, 7 (3):355–356, May 1997. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/6B5889E32C8C4D0A58E9E0965C209ED1.

Morihata:2012:MAF

[MKHT12] Akimasa Morihata, Kazuhiko Kakehi, Zhenjiang Hu, and Masato Takeichi. Manipulating accumulative functions by swapping calltime and return-time computations. Journal of Functional Programming, 22(3):275–299, May 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/CDE3159D777DD10FB5058D693E6CF294.

Mu:2009:APA

[MKJ09] Shin-Cheng Mu, Hsiang-Shang Ko, and Patrik Jansson. Algebra of programming in Agda: Dependent types for relational program derivation. Journal of Functional Programming, 19(5):545– 579, September 2009. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/ACA0C08F29621A892FB0C0B745254D15.

Major:1991:DGF

[MLC91] François Major, Guy Lapalme, and Robert Cedergren. Domain generating functions for solving constraint satisfaction problems. Journal of Functional Programming, 1(2):213–227, April 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B527101ADDFD765B2EAA84F0CC380443.

McIver:2003:PM

[MM03] Annabelle McIver and Carroll Morgan, editors. Programming methodology. Monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95349-3. xvii + 469 pp. LCCN QA76.6.P76 2003.

Mcbride:2004:VL

[MM04] Conor Mcbride and James Mckinna. The view from the left. Journal of Functional Programming, 14(1):69–111, January 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F8A44CAC27CCA178AF69DD84BC585A2D.

Madhavapeddy:2012:CWR

[MME12] Anil Madhavapeddy, Yaron Minsky, and Marius Eriksen. CUFP 2011 workshop report. Journal of Functional Programming, 22(1): 1–8, January 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/F22A5B087C6DD9A382D518F6DE08477A.

Mokhov:2020:BSC

[MMJ20] Andrey Mokhov, Neil Mitchell, and Simon Peyton Jones. Build systems à la carte: Theory and practice. Journal of Functional Programming, 30(??):e11, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/build-systems-a-la-carte-theory-and-
practice/097CE52C750E69BD16B78C318754C7A4.

Muller:1999:HHL

[MNVS99] Olaf Müller, Tobias Nipkow, David Von Oheimb, and Oscar Slotosch. HOLCF = HOL + LCF. Journal of Functional Programming, 9(2):191–223, March 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/DF841795DC24D986945EB72EBBE31939.

Myreen:2014:PPT

[MO14] Magnus O. Myreen and Scott Owens. Proof-producing translation of higher-order logic into pure and stateful ML. Journal of Functional Programming, 24(2–3):284–315, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4836EEA7A25F733339A9D23CD6C9F208.

Mogensen:1992:ESI

[Mog92] Torben Æ. Mogensen. Efficient self-interpretation in lambda calculus. Journal of Functional Programming, 2(3):345–364, July 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4727ED102FD8F5C5950EBBFE776E8C69.

Molyneux:1993:FPB

[Mol93] Phil Molyneux. Functional programming for business students. Journal of Functional Programming, 3(1):35–48, January 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
22523DD5F1EA13510D348DE4E9E6B226.

Morazan:2018:IHB

[Mor18] Marco T. Morazán. Infusing an HtDP-based CS1 with distributed programming using functional video games. Journal of Functional Programming, 28:e5, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
infusing-an-htdpbased-cs1-with-distributed-programming-
using-functional-video-games/22DB3A4087A49FF56A77354300B746AC.

Morihata:2021:LCA

[Mor21] Akimasa Morihata. Lambda calculus with algebraic simplification for reduction parallelisation: Extended study. Journal of Functional Programming, 31(??):e7, ???? 2021. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/lambda-calculus-with-algebraic-simplification-
for-reduction-parallelisation-extended-study/6B54DF4F6E295C414DE8A4029FE8B1F1.

Maraist:1998:CNL

[MOW98] John Maraist, Martin Odersky, and Philip Wadler. The call-by-need lambda calculus. Journal of Functional Programming, 8(3): 275–317, May 1998. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/7EDF4164D2F6EFBB5D36544D5390151A.

Moy:2024:KMP

[Moy24] Cameron Moy. Knuth–Morris–Pratt illustrated. Journal of Functional Programming, 34:??, ???? 2024. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/knuthmorrispratt-illustrated/8EFA77D663D585B68630E372BCE1EBA4.

Marlow:2006:MFC

[MP06] Simon Marlow and Simon Peyton Jones. Making a fast curry: push/ enter vs. eval/apply for higher-order languages. Journal of Functional Programming, 16(4–5):415–449, July 2006. CODEN JFPRES.

ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/02447DB613E94DC35ACDCB24DB39F085.

Mcbride:2008:APE

[MP08] Conor Mcbride and Ross Paterson. Applicative programming with effects. Journal of Functional Programming, 18(1):1–13, January 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C80616ACD5687ABDC86D2B341E83D298.

Moss:2001:IBP

[MR01] Graeme E. Moss and Colin Runciman. Inductive benchmarking for purely functional data structures. Journal of Functional Programming, 11(5):525–556, September 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/65B673660F49C5C993B66CA4397A6189.

Michaelson:1995:PPV

[MS95] Greg Michaelson and Norman Scaife. Prototyping a parallel vision system in Standard ML. Journal of Functional Programming, 5(3): 345–382, July 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/6EDFE4ACC037662AA4B15F7D0FBFD819.

Moggi:2001:MEE

[MS01] E. Moggi and Amr Sabry. Monadic encapsulation of effects: a revised approach (extended version). Journal of Functional Programming, 11(6):591–627, November 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/D31AEC33C1E8291B309578FA84B4664D.

Monnier:2003:ISC

[MS03] Stefan Monnier and Zhong Shao. Inlining as staged computation. Journal of Functional Programming, 13(3):647–676, May 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
AC9DC073D315A61870C3D85FEBF23657.

Mcnamara:2004:FPF

[MS04] Brian Mcnamara and Yannis Smaragdakis. Functional programming with the FC++ library. Journal of Functional Programming, 14(4): 429–472, July 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/1EC78833E953D6798C39C2E4B15EE4B4.

Might:2008:ERC

[MS08] Matthew Might and Olin Shivers. Exploiting reachability and cardinality in higher-order flow analysis. Journal of Functional Programming, 18(5):821–864, September 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/5B6F7BCB241D1B0701DB312E12AE6769.

Mason:1991:EFL

[MT91] Ian Mason and Carolyn Talcott. Equivalence in functional languages with effects. Journal of Functional Programming, 1(3):287–327, July 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
AC91F676E376BD48AA9DB0423E179249.

McBride:2009:P

[MU09] Conor McBride and Tarmo Uustalu. Preface. Journal of Functional Programming, 19(3–4):263–264, July 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/EE6173DB19BE4F4751D225EB6AF770A4.

Minsky:2008:CTE

[MW08] Yaron Minsky and Stephen Weeks. Caml trading — experiences with functional programming on Wall Street. Journal of Functional Programming, 18(4):553–564, July 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/02F18023B4C43BF6E53512AA7062A9A5.

Matsuda:2018:ABP

[MW18] Kazutaka Matsuda and Meng Wang. Applicative bidirectional programming: Mixing lenses and semantic bidirectionalization. Journal of Functional Programming, 28:e15, ???? 2018. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/applicative-bidirectional-programming/
82D617027973B92A50FE977FCAE72FAE.

Matsuda:2024:SLP

[MW24] Kazutaka Matsuda and Meng Wang. Sparcl: A language for partially invertible computation. Journal of Functional Programming, 34:??, ???? 2024. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/

sparcl-a-language-for-partially-invertible-computation/
809BDECF87B3748ED960FEFD42498BBE.

Nakano:2013:MJ

[Nak13] Keisuke Nakano. Metamorphism in jigsaw. Journal of Functional Programming, 23(2):161–173, March 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/15F720F604C69433542883DF43EE88BC.

Narbel:2007:TSC

[Nar07] Philippe Narbel. Type sharing constraints and undecidability. Journal of Functional Programming, 17(2):207–214, March 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
032BC34F95AB3CD3CB4DFE04DA560208.

Neis:2011:NPP

[NDR11] Georg Neis, Derek Dreyer, and Andreas Rossberg. Non-parametric parametricity. Journal of Functional Programming, 21(4–5):497– 562, September 2011. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/C81A547CCE117F612341364FB10FDEEF.

Neergaard:2005:TPB

[Nee05] Peter Møller Neergaard. Theoretical pearls: A bargain for intersection types: a simple strong normalization proof. Journal of Functional Programming, 15(5):669–677, September 2005. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/498F1ED0E87B54F3CD97E8EE41BA221E.

Nilsson:1994:ADL

[NF94] Henrik Nilsson and Peter Fritzson. Algorithmic debugging for lazy functional languages. Journal of Functional Programming, 4(3): 337–369, July 1994. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/2A77B114772CC3E74E0244E4586B60A5.

New:2017:FEC

[NFFM17] Max S. New, Burke Fetscher, Robert Bruce Findler, and Jay Mccarthy. Fair enumeration combinators. Journal of Functional Programming, 27:e19, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
fair-enumeration-combinators/CCEC7CBA286B2F792B073586EFCD6F3A.

Nakata:2009:SSB

[NH09] Keiko Nakata and Masahito Hasegawa. Small-step and big-step semantics for call-by-need. Journal of Functional Programming, 19 (6):699–722, November 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/B3210592CE62AB92DF78D9B055DFC04A.

Niehren:2000:UCC

[Nie00] Joachim Niehren. Uniform confluence in concurrent computation. Journal of Functional Programming, 10(5):453–499, September 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
271AB639DACFA4796A3E1A01EA9A3E1B.

Nilsson:2001:HLB

[Nil01] Henrik Nilsson. How to look busy while being as lazy as ever: the implementation of a lazy functional debugger. Journal of Functional Programming, 11(6):629–671, November 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/CEE05894A09DA11ECD1FC9952AD4C008.

New:2021:GTT

[NLA21] Max S. New, Daniel R. Licata, and Amal Ahmed. Gradual type theory. Journal of Functional Programming, 31(??):e21, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/gradual-type-
theory/2D5DC0E87301B1724C42B7E31F90DD6B.

Nanevski:2008:HTT

[NMB08] Aleksandar Nanevski, Greg Morrisett, and Lars Birkedal. Hoare type theory, polymorphism and separation. Journal of Functional Programming, 18(5):865–911, September 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D6B10CE5025B4C895C2FC7438393195E.

Nielson:1991:UTI

[NN91] Hanne Riis Nielson and Flemming Nielson. Using transformations in the implementation of higher-order functions. Journal of Functional Programming, 1(4):459–494, October 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/EA710E7904CD94841A606FBEBA8EA1B6.

Nishimura:1999:PFP

[NO99] Susumu Nishimura and Atsushi Ohori. Parallel functional programming on recursively defined data via data-parallel recursion. Journal of Functional Programming, 9(4):427–462, July 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B67FA8F9DD3686BA2DC7B44F277A13A0.

Normark:2005:WPS

[Nør05] Kurt Nørmark. Web programming in Scheme with LAML. Journal of Functional Programming, 15(1):53–65, January 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/4C8A275E2A7AAE2AB4A24977EF1EEB8F.

Nipkow:1995:TRT

[NP95] Tobias Nipkow and Christian Prehofer. Type reconstruction for type classes. Journal of Functional Programming, 5(2):201–224, April 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
073D53A7CFC6DB5304EC3FE4469B61C2.

Nanevski:2005:SCN

[NP05] Aleksandar Nanevski and Frank Pfenning. Staged computation with names and necessity. Journal of Functional Programming, 15 (6):893–939, November 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/13E2C64B0C6D91711C2DE0F3CD03C002.

Naylor:2012:RRR

[NR12] Matthew Naylor and Colin Runciman. The Reduceron reconfigured and re-evaluated. Journal of Functional Programming, 22(4–5):574– 613, September 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/9818E081664ADAFE9F61F1AEDAD0B043.

Nocker:1993:PSN

[NS93] Eric Nöcker and Sjaak Smetsers. Partially strict non-recursive data types. Journal of Functional Programming, 3(2):191–215, April 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
627109C103B7B5752E1C191699FC4715.

Nordin:2001:MLS

[NT01] Thomas Nordin and Andrew Tolmach. Modular lazy search for constraint satisfaction problems. Journal of Functional Programming, 11(5):557–587, September 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/0BD2B81EEF6550BD9DF9EC6BB5E1F079.

Nguyen:2017:HOS

[NTHV17] Phúc C. Nguyên, Sam Tobin-Hochstadt, and David Van Horn. Higher order symbolic execution for contract verification and refutation. Journal of Functional Programming, 27:e3, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/higher-order-
symbolic-execution-for-contract-verification-and-refutation/
129E10B4113242ABB0B7D6890CCFDBC9.

Nykanen:2011:NGS

[Nyk11] Matti Nykänen. A note on the genuine Sieve of Eratosthenes. Journal of Functional Programming, 21(6):563–572, November 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E71C1ECF451F35342D9CF2F33F8FEF66. See [O’N09].

ONeill:1997:NMF

[OB97] Melissa E. O’Neill and F. Warren Burton. A new method for functional arrays. Journal of Functional Programming, 7(5):487– 513, September 1997. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/7BBFBB58A34DFC28CBF0E758D989D622.

OConnor:2021:CUT

[OCR^{+}21] Liam O’Connor, Zilin Chen, Christine Rizkallah, Vincent Jackson, Sidney Amani, Gerwin Klein, Toby Murray, Thomas Sewell, and Gabriele Keller. Cogent: uniqueness types and certifying compilation. Journal of Functional Programming, 31(??):e25, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/cogent-uniqueness-types-
and-certifying-compilation/47AC86F02534818B95A56FA1A283A0A6.

ODonnell:2001:BRC

[O’D01] John O’Donnell. Book review: Concurrent Programming in ML, by John H. Reppy, Cambridge University Press, 1999, ISBN 0-521-48089-2, xv+308 pp. Journal of Functional Programming, 11(2): 253–262, March 2001. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/621FF1E119546D900CCFFE0DBDB3A098.

Oliveira:2010:SGP

[OG10] Bruno C. D. S. Oliveira and Jeremy Gibbons. Scala for generic programmers: Comparing Haskell and Scala support for generic programming. Journal of Functional Programming, 20(3–4):303–352, July 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
223EB37E77EA36B27AE33A644DA70926.

OHearn:1996:NAC

[O’H96] Peter W. O’Hearn. Note on Algol and conservatively extending functional programming. Journal of Functional Programming, 6(1):171– 180, January 1996. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/050F6652EB31D730ECC20A16E5B0E8EE.

OHearn:2003:BT

[O’H03] Peter O’Hearn. On bunched typing. Journal of Functional Programming, 13(4):747–796, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/4AFEBD081F0F0AFFCF03C8415DDE53A4.

Okasaki:1995:SEP

[Oka95] Chris Okasaki. Simple and efficient purely functional queues and deques. Journal of Functional Programming, 5(4):583–592, October 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7B3036772616B39E87BF7FBD119015AB.

Okasaki:1997:TAB

[Oka97] Chris Okasaki. Three algorithms on Braun trees. Journal of Functional Programming, 7(6):661–666, November 1997. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/59CF6B17F75EC216EBCA8A08E8780DFF.

Okasaki:1998:EHO

[Oka98] Chris Okasaki. Even higher-order functions for parsing or why would anyone ever want to use a sixth-order function? Journal of Functional Programming, 8(2):195–199, March 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/AAAA5C5E29889CEBC5E944CC1080FE8D.

Okasaki:1999:BRC

[Oka99a] Chris Okasaki. Book review: Computability and Complexity: From a Programming Perspective, by Neil D. Jones, MIT Press, 1997. Journal of Functional Programming, 9(4):479–482, July 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1373D4359E7AEA4AFAD334F84129F521.

Okasaki:1999:RBT

[Oka99b] Chris Okasaki. Red-black trees in a functional setting. Journal of Functional Programming, 9(4):471–477, July 1999. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/62BC5EA75A2C95E3F6EE95AE3DADF0E5.

Okasaki:2001:SIA

[Oka01] Chris Okasaki. Special issue on algorithmic aspects of functional programming languages. Journal of Functional Programming, 11 (5):439–440, September 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/504CECC3DCAF5F29E09AC217692C4DE6.

Okasaki:2003:TPF

[Oka03] Chris Okasaki. Theoretical pearls: Flattening combinators: surviving without parentheses. Journal of Functional Programming, 13 (4):815–822, July 2003. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/3E99993FE5464986AD94D292FF5EA275.

ONeill:2009:GSE

[O’N09] Melissa E. O’Neill. The genuine Sieve of Eratosthenes. Journal of Functional Programming, 19(1):95–106, January 2009. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/FD3E90871269020CA6C64C25AB8A4FBD.

See note [Nyk11].

Orbaek:1997:TC

[ØP97] P. Ørbæk and J. Palsberg. Trust in the λ-calculus. Journal of Functional Programming, 7(6):557–591, November 1997. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/302CC0487CED79DF534733D937EDE33C.

ODonnell:2004:FPD

[OR04] John T. O’Donnell and Gudula Rünger. Functional pearl: Derivation of a logarithmic time carry lookahead addition circuit. Journal of Functional Programming, 14(6):697–713, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
EE7E2A6B541526A0472DD180FDF6789F.

Orchard:2014:BRC

[Orc14] Dominic Orchard. Book review: Computational Semantics with Functional Programming, by Jan van Eijck and Christina Unger. Journal of Functional Programming, 24(4):524–527, July 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D64D86CC51BD935F57EB64AE2A9503FA.

Owens:2009:RED

[ORT09] Scott Owens, John Reppy, and Aaron Turon. Regular-expression derivatives re-examined. Journal of Functional Programming, 19(2): 173–190, March 2009. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/E5734B86DEB96C61C69E5CF3C4FB0AFA.

Oliveira:2012:MMR

[OSC12] Bruno C. D. S. Oliveira, Tom Schrijvers, and William R. Cook. MRI: Modular reasoning about interference in incremental programming. Journal of Functional Programming, 22(6):797–852, November 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
34D74AA93EA4EAAE23A30318EFBFF49D.

Padovani:2017:SLI

[Pad17] Luca Padovani. A simple library implementation of binary sessions. Journal of Functional Programming, 27:e4, ???? 2017. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-

programming/article/simple-library-implementation-of-binary-
sessions/D4F6DF42D95358937A5269A60193AAE6.

Page:1997:BRA

[Pag97] Rex Page. Book review: Applications of Functional Programming, by Colin Runciman and David Wakeling, editors, UCL Press, 1995. Journal of Functional Programming, 7(5):555–556, September 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
933AD166F65F8100385624F43C2A706E.

Page:1998:BRA

[Pag98] Rex Page. Book review: Applications of Functional Programming edited by Colin Runciman and David Wakeling, UCL Press Limited, 1995. Journal of Functional Programming, 8(4):445–446, July 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
73E7FBA126A66A293E0B751CCFDFD75D.

Page:2007:ESC

[Pag07] Rex Page. Engineering software correctness. Journal of Functional Programming, 17(6):675–686, November 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/FECD1C78F3965973EC5BA0BFE1ABAC06.

Palsberg:1993:CBT

[Pal93] Jens Palsberg. Correctness of binding-time analysis. Journal of Functional Programming, 3(3):347–363, July 1993. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/BF52D99DB03BE9722F476CC4295D7556.

Parr:2010:LIP

[Par10] Terence (Terence John) Parr. Language implementation patterns: create your own domain-specific and general programming languages. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2010. ISBN 1-934356-45-X (paperback). xix + 374 pp. LCCN QA76.7 .P375 2010.

Parys:2016:CLT

[Par16] Pawel Parys. A characterization of lambda-terms transforming numerals. Journal of Functional Programming, 26:e12, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
characterization-of-lambdaterms-transforming-numerals/852D33A97D7A849191B3EFA5ECE7E458.

Patryshev:2014:BRI

[Pat14] Vlad Patryshev. Book review: Introduction to the art of programming using Scala, by Mark C. Lewis, Chapman and Hall/CRC Press, 2012, £46.99 (paperback) ISBN-10: 1-4398-9666-

6. Journal of Functional Programming, 24(5):608–609, September 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F8E004AA7CFDA2440A6A9FCA9BD81CCF.

Paulson:1991:MWP

[Pau91] Lawrence C. Paulson. ML for the working programmer. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-39022-2. x + 429 pp. LCCN QA76.73.M6 P38 1991.

Petricek:2021:CDV

[Pet21] Tomas Petricek. Composable data visualizations. Journal of Functional Programming, 31(??):e13, ???? 2021. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/composable-data-visualizations/CFC3E7AFACBEE62AE3AC70AD6DF4F3D5.

PeytonJones:1992:ILF

[Pey92] Simon L. Peyton Jones. Implementing lazy functional languages on stock hardware: the Spineless Tagless G-machine. Journal of Functional Programming, 2(2):127–202, April 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/354FFB29102309CCD2A3824F894A2799.

PeytonJones:2003:I

[Pey03a] Simon Peyton Jones. 1 introduction. Journal of Functional Programming, 13(1):i–6, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C418D095ECDDA0E2C777BA68D09C2592.

PeytonJones:2003:SDI

[Pey03b] Simon Peyton Jones. 10 specification of derived instances. Journal of Functional Programming, 13(1):139–144, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
81CAE922465978B6F3410EDC53BBCB0D.

PeytonJones:2003:CP

[Pey03c] Simon Peyton Jones. 11 compiler pragmas. Journal of Functional Programming, 13(1):145–146, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/B57AE39A32A5A12C2482FBBBF1053AE5.

PeytonJones:2003:RNa

[Pey03d] Simon Peyton Jones. 12 rational numbers. Journal of Functional Programming, 13(1):149–152, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/AC5742F18F56DDBC1A1F9E0183B8A215.

PeytonJones:2003:CN

[Pey03e] Simon Peyton Jones. 13 complex numbers. Journal of Functional Programming, 13(1):153–156, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/8C55E08669D5EB6F40DBED820B3CFA85.

PeytonJones:2003:NF

[Pey03f] Simon Peyton Jones. 14 numeric functions. Journal of Functional Programming, 13(1):157–168, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/6D11CE287442484A58F189C8619B1A1F.

PeytonJones:2003:IOa

[Pey03g] Simon Peyton Jones. 15 indexing operations. Journal of Functional Programming, 13(1):169–172, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/47AC0A8BD9F0D447F15145CBF341FF05.

PeytonJones:2003:A

[Pey03h] Simon Peyton Jones. 16 arrays. Journal of Functional Programming, 13(1):173–178, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/6CCA8C69B7ADF5D795FBFB3C12F15244.

PeytonJones:2003:LU

[Pey03i] Simon Peyton Jones. 17 list utilities. Journal of Functional Programming, 13(1):179–190, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/E46CD6A1ABCDF6E08170B8255C58EACE.

PeytonJones:2003:MUa

[Pey03j] Simon Peyton Jones. 18 Maybe utilities. Journal of Functional Programming, 13(1):191–192, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/F01A9571ED93426312B03C95381F852B.

PeytonJones:2003:CU

[Pey03k] Simon Peyton Jones. 19 character utilities. Journal of Functional Programming, 13(1):193–198, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/28AB9D1B1C5DB1C431114A9BACCBC886.

PeytonJones:2003:LS

[Pey03l] Simon Peyton Jones. 2 lexical structure. Journal of Functional Programming, 13(1):7–16, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/52807D0D55B94769840471E2A39864C7.

PeytonJones:2003:MUb

[Pey03m] Simon Peyton Jones. 20 monad utilities. Journal of Functional Programming, 13(1):199–204, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/D74BCFCB041AFF5AF5A799451999A8EB.

PeytonJones:2003:IOb

[Pey03n] Simon Peyton Jones. 21 input/output. Journal of Functional Programming, 13(1):205–218, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/032C8D62ACDD94C0D3EA86900DC9843B.

PeytonJones:2003:DF

[Pey03o] Simon Peyton Jones. 22 directory functions. Journal of Functional Programming, 13(1):219–222, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/F1C2EF26D6D98A1F16C6677151D5AD19.

PeytonJones:2003:SF

[Pey03p] Simon Peyton Jones. 23 system functions. Journal of Functional Programming, 13(1):223–224, January 2003. CODEN JFPRES.

ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/01D58F1F5AE939E537158458F973F412.

PeytonJones:2003:DT

[Pey03q] Simon Peyton Jones. 24 dates and times. Journal of Functional Programming, 13(1):225–230, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7DB9AC52468CE97307DF137FAEB47224.

PeytonJones:2003:L

[Pey03r] Simon Peyton Jones. 25 locales. Journal of Functional Programming, 13(1):231–232, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/3C08562BE49F3903F9FDF9204E027404.

PeytonJones:2003:CT

[Pey03s] Simon Peyton Jones. 26 CPU time. Journal of Functional Programming, 13(1):233–234, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/462582A0E2C7C43B347A1FBDCDB0050F.

PeytonJones:2003:RNb

[Pey03t] Simon Peyton Jones. 27 random numbers. Journal of Functional Programming, 13(1):235–240, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/E141AFF92D015912D4A5CC313B0D6A9F.

PeytonJones:2003:E

[Pey03u] Simon Peyton Jones. 3 expressions. Journal of Functional Programming, 13(1):17–38, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3994C7B69430EBECD0C9127FBCAAE2EA.

PeytonJones:2003:DB

[Pey03v] Simon Peyton Jones. 4 declarations and bindings. Journal of Functional Programming, 13(1):39–66, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A525696031147F9B7C9452E12DE4422B.

PeytonJones:2003:M

[Pey03w] Simon Peyton Jones. 5 modules. Journal of Functional Programming, 13(1):67–80, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/6363089C539D816849F92D67B1D2374B.

PeytonJones:2003:PTC

[Pey03x] Simon Peyton Jones. 6 predefined types and classes. Journal of Functional Programming, 13(1):81–96, January 2003. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/A512C0BB77D425E55066CBF97E68C153.

PeytonJones:2003:BIO

[Pey03y] Simon Peyton Jones. 7 basic input/output. Journal of Functional Programming, 13(1):97–102, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/43840B587166A4609773DA629DD318CD.

PeytonJones:2003:SP

[Pey03z] Simon Peyton Jones. 8 standard prelude. Journal of Functional Programming, 13(1):103–124, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/52BD5428223DA027C080DD24A25ECF60.

PeytonJones:2003:SR

[Pey03-27] Simon Peyton Jones. 9 syntax reference. Journal of Functional Programming, 13(1):125–138, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/F5BC1846746FDDEDC06E0A0FEF8408C9.

PeytonJones:2003:B

[Pey03-28] Simon Peyton Jones. Bibliography. Journal of Functional Programming, 13(1):241–255, January 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/99DCBB6B2FF8F852518493D2A27FB983.

Perez:2020:FTF

[PG20] Ivan Perez and Alwyn E. Goodloe. Fault-tolerant functional reactive programming (extended version). Journal of Functional Programming, 30(??):e12, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
faulttolerant-functional-reactive-programming-extended-
version/F0C270C83E218FA5627D96A7FD6C56E9.

Park:2009:TSH

[PI09] Sungwoo Park and Hyeonseung Im. Type-safe higher-order channels with channel locality. Journal of Functional Programming, 19(1): 107–142, January 2009. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/D351717817CBFA2884DF0234F29A1778.

Park:2011:CHD

[PI11] Sungwoo Park and Hyeonseung Im. A calculus for hardware description. Journal of Functional Programming, 21(1):21–58, January 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E755E501B667D2B9F3F4D2F80F256464.

Pickering:2007:FF

[Pic07] Robert Pickering. Foundations of F#. Expert’s voice in .NET. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-757-5. xxi + 360

pp. LCCN QA76.62 .P53 2007. URL http://www.loc.gov/catdir/
enhancements/fy1502/2008297942-b.html;   http://www.loc.
gov/catdir/enhancements/fy1502/2008297942-d.html;   http:
//www.loc.gov/catdir/enhancements/fy1502/2008297942-t.html.

Pientka:2013:ILL

[Pie13] Brigitte Pientka. An insider’s look at LF type reconstruction: everything you (n)ever wanted to know. Journal of Functional Programming, 23(1):1–37, January 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/154A57699A72B0B440CD95B717908348.

Pitts:2011:SRL

[Pit11] Andrew M. Pitts. Structural recursion with locally scoped names. Journal of Functional Programming, 21(3):235–286, May 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3BA6B0526749AAA53EE4C9067135BA52.

Papaspyrou:2000:SEO

[PM00] Nikolaos S. Papaspyrou and Dragan Maćos. A study of evaluation order semantics in expressions with side effects. Journal of Functional Programming, 10(3):227–244, May 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/E7D20CA2755BFF9927644E34E20D8FB6.

PeytonJones:2002:SGH

[PM02] Simon Peyton Jones and Simon Marlow. Secrets of the Glasgow Haskell Compiler inliner. Journal of Functional Programming, 12(4– 5):393–434, July 2002. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/8DD9A82FF4189A0093B7672193246E22.

Perez:2020:RVV

[PN20] Ivan Perez and Henrik Nilsson. Runtime verification and validation of functional reactive systems. Journal of Functional Programming, 30(??):e28, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/runtime-verification-and-validation-
of-functional-reactive-systems/875DE7B51D38C418739B441874EB23D2.

Potanin:2006:FGC

[PNCB06] Alex Potanin, James Noble, Dave Clarke, and Robert Biddle. Featherweight generic confinement. Journal of Functional Programming, 16(6):793–811, November 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/65F0C15E9EA9DFC113D2F8529D9171CE.

Poll:1998:EPN

[Pol98] Erik Poll. Expansion postponement for normalising pure type systems. Journal of Functional Programming, 8(1):89–96, January 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7012C229AB8920E2241988163C817B51.

Pottier:2013:SSP

[Pot13] François Pottier. Syntactic soundness proof of a type-and-capability system with hidden state. Journal of Functional Programming, 23(1):38–144, January 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/C389E0D461C51F7D20E871BF2DDF4B21.

Palsberg:2001:PFI

[PP01] Jens Palsberg and Christina Pavlopoulou. From polyvariant flow information to intersection and union types. Journal of Functional Programming, 11(3):263–317, May 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/124B92D33BB601EA71CEE2916255E782.

Pouillard:2012:UTS

[PP12] Nicolas Pouillard and François Pottier. A unified treatment of syntax with binders. Journal of Functional Programming, 22(4–5):614– 704, September 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/A12B7450ED7933F225D4C82195FD667A.

Pruiksma:2022:BF

[PP22] Klaas Pruiksma and Frank Pfenning. Back to futures. Journal of Functional Programming, 32(??):e6, ???? 2022. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/back-to-futures/9FA8FD46A69EDCD889CEE58756418CC6.

Pieters:2020:GME

[PRS20] Ruben P. Pieters, Exequiel Rivas, and Tom Schrijvers. Generalized monoidal effects and handlers. Journal of Functional Programming, 30(??):e23, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
generalized-monoidal-effects-and-handlers/E11A8F2693A1762295769026C91E2D1D.

Pena:2005:NDA

[PS05] Ricardo Peña and Clara Segura. Non-determinism analyses in a parallel-functional language. Journal of Functional Programming, 15(1):67–100, January 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/8D69E37F6547488134C6DE1E724CABB6.

Pirog:2017:BCD

[PS17] Maciej Piróg and Sam Staton. Backtracking with cut via a distributive law and left-zero monoids. Journal of Functional Programming, 27:e17, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
backtracking-with-cut-via-a-distributive-law-and-leftzero-
monoids/9B7B1620EEFD293888B7B1E0F805156B.

Pieters:2020:FCP

[PS20] Ruben P. Pieters and Tom Schrijvers. Faster coroutine pipelines: A reconstruction. Journal of Functional Programming, 30 (??):e22, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
faster-coroutine-pipelines-a-reconstruction/BE694B384B35DF1DC03684FF5AE3FBDE.

Pierce:1994:STT

[PT94] Benjamin C. Pierce and David N. Turner. Simple type-theoretic foundations for object-oriented programming. Journal of Functional Programming, 4(2):207–247, April 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/5C18E2E055B028F7214FBB183701830E.

Poll:1999:BRA

[PT99] Erik Poll and Simon Thompson. Book review: Algebra of Programming, by Richard Bird and Oege de Moor, Prentice Hall, 1996 (dated 1997). Journal of Functional Programming, 9(3):347–354, May 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7FE8E8A9A0B49E557E894D78DC085FF2.

Perna:2022:IMC

[PTW22] Stefano Perna, Val Tannen, and Limsoon Wong. Iterating on multiple collections in synchrony. Journal of Functional Programming, 32: e9, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/iterating-on-
multiple-collections-in-synchrony/E1E868AA95761C15F9896238442D790B.

PeytonJones:2007:PTI

[PVWS07] Simon Peyton Jones, Dimitrios Vytiniotis, Stephanie Weirich, and Mark Shields. Practical type inference for arbitrary-rank types. Journal of Functional Programming, 17(1):1–82, January 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
5339FB9DAB968768874D4C20FA6F8CB6.

Partridge:1996:PPC

[PW96] Andrew Partridge and David Wright. Predictive parser combinators need four values to report errors. Journal of Functional Programming, 6(2):355–364, March 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/8387A6EE9429D4D1EA84CE644C0E57CA.

PeytonJones:1998:E

[PW98] Simon Peyton Jones and Phil Wadler. EDITORIAL. Journal of Functional Programming, 8(4):319–321, July 1998. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/24AAB2F942D5ED6691EA2293DC87349C.

Palsberg:2003:CTF

[PW03a] Jens Palsberg and Mitchell Wand. CPS transformation of flow information. Journal of Functional Programming, 13(5):905–923, September 2003. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/1C74FEADD1A7C0BF6F85CFEEB39077DF.

PeytonJones:2003:EPC

[PW03b] Simon Peyton Jones and Phil Wadler. The educational pearls column. Journal of Functional Programming, 13(5):833–834, September 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6A30044075FC0B38B32122D6666E9B41.

Qu:2021:RCA

[QGG21] Weihao Qu, Marco Gaboardi, and Deepak Garg. Relational cost analysis in a functional-imperative setting. Journal of Functional Programming, 31(??):e27, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/relational-cost-analysis-in-a-functionalimperative-
setting/048CDEF0960FBB8C42704284E684CB1E.

Rompf:2019:SCC

[RA19] Tiark Rompf and Nada Amin. A SQL to C compiler in 500 lines of code. Journal of Functional Programming, 29:e9, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/sql-to-c-compiler-in-500-
lines-of-code/C38B40C78B6A9C55232D4A850587FC64.

Ragde:2013:BRO

[Rag13] Prabhakar Ragde. Book review: OCaml from the Very Beginning, by John Whitington, Coherent Press, 2013, £25.99, US $37.99. ISBN-10: 0-9576711-0-5 (paperback), 204 pp. Journal of Functional Programming, 23(3):352–354, May 2013. CODEN JFPRES.

ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/43ACFEB5D85B716D2176FAFE213F7A5C.

Ragde:2020:RAF

[Rag20] Prabhakar Ragde. Review of “Algorithms for Functional Programming” by John David Stone, Springer-Verlag, 2018. Journal of Functional Programming, 30(??):e19, ???? 2020. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/review-of-algorithms-for-functional-
programming-by-john-david-stone-springerverlag-2018/8FBE435BA79EC4F35CAAB64FAFCC183D.

Ramsey:2003:PAR

[Ram03] Norman Ramsey. Pragmatic aspects of reusable program generators. Journal of Functional Programming, 13(3):601–646, May 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C9BA9EBBE693BDD5323B2AC862F14A3E.

Ramsey:2011:EIL

[Ram11] Norman Ramsey. Embedding an interpreted language using higherorder functions and types. Journal of Functional Programming, 21 (6):585–615, November 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/922400659B6E10D7C4EE3F5DEF75A4B1.

Ranta:2004:GF

[Ran04] Aarne Ranta. Grammatical framework. Journal of Functional Programming, 14(2):145–189, March 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/09C4B3F43447ADDB0F632C64F195BC9C.

Reade:1993:TC

[Rea93] Chris Reade. Terminating comprehensions. Journal of Functional Programming, 3(2):247–250, April 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/988B876439491F61A3E72DBA3AD2F7D8.

Reade:1997:BRM

[Rea97] Chris Reade. Book reviews: ML for the Working Programmer (2nd edition) by L. C. Paulson, Cambridge University Press, 1996. A Practical Course in Functional Programming Using Standard ML, by R. Bosworth, McGraw Hill, 1996. Journal of Functional Programming, 7(4):447–449, July 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/D48ECEDA4961D4B7D37EC668939F6DE9.

Reade:2004:BRP

[Rea04] Chris Reade. Book review: Proof, Language and Interaction, Essays in Honour of Robin Milner, eds. Gordon Plotkin, Colin Sterling, Mads Tofte, Foundations of Computing Series, MIT Press 2000. Journal of Functional Programming, 14(1):125–128, January 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
651F6FDC2B8DDAD904D778B7889B76C2.

Reppy:1999:CPM

[Rep99] John H. Reppy. Concurrent programming in ML. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-48089-2 (hardcover). xiv + 308 pp. LCCN QA76.73.M6 R47 1999. URL http://www.

loc.gov/catdir/description/cam029/99020465.html;   http:
//www.loc.gov/catdir/toc/cam023/99020465.html.

Rockl:2003:FAS

[RH03] Christine Röckl and Daniel Hirschkoff. A fully adequate shallow embedding of the π-calculus in Isabelle/HOL with mechanized syntax analysis. Journal of Functional Programming, 13(2):415–451, March 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1754542B03E4A446E0A6686D6C2E5AF8.

Rhiger:2009:TSP

[Rhi09] Morten Rhiger. Type-safe pattern combinators. Journal of Functional Programming, 19(2):145–156, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/1E3D0890F2ED1B70F80722A732756910.

Richards:1998:BRH

[Ric98] Hamilton Richards. Book reviews: Haskell: The Craft of Functional Programming, by Simon Thompson, Addison–Wesley, 1996. Miranda 81: The Craft of Functional Programming, by Simon Thompson, Addison–Wesley, 1995. Journal of Functional Programming, 8(6):633–637, November 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/31DDA65AA3AF51BBEEA5B3C4FC66F32A.

Rittri:1991:UTS

[Rit91] Mikael Rittri. Using types as search keys in function libraries. Journal of Functional Programming, 1(1):71–89, January 1991. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BA56BB3061DB73396847E37048544E19.

Rivas:2017:NCM

[RJ17] Exequiel Rivas and Mauro Jaskelioff. Notions of computation as monoids. Journal of Functional Programming, 27:e21, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
notions-of-computation-as-monoids/70019FC0F2384270E9F41B9719042528.

Rabhi:1999:AFP

[RL99] Fethi Rabhi and Guy Lapalme. Algorithms: a functional programming approach. International computer science series. Addison-Wes-ley, Reading, MA, USA, second edition, 1999. ISBN 0-201-59604-0 (paperback). xi + 235 pp. LCCN QA76.9.A43 R344 1999.

Rossberg:2018:CMU

[Ros18] Andreas Rossberg. 1ML — core and modules united. Journal of Functional Programming, 28:e22, ???? 2018. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/1ml-core-and-modules-united/47B10882829E4B32F98FBA93B28CEF30.

Runciman:1996:NDH

[RR96] Colin Runciman and Niklas Röjemo. New dimensions in heap profiling. Journal of Functional Programming, 6(4):587–620, July 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
15780E8F8EE200809B26A6025AFF4027.

Rossberg:2014:FIM

[RRD14] Andreas Rossberg, Claudio Russo, and Derek Dreyer. F-ing modules. Journal of Functional Programming, 24(5):529–607, September 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B573FA00832D55D4878863DE1725D90B.

Runciman:1991:RRS

[RT91] Colin Runciman and Ian Toyn. Retrieving reusable software components by polymorphic type. Journal of Functional Programming, 1 (2):191–211, April 1991. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/A7B301B2E32F01136F9EC29DD70B0881.

Runciman:1993:IET

[RTF93] Colin Runciman, Ian Toyn, and Mike Firth. An incremental, exploratory and transformational environment for lazy functional programming. Journal of Functional Programming, 3(1):93–115, January 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BFF5A52A3D603F5BA23B7821E2C5E6E4.

Runciman:1997:FPL

[Run97] Colin Runciman. Functional pearl: Lazy wheel sieves and spirals of primes. Journal of Functional Programming, 7(2):219–225, March 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
24D8441CCD9E7CADB05EAFD47540D69C.

Russell:2003:BRA

[Rus03] D. Russell. Book review: Algorithms: A Functional Programming Approach, by Fethi Rabhi and Guy Lapalme, Addison–Wesley, 1999, ISBN 0-201-59604-0, xi + 235 pp. Journal of Functional Programming, 13(4):828–829, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/56B09E23C4B74ABBC5F3302C331F4795.

Ruf:1993:SOP

[RW93a] Erik Ruf and Daniel Weise. On the specialization of online program specializers. Journal of Functional Programming, 3(3):251–281, July 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
83985B7CE2B4E6CAC5DEAEF12B9B0000.

Runciman:1993:HPL

[RW93b] Colin Runciman and David Wakeling. Heap profiling of lazy functional programs. Journal of Functional Programming, 3(2):217–245, April 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E4DC90589F89A5A0C05CE41033780760.

Rondogiannis:1997:FOF

[RW97] P. Rondogiannis and W. W. Wadge. First-order functional languages and intensional logic. Journal of Functional Programming, 7(1):73– 101, January 1997. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/CDA5800533BC35832DDC9587E15EFCE0.

Rondogiannis:1999:HOF

[RW99] P. Rondogiannis and W. W. Wadge. Higher-order functional languages and intensional logic. Journal of Functional Programming, 9 (5):527–564, September 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/A768CE60E432D7F635C82C0F3BD64E4C.

Sperber:2014:SIR

[SA14] Michael Sperber and Lennart Augustsson. Special issue on runtime systems and target platforms for functional languages: Editorial. Journal of Functional Programming, 24(6):612, November 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
5F2B2C5D91A7121E11600E08CEA24BF4.

Sabry:1998:WPF

[Sab98] Amr Sabry. What is a purely functional language? Journal of Functional Programming, 8(1):1–22, January 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/3A39D50DA48F628D17D9A768A1FA39C3.

Sankel:2006:BRH

[San06] David Sankel. Book review: The Haskell Road to Logic, Maths and Programming, by Kees Doets and Jan van Eijck, King’s College Publications, 2004, ISBN 0-9543006-9-6. Journal of Functional Programming, 16(2):249–250, March 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4C3FA37728FBCF8FC70B0B38DCD9AB94.

Sankar:2009:BRP

[San09] Krishna Sankar. Book review: Programming Erlang — Software for a Concurrent World, by Armstrong Joe, Pragmatic Bookshelf, 2007, pp. 536. ISBN-10: 1-934356-00-X. Journal of Functional Programming, 19(2):259–261, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/51B42E4D3FFBB9A9D068ACA39B1081C0.

Statman:1999:APT

[SB99] Richard Statman and Henk Barendregt. Applications of plotkinterms: partitions and morphisms for closed terms. Journal of Functional Programming, 9(5):565–575, September 1999. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/6BCD6ACD4331B7DA55CC1987D99A49DB.

Spoonhower:2010:SPP

[SBHG10] Daniel Spoonhower, Guy E. Blelloch, Robert Harper, and Phillip B. Gibbons. Space profiling for parallel functional programs. Journal of Functional Programming, 20(5–6):417–461, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
31C0C87A456BBE21B5FCFEA19F600EF7.

Swierstra:2009:LBF

[SC09] S. Doaitse Swierstra and Olaf Chitil. Linear, bounded, functional pretty-printing. Journal of Functional Programming, 19(1):1–16, January 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0800ABD4B4A45914B02F9CC54213D040.

Siek:2021:PCC

[SC21] Jeremy G. Siek and Tianyu Chen. Parameterized cast calculi and reusable meta-theory for gradually typed lambda calculi. Journal of Functional Programming, 31(??):e30, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/parameterized-
cast-calculi-and-reusable-metatheory-for-gradually-typed-
lambda-calculi/7A6E772C5AEB832ED157E80C5D2085D4.

Swamy:2013:SDP

[SCF^{+}13] Nikhil Swamy, Juan Chen, Cédric Fournet, Pierre-Yves Strub, Karthikeyan Bhargavan, and Jean Yang. Secure distributed programming with value-dependent types. Journal of Functional Programming, 23(4):402–451, July 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/14288BBFE9B373738DC8F60C3E89E3E3.

Schmid:2003:ISF

[Sch03a] Ute Schmid, editor. Inductive Synthesis of Functional Programs: Universal Planning, Folding of Finite Programs, and Schema Abstraction by Analogical Reasoning, volume 2654 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CO-DEN LNCSD9. ISBN 3-540-40174-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xxii + 398 pp. LCCN QA76.6

.S3855 2003. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t2654.htm;   http://www.loc.gov/catdir/
enhancements/fy0818/2003059058-d.html;   http://www.springerlink.
com/content/978-3-540-40174-2;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=2654;
 http://www.springerlink.com/openurl.asp?genre=volume&id=
doi:10.1007/b12055.

Scholz:2003:SAC

[Sch03b] Sven-Bodo Scholz. Single Assignment C: efficient support for high-level array operations in a functional setting. Journal of Functional Programming, 13(6):1005–1059, November 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D0DDF8167C20487D81F84A2310838594.

Schneider:2004:VRS

[Sch04] Klaus Schneider. Verification of Reactive Systems: Formal Methods and Algorithms. Texts in theoretical computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-00296-0. xiv + 600 pp.

LCCN QA76.76.V47 S34 2004. URL http://www.loc.gov/catdir/
enhancements/fy0817/2003066407-b.html;   http://www.loc.
gov/catdir/enhancements/fy0817/2003066407-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0817/2003066407-t.html.

Schwinghammer:2009:CSM

[Sch09] Jan Schwinghammer. Coherence of subsumption for monadic types. Journal of Functional Programming, 19(2):157–172, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E9C237B07B01085E91762BAF14B0698E.

Schaathun:2015:ESP

[Sch15] Hans Georg Schaathun. Evaluation of splittable pseudo-random generators. Journal of Functional Programming, 25:e6, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-

functional-programming/article/evaluation-of-splittable-
pseudorandom-generators/3EBAA9F14939C5BB5560E32D1A132637.

Su:2024:SBG

[SCYZ24] Chenghao Su, Lin Chen, Li Yanhui, and Yuming Zhou. Static Blame for gradual typing. Journal of Functional Programming, 34: ??, ???? 2024. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/static-blame-
for-gradual-typing/A2CADE75A6BB303599A9FF8367DF9628.

Swierstra:2017:SIP

[SD17] Wouter Swierstra and Peter Dybjer. Special issue on programming with dependent types editorial. Journal of Functional Programming, 27:e15, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
special-issue-on-programming-with-dependent-types-editorial/
B64163F12FCF62513E2148A6F0E52EE0.

Skorstengaard:2021:SEW

[SDB21] Lau Skorstengaard, Dominique Devriese, and Lars Birkedal. Stk-Tokens: Enforcing well-bracketed control flow and stack encapsulation using linear capabilities. Journal of Functional Programming, 31(??):e9, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
stktokens-enforcing-wellbracketed-control-flow-and-stack-
encapsulation-using-linear-capabilities/10DE9942E61FC9D4B6B123560C5801AD.

Sperber:2009:RRA

[SDF^{+}09] Michael Sperber, R. Kent Dybvig, Matthew Flatt, Anton Van Straaten, Robby Findler, and Jacob Matthews. Revised 6 report on the algorithmic language Scheme. Journal of Functional Programming, 19(S1):1–301, August 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4896B66B864A3D40FFA508528F2A3FD7.

Shaikhha:2018:PVP

[SDK18] Amir Shaikhha, Mohammad Dashti, and Christoph Koch. Push versus pull-based loop fusion in query engines. Journal of Functional Programming, 28:e10, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
push-versus-pullbased-loop-fusion-in-query-engines/D67AE4899E87F4B5102F859B0FC02045.

Sulzmann:2007:UFD

[SDPJS07] Martin Sulzmann, Gregory J. Duck, Simon Peyton-Jones, and Peter J. Stuckey. Understanding functional dependencies via constraint handling rules. Journal of Functional Programming, 17(1): 83–129, January 2007. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/49E533CD7975431B5339456255DA9BE5.

Selinger:2002:LCA

[Sel02] Peter Selinger. The lambda calculus is algebraic. Journal of Functional Programming, 12(6):549–566, November 2002. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/6F42B3B30AD21B0E0E026BAFFFB79C8B.

Serrano:2000:BID

[Ser00] Manuel Serrano. Bee: an integrated development environment for the Scheme programming language. Journal of Functional Programming, 10(4):353–395, July 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/BED154F41D9CB0F01F98C08717EC287B.

Sestoft:1997:DLA

[Ses97] Peter Sestoft. Deriving a lazy abstract machine. Journal of Functional Programming, 7(3):231–264, May 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A1CF974BD4A92A2A9B47287F55B68DB6.

Shivers:2006:MRF

[SF06] Olin Shivers and David Fisher. Multi-return function call. Journal of Functional Programming, 16(4–5):547–582, July 2006. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/0AF24BFFCE5BE1B56F58D824AFD81A9D.

Singh:2014:SID

[SF14] Satnam Singh and Robert Bruce Findler. Special issue dedicated to ICFP 2012: Editorial. Journal of Functional Programming, 24(2– 3):131–132, May 2014. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/7CA735936D1B06869969CEA30A25BBB0.

Stump:2016:ELE

[SF16] Aaron Stump and Peng Fu. Efficiency of lambda-encodings in total type theory. Journal of Functional Programming, 26:e3, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/efficiency-of-
lambdaencodings-in-total-type-theory/61BB015E068EAC16C6C31D5F7654F3AD.

Sculthorpe:2014:KUR

[SFG14] Neil Sculthorpe, Nicolas Frisby, and Andy Gill. The Kansas University rewrite engine: A Haskell-embedded strategic programming language with custom closed universes. Journal of Functional Programming, 24(4):434–473, July 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/B188BD6CAC68A121C219FA5315AF0C21.

Seco:2017:CBL

[SFL17] João Costa Seco, Paulo Ferreira, and Hugo Lourenço. Capabilitybased localization of distributed and heterogeneous queries. Journal of Functional Programming, 27:e26, ???? 2017. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/capabilitybased-localization-of-distributed-
and-heterogeneous-queries/C9401FAD261D3785552F1038A59E980C.

Sorensen:1996:PS

[SGJ96] M. H. Sørensen, R. Glück, and N. D. Jones. A positive supercompiler. Journal of Functional Programming, 6(6):811–838, November 1996. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4EEE2EBC972AA2FDC861EF7A713EE898.

Smith:2003:CTB

[SGM^{+}03] Frederick Smith, Dan Grossman, Greg Morrisett, Luke Hornof, and Trevor Jim. Compiling for template-based run-time code generation. Journal of Functional Programming, 13(3):677–708, May 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
D903A3D6D54DBD0C84983335FA252E1C.

Siles:2012:PTS

[SH12] Vincent Siles and Hugo Herbelin. Pure type system conversion is always typable. Journal of Functional Programming, 22(2):153–180, March 2012. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0BFD4C10E4EBB7884E906982CD1B017F.

Sculthorpe:2014:WIW

[SH14] Neil Sculthorpe and Graham Hutton. Work it, wrap it, fix it, fold it. Journal of Functional Programming, 24(1):113–127, January 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
FF151CC992C1B18767B45B021461BD17.

Shah:2004:BRK

[Sha04a] Nimish Shah. Book review: Knowledge Representation, Reasoning and Declarative Problem Solving, by C. Baral, Cambridge University Press, 2003. Journal of Functional Programming, 14(5):588– 589, September 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/8A7204BC9994101D4DD3DA81CFC95B3E.

Shah:2004:BRP

[Sha04b] Nimish Shah. Book review: Program Construction: Calculating Implementations from Specifications, by R. C. Backhouse, John Wiley & Sons, 2004. Journal of Functional Programming, 14(5):598– 600, September 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0B56011F9052E7930B48915C56C91207.

Shah:2006:BRA

[Sha06] Nimish Shah. Book review: Advanced Topics in Term Rewriting, by Enno Ohlebusch, Springer Verlag, 2002. Journal of Functional Programming, 16(2):247–249, March 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/C1C10937545A6B3FF1D4D60BCC8C47CD.

Sheeran:2011:FDP

[She11] Mary Sheeran. Functional and dynamic programming in the design of parallel prefix networks. Journal of Functional Programming, 21(1):59–114, January 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/BF6065585BE32438CDF081A44741A5B8.

Scaife:2005:PSC

[SHMB05] Norman Scaife, Susumi Horiguchi, Greg Michaelson, and Paul Bristow. A parallel SML compiler based on algorithmic skeletons. Journal of Functional Programming, 15(4):615–650, July 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
A9964145FED1C19F4BBE55465488186A.

Sivaramakrishnan:2016:CSA

[SHMJ16] K. C. Sivaramakrishnan, Tim Harris, Simon Marlow, and Simon Peyton Jones. Composable scheduler activations for Haskell. Journal of Functional Programming, 26:e9, ???? 2016. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/composable-scheduler-activations-for-
haskell/5F4A86F27A77CAC76080B5F79667E619.

Simpson:2007:BRP

[Sim07] Alex Simpson. Book review: Programming Languages and Operational Semantics, by Maribel Fernández, King’s College Publications, 2004, ISBN 0-9543006-3-7. Journal of Functional Programming, 17(3):424–426, May 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/B29192BFFCB76F9C7BD0430579EB9317.

Saito:2008:LFP

[SIV08] Chieri Saito, Atsushi Igarashi, and Mirko Viroli. Lightweight family polymorphism. Journal of Functional Programming, 18(3):285–331, May 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
60B7A0962244CC5314A48B43A7B5B953.

Silbermann:1992:DTA

[SJ92] Frank S. K. Silbermann and Bharat Jayaraman. A domain-theoretic approach to functional and logic programming. Journal of Functional Programming, 2(3):273–321, July 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A57134D985A2C77052D4B796F290F675.

Smallbone:2017:QSB

[SJCA17] Nicholas Smallbone, Moa Johansson, Koen Claessen, and Maximilian Algehed. Quick specifications for the busy programmer. Journal of Functional Programming, 27:e18, ???? 2017. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/quick-specifications-for-
the-busy-programmer/1E1E8A11F91A927DD56AF08280A9F58D.

Seidel:2018:DWS

[SJW18] Eric L. Seidel, Ranjit Jhala, and Westley Weimer. Dynamic witnesses for static type errors (or, ill-typed programs usually go wrong). Journal of Functional Programming, 28:e13, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
dynamic-witnesses-for-static-type-errors-or-illtyped-programs-
usually-go-wrong/C9525689DAC80D153B94FAD292774444.

Sloane:2009:BRS

[Slo09] Anthony M. Sloane. Book review: Software Abstractions: Logic, Language, and Analysis, by Jackson Daniel, The MIT Press, 2006, 366 pp., ISBN 978-0-262-10114-1. Journal of Functional Programming, 19(2):253–254, March 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/0897369CCD962A22EC15B4E53DCBA3FA.

Sewell:2007:AHL

[SLW^{+}07] Peter Sewell, James J. Leifer, Keith Wansbrough, Francesco Zappa Nardelli, Mair Allen-Williams, Pierre Habouzit, and Viktor Vafeiadis. Acute: High-level programming language design for distributed computation. Journal of Functional Programming, 17(4–5): 547–612, July 2007. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/095AA356E4FBE1417C18C0B8338DD3F3.

Sperber:2013:CUF

[SM13] Michael Sperber and Anil Madhavapeddy. Commercial users of functional programming workshop report. Journal of Functional Programming, 23(6):701–712, November 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7B8E5D99E6C0D40D45B37D972B82598D.

Sison:2021:VSC

[SM21] Robert Sison and Toby Murray. Verified secure compilation for mixed-sensitivity concurrent programs. Journal of Functional Programming, 31(??):e18, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/verified-secure-compilation-for-mixedsensitivity-
concurrent-programs/A46AD098BDBBE50665D4094BE8F6D8DD.

Stefan:2017:FDI

[SMMR17] Deian Stefan, David Mazières, John C. Mitchell, and Alejandro Russo. Flexible dynamic information flow control in the presence of exceptions. Journal of Functional Programming, 27:e5, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
flexible-dynamic-information-flow-control-in-the-presence-
of-exceptions/DA92A6BA5FBBDC534F1B498217D160A8.

Stewart:2016:TFT

[SMT16] Robert Stewart, Patrick Maier, and Phil Trinder. Transparent fault tolerance for scalable functional computation. Journal of Functional Programming, 26:e5, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
transparent-fault-tolerance-for-scalable-functional-computation/
48B954EF9ADDC9F5244904061DF1123C.

Shivkumar:2021:RTM

[SMZ21] Bhargav Shivkumar, Jeffrey Murphy, and Lukasz Ziarek. Realtime MLton: a standard ML runtime for real-time functional programs. Journal of Functional Programming, 31(??):e19, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
realtime-mlton-a-standard-ml-runtime-for-realtime-functional-
programs/2C7EF6A583AB68662F4DF3BF12844A53.

Sewell:2010:OET

[SNO^{+}10] Peter Sewell, Francesco Zappa Nardelli, Scott Owens, Gilles Peskine, Thomas Ridge, Susmit Sarkar, and Rok Strnisa. Ott: Effective tool support for the working semanticist. Journal of Functional Programming, 20(1):71–122, January 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/4F02FB206B5209D42B3BF2ED2240A802.

Svensson:2016:LHD

[SNS16] Bo Joel Svensson, Ryan R. Newton, and Mary Sheeran. A language for hierarchical data parallel design-space exploration on GPUs. Journal of Functional Programming, 26:e6, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/language-for-
hierarchical-data-parallel-designspace-exploration-on-gpus/
C406E732CBFFD3AF80E3BECBBE7F8B7B.

Schrijvers:2019:CSC

[SOWM19] Tom Schrijvers, Bruno C. D. S. Oliveira, Philip Wadler, and Koar Marntirosian. COCHIS: Stable and coherent implicits. Journal of Functional Programming, 29:e3, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic).  URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/cochis-stable-
and-coherent-implicits/F0127B641DC6AFBA1488A0C64A25A289.

Sheard:2004:TLT

[SP04] Tim Sheard and Emir Pasalic. Two-level types and parameterized modules. Journal of Functional Programming, 14(5):547–587, September 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/3118E358DBB8D5AA599450E77151CAA5.

Sperber:2014:BRS

[Spe14] Michael Sperber. Book review: Systematic Program Design: From Clarity to Efficiency, by Yanhong Annie Liu, Cambridge University Press, 2013, ISBN: 978-1-107-03660-4. Journal of Functional Programming, 24(1):128–130, January 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/9072682693050A26B2720A6443FB3CCF.

Sperber:2020:RFR

[Spe20] Michael Sperber. Review of “Functional and Reactive Domain Modeling” by Debasish Ghosh, Manning, 2017, ISBN 978-1-617-29224-8. Journal of Functional Programming, 30(??):e21, ???? 2020. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/review-of-functional-and-
reactive-domain-modeling-by-debasish-ghosh-manning-2017-
isbn-9781617292248/EDC2C65147D376E3763D0177C74FE23F.

Sperber:2024:RJF

[Spe24] Michael Sperber. A review for the Journal of Functional Programming of Sandy Maguire, Algebra-Driven Design (Leanpub, 2020).

Journal of Functional Programming, 34:??, ???? 2024. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/review-for-the-journal-of-functional-
programming-of-sandy-maguire-algebradriven-design-leanpub-
2020/31F9331736F7EE664C017E5C340DD0D6.

Spivey:2000:CBF

[Spi00] Michael Spivey. Combinators for breadth-first search. Journal of Functional Programming, 10(4):397–408, July 2000. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/60383337C85657F3F6549C18F4E345BA.

Spivey:2009:ACS

[Spi09] J. Michael Spivey. Algebras for combinatorial search. Journal of Functional Programming, 19(3–4):469–487, July 2009. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/AB57FF99CEA76C1C31A336B560D6FD3C.

Spivey:2012:WMG

[Spi12] Michael Spivey. When Maybe is not good enough. Journal of Functional Programming, 22(6):747–756, November 2012. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/A6D1423D2D8B62105F04FDC012E7DDB9.

Streicher:1998:CLC

[SR98] Th. Streicher and B. Reus. Classical logic, continuation semantics and abstract machines. Journal of Functional Programming, 8(6):543–572, November 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/CCC69E5E233D89BE6378EB7BF50E3AB9.

Sulzmann:2008:HXT

[SS08] Martin Sulzmann and Peter J. Stuckey. HM(X) type inference is CLP(X) solving. Journal of Functional Programming, 18(2):251– 283, March 2008. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/E43F57EC1FF8C202A3F64C55C436930E.

Sewell:2008:DRM

[SSH^{+}08] Peter Sewell, Gareth Stoyle, Michael Hicks, Gavin Bierman, and Keith Wansbrough. Dynamic rebinding for marshalling and update, via redex-time and destruct-time reduction. Journal of Functional Programming, 18(4):437–502, July 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/146BCC9FFC7784B25BBED05CDAD3BE42.

Schmidt-Schauss:2008:SNS

[SSSS08] Manfred Schmidt-Schauss, David Sabel, and Marko Schütz. Safety of Nöcker’s strictness analysis. Journal of Functional Programming, 18 (4):503–551, July 2008. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/0B4B565CDCE0159FE6F8FBF801753FF3.

Swords:2018:EAC

[SSTH18] Cameron Swords, Amr Sabry, and Sam Tobin-Hochstadt. An extended account of contract monitoring strategies as patterns of communication. Journal of Functional Programming, 28: e4, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
an-extended-account-of-contract-monitoring-strategies-as-
patterns-of-communication/D410A73D44AE4F8126C44C558672F875.

Skalka:2008:TTE

[SSV08] Christian Skalka, Scott Smith, and David Van Horn. Types and trace effects of higher order programs. Journal of Functional Programming, 18(2):179–249, March 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/2A66C1E43EA859C82419830ED04C8E99.

Schrijvers:2009:MCP

[SSW09] Tom Schrijvers, Peter Stuckey, and Philip Wadler. Monadic constraint programming. Journal of Functional Programming, 19(6): 663–697, November 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/213541C635A4EEC9ADA73B345B731E9D.

Stark:1998:WCU

[Stä98] Robert F. Stärk. Why the constant ‘undefined’? logics of partial terms for strict and non-strict functional programming languages. Journal of Functional Programming, 8(2):97–129, March 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3351BE72FC143DEEC339C155D7910A0B.

Sterling:2021:HOF

[Ste21] Jonathan Sterling. Higher order functions and Brouwer’s thesis. Journal of Functional Programming, 31(??):e11, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic).  URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/higher-order-
functions-and-brouwers-thesis/938CB9864C34CA33A41F8202620918A3.

Sekiyama:2024:SRP

[STI24] Taro Sekiyama, Takeshi Tsukada, and Atsushi Igarashi. Signature restriction for polymorphic algebraic effects. Journal of Functional Programming, 34:??, ???? 2024. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
signature-restriction-for-polymorphic-algebraic-effects/
1FCD90F7590C031791DBE08DCD65CED5.

Stratford:2011:CLD

[Str11] Barney Stratford. A combinator library for the design of railway track layouts. Journal of Functional Programming, 21(3):309–329, May 2011. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4D4C3D3AFCF65BBA7B40BC49F3E08951.

Stump:2017:CDL

[Stu17] Aaron Stump. The calculus of dependent lambda eliminations. Journal of Functional Programming, 27:e14, ???? 2017. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/calculus-of-dependent-lambda-eliminations/
1D0BDA070E9273AC56C108D8F6F2B078.

Siek:2021:BCT

[STW21] Jeremy G. Siek, Peter Thiemann, and Philip Wadler. Blame and coercion: Together again for the first time. Journal of Functional Programming, 31(??):e20, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/blame-and-coercion-together-again-for-
the-first-time/657CEF36278FE66B039DEF1D06DF8178.

Suranyi:2007:OOC

[Sur07] Gábor M. Surányi. An object-oriented calculus with term constraints. Journal of Functional Programming, 17(3):353–386, May 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9D41BDA554D63884652B3E28E1B6405F.

Swierstra:2013:LPD

[SV13] Wouter Swierstra and Thomas Van Noort. A library for polymorphic dynamic typing. Journal of Functional Programming, 23(3): 229–248, May 2013. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/BF1DA1BFB6EABDB80D36104D70427E85.

Sergey:2017:MHO

[SVJB17] Ilya Sergey, Dimitrios Vytiniotis, Simon L. Peyton Jones, and Joachim Breitner. Modular, higher order cardinality analysis in theory and practice. Journal of Functional Programming, 27:e11, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
modular-higher-order-cardinality-analysis-in-theory-and-
practice/5D815BD54F43FD49146B2F4154565DE4.

Stansifer:2016:RSM

[SW16] Paul Stansifer and Mitchell Wand. Romeo: A system for more flexible binding-safe programming. Journal of Functional Programming, 26:e13, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
romeo-a-system-for-more-flexible-bindingsafe-programming/
F9C128D1D92B6BCD489E0FE41D255AE0.

Schultz:2017:ADI

[SW17] Patrick Schultz and Ryan Wisnesky. Algebraic data integration. Journal of Functional Programming, 27:e24, ???? 2017. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/algebraic-data-integration/AC36B1C39D4C76A18CB54F75B0D26796.

Sulzmann:2023:TDD

[SW23] Martin Sulzmann and Stefan Wehr. A type-directed, dictionarypassing translation of method overloading and structural subtyping in Featherweight Generic Go. Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
typedirected-dictionarypassing-translation-of-method-overloading-
and-structural-subtyping-in-featherweight-generic-go/993459B2AFE973B82F93D29C8A16C774.

Sarkar:2005:EPN

[SWD05] Dipanwita Sarkar, Oscar Waddell, and R. Kent Dybvig. Educational pearl: A Nanopass framework for compiler education. Journal of Functional Programming, 15(5):653–667, September 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
1E378B9B451270AF6A155FA0C21C04A3.

Swierstra:2008:DTC

[Swi08] Wouter Swierstra. Data types à la carte. Journal of Functional Programming, 18(4):423–436, July 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/14416CB20C4637164EA9F77097909409.

Swierstra:2011:SVP

[Swi11] Wouter Swierstra. Sorted: Verifying the problem of the Dutch National Flag in Agda. Journal of Functional Programming, 21(6):573– 583, November 2011. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/B94802514DC2786CCF323154AEB91953.

Swierstra:2020:HBR

[Swi20] Wouter Swierstra. Heterogeneous binary random-access lists. Journal of Functional Programming, 30(??):e10, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic).  URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/heterogeneous-
binary-randomaccess-lists/CC82B2E79DC5CCAD57E0AC5DF0D43DEC.

Swierstra:2022:WKR

[Swi22] Wouter Swierstra. A well-known representation of monoids and its application to the function ‘vector reverse’. Journal of Functional Programming, 32:e10, ???? 2022. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
wellknown-representation-of-monoids-and-its-application-
to-the-function-vector-reverse/929D15452762F6BB0A4A68503026D1C4.

Swierstra:2023:CCC

[Swi23] Wouter Swierstra. A correct-by-construction conversion from lambda calculus to combinatory logic. Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
correctbyconstruction-conversion-from-lambda-calculus-to-
combinatory-logic/B2A7BCF12A7A5D52B294C78A6F04191F.

Sussman:2001:SIC

[SWM01] Gerald Jay Sussman, Jack Wisdom, and Meinhard E. (Meinhard Edwin) Mayer. Structure and interpretation of classical mechanics. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-19455-4. xxi + 534 pp. LCCN QC125.2 .S895 2001.

Sivaramakrishnan:2014:MMA

[SZJ14] K. C. Sivaramakrishnan, Lukasz Ziarek, and Suresh Jagannathan. MultiMLton: A multicore-aware runtime for standard ML. Journal of Functional Programming, 24(6):613–674, November 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
AE2EE448CF8C677D64B47E657F8BBB9A.

Tolmach:1995:DSM

[TA95] Andrew Tolmach and Andrew W. Appel. A debugger for Standard ML. Journal of Functional Programming, 5(2):155–200, April 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
BF54A954B1F2E6447C9B404627F87EC4.

Taha:2003:BRE

[Tah03a] Walid Taha. Book review: Essentials of Programming Languages (2nd ed), by Daniel P. Friedman, Mitchell Wand and Christopher T. Haynes, MIT Press, ISBN 0-262-06217-8, 2001. Journal of Functional Programming, 13(4):829–831, July 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/C297177823411C9D82A860D4534B548C.

Taha:2003:SAI

[Tah03b] Walid Taha. ‘semantics, applications, and implementation of program generation’. Journal of Functional Programming, 13(3):453– 454, May 2003. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/13996AAD17C58EF97A28E2CA9424EC14.

Tanter:2020:RLP

[Tan20] Éric Tanter. Review of “The Little Prover” by Daniel P. Friedman and Carl Eastlund, MIT Press, 2015. Journal of Functional Programming, 30(??):e6, ???? 2020. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
review-of-the-little-prover-by-daniel-p-friedman-and-carl-
eastlund-mit-press-2015/F85BF92A2EBE5A0B46063D3E6EA9D457.

Tennent:1997:BRF

[Ten97] R. D. Tennent. Book review: Foundations for Programming Languages, by John C. Mitchell, MIT Press, 1996. Journal of Functional Programming, 7(6):667–668, November 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/53D17715F45FC137A4D4C2B2488A0209.

Tennent:2000:BRT

[Ten00] R. D. Tennent. Book review: Theories of Programming Languages, by John C. Reynolds, Cambridge University Press, 1998. Journal of Functional Programming, 10(5):501–508, September 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0E17EB1A328285CF70B73D613981D43C.

Thiemann:1999:CPG

[Thi99] Peter Thiemann. Combinators for program generation. Journal of Functional Programming, 9(5):483–525, September 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/7D7702570EFD837B154B23859577D701.

Thiemann:2002:TRH

[Thi02] Peter Thiemann. A typed representation for HTML and XML documents in Haskell. Journal of Functional Programming, 12(4–5): 435–468, July 2002. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/BF4BC3BA84763A9D4FB3007719BB09FE.

Thiemann:2003:PSE

[Thi03] Peter Thiemann. Program specialization for execution monitoring. Journal of Functional Programming, 13(3):573–600, May 2003. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
DC96972547AD702BC33BCC8D5989EA9C.

Thielecke:2009:CEM

[Thi09] Hayo Thielecke. Control effects as a modality. Journal of Functional Programming, 19(1):17–26, January 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/6B84CDBDD98C3D9A4F368B357ABA9FCE.

Trinder:1998:ASP

[THLJ98] P. W. Trinder, K. Hammond, H.-W. Loidl, and S. L. Peyton Jones. Algorithm + strategy = parallelism. Journal of Functional Programming, 8(1):23–60, January 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/24CE696A9E76AEA63C2D6132BE25FC09.

Thompson:1999:BRA

[Tho99] Simon Thompson. Book review: Advanced Functional Programming, edited by John Lauchbury, Erik Meijer and Tim Sheard, Lecture Notes in Computer Science 1129, Springer-Verlag, 1996. Journal of Functional Programming, 9(1):105–111, January 1999. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
596D4554B216475A6A957DD39E071924.

Thompson:2000:FRA

[Tho00] Simon Thompson. A functional reactive animation of a lift using fran. Journal of Functional Programming, 10(3):245–268, May 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C828ACD945F6DF45CE9B5DFDF0B67C76.

Thomas:2002:BRP

[Tho02] Mike Thomas. Book review: Principles of Programming Languages Design, Evaluation, and Implementation (3rd ed.), by Bruce J. MacLennan, Oxford University Press, 1999, ISBN 0-19-511306-3. Journal of Functional Programming, 12(1):89–90, January 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6891FE9D72DB2727931CBC75ABA8C591.

Thomas:2004:BRI

[Tho04a] Mike Thomas. Book review: An Introduction to Computing with Haskell, by Manuel M. T. Chakravarty and Gabrielle C. Keller, Pearson SprintPrint, 2002, ISBN 1-74009-404-2. Price $AU 39.95, Pp. 150. Journal of Functional Programming, 14(2):247–249, March 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
9889C1B2F32685F1A2525195061F623D.

Thomas:2004:BRR

[Tho04b] Mike Thomas. Book review: Reasoning About Program Transformations: Imperative Programming and Flow of Data, by Jean-François Collard, Springer-Verlag, 2003, ISBN 0-387-95391-4. Journal of Functional Programming, 14(5):589–592, September 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
71E492B42EE8F88447D088C4F9D223CE.

Thompson:2016:RSI

[Tho16] Simon Thompson. Review of Spreadsheet implementation technology: Basics and extensions, by Peter Sestoft, MIT Press, 2014, ISBN 978-0-262-52664-7. Journal of Functional Programming, 26:e17, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
review-of-spreadsheet-implementation-technology-basics-
and-extensions-by-peter-sestoft-mit-press-2014-isbn-9780262526647/
B1F6E62B5E20014B4BBC9BF5135DD0BF.

Tirronen:2015:TTC

[TI15] Ville Tirronen and Ville Isomöttönen. Teaching types with a cognitively effective worked example format. Journal of Functional Programming, 25:e23, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
teaching-types-with-a-cognitively-effective-worked-example-
format/D793285CAC3FDC1FAB3B1C846EEEFF54.

Talpin:1992:PTR

[TJ92] Jean-Pierre Talpin and Pierre Jouvelot. Polymorphic type, region and effect inference. Journal of Functional Programming, 2(3): 245–271, July 1992. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/CC658A8BA695A47E0113BE4DD3BBB4CB.

Tolmach:2011:SID

[TL11] Andrew Tolmach and Xavier Leroy. Special issue dedicated to ICFP 2009 editorial. Journal of Functional Programming, 21(4–5):331– 332, September 2011. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/B427F1324C994F63AACC5CE678D8F705.

Thompson:2013:RTF

[TL13] Simon Thompson and Huiqing Li. Refactoring tools for functional languages. Journal of Functional Programming, 23(3):293–350, May 2013. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F78282D0AE831BD11AD5531826892139.

Turcotte:1995:ECN

[TLM95] Marcel Turcotte, Guy Lapalme, and François Major. Exploring the conformations of nucleic acids. Journal of Functional Programming, 5(3):443–460, July 1995. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/32294B1EE1AD1DA167ACCF998E445355.

Trinder:2002:PDH

[TLP02] P. W. Trinder, H.-W. Loidl, and R. F. Pointon. Parallel and distributed Haskells. Journal of Functional Programming, 12(4–5): 469–510, July 2002. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/D24021D5389F8A8CBA96C15FBBB5EFC9.

Tan:2019:VCC

[TMK^{+}19] Yong Kiam Tan, Magnus O. Myreen, Ramana Kumar, Anthony Fox, Scott Owens, and Michael Norrish. The verified CakeML compiler backend. Journal of Functional Programming, 29:e2, ???? 2019. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic).

URL https://www.cambridge.org/core/journals/journal-of-
functional-programming/article/verified-cakeml-compiler-
backend/E43ED3EA740D2DF970067F4E2BB9EF7D.

Thiemann:2010:SID

[TN10] Peter Thiemann and Henrik Nilsson. Special issue dedicated to ICFP 2008 editorial. Journal of Functional Programming, 20(5–6):415– 416, November 2010. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/8855148164EBA43E91523E339067EB40.

Traytel:2015:VDP

[TN15] Dmitriy Traytel and Tobias Nipkow. Verified decision procedures for MSO on words based on derivatives of regular expressions. Journal of Functional Programming, 25:e18, ???? 2015. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/journals/journal-of-functional-
programming/article/verified-decision-procedures-for-mso-
on-words-based-on-derivatives-of-regular-expressions/18DCED718D5D525252C97EFA3501B4A4.

Tolmach:1998:MAS

[TO98] Andrew Tolmach and Dino P. Oliva. From ML to Ada: Stronglytyped language interoperability via source translation. Journal of Functional Programming, 8(4):367–412, July 1998. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/9ACB018A7F71ECF8A0FEE0DFEDAC4E84.

Tofte:1994:PSH

[Tof94] Mads Tofte. Principal signatures for higher-order program modules. Journal of Functional Programming, 4(3):285–335, July 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
FF20CCEF92F3E5D41183EDC2F68FDCDB.

Trinder:2005:SIH

[Tri05] P. W. Trinder. Special issue high performance parallel functional programming. Journal of Functional Programming, 15(3):351–352, May 2005. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
B3FB7A343B35AD624A2A707032FB9241.

Tronci:1995:DDS

[Tro95] Enrico Tronci. Defining data structures via Böhm-out. Journal of Functional Programming, 5(1):51–64, January 1995. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/F224D11499AB9872806C8D3F0467995B.

Tirronen:2015:UBM

[TUmI15] Ville Tirronen, Samuel Uusi-mäkelä, and Ville Isomöttönen. Understanding beginners’ mistakes with Haskell. Journal of Functional Programming, 25:e11, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
understanding-beginners-mistakes-with-haskell/244DB6807F3BD77E14CE7D627514D6D3.

Turchin:1993:PTM

[Tur93] Valentin F. Turchin. Program transformation with metasystem transitions. Journal of Functional Programming, 3(3):283–313, July 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
F0D67B654FB2C053A961551C357B92E9.

Thompson:1993:FPE

[TW93] Simon Thompson and Philip Wadler. Functional programming in education — introduction. Journal of Functional Programming, 3 (1):3–4, January 1993. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/820812F91668A599611A4F906BA8F5DB.

Taha:2000:SIS

[TW00] Walid Taha and Peter Wadler. Special issue on semantics, applications, and implementation of program generation. Journal of Functional Programming, 10(6):627, November 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/A70CD48E1D425AE54F0DD70295E903DE.

Ullman:1994:EMP

[Ull94] Jeffrey D. (Jeffrey David) Ullman. Elements of ML programming. Prentice Hall International Editions. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-288788-6 (paperback), 0-13-184854-2 (hardcover). xiii + 320 pp. LCCN QA76.73.M6 U45 1994. URL http://books.google.com/

books?id=XalQAAAAMAAJ;   http://catalog.hathitrust.org/
api/volumes/oclc/28844308.html;   http://www.gbv.de/dms/
bowker/toc/9780131848542.pdf;   http://www.gbv.de/dms/bowker/
toc/9780132887885.pdf;   http://www.zentralblattmath.org/
zmath/en/search/?an=0821.68025.

Uustalu:2017:FRS

[UV17] Tarmo Uustalu and Niccolò Veltri. Finiteness and rational sequences, constructively. Journal of Functional Programming, 27:e13, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
finiteness-and-rational-sequences-constructively/FCC024B602F850DDCEB2E9092E72369B.

Vasconcelos:2005:LPC

[Vas05] Vasco Thudichum Vasconcelos. Lambda and pi calculi, CAM and SECD machines. Journal of Functional Programming, 15(1):101– 127, January 2005. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/29F2716CAC0EFCEEC02A3DF9D1FD2250.

Verbruggen:2010:FPP

[VDH10] Wendy Verbruggen, Edsko De Vries, and Arthur Hughes. Formal polytypic programs and proofs. Journal of Functional Programming, 20(3–4):213–270, July 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/AFAB9C723CCC9D4118727BC0C8CEDF63.

Vree:1995:CLF

[VH95] Willem G. Vree and Pieter H. Hartel. Communication lifting: fixed point computation for parallelism. Journal of Functional Programming, 5(4):549–581, October 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/ADFDDE42A79B551BAE75E34CDBEC3AE0.

VanRoy:2004:CTM

[VH04] Peter Van Roy and Seif Haridi. Concepts, techniques, and models of computer programming. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-22069-5. xxvii + 900 pp. LCCN QA76.6 .V36 2004.

Voigtlander:2013:ESB

[VHMW13] Janis Voigtländer, Zhenjiang Hu, Kazutaka Matsuda, and Meng Wang. Enhancing semantic bidirectionalization via shape bidirectionalizer plug-ins. Journal of Functional Programming, 23(5):515– 551, September 2013. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/C8569C350E608F454B02EA7A4EE42E82.

Villanueva:2007:BRV

[Vil07] Alicia Villanueva. Book review: Verification of Reactive Systems, by Schneider, Klaus, Springer-Verlag, 2003, 600 pp., ISBN 3-540-00296-0. Journal of Functional Programming, 17(6):813–814, November 2007. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
E12CCF75D145F46FC6A8AD0EB836880B.

Voigtlander:2004:CFA

[VK04] Janis Voigtländer and Armin Kühnemann. Composition of functions with accumulating parameters. Journal of Functional Programming, 14(3):317–363, May 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/D79D8E3BBE66826BEE0B1DDB14FED8B5.

VanHorn:2012:SAA

[VM12] David Van Horn and Matthew Might. Systematic abstraction of abstract machines. Journal of Functional Programming, 22(4–5):705– 746, September 2012. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/8EF3C3BB59F64000F40CDB5F6CCD48DD.

Vezzosi:2021:CAD

[VMA21] Andrea Vezzosi, Anders MÖRtberg, and Andreas Abel. Cubical Agda: a dependently typed programming language with univalence and higher inductive types. Journal of Functional Programming, 31(??):e8, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
cubical-agda-a-dependently-typed-programming-language-with-
univalence-and-higher-inductive-types/839F14B5227969B039D7B57AA8272C6B.

Strydonck:2021:LCF

[VPD21] Thomas Van Strydonck, Frank Piessens, and Dominique Devriese. Linear capabilities for fully abstract compilation of separation-logic-verified code. Journal of Functional Programming, 31(??):e6, ???? 2021. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
linear-capabilities-for-fully-abstract-compilation-of-separationlogicverified-
code/EA316E0FBA47C39AA18EA4D85A0E3CD8.

Vytiniotis:2011:OXM

[VPSS11] Dimitrios Vytiniotis, Simon Peyton Jones, Tom Schrijvers, and Martin Sulzmann. OutsideIn(X) modular type inference with local assumptions. Journal of Functional Programming, 21(4–5):333– 412, September 2011. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/65110D74CF75563F91F9C68010604329.

vonSydow:2000:BRH

[vS00] Björn von Sydow. Book review: The Haskell School of Expression: Learning Functional Programming Through Multimedia, by Paul Hudak, Cambridge University Press, 2000, 363 pp., ISBN 0-521-64408-9. Journal of Functional Programming, 10(5):501–508, September 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6215F433B7EF37CCC59F36B888AB9BDA.

Vytiniotis:2010:PTE

[VW10] Dimitrios Vytiniotis and Stephanie Weirich. Parametricity, type equality, and higher-order polymorphism. Journal of Functional Programming, 20(2):175–210, March 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/94D816825934AEAC3E556D5CA0614B83.

VanNoort:2010:LAD

[VYH^{+}10] Thomas Van Noort, Alexey Rodriguez Yakushev, Stefan Holdermans, Johan Jeuring, Bastiaan Heeren, and José Pedro Magalhães. A lightweight approach to datatype-generic rewriting. Journal of Functional Programming, 20(3–4):375–413, July 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
AA248FC84F0A459CE7CF073FEA985C59.

Wadler:1994:TSO

[Wad94] Philip Wadler. Type systems for object-oriented programming. Journal of Functional Programming, 4(2):125, April 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
4E7FE0F6F61ADF67DE151A40A450CD37.

Wadler:1997:EHO

[Wad97] Philip Wadler. EDITORIAL: A HOT opportunity. Journal of Functional Programming, 7(2):127–128, March 1997. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/7C024361430875DD5C5C29B551D5E18D.

Wadler:2014:PS

[Wad14] Philip Wadler. Propositions as sessions. Journal of Functional Programming, 24(2–3):384–418, May 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/0985539E5D607AC00FB00FF900BA1C86.

Wakeling:1998:DCL

[Wak98] David Wakeling. The dynamic compilation of lazy functional programs. Journal of Functional Programming, 8(1):61–81, January 1998. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
3CBEE8F4DB9C6F3446FA3E33AB9951D1.

Wakeling:1999:CLF

[Wak99] David Wakeling. Compiling lazy functional programs for the Java virtual machine. Journal of Functional Programming, 9(6):579– 603, November 1999. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/AC59FAFC68063D9EC71E4FCF283E7EFC.

Wakeling:2006:EPB

[Wak06] David Wakeling. Educational pearl: Biological sequence similarity. Journal of Functional Programming, 16(1):1–12, January 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
581652C1964C346737F43D0BE46295B5.

Wakeling:2007:SFP

[Wak07] David Wakeling. Spreadsheet functional programming. Journal of Functional Programming, 17(1):131–143, January 2007. CODEN JF-PRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

//www.cambridge.org/core/product/ADAC2F790AA176B3CE72E40412C63F06.

Wallingford:2004:BRP

[Wal04] Eugene Wallingford. Book review: Practical Aspects of Declarative Languages, by Shriram Krishnamurthi and C. R. Ramakrishnan (eds.), LNCS 2257, Springer-Verlag, 2002. ISBN 3-540-43092-X Price £31.50, pp. 359. Journal of Functional Programming, 14(2): 249–251, March 2004. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/2DD6F82F2720F922BE9C81BE10BF5274.

Wampler:2011:FPJ

[Wam11] Dean Wampler. Functional programming for Java developers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 1-4493-1265-9, 1-4493-1103-2. xi + 72 pp. LCCN QA76.62 .W36 2011. URL http://proquest.

safaribooksonline.com/9781449312657.

Wand:1993:SCB

[Wan93] Mitchell Wand. Specifying the correctness of binding-time analysis. Journal of Functional Programming, 3(3):365–387, July 1993. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
0B598D301F7742EE5F3159340ADD58EE.

Wand:1994:I

[Wan94] Mitchell Wand. Introduction. Journal of Functional Programming, 4 (2):125–126, April 1994. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/DB5982C1FC7A275391C526B396D61461.

Watt:2014:BRP

[Wat14] David Watt. Book review: Practical Foundations for Programming Languages, by Robert Harper, Cambridge University Press, New York, December 2012, English, ISBN-10: 1-107-02957-0, Hardcover, 487 pp. Journal of Functional Programming, 24(4):523, July 2014. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
48A4960F719F454E396AEB8C188461FE.

Wazny:2010:ERF

[Waz10] Jeremy Wazny. Experience report: Functional programming in C-Rules. Journal of Functional Programming, 20(2):123–136, March 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
7504B9AB7F18AB7E026DFCDC30486DB8.

Wand:2001:SCD

[WC01] Mitchell Wand and William D. Clinger. Set constraints for destructive array update optimization. Journal of Functional Programming, 11(3):319–346, May 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/7FB3D17C44EC7B1C98EE99CD6825A303.

Walukiewicz-Chrzaszcz:2003:TRC

[WC03] Daria Walukiewicz-Chrzaszcz. Termination of rewriting in the calculus of constructions. Journal of Functional Programming, 13(2): 339–414, March 2003. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/A9C59D53BDF36F535AB3DB938140C0D3.

Wells:2002:CPP

[WDMT02] J. B. Wells, Allyn Dimock, Robert Muller, and Franklyn Turbak. A calculus with polymorphic and polyvariant flow types. Journal of Functional Programming, 12(3):183–227, May 2002. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
5A4E09535BC9D305098BB062655407AB.

Walkingshaw:2009:DSL

[WE09] Eric Walkingshaw and Martin Erwig. A domain-specific language for experimental game theory. Journal of Functional Programming, 19(6):645–661, November 2009. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/7D23DA54801AFA0FD239DEF3B0C05E26.

Weirich:2004:FPT

[Wei04] Stephanie Weirich. Functional pearl: Type-safe cast. Journal of Functional Programming, 14(6):681–695, November 2004. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL

https://www.cambridge.org/core/product/5C09BBECAB4721CE5E20DCCF2961783E.

Weirich:2006:TSR

[Wei06] Stephanie Weirich. Type-safe run-time polytypic programming. Journal of Functional Programming, 16(6):681–710, November 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
6461794C150F8570D90E67ED67324308.

Whitington:2013:OVB

[Whi13] John Whitington. OCaml from the very beginning. Coherent Press, Cambridge, UK, 2013. ISBN 0-9576711-0-5 (paperback). ix + 194 pp. LCCN QA76.73.O12 W5 2013.

Wong:2000:KFQ

[Won00] Limsoon Wong. Kleisli, a functional query system. Journal of Functional Programming, 10(1):19–56, January 2000. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/BDB5049BBB1F2263B2431C5DAF951DC3.

Wood:2001:BRS

[Woo01] Alan Wood. Book review: Structure and Interpretation of Computer Programs, 2nd Ed by Abelson and Sussman, with Sussman, MIT Press, 1996, ISBN 0-262-51087-1, 657 pp. Journal of Functional Programming, 11(2):253–262, March 2001. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/DF5CF9995BD04B6916A407D09EDAB6F2.

Wainwright:1992:SSM

[WS92] Roger L. Wainwright and Marian E. Sexton. A study of sparse matrix representations for solving linear systems in a functional language. Journal of Functional Programming, 2(1):61–72, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
C4E6B889BD52AFE1F5CE4F469FCCD973.

Washburn:2008:BGB

[WW08] Geoffrey Washburn and Stephanie Weirich. Boxes go bananas: Encoding higher-order abstract syntax with parametric polymorphism. Journal of Functional Programming, 18(1):87–140, January 2008. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
06272FA70BF929DF103CF92700200CB7.

Xi:2007:DMA

[Xi07] Hongwei Xi. Dependent ML an approach to practical programming with dependent types. Journal of Functional Programming, 17(2): 215–286, March 2007. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/4A1FC643ACD49EF31DAF5EB955D2CCC7.

Yi:2006:EPP

[Yi06] Kwangkeun Yi. Educational pearl: ‘proof-directed debugging’ revisited for a first-order version. Journal of Functional Programming, 16(6):663–670, November 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/product/F7CC0A759398A52C35F21F13236C0E00.

Ziliani:2015:MMT

[ZDK^{+}15] Beta Ziliani, Derek Dreyer, Neelakantan R. Krishnaswami, Aleksandar Nanevski, and Viktor Vafeiadis. Mtac: A monad for typed tactic programming in Coq. Journal of Functional Programming, 25:e12, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/mtac-a-monad-
for-typed-tactic-programming-in-coq/75B49F20037D8A0F718EAB21C662ABA0.

Zeilberger:2016:LLT

[Zei16] Noam Zeilberger. Linear lambda terms as invariants of rooted trivalent maps. Journal of Functional Programming, 26:e21, ???? 2016. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/journals/
journal-of-functional-programming/article/linear-lambda-
terms-as-invariants-of-rooted-trivalent-maps/12B63B0C31472A26E4D1173016D233C7.

Zhu:1994:HPF

[Zhu94] Hong Zhu. How powerful are folding/unfolding transformations? Journal of Functional Programming, 4(1):89–112, January 1994. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
851C3EA436ACA3556B782F2E9F499D92.

Ziarek:2010:LCC

[ZJ10] Lukasz Ziarek and Suresh Jagannathan. Lightweight checkpointing for concurrent ML. Journal of Functional Programming, 20(2): 137–173, March 2010. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic). URL https://www.cambridge.org/core/
product/A8CBF8766727B44869F7C7C5D01B9EC5.

Zakian:2018:GTS

[ZMCN18] Timothy A. K. Zakian, Trevor L. Mcdonell, Matteo Cimini, and Ryan R. Newton. Ghostbuster: A tool for simplifying and converting GADTs. Journal of Functional Programming, 28:e16, ???? 2018. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
ghostbuster-a-tool-for-simplifying-and-converting-gadts/
D8FEC9439E1D550CC05CFEC7AD5F1E1D.

Zhao:2006:TBC

[ZPV06] Tian Zhao, Jens Palsberg, and Jan Vitek. Type-based confinement. Journal of Functional Programming, 16(1):83–128, January 2006. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653

(electronic). URL https://www.cambridge.org/core/product/
67728DD6A4F34364B042ED17BB7296AD.

Ziliani:2017:CGN

[ZS17] Beta Ziliani and Matthieu Sozeau. A comprehensible guide to a new unifier for CIC including universe polymorphism and overloading. Journal of Functional Programming, 27:e10, ???? 2017. CODEN JFPRES. ISSN 0956-7968 (print),

1469-7653 (electronic).  URL https://www.cambridge.org/
core/journals/journal-of-functional-programming/article/
comprehensible-guide-to-a-new-unifier-for-cic-including-
universe-polymorphism-and-overloading/19A095CA0645F89A772B7E2B7B3D92B2.

Ziff:1995:FFS

[ZSW95] Donald A. Ziff, Stephen P. Spackman, and Keith Waclena. Funser: a functional server for textual information retrieval. Journal of Functional Programming, 5(3):317–343, July 1995. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.

cambridge.org/core/product/BEA93BBFBCEE3DF98C778D74691CF9D3.

Zhang:2023:CLR

[ZTX^{+}23] Hanliang Zhang, Wenhao Tang, Ruifeng Xie, Meng Wang, and Zhenjiang Hu. Contract lenses: Reasoning about bidirectional programs via calculation. Journal of Functional Programming, 33:??, ???? 2023. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https://www.cambridge.org/

core/journals/journal-of-functional-programming/article/
contract-lenses-reasoning-about-bidirectional-programs-
via-calculation/43F612938DAA399A9D35193FB6278F56.